### 11 Mar 2026 v1

# Primitive-Root Determinant Densities over Prime Fields and Implications for PRIM-LWE

Abstract
The prim-lwe problem, introduced by Sehrawat, Yeo, and Desmedt (Theoretical Computer Science, 886 (2021)), is a variant of the Learning with Errors problem in which the secret matrix is required to have a primitive-root determinant. The dimension-uniform reduction constant is $c(p)=\inf_{m\geq 1}c_{m}(p)$, where $c_{n}(p)$ is the exact density of $n\times n$ matrices over $\mathbb{F}_{p}$ with primitive-root determinant. Sehrawat, Yeo, and Desmedt asked whether $\inf_{p\text{ prime}}c(p)=0$, observing that an affirmative answer would follow from the conjectural infinitude of primorial primes.
We resolve this question unconditionally using only Dirichlet’s theorem and Mertens’ product formula, entirely bypassing the primorial-prime hypothesis. We further establish the sharp order
$\min_{p\leq x}c(p)\asymp\frac{1}{\log\log x}\qquad(x\to\infty),$
prove that $c(p)$ possesses a continuous purely singular limiting distribution over the primes with support exactly $[0,1/2]$, and derive explicit lower bounds on $c(q)$ for primes of cryptographic interest, parameterized solely by the number $\omega(q-1)$ of distinct prime factors of q1q-1. These bounds apply to every prime $q$ whose predecessor q1q-1 has controlled factorization structure, as measured by the number $\omega(q-1)$ of distinct prime factors; this includes many common NTT-friendly moduli, though NTT-friendliness alone does not imply the needed bound. For the NIST-standardized moduli $q=3329$ and $q=8380417$, the dimension-uniform expected rejection-sampling overhead $1/c(q)$ in the prim-lwe reduction is at most $2.17$ and $3.42$, respectively. As a simple conservative explicit bound, for any prime $q>2^{30}$ the dimension-uniform expected rejection-sampling overhead $1/c(q)$ is at most $1.79\log q$. On the other hand, our results show that the worst-case overhead among primes $p\leq x$ is of order $\Theta(\log\log x)$, and in particular $1/c(q)=O(\log\log q)$ pointwise.
1 Introduction
Let $p$ be a prime and $M_{n}(\mathbb{F}_{p})$ denote the ring of $n\times n$ matrices over $\mathbb{F}_{p}$. Following [1], a matrix $M\in M_{n}(\mathbb{F}_{p})$ has primitive-root determinant if $\det(M)$ generates $\mathbb{F}_{p}^{*}$. The fraction of such matrices is
$c_{n}(p)\;=\;\frac{\varphi(p-1)}{p-1}\;\prod_{j=1}^{n}\Bigl(1-\frac{1}{p^{j}}\Bigr),$
and since $c_{m}(p)$ is decreasing in $m$, the dimension-uniform constant governing the reduction from decision-LWE to decision-prim-lwe is
$c(p)\;:=\;\inf_{m\geq 1}c_{m}(p)\;=\;\frac{\varphi(p-1)}{p-1}\;\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{p^{j}}\Bigr).$
Sehrawat, Yeo, and Desmedt [1, Remark 3(ii)] observed that $\inf_{p}c(p)=0$ would follow from the conjectural infinitude of primorial primes and asked whether this can be proved unconditionally.
Main results
We resolve this question and establish several natural refinements.
1. Unconditional vanishing (Theorem 3). We prove $\inf_{p\text{ prime}}c(p)=0$ using only Dirichlet’s theorem on primes in arithmetic progressions and Mertens’ third theorem. The key observation is that for each primorial $N_{k}=\prod_{i=1}^{k}p_{i}$, Dirichlet’s theorem supplies infinitely many primes $p\equiv 1\pmod{N_{k}}$, and for any such prime the totient ratio $\varphi(p-1)/(p-1)$ is bounded by the Mertens product $\prod_{i=1}^{k}(1-1/p_{i})\sim e^{-\gamma}/\!\log p_{k}\to 0$. This bypasses the primorial-prime hypothesis entirely.
2. Sharp order of the minimum (Proposition 1, Corollary 6). By additionally invoking Linnik’s theorem on the least prime in an arithmetic progression, together with standard prime-number-theorem consequences for primorials, we establish
$\min_{p\leq x}c(p)\;\asymp\;\frac{1}{\log\log x}\qquad(x\to\infty).$
The matching lower bound follows from two observations: if $p\leq x$ is an odd prime, then $\omega(p-1)\leq m(x):=\max\{k:N_{k}\leq x\}$, and Mertens’ theorem converts this into a $1/\!\log\log x$ lower bound for $\varphi(p-1)/(p-1)$; the additional factor $\prod_{j\geq 1}(1-p^{-j})\geq P_{3}$ is handled by Theorem 8. Hence the decay to zero is completely quantified and extremely slow. To put this rate in perspective, $\log\log(2^{128})\approx 4.5$ and $\log\log(2^{4096})\approx 7.95$, so the asymptotic lower bound $\min_{p\leq x}c(p)\gg 1/\!\log\log x$ already guarantees that the minimum cannot drop faster than the reciprocal of a slowly growing function. Nonetheless, the implied constant is modest (roughly $P_{3}e^{-\gamma}\approx 0.31$), so primes $p$ whose predecessor p1p-1 is divisible by many small primes can still have $c(p)<0.1$ well within practical ranges. The concrete impact on prim-lwe therefore depends on the factorization structure of the chosen modulus; the explicit bounds in Section 6 show that standard cryptographic moduli with controlled factorization incur only a small constant-factor overhead. These implied constants should in principle be capable of being made explicit by combining explicit estimates for the Mertens product and for $\vartheta(x)$ with an explicit least-prime-in-an-arithmetic-progression bound; see Remark 3. We do not pursue that bookkeeping here.
3. Limiting distribution (Theorems 6 and 7, Corollary 2). The quantity $c(p)$ possesses a continuous purely singular limiting distribution over the primes, with support exactly $0,\tfrac{1}{2}]$. If $G$ denotes the distribution function of this law, then
$0<G(\tau)<1\qquad(0<\tau<1/2),$
the set of primes with $c(p)\leq\tau$ has relative density $G(\tau)$, and $G(\tau)\to 0$ as $\tau\downarrow 0$.
The existence and continuity of $G$ follow from the shifted-prime Erdős–Wintner framework for additive functions evaluated at p1p-1; for the relevant classical background see Tenenbaum [2] and Deshouillers–Hassani [3]. The full-support conclusion on $0,\tfrac{1}{2}]$, which to our knowledge has not appeared explicitly in the literature, combines the left non-differentiability result of [3, Theorem 1.1] with the density of $\{\varphi(m)/m:m\text{ even}\}$ in $[0,1/2]$ (Lemma 3). Our novel contribution is the transfer of this classical law to $c(p)$ via a transport lemma (Lemma 4), which establishes that the dimension-uniform prim-lwe reduction constant shares the same limiting distribution.
4. Explicit bounds for cryptographic moduli (Theorem 8, Table 1). We give unconditional lower bounds on $c(q)$ for primes $q$ parameterized by $\omega(q-1)$, the number of distinct prime factors of q1q-1, yielding concrete bounds for the NIST-standardized primes and other moduli whose predecessors have controlled factorization structure. The following consequence singles out the cryptographic implication.
Corollary 1 (Cryptographic consequence).
Let
$P_{3}:=\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{3^{j}}\Bigr)\approx 0.5601260779.$
Let $q\geq 3$ be a prime modulus with $\omega(q-1)\leq K$. Then, in the reduction from decision-LWE to decision-prim-lwe over $\mathbb{F}_{q}$, the dimension-uniform expected rejection-sampling overhead $1/c(q)$ satisfies
$\frac{1}{c(q)}\;\leq\;\frac{1}{P_{3}}\prod_{i=1}^{K}\frac{p_{i}}{p_{i}-1}.$
For each fixed dimension $n$, the exact expected overhead is $1/c_{n}(q)\leq 1/c(q)$.
The general $\omega(q-1)$-based bound is immediate from Theorem 8 and gives factors at most $5.36$ and $7.82$ for the NIST-standardized moduli $q=3329$ used in ML-KEM [4] and $q=8380417$ used in ML-DSA [5], respectively. The sharper NIST figures use the exact factorizations recorded in Section 6.
Remark 1.
The constant $P_{3}$ may also be written in standard $q$-series notation as
$P_{3}=\left(\frac{1}{3};\frac{1}{3}\right)_{\infty},\qquad(a;q)_{\infty}:=\prod_{m=0}^{\infty}(1-aq^{m}).$
Related work
A related distribution problem was studied by Aivazidis and Sofos [6], who considered the density of maximal-order elements in $\mathrm{GL}_{n}(q)$ as $q$ varies over prime powers and proved corresponding distribution laws. Their setting is different from ours: we work over prime fields and with the primitive-root determinant density arising in prim-lwe. On the distributional side, Deshouillers and Hassani [3, Proposition 2.1 and equation (2.1)] proved limiting-distribution laws for $\varphi(p-1)/(p-1)$ in the residue classes $p\equiv 1\pmod{m}$, including continuity, and [3, Theorem 1.1] established left non-differentiability of the distribution function at every value $\varphi(m)/m$ with $m$ even. The analogous support question for the classical distribution of $\varphi(n)/n$ over the integers remains open; see [3].
Organization
Section 2 records the precise definitions and the open question from [1]. Section 3 gives the unconditional proof that $\inf_{p}c(p)=0$. Section 4 proves the upper bound $\min_{p\leq x}c(p)\ll 1/\log\log x$. Section 5 develops the limiting distribution theory. Section 6 derives explicit lower bounds for primes of cryptographic interest and, in particular, supplies the matching lower bound $\min_{p\leq x}c(p)\gg 1/\log\log x$. Section 7 discusses the implications for prim-lwe.
Notation
We write $\omega(n)$ for the number of distinct prime factors of an integer $n$, $\pi(x)$ for the number of primes up to $x$, and adopt the convention that $p_{1}=2<p_{2}=3<p_{3}=5<\cdots$ denotes the sequence of all primes in increasing order. All limits are taken as $x\to\infty$ unless stated otherwise. We use the standard Vinogradov notation $A\ll B$ and $A\gg B$ synonymously with $A=O(B)$ and $B=O(A)$, respectively. $A\asymp B$ means both $A\ll B$ and $A\gg B$. We write $\operatorname{Li}(x)=\int_{2}^{x}\mathrm{d}t/\!\log t$ for the logarithmic integral, and throughout $\log$ denotes the natural logarithm. We use the notation $A=O^{*}(B)$ to mean $|A|\leq B$, i.e. an implied constant of $1$.
2 Background and the Open Question
We now record the precise setup from [1] and state the open question. Following [1], define
$M_{n}^{\mathrm{prim}}(\mathbb{F}_{p})\;=\;\bigl\{\,M\in M_{n}(\mathbb{F}_{p}):\det(M)\text{ generates }\mathbb{F}_{p}^{*}\,\bigr\}.$
The determinant map
$\det:\mathrm{GL}_{n}(\mathbb{F}_{p})\longrightarrow\mathbb{F}_{p}^{\times}$
is a surjective homomorphism, so all its fibres have the same cardinality. Therefore a fraction $\varphi(p-1)/(p-1)$ of invertible matrices have primitive-root determinant. Since
$|\mathrm{GL}_{n}(\mathbb{F}_{p})|=p^{n^{2}}\prod_{j=1}^{n}\Bigl(1-\frac{1}{p^{j}}\Bigr),$
it follows that
$c_{n}(p)\;:=\;\frac{|M_{n}^{\mathrm{prim}}(\mathbb{F}_{p})|}{|M_{n}(\mathbb{F}_{p})|}\;=\;\frac{\varphi(p-1)}{p-1}\;\prod_{j=1}^{n}\Bigl(1-\frac{1}{p^{j}}\Bigr),$
recovering [1, Corollary 4]. Consequently, the limit
$c(p)\;:=\;\lim_{n\to\infty}c_{n}(p)\;=\;\frac{\varphi(p-1)}{p-1}\;\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{p^{j}}\Bigr)$
exists and is strictly positive for each fixed $p$. The quantity $c_{n}(p)$ is the exact acceptance probability in dimension $n$. Since $c_{m}(p)$ is decreasing in $m$, the dimension-uniform constant that enters the prim-lwe reduction is
$\inf_{m\geq 1}c_{m}(p)=\lim_{m\to\infty}c_{m}(p)=c(p).$
Thus $c(p)$ is not the exact finite-dimensional density but the limiting, dimension-uniform reduction constant.
Recall that one may equivalently bundle $n$ standard LWE secret vectors into a matrix and view decision-LWE in matrix form: given $\mathbf{A}\in\mathbb{Z}_{q}^{w\times n}$, one must distinguish
$(\mathbf{A},\,\mathbf{A}\mathbf{S}+\mathbf{E})$
from uniform over $\mathbb{Z}_{q}^{w\times n}\times\mathbb{Z}_{q}^{w\times n}$, where $\mathbf{S}\in\mathbb{Z}_{q}^{n\times n}$ is the secret matrix and $\mathbf{E}\in\mathbb{Z}_{q}^{w\times n}$ is the error matrix [7]. Via the identification $\mathbb{Z}_{q}\cong\mathbb{F}_{q}$ for prime $q$, decision-prim-lwe [1] is the variant in which the secret matrix is additionally required to lie in $M_{n}^{\mathrm{prim}}(\mathbb{F}_{q})$.
In [1, Remark 3(ii)], Sehrawat, Yeo, and Desmedt observed that if there exist infinitely many primorial primes
$p=\prod_{i=1}^{k}p_{i}+1,$
then
$\frac{\varphi(p-1)}{p-1}=\prod_{i=1}^{k}\frac{p_{i}-1}{p_{i}}\longrightarrow 0$
by Mertens’ theorem, which would imply $\inf_{p}c(p)=0$. They left open whether this can be proved unconditionally.
3 Unconditional Proof that $\inf_{p}c(p)=0$
The argument requires only two classical results.
Theorem 1 (Dirichlet [8], 1837).
For coprime integers $a$ and $q$, there are infinitely many primes $p\equiv a\pmod{q}$.
Theorem 2 (Mertens’ Third Theorem [9], 1874).
As $x\to\infty$,
$\prod_{\begin{subarray}{c}q\leq x\\ q\text{ prime}\end{subarray}}\Bigl(1-\frac{1}{q}\Bigr)\;\sim\;\frac{e^{-\gamma}}{\log x},$
where $\gamma\approx 0.5772$ is the Euler–Mascheroni constant.
We also record the following simple observation.
Lemma 1 (Lower bound for the matrix-count Euler product).
For all primes $p\geq 3$,
$\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{p^{j}}\Bigr)\;\geq\;\frac{1}{2}.$
Proof.
For each $J\geq 1$, set $a_{j}:=p^{-j}$. Then $0<a_{j}\leq 1$ for all $j$, so the finite-product inequality
$\prod_{j=1}^{J}(1-a_{j})\geq 1-\sum_{j=1}^{J}a_{j}$
gives
$\prod_{j=1}^{J}\Bigl(1-\frac{1}{p^{j}}\Bigr)\;\geq\;1-\sum_{j=1}^{J}\frac{1}{p^{j}}.$
Letting $J\to\infty$ yields
$\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{p^{j}}\Bigr)\;\geq\;1-\sum_{j=1}^{\infty}\frac{1}{p^{j}}=1-\frac{1}{p-1}\geq\frac{1}{2}\qquad(p\geq 3).$
Theorem 3 (Unconditional vanishing of the infimal PRIM-LWE density).
$\inf_{p\textup{ prime}}c(p)=0,\qquad\liminf_{\begin{subarray}{c}p\to\infty\\ p\textup{ prime}\end{subarray}}c(p)=0.$
Proof.
For each $k\geq 1$ define the $k$th primorial
$N_{k}\;=\;\prod_{i=1}^{k}p_{i}.$
Since $\gcd(1,N_{k})=1$, Theorem 1 guarantees infinitely many primes in the arithmetic progression $\{1+j\,N_{k}:j\geq 1\}$. Fix any such prime $p\equiv 1\pmod{N_{k}}$, so that $N_{k}\mid(p-1)$.
Because every prime $q\leq p_{k}$ divides p1p-1, the multiplicative formula for the totient gives
$\frac{\varphi(p-1)}{p-1}\;=\;\prod_{q\mid(p-1)}\Bigl(1-\frac{1}{q}\Bigr)\;\leq\;\prod_{i=1}^{k}\Bigl(1-\frac{1}{p_{i}}\Bigr),$
where the inequality holds because each factor $\bigl(1-\frac{1}{q}\bigr)$ lies in $(0,1)$, so the full product over all prime divisors $q\mid(p-1)$ is no larger than the partial product over the subset $p_{1},\dots,p_{k}$.
By Mertens’ third theorem (Theorem 2),
$\prod_{i=1}^{k}\Bigl(1-\frac{1}{p_{i}}\Bigr)\;\sim\;\frac{e^{-\gamma}}{\log p_{k}}\;\xrightarrow{k\to\infty}\;0.$
Since the infinite product satisfies $\prod_{j=1}^{\infty}(1-p^{-j})\leq 1$ for every prime $p$, we obtain
$c(p)\;=\;\frac{\varphi(p-1)}{p-1}\;\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{p^{j}}\Bigr)\;\leq\;\prod_{i=1}^{k}\Bigl(1-\frac{1}{p_{i}}\Bigr)\;\sim\;\frac{e^{-\gamma}}{\log p_{k}}\;\xrightarrow{k\to\infty}\;0.$
Since for every $k$ there are infinitely many primes $p\equiv 1\pmod{N_{k}}$, the displayed upper bound holds for arbitrarily large primes $p$. Letting $k\to\infty$ gives
$\liminf_{\begin{subarray}{c}p\to\infty\\ p\textup{ prime}\end{subarray}}c(p)=0,$
and the infimum statement follows immediately. ∎
Remark 2.
The key advantage of this argument over the approach suggested in [1] is that we require only a prime $p\equiv 1\pmod{N_{k}}$—an infinite supply of which is guaranteed by Dirichlet’s theorem—rather than a primorial prime $p=N_{k}+1$, whose infinitude remains a longstanding open problem; see [10] for computational results on primorial primes.
4 Quantitative Refinements
We first record the form of Linnik’s theorem used in the proof.
Theorem 4 (Least prime in an arithmetic progression; Linnik [11]).
There exist absolute constants $C_{0},L>0$ such that for every modulus $q\geq 1$ and every residue class $a\pmod{q}$ with $\gcd(a,q)=1$, the least prime $p(a,q)\equiv a\pmod{q}$ satisfies
$p(a,q)\leq C_{0}q^{L}.$
Lemma 2 (Prime-number-theorem consequences for primorials; cf. [12]).
Let
$N_{k}=\prod_{i=1}^{k}p_{i},\qquad\vartheta(x):=\sum_{\begin{subarray}{c}p\leq x\\ p\text{ prime}\end{subarray}}\log p.$
Then
$\log N_{k}=\vartheta(p_{k})\sim p_{k},\qquad p_{k+1}\sim p_{k}\qquad(k\to\infty).$
Consequently, if
$m(x):=\max\{k\geq 1:N_{k}\leq x\},$
then
$\log N_{m(x)}\sim\log x,\qquad p_{m(x)}\sim\log x\qquad(x\to\infty).$
Proof.
The first two assertions are standard consequences of the prime number theorem. Since
$N_{m(x)}\leq x<N_{m(x)+1}=N_{m(x)}\,p_{m(x)+1},$
we have
$\log N_{m(x)}\leq\log x<\log N_{m(x)}+\log p_{m(x)+1}.$
Now
$\log p_{m(x)+1}=o\!\bigl(p_{m(x)}\bigr)=o\!\bigl(\log N_{m(x)}\bigr),$
because $\log N_{m(x)}=\vartheta(p_{m(x)})\sim p_{m(x)}$ and $p_{m(x)+1}\sim p_{m(x)}$. Hence $\log x\sim\log N_{m(x)}$, and therefore $p_{m(x)}\sim\log x$. ∎
Proposition 1 (Upper bound for the minimum density up to $x$).
There exists an absolute constant $C>0$ such that for all sufficiently large $x$,
$\min_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}c(p)\;\leq\;\frac{C}{\log\log x}.$
Proof.
By Theorem 4, fix absolute constants $C_{0},L>0$ such that the least prime in any reduced residue class modulo $q$ is at most $C_{0}q^{L}$. For each $k\geq 1$, let $p^{*}(k)$ denote the least prime satisfying $p^{*}(k)\equiv 1\pmod{N_{k}}$. Given $x$, let $k$ be the largest positive integer such that
$C_{0}N_{k}^{\,L}\leq x.$
Then
$p^{*}(k)\leq C_{0}N_{k}^{\,L}\leq x.$
Such a $k$ exists for all sufficiently large $x$, and $k\to\infty$ as $x\to\infty$.
Since $\log N_{k}=\sum_{i=1}^{k}\log p_{i}\sim p_{k}$ by the prime number theorem, the inequality $N_{k}^{\,L}\leq x$ gives
$L\,\log N_{k}\leq\log x,$
and hence $p_{k}\ll\log x$. Therefore $\log p_{k}=O(\log\log x)$.
Conversely, the maximality of $k$ forces $C_{0}N_{k+1}^{\,L}>x$, so
$L\log N_{k+1}>\log x-\log C_{0}.$
Since $\log N_{k+1}=\log N_{k}+\log p_{k+1}$, while
$\log N_{k}=\sum_{i=1}^{k}\log p_{i}\sim p_{k}\quad\text{and}\quad\log p_{k+1}\sim\log p_{k}=o(p_{k})=o(\log N_{k}),$
we have $\log N_{k+1}\sim\log N_{k}$. Hence $\log N_{k}\gg\log x$, so $p_{k}\gg\log x$ and therefore $\log p_{k}=\Omega(\log\log x)$. Combined with the upper bound just proved, this yields
$\log p_{k}\asymp\log\log x.$
By construction the least prime $p^{*}(k)\equiv 1\pmod{N_{k}}$ satisfies $p^{*}(k)\leq x$, and the argument in the proof of Theorem 3 gives
$c\bigl(p^{*}(k)\bigr)\;\leq\;\prod_{i=1}^{k}\Bigl(1-\frac{1}{p_{i}}\Bigr)\;\sim\;\frac{e^{-\gamma}}{\log p_{k}}\;\asymp\;\frac{1}{\log\log x}.$
Therefore
$\min_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}c(p)\;\leq\;c\bigl(p^{*}(k)\bigr)\;\ll\;\frac{1}{\log\log x},$
as claimed. ∎
5 The Limiting Distribution of $c(p)$
We now turn to the distributional theory of $c(p)$ over the primes. A related density problem in $\mathrm{GL}_{n}(q)$, with $q$ ranging over prime powers, was studied by Aivazidis and Sofos [6], who proved a distribution law for the density of maximal-order elements. Our setting is different: we work over prime fields and with the prim-lwe density $c(p)$. Recall that
$f(n):=\log\frac{n}{\varphi(n)}=\sum_{\ell\mid n}\log\frac{\ell}{\ell-1},\qquad\frac{\varphi(p-1)}{p-1}=e^{-f(p-1)},$
and hence the problem sits inside the shifted-prime Erdős–Wintner framework [13] for additive functions on $p+a$ with fixed $a\neq 0$—for this viewpoint, see Tenenbaum [2]. Its singularity is also essentially classical, ultimately resting on the Jessen–Wintner theory of infinite convolutions [14]: in the integer setting, pure singularity under the condition
$\sum_{\ell\ \text{prime}}\frac{f(\ell)^{2}}{\ell}<\infty$
is due to Erdős [15]; see also [16, Exercise 256]. Tenenbaum [2, p. 258] observes that the same conclusion carries over to the shifted-prime law by adapting the same argument to the shifted-prime characteristic function. The full-support conclusion on $0,\tfrac{1}{2}]$ then follows from Deshouillers and Hassani [3, Theorem 1.1] together with Lemma 3. Our novel contribution in this section is the transport of this classical law to $c(p)$ via Lemma 4, which establishes that the dimension-uniform prim-lwe reduction constant shares the same limiting distribution.
5.1 The distribution of $\varphi(p-1)/(p-1)$
Historically, the distribution of additive functions on shifted primes $p+a$, with fixed $a\neq 0$, is governed by the shifted-prime analogue of the Erdős–Wintner theorem; see [17, 18, 2] for a concise overview and references. For the specific function needed here, however, one can bypass that general machinery and use the following direct p1p-1-specific input from Deshouillers and Hassani.
Proposition 2 (Deshouillers–Hassani; special case of [3, Proposition 2.1 and equation (2.1)]).
For each integer $m\geq 1$, there exists a continuous distribution function $G_{m}\colon\mathbb{R}\to[0,1]$ such that for every $\alpha\in\mathbb{R}$,
$\lim_{x\to\infty}\frac{1}{\pi(x;m,1)}\#\left\{p\leq x:p\text{ prime},\ p\equiv 1\pmod{m},\;\frac{\varphi(p-1)}{p-1}\leq\alpha\right\}=G_{m}(\alpha).$
Moreover, the limiting measure is supported on $0,\varphi(m)/m]$.
Proof.
Apply Deshouillers and Hassani [3, Proposition 2.1] to
$g(n):=\frac{\varphi(n)}{n}.$
For a prime $\ell$,
$\log g(\ell)=\log\!\left(1-\frac{1}{\ell}\right)=-\frac{1}{\ell}+O\!\left(\frac{1}{\ell^{2}}\right),$
so, writing $h(\ell):=\log g(\ell)$ for each prime $\ell$, we have $|h(\ell)|=|\log(1-1/\ell)|\leq\log 2<1$ for every prime $\ell$. Therefore the third Erdős–Wintner series $\sum_{|h(\ell)|>1}1/\ell$ is vacuous, while
$\frac{h(\ell)}{\ell}=-\frac{1}{\ell^{2}}+O\!\left(\frac{1}{\ell^{3}}\right),\qquad\frac{h(\ell)^{2}}{\ell}=\frac{1}{\ell^{3}}+O\!\left(\frac{1}{\ell^{4}}\right).$
Hence
$\sum_{\ell}\frac{h(\ell)}{\ell}\quad\text{converges absolutely, and}\quad\sum_{\ell}\frac{h(\ell)^{2}}{\ell}\quad\text{converges}.$
Thus the three convergence hypotheses of [3, Proposition 2.1] are satisfied, and there exists a distribution function $G_{m}$ such that the displayed limit holds at every continuity point of $G_{m}$.
By the continuity criterion in the same proposition, $G_{m}$ is continuous if and only if
$\sum_{\begin{subarray}{c}\ell\equiv 1\pmod{m}\\ g(\ell)\neq 1\end{subarray}}\frac{1}{\ell}$
diverges. Here $g(\ell)=1-\frac{1}{\ell}\neq 1$ for every prime $\ell$, and
$\sum_{\ell\equiv 1\pmod{m}}\frac{1}{\ell}=\infty$
by the divergence of the prime harmonic series in every reduced residue class. Thus $G_{m}$ is continuous, so the displayed limit holds for every $\alpha\in\mathbb{R}$.
If $p\equiv 1\pmod{m}$, then every prime divisor of $m$ also divides p1p-1, so
$\frac{\varphi(p-1)}{p-1}=\prod_{q\mid(p-1)}\Bigl(1-\frac{1}{q}\Bigr)\leq\prod_{q\mid m}\Bigl(1-\frac{1}{q}\Bigr)=\frac{\varphi(m)}{m},$
which yields the support claim. ∎
We shall also use the following standard Siegel–Walfisz (see [12, 16]) estimate in a range of polylogarithmic moduli.
Theorem 5 (Siegel–Walfisz).
For every fixed $A>0$ there exists $c_{A}>0$ such that uniformly for integers $m\leq(\log x)^{A}$ and reduced residue classes $a\pmod{m}$,
$\pi(x;m,a)=\frac{\operatorname{Li}(x)}{\varphi(m)}+O_{A}\!\bigl(xe^{-c_{A}\sqrt{\log x}}\bigr),$
where
$\pi(x;m,a)=\#\{p\leq x:p\text{ prime},\ p\equiv a\pmod{m}\}$.
For an independent, self-contained proof (based on Siegel–Walfisz) that intervals near the endpoint $1/2$ carry positive limiting measure, see Lemma 5 in Appendix A. That result is not used in what follows.
Lemma 3 (Density of even totient ratios).
The set
$\left\{\frac{\varphi(m)}{m}:m\geq 1\text{ even}\right\}$
is dense in $\left[0,\tfrac{1}{2}\right]$.
Proof.
Since $\varphi(m)/m$ depends only on the distinct prime divisors of $m$, we have
$\left\{\frac{\varphi(m)}{m}:m\geq 1\text{ even}\right\}=\left\{\frac{1}{2}\prod_{p\in S}\left(1-\frac{1}{p}\right):S\subset\{\text{odd primes}\},\ S\text{ finite}\right\}.$
Enumerate the odd primes as $r_{1}<r_{2}<\cdots$, and set
$a_{j}:=\log\frac{r_{j}}{r_{j}-1}.$
Then $a_{j}>0$, $a_{j}\downarrow 0$, and
$a_{j}=\frac{1}{r_{j}}+O\!\left(\frac{1}{r_{j}^{2}}\right).$
Since $\sum_{j}1/r_{j}$ over the odd primes diverges, it follows that
$\sum_{j=1}^{\infty}a_{j}=+\infty.$
We claim that the set of finite subsums
$\left\{\sum_{j\in J}a_{j}:J\subset\mathbb{N}\ \text{finite}\right\}$
is dense in $0,\infty)$. Indeed, fix $y\geq 0$ and $\varepsilon>0$. Choose $J_{0}$ so large that $a_{j}<\varepsilon$ for all $j\geq J_{0}$. Since $\sum_{j\geq J_{0}}a_{j}=\infty$, there exists $N\geq J_{0}$ minimal such that
$\sum_{j=J_{0}}^{N}a_{j}>y.$
By minimality,
$\sum_{j=J_{0}}^{N-1}a_{j}\leq y<\sum_{j=J_{0}}^{N}a_{j}=\sum_{j=J_{0}}^{N-1}a_{j}+a_{N}.$
Because $a_{N}<\varepsilon$, it follows that
$y-\varepsilon<\sum_{j=J_{0}}^{N-1}a_{j}\leq y.$
Thus a finite subsum lies within $\varepsilon$ of $y$, proving the claim.
Exponentiating the negative of these finite subsums therefore shows that
$\left\{\frac{1}{2}\exp\!\left(-\sum_{j\in J}a_{j}\right):J\subset\mathbb{N}\ \text{finite}\right\}=\left\{\frac{1}{2}\prod_{j\in J}\left(1-\frac{1}{r_{j}}\right):J\subset\mathbb{N}\ \text{finite}\right\}$
is dense in $(0,1/2]$. Since the finite subsums can be arbitrarily large, 0 is also a limit point. Hence the displayed set is dense in $[0,1/2]$, as claimed. ∎
Theorem 6 (Continuous limiting distribution and full support of the totient ratio over primes).
There exists a continuous distribution function $G\colon\mathbb{R}\to[0,1]$ such that for every $\alpha\in\mathbb{R}$,
$\lim_{x\to\infty}\;\frac{1}{\pi(x)}\,\#\bigl\{p\leq x:p\text{ prime},\;\varphi(p-1)/(p-1)\leq\alpha\bigr\}\;=\;G(\alpha).$
The limiting law has support exactly $0,\tfrac{1}{2}]$, with $G(\alpha)=0$ for $\alpha\leq 0$ and $G(\alpha)=1$ for $\alpha\geq 1/2$. In particular, $G$ is strictly increasing on $0,\tfrac{1}{2}]$.
Proof.
Apply Proposition 2 with $m=2$. Since
$\pi(x;2,1)=\pi(x)-1\sim\pi(x),$
there exists a continuous distribution function $G\colon\mathbb{R}\to[0,1]$ such that for every $\alpha\in\mathbb{R}$,
$\lim_{x\to\infty}\frac{1}{\pi(x)}\#\bigl\{p\leq x:p\text{ prime},\;\varphi(p-1)/(p-1)\leq\alpha\bigr\}=G(\alpha).$
For every odd prime $p$, the integer p1p-1 is even, so
$0<\frac{\varphi(p-1)}{p-1}\leq\frac{1}{2}.$
The exceptional prime $p=2$ is negligible in the limit. Hence the limiting law is carried by $0,\tfrac{1}{2}]$, and therefore
$G(\alpha)=0\qquad(\alpha\leq 0),\qquad G(\alpha)=1\qquad(\alpha\geq 1/2).$
Let $\mu_{\varphi}$ denote the limiting measure with distribution function $G$. Fix an even integer $m\geq 2$, and write
$x_{m}:=\frac{\varphi(m)}{m}.$
By Deshouillers and Hassani [3, Theorem 1.1], the distribution function $G$ is not left-differentiable at $x_{m}$. If $x_{m}\notin\operatorname{supp}(\mu_{\varphi})$, then some open interval $I$ containing $x_{m}$ has $\mu_{\varphi}(I)=0$. Hence $G$ is locally constant on $I$, and in particular left-differentiable at $x_{m}$, a contradiction. Therefore $x_{m}\in\operatorname{supp}(\mu_{\varphi})$ for every even $m$.
By Lemma 3, the support contains a dense subset of $0,\tfrac{1}{2}]$. Since the support is closed and already lies inside $0,\tfrac{1}{2}]$, it follows that
$\operatorname{supp}(\mu_{\varphi})=[0,\tfrac{1}{2}].$
Now let $0\leq a<b\leq 1/2$. Choose any $x\in(a,b)$. Since $x\in\operatorname{supp}(\mu_{\varphi})$, every neighbourhood of $x$ has positive $\mu_{\varphi}$-measure. Taking $r>0$ so small that $(x-r,x+r)\subset(a,b)$, we obtain
$\mu_{\varphi}((a,b])\geq\mu_{\varphi}((x-r,x+r))>0.$
Hence
$G(b)-G(a)=\mu_{\varphi}((a,b])>0,$
so $G$ is strictly increasing on $0,\tfrac{1}{2}]$. The displayed bounds $0<G(\tau)<1$ for $0<\tau<1/2$ follow immediately from this and the endpoint values $G(0)=0$, $G(1/2)=1$. ∎
Theorem 6 completely determines the shifted-prime support: it is exactly $0,\tfrac{1}{2}]$. This does not settle the classical support question for the distribution of $\varphi(n)/n$ over the integers, which remains open; see [3]. The residue-class laws in [3, Proposition 2.1 and equation (2.1)] are consistent with this picture, since for primes $p\equiv 1\pmod{m}$ the corresponding limiting measure is supported in $0,\varphi(m)/m]$.
5.2 From $\varphi(p-1)/(p-1)$ to $c(p)$
Lemma 4 (Transport lemma).
Let $X(p)\geq 0$ and $a(p)>0$ be functions of the primes. Assume that $X(p)$ has a continuous limiting distribution function $G$ over the primes, i.e.
$\lim_{x\to\infty}\frac{1}{\pi(x)}\#\bigl\{p\leq x:p\text{ prime},\ X(p)\leq\alpha\bigr\}=G(\alpha)\qquad(\alpha\in\mathbb{R}).$
If $a(p)\to 1$ as $p\to\infty$ through the primes, then $a(p)X(p)$ has the same limiting distribution $G$.
Proof.
For $\alpha<0$ there is nothing to prove, since both distribution functions vanish. Fix $\alpha\geq 0$ and $0<\varepsilon<1$. Since $a(p)\to 1$, there exists $P_{0}=P_{0}(\varepsilon)$ such that
$1-\varepsilon<a(p)<1+\varepsilon\qquad(p>P_{0},\ p\text{ prime}).$
For such primes $p$,
$X(p)\leq\frac{\alpha}{1+\varepsilon}\;\Longrightarrow\;a(p)X(p)\leq\alpha\;\Longrightarrow\;X(p)\leq\frac{\alpha}{1-\varepsilon}.$
Hence
$\frac{1}{\pi(x)}\#\left\{p\leq x:p\text{ prime},\ X(p)\leq\frac{\alpha}{1+\varepsilon}\right\}\leq\frac{1}{\pi(x)}\#\bigl\{p\leq x:p\text{ prime},\ a(p)X(p)\leq\alpha\bigr\}+\frac{\pi(P_{0})}{\pi(x)},$
and
$\frac{1}{\pi(x)}\#\bigl\{p\leq x:p\text{ prime},\ a(p)X(p)\leq\alpha\bigr\}\leq\frac{1}{\pi(x)}\#\left\{p\leq x:p\text{ prime},\ X(p)\leq\frac{\alpha}{1-\varepsilon}\right\}+\frac{\pi(P_{0})}{\pi(x)}.$
Taking $\liminf$ and $\limsup$ as $x\to\infty$, and then letting $\varepsilon\downarrow 0$, continuity of $G$ gives the claim. ∎
Theorem 7 (Transfer of the limiting distribution to the PRIM-LWE density).
The quantity $c(p)$ has the same limiting distribution as $\varphi(p-1)/(p-1)$. That is, for every $\alpha\in\mathbb{R}$,
$\lim_{x\to\infty}\;\frac{1}{\pi(x)}\,\#\bigl\{p\leq x:p\text{ prime},\;c(p)\leq\alpha\bigr\}\;=\;G(\alpha),$
where $G$ is the distribution function from Theorem 6.
Proof.
Write
$X(p):=\frac{\varphi(p-1)}{p-1},\qquad P(p):=\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{p^{j}}\Bigr),$
so that $c(p)=X(p)P(p)$.
For every prime $p\geq 2$,
[…]
hence
$P(p)=1+O\!\left(\frac{1}{p}\right)\qquad(p\to\infty\text{ through the primes}).$
By Theorem 6, $X(p)$ has continuous limiting distribution function $G$. Therefore Lemma 4 applies with $a(p)=P(p)$, and shows that $c(p)=a(p)X(p)$ has the same limiting distribution $G$. ∎
Proposition 3 (Shifted-prime limiting law for $f$).
Let
$f(n):=\log\frac{n}{\varphi(n)}=\sum_{\begin{subarray}{c}\ell\mid n\\ \ell\textup{ prime}\end{subarray}}\log\frac{\ell}{\ell-1}.$
Then the values f(p1)f(p-1), as $p$ ranges over the primes, possess a continuous purely singular limiting distribution. More precisely, the empirical measures
$\mu_{x}:=\frac{1}{\pi(x)}\sum_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}\delta_{f(p-1)}$
converge weakly, as $x\to\infty$, to a Borel probability measure $\mu_{f}$ on $0,\infty)$ that is continuous and purely singular with respect to Lebesgue measure.
Proof.
The shifted-prime analogue of the Erdős–Wintner theorem, due independently to Kátai [17] and Hildebrand [18], shows that a strongly additive function $h$ possesses a limiting distribution on the shifted primes p1p-1 if and only if the three classical Erdős–Wintner series
$\sum_{\begin{subarray}{c}\ell\ \textup{prime}\\ |h(\ell)|\leq 1\end{subarray}}\frac{h(\ell)}{\ell},\qquad\sum_{\begin{subarray}{c}\ell\ \textup{prime}\\ |h(\ell)|\leq 1\end{subarray}}\frac{h(\ell)^{2}}{\ell},\qquad\sum_{\begin{subarray}{c}\ell\ \textup{prime}\\ |h(\ell)|>1\end{subarray}}\frac{1}{\ell}$
converge; see Tenenbaum [2, pp. 257–258]. In the same reference it is noted that the shifted-prime limiting law is of pure type and is continuous if and only if
$\sum_{\begin{subarray}{c}\ell\ \textup{prime}\\ h(\ell)\neq 0\end{subarray}}\frac{1}{\ell}=\infty.$
We apply this to the strongly additive function
$f(n)=\sum_{\begin{subarray}{c}\ell\mid n\\ \ell\textup{ prime}\end{subarray}}\log\frac{\ell}{\ell-1}.$
For every prime $\ell$,
$0<f(\ell)=\log\frac{\ell}{\ell-1}\leq\log 2<1,\qquad f(\ell)=\frac{1}{\ell}+O\!\left(\frac{1}{\ell^{2}}\right).$
Hence
$\sum_{\ell}\frac{f(\ell)}{\ell}<\infty,\qquad\sum_{\ell}\frac{f(\ell)^{2}}{\ell}<\infty,\qquad\sum_{|f(\ell)|>1}\frac{1}{\ell}=0.$
Therefore the values f(p1)f(p-1), as $p$ ranges over the primes, possess a limiting distribution $\mu_{f}$.
Since $f(\ell)\neq 0$ for every prime $\ell$ and $\sum_{\ell}1/\ell=\infty$, the limiting law $\mu_{f}$ is continuous.
Invoking the shifted-prime extension of Erdős’ singularity criterion (Tenenbaum [2, p. 258]; see also [3, p. 77]): for a strongly additive function $h$, once the limiting law of h(p1)h(p-1) exists, the condition
$\sum_{\ell\ \textup{prime}}\frac{h(\ell)^{2}}{\ell}<\infty$
forces that limiting law to be singular. Applying this with $h=f$, and using the already verified convergence of $\sum_{\ell}f(\ell)^{2}/\ell$, we conclude that $\mu_{f}$ is singular. Since $\mu_{f}$ is also continuous, it is purely singular. ∎
Corollary 2 (Structure of the limiting PRIM-LWE law).
The density $c(p)$ has a continuous, purely singular limiting distribution with support exactly $0,\tfrac{1}{2}]$. Its distribution function $G$ is strictly increasing on $0,\tfrac{1}{2}]$.
Proof.
Write
$Y(p):=\frac{\varphi(p-1)}{p-1}.$
By Theorem 7, $c(p)$ has the same limiting law as $Y(p)$. By Theorem 6, this common law is continuous, has support exactly $0,\tfrac{1}{2}]$, and its distribution function is strictly increasing on $0,\tfrac{1}{2}]$. It remains only to justify pure singularity.
Set
$f(n):=\log\frac{n}{\varphi(n)}=\sum_{\ell\mid n}\log\frac{\ell}{\ell-1}.$
Then $f$ is strongly additive and
$f(\ell)=\log\frac{\ell}{\ell-1}\sim\frac{1}{\ell}\qquad(\ell\to\infty,\ \ell\ \text{prime}),$
while
$Y(p)=e^{-f(p-1)}.$
By Proposition 3, the empirical measures of f(p1)f(p-1) converge weakly to a continuous purely singular law $\mu_{f}$. Since $T(u):=e^{-u}$ is continuous on $0,\infty)$, the continuous mapping theorem implies that
Y(p)=T(f(p1))Y(p)=T(f(p-1))
has limiting law $T_{\#}\mu_{f}$.
Choose a Borel set $E\subset[0,\infty)$ with Lebesgue measure zero and $\mu_{f}(E)=1$. The map $T$ is strictly monotone and Lipschitz on each compact interval $[m,m+1]$, $m\geq 0$. Writing $E=\bigcup_{m\geq 0}(E\cap[m,m+1])$, each image $T(E\cap[m,m+1])$ has Lebesgue measure zero (since a Lipschitz map on a compact interval preserves null sets), and therefore $T(E)$ has Lebesgue measure zero by countable subadditivity. Therefore $T_{\#}\mu_{f}$ is supported on the null set $T(E)$ and is therefore purely singular.
By Theorem 7, the same is true for $c(p)$. Since the limiting law has support containing more than one point, it is not a point mass, and therefore $c(p)$ does not converge to a constant along the primes. ∎
Note that $c(p)<1/2$ for every fixed prime $p$. Indeed, for $p=2$ one has
$\frac{\varphi(p-1)}{p-1}=\frac{\varphi(1)}{1}=1,$
so
$c(2)=\prod_{j\geq 1}\Bigl(1-2^{-j}\Bigr)=\left(\frac{1}{2};\frac{1}{2}\right)_{\infty}\approx 0.288788<\frac{1}{2}.$
For odd $p$, one has $\varphi(p-1)/(p-1)\leq 1/2$ together with $\prod_{j\geq 1}(1-p^{-j})<1$. Thus $1/2$ is only a boundary point of the support of the limiting law for $c(p)$, not an attained value.
Corollary 3 (Unconditional limsup).
$\limsup_{\begin{subarray}{c}p\to\infty\\ p\textup{ prime}\end{subarray}}c(p)=\frac{1}{2}.$
Moreover, for every $\varepsilon>0$, the set of primes with $c(p)>1/2-\varepsilon$ has relative density $1-G(1/2-\varepsilon)$ among the primes.
Proof.
Since $c(p)<1/2$ for every prime, $\limsup c(p)\leq 1/2$. For the reverse inequality, fix $\varepsilon>0$. By Corollary 2, the limiting distribution of $c(p)$ has support exactly $0,\tfrac{1}{2}]$, so the limiting measure assigns positive mass to the interval $(1/2-\varepsilon,\,1/2]$. Hence
$\lim_{x\to\infty}\frac{1}{\pi(x)}\#\bigl\{p\leq x:c(p)>1/2-\varepsilon\bigr\}=1-G(1/2-\varepsilon)>0.$
In particular, infinitely many primes satisfy $c(p)>1/2-\varepsilon$, so $\limsup c(p)\geq 1/2-\varepsilon$. Letting $\varepsilon\downarrow 0$ yields the claim. ∎
Corollary 4 (Small-density moduli).
For every fixed $\tau\in(0,1/2)$, both sets $\{p:c(p)\leq\tau\}$ and $\{p:c(p)>\tau\}$ have positive relative density among the primes, equal to $G(\tau)$ and $1-G(\tau)$, respectively.
Proof.
By Theorem 7 with $\alpha=\tau$, the limiting density of $\{p:c(p)\leq\tau\}$ is $G(\tau)$. Since $G$ is strictly increasing on $0,\tfrac{1}{2}]$ by Theorem 6, we have $0<G(\tau)<1$, giving positive density for both sets. ∎
Corollary 5 (Fixed-dimension analogue).
For each fixed $n\geq 1$, the exact density $c_{n}(p)$ has the same continuous limiting distribution as $\varphi(p-1)/(p-1)$; in particular, its limiting law has support exactly $0,\tfrac{1}{2}]$. Moreover,
$\min_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}c_{n}(p)\;\asymp\;\frac{1}{\log\log x}.$
Proof.
Write
$c_{n}(p)=\frac{\varphi(p-1)}{p-1}\,P_{n}(p),\qquad P_{n}(p):=\prod_{j=1}^{n}\Bigl(1-\frac{1}{p^{j}}\Bigr).$
Since $P_{n}(p)\to 1$ as $p\to\infty$ through the primes, Lemma 4 shows that $c_{n}(p)$ has the same limiting law as $\varphi(p-1)/(p-1)$.
Since $c_{n}(2)>0$, it suffices for large $x$ to consider odd primes $p\leq x$. Let
$m(x):=\max\{k\geq 1:N_{k}\leq x\}.$
For each odd prime $p\leq x$, we have
$N_{\omega(p-1)}\leq\prod_{r\mid(p-1)}r\leq p-1\leq x,$
so $\omega(p-1)\leq m(x)$. Let
$r_{1}<r_{2}<\cdots<r_{d}$
be the distinct prime divisors of p1p-1, where $d=\omega(p-1)$. Then $d\leq m(x)$, and $r_{i}\geq p_{i}$ for $1\leq i\leq d$. Hence
$\frac{\varphi(p-1)}{p-1}=\prod_{i=1}^{d}\Bigl(1-\frac{1}{r_{i}}\Bigr)\geq\prod_{i=1}^{d}\Bigl(1-\frac{1}{p_{i}}\Bigr)\geq\prod_{i=1}^{m(x)}\Bigl(1-\frac{1}{p_{i}}\Bigr),$
because each factor $1-\frac{1}{p_{i}}$ lies in $(0,1)$. Also, by the same elementary estimate used in Lemma 1,
$P_{n}(p)\geq 1-\sum_{j=1}^{n}\frac{1}{p^{j}}\geq 1-\frac{1}{p-1}\geq\frac{1}{2}.$
Therefore
$c_{n}(p)\geq\frac{1}{2}\prod_{i=1}^{m(x)}\Bigl(1-\frac{1}{p_{i}}\Bigr).$
Now $N_{m(x)}\leq x<N_{m(x)+1}=N_{m(x)}p_{m(x)+1}$, so $\log x\sim\log N_{m(x)}$. Since
$\log N_{m(x)}=\sum_{i=1}^{m(x)}\log p_{i}\sim p_{m(x)},$
we get $p_{m(x)}\sim\log x$. By Mertens’ theorem,
$\prod_{i=1}^{m(x)}\Bigl(1-\frac{1}{p_{i}}\Bigr)\sim\frac{e^{-\gamma}}{\log p_{m(x)}}\asymp\frac{1}{\log\log x},$
and therefore
$\min_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}c_{n}(p)\gg\frac{1}{\log\log x}.$
For the upper bound, the argument of Proposition 1 applies verbatim, with the additional observation that $P_{n}(p)\leq 1$ for every prime $p$. If $p^{*}(k)\leq x$ is the least prime congruent to $1\pmod{N_{k}}$ arising there, then
$c_{n}\bigl(p^{*}(k)\bigr)=\frac{\varphi(p^{*}(k)-1)}{p^{*}(k)-1}\,P_{n}\bigl(p^{*}(k)\bigr)\leq\frac{\varphi(p^{*}(k)-1)}{p^{*}(k)-1}\leq\prod_{i=1}^{k}\Bigl(1-\frac{1}{p_{i}}\Bigr)\ll\frac{1}{\log\log x}.$
Hence
$\min_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}c_{n}(p)\ll\frac{1}{\log\log x},$
which proves the claim. ∎
6 Explicit Lower Bounds for Primes of Cryptographic Interest
We now give explicit lower bounds on $c(q)$ for primes $q$ whose predecessor q1q-1 has few distinct prime factors—a condition satisfied by many primes commonly used in lattice-based cryptography, especially Number Theoretic Transform (NTT)-friendly ones.
Theorem 8 (Explicit lower bound on $c(q)$ via the number of distinct prime factors).
Let $q\geq 3$ be a prime, and let $\omega=\omega(q-1)$ denote the number of distinct prime factors of q1q-1. Then
$c(q)\;\geq\;P_{3}\prod_{i=1}^{\omega}\Bigl(1-\frac{1}{p_{i}}\Bigr),\qquad P_{3}:=\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{3^{j}}\Bigr)\approx 0.5601260779,$
where $p_{1}<p_{2}<\cdots<p_{\omega}$ are the first $\omega$ primes.
Proof.
Write the distinct prime factors of q1q-1 as $r_{1}<r_{2}<\cdots<r_{\omega}$. Then
$\frac{\varphi(q-1)}{q-1}=\prod_{i=1}^{\omega}\Bigl(1-\frac{1}{r_{i}}\Bigr)\geq\prod_{i=1}^{\omega}\Bigl(1-\frac{1}{p_{i}}\Bigr),$
since the product is minimized by choosing the smallest possible distinct primes.
Now define
$P(t):=\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{t^{j}}\Bigr)\qquad(t>1).$
Each factor $1-t^{-j}$ is increasing in $t$, so $P(t)$ is increasing on $(1,\infty)$. Since $q\geq 3$ is prime, we have $P(q)\geq P(3)=P_{3}$. Therefore
$c(q)=\frac{\varphi(q-1)}{q-1}\,P(q)\geq P_{3}\prod_{i=1}^{\omega}\Bigl(1-\frac{1}{p_{i}}\Bigr),$
as claimed. ∎
Corollary 6 (Matching lower bound).
There exists an absolute constant $c>0$ such that for all sufficiently large $x$,
$\min_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}c(p)\;\geq\;\frac{c}{\log\log x}.$
Consequently,
$\min_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}c(p)\;\asymp\;\frac{1}{\log\log x}.$
Proof.
The lower-bound argument is the same primorial/Mertens argument used in the proof of Corollary 5, but now with Theorem 8 replacing the crude bound $P_{n}(p)\geq\tfrac{1}{2}$.
Let
$m(x):=\max\{k\geq 1:N_{k}\leq x\}.$
Since $c(2)>0$, it suffices for large $x$ to consider odd primes $p\leq x$. For such $p$,
$N_{\omega(p-1)}\leq\prod_{r\mid(p-1)}r\leq p-1\leq x,$
so $\omega(p-1)\leq m(x)$. Therefore Theorem 8 gives
$c(p)\geq P_{3}\prod_{i=1}^{\omega(p-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr)\geq P_{3}\prod_{i=1}^{m(x)}\Bigl(1-\frac{1}{p_{i}}\Bigr).$
As in the proof of Corollary 5, the relation
$N_{m(x)}\leq x<N_{m(x)+1}=N_{m(x)}\,p_{m(x)+1}$
implies $p_{m(x)}\sim\log x$. Hence, by Mertens’ theorem,
$\prod_{i=1}^{m(x)}\Bigl(1-\frac{1}{p_{i}}\Bigr)\sim\frac{e^{-\gamma}}{\log p_{m(x)}}\asymp\frac{1}{\log\log x}.$
Therefore
$\min_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}c(p)\gg\frac{1}{\log\log x}.$
The matching upper bound is Proposition 1. ∎
The lower bound in Corollary 6 may be viewed as the prime-shifted specialization n=p1n=p-1 of the classical universal estimate
$\frac{\varphi(n)}{n}\gg\frac{1}{\log\log n},$
(see, e.g., Rosser and Schoenfeld [19]) together with the uniform bound $\prod_{j\geq 1}(1-p^{-j})\geq\tfrac{1}{2}$. The proof above keeps the primorial structure explicit, which makes it parallel to the upper-bound construction of Proposition 1.
Corollary 7 (Pointwise universal lower bound).
Let
$m(n):=\max\{k\geq 1:N_{k}\leq n\}.$
Then for every prime $q\geq 3$,
$c(q)\;\geq\;P_{3}\prod_{i=1}^{m(q-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr).$
In particular, there exists an absolute constant $c_{0}>0$ such that, for all sufficiently large primes $q$,
$c(q)\geq\frac{c_{0}}{\log\log q}.$
Proof.
By Theorem 8,
$c(q)\geq P_{3}\prod_{i=1}^{\omega(q-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr).$
Also,
$N_{\omega(q-1)}\leq\prod_{r\mid(q-1)}r\leq q-1,$
so $\omega(q-1)\leq m(q-1)$. Since each factor $1-\frac{1}{p_{i}}$ lies in $(0,1)$,
$\prod_{i=1}^{\omega(q-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr)\geq\prod_{i=1}^{m(q-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr),$
which proves the first claim.
Finally, by Mertens’ theorem and Lemma 2,
$\prod_{i=1}^{m(q-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr)\asymp\frac{1}{\log p_{m(q-1)}}\asymp\frac{1}{\log\log q},$
as $q\to\infty$ through the primes. This gives the second claim; equivalently, $1/c(q)\ll\log\log q$. ∎
Remark 3 (On effectivity).
The comparison
$\min_{p\leq x}c(p)\asymp\frac{1}{\log\log x}$
should in principle be capable of being made fully explicit by combining explicit estimates for the Mertens product and for $\vartheta(x)$ (see, e.g., Rosser–Schoenfeld [19] and Dusart [20]) with an explicit least-prime-in-an-arithmetic-progression bound of Linnik type; see also Xylouris [21] for an admissible Linnik exponent. We do not carry out that bookkeeping here, since the present note only needs the qualitative asymptotic order.
Remark 4.
The bound $1/2$ for the infinite product can be sharpened for specific primes. For instance, $\prod_{j\geq 1}(1-q^{-j})\geq 1-1/(q-1)$, and this lower bound exceeds $2/3$ for $q\geq 5$ and is at least $0.99$ for $q\geq 101$. For large implementation-oriented NTT primes (for example $q>2^{30}$), this factor is numerically indistinguishable from $1$. The current NIST-standardized lattice moduli $q=3329$ and $q=8380417$ are much smaller [4, 5], but even there the same elementary bound already gives $\prod_{j\geq 1}(1-q^{-j})>0.9996$ and $\prod_{j\geq 1}(1-q^{-j})>0.9999998$, respectively.
We now instantiate the general $\omega(q-1)$-based bound in several parameter regimes relevant to cryptographic applications, including many NTT-friendly moduli.
Proposition 4 (Bounds for primes with bounded $\omega(q-1)$).
Let $q\geq 3$ be a prime. Define
$B(\omega)=P_{3}\prod_{i=1}^{\omega}\Bigl(1-\frac{1}{p_{i}}\Bigr).$
Then $c(q)\geq B(\omega(q-1))$. Explicit values:
  1. (a)
    $\omega(q-1)=1$: $B(1)=P_{3}/2\approx 0.280063$.
  2. (b)
    $\omega(q-1)\leq 5$: $B(5)\approx 0.116390$.
  3. (c)
    $\omega(q-1)\leq 10$: $B(10)\approx 0.0884704$.
  4. (d)
    $\omega(q-1)\leq 15$: $B(15)\approx 0.0776918$.
  5. (e)
    $\omega(q-1)\leq 20$: $B(20)\approx 0.0715828$.
Proof.
The bound $c(q)\geq B(\omega(q-1))$ is immediate from Theorem 8. Direct evaluation gives
$B(1)=\frac{P_{3}}{2}\approx 0.280063,\qquad B(5)=\frac{16P_{3}}{77}\approx 0.116390,$
$B(10)\approx 0.0884704,\qquad B(15)\approx 0.0776918,\qquad B(20)\approx 0.0715828.$
This yields the stated bounds. ∎
Remark 5 (On the case $\omega(q-1)=1$).
If $\omega(q-1)=1$ and $q$ is an odd prime, then q1q-1 is an even prime power, hence $q-1=2^{m}$ for some $m\geq 1$; equivalently, $q$ is a Fermat prime. In this case
$\frac{\varphi(q-1)}{q-1}=\frac{1}{2},\qquad c(q)=\frac{1}{2}\,P(q),\qquad P(q):=\prod_{j\geq 1}\Bigl(1-\frac{1}{q^{j}}\Bigr).$
Hence $c(q)$ is close to $1/2$ for large Fermat primes; so if infinitely many Fermat primes exist, then $c(q)\to 1/2$ along that subsequence. Thus the universal bound $B(1)=P_{3}/2\approx 0.280063$ is still not tight in this special case.
Corollary 8 (Practical bounds for cryptographic primes).
Let $q$ be a prime with $q>2^{30}$. Then
$c(q)>\frac{0.56}{\log q}.$
Moreover, if $\omega(q-1)\leq 10$, then
$c(q)>0.088.$
In particular, this applies to cryptographic-size primes $q\in[2^{64},2^{128}]$ satisfying $\omega(q-1)\leq 10$.
Proof.
The claim under $\omega(q-1)\leq 10$ follows from Proposition 4(c), since
$B(10)=P_{3}\prod_{i=1}^{10}\Bigl(1-\frac{1}{p_{i}}\Bigr)\approx 0.08847>0.088.$
For the general bound, since every prime divisor of q1q-1 is at most q1q-1,
$\frac{\varphi(q-1)}{q-1}=\prod_{r\mid(q-1)}\Bigl(1-\frac{1}{r}\Bigr)\;\geq\;\prod_{\begin{subarray}{c}r\leq q-1\\ r\text{ prime}\end{subarray}}\Bigl(1-\frac{1}{r}\Bigr),$
where the inequality holds because the right-hand side includes additional factors in $(0,1)$.
By Dusart’s explicit product estimate [20, Theorem 5.9], which we use conservatively with threshold $x\geq 2\,278\,383$,
$\prod_{\begin{subarray}{c}r\leq x\\ r\text{ prime}\end{subarray}}\Bigl(1-\frac{1}{r}\Bigr)\;=\;\frac{e^{-\gamma}}{\log x}\,\Bigl(1+O^{*}\!\Bigl(\frac{0.2}{\log^{3}x}\Bigr)\Bigr).$
Since $q>2^{30}$, we have q1>2 278 383q-1>2\,278\,383 and $\log(q-1)>20$, so the parenthetical factor lies in $(1-2.5\times 10^{-5},\;1+2.5\times 10^{-5})$.
Also,
$\prod_{j\geq 1}\Bigl(1-\frac{1}{q^{j}}\Bigr)\geq 1-\sum_{j\geq 1}\frac{1}{q^{j}}=1-\frac{1}{q-1}>0.9999.$
Therefore
$c(q)=\frac{\varphi(q-1)}{q-1}\prod_{j\geq 1}\Bigl(1-\frac{1}{q^{j}}\Bigr)\;>\;\Bigl(1-\frac{0.2}{\log^{3}(q-1)}\Bigr)\cdot 0.9999\cdot\frac{e^{-\gamma}}{\log(q-1)}\;>\;\frac{0.56}{\log q},$
using $e^{-\gamma}>0.5614$, $\log(q-1)<\log q$, and $0.2/\log^{3}(q-1)<2.5\times 10^{-5}$ for $q>2^{30}$. ∎
The bound $c(q)>0.56/\!\log q$ of Corollary 8 is a simple, fully explicit estimate obtained by taking the Mertens product over all primes up to q1q-1. However, the sharper universal pointwise bound
$c(q)\;\geq\;P_{3}\prod_{i=1}^{\omega(q-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr)\;\geq\;P_{3}\prod_{i=1}^{m(q-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr)\;\gg\;\frac{1}{\log\log q}$
follows directly from Theorem 8 together with $\omega(q-1)\leq m(q-1)$ and Mertens’ theorem, where $m(n):=\max\{k:N_{k}\leq n\}$. This matches the sharp order established in Corollary 6 and shows that $1/\!\log\log q$, not $1/\!\log q$, is the correct worst-case scale for the rejection-sampling overhead.
Remark 6 (Typical size of $\omega(q-1)$).
Let $\Phi$ denote the standard normal distribution function. The shifted-prime analogue of the Erdős–Kac theorem (due to Barban [22]; see also Halberstam [23] and Elliott [24, Ch. 12]) asserts that for every real $z$,
$\lim_{x\to\infty}\frac{1}{\pi(x)}\#\left\{q\leq x:q\text{ prime and }\frac{\omega(q-1)-\log\log x}{\sqrt{\log\log x}}\leq z\right\}=\Phi(z).$
Consequently, for every fixed $\delta>0$,
$\frac{1}{\pi(x)}\#\left\{q\leq x:q\text{ prime and }\omega(q-1)>(1+\delta)\log\log x\right\}\longrightarrow 0.$
Thus, for large $x$, almost all primes $q\leq x$ satisfy
$\omega(q-1)\leq(1+\delta)\log\log x,$
which explains why bounds depending only on $\omega(q-1)$ are relevant in cryptographic parameter ranges.
Scheme $q$ factorization of q1q-1 general lower bound for $c(q)$ exact-factorization lower bound
ML-KEM $3329$ $2^{8}\cdot 13$ $B(2)=P_{3}/3\approx 0.1867$ $\dfrac{6}{13}\Bigl(1-\dfrac{1}{3328}\Bigr)\approx 0.4614$
ML-DSA $8380417$ $2^{13}\cdot 3\cdot 11\cdot 31$ $B(4)=8P_{3}/35\approx 0.1280$ $\dfrac{100}{341}\Bigl(1-\dfrac{1}{8380416}\Bigr)\approx 0.2933$
Table 1: The prime moduli appearing in the current NIST standards for ML-KEM [4] and ML-DSA [5]. The fourth column gives the general $\omega(q-1)$-based lower bound for $c(q)$ from Theorem 8, while the fifth gives the sharper lower bound obtained from the exact factorization of q1q-1 together with $P(q)\geq 1-\frac{1}{q-1}$. The exact-factorization bounds imply expected rejection-sampling overheads at most $2.17$ and $3.42$, respectively.
For the two prime moduli appearing, across all parameter sets, in the current NIST standards for ML-KEM and ML-DSA [4, 5], Table 1 records both the general $\omega(q-1)$-based bounds and the sharper exact-factorization bounds. The fourth column records the density lower bound $B(\omega(q-1))$, so the corresponding general rejection-sampling overhead upper bound is its reciprocal $1/B(\omega(q-1))$; for the two rows of Table 1 these are $1/B(2)\approx 5.36$ and $1/B(4)\approx 7.82$, respectively. The exact factorizations give
$\frac{\varphi(3328)}{3328}=\frac{6}{13},\qquad\frac{\varphi(8380416)}{8380416}=\frac{100}{341},$
so together with $P(q)\geq 1-\frac{1}{q-1}$ one obtains
$c(3329)\geq\frac{6}{13}\Bigl(1-\frac{1}{3328}\Bigr)\approx 0.4614,\qquad c(8380417)\geq\frac{100}{341}\Bigl(1-\frac{1}{8380416}\Bigr)\approx 0.2933.$
Direct evaluation of the defining product shows that these lower bounds are already essentially exact:
$c(3329)=\frac{6}{13}\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{3329^{j}}\Bigr)\approx 0.46139978,\qquad c(8380417)=\frac{100}{341}\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{8380417^{j}}\Bigr)\approx 0.29325510.$
Equivalently, the corresponding dimension-uniform expected rejection-sampling overheads $1/c(q)$ are approximately $2.1674$ and $3.4101$. For each fixed dimension $n$, the exact overheads $1/c_{n}(q)$ are no larger.
Remark 7 (Concrete example).
Consider $q=2^{64}-2^{32}+1=18\,446\,744\,069\,414\,584\,321$, a well-known NTT-friendly prime. This modulus is not the native modulus of the standardized NIST-PQC schemes, but it is used as a 64-bit NTT-friendly ambient field in implementation-oriented work on lattice cryptography, so it is a natural cryptographic test case [25]. One verifies that $q-1=2^{32}\cdot 3\cdot 5\cdot 17\cdot 257\cdot 65537$, so $\omega(q-1)=6$ and
$\frac{\varphi(q-1)}{q-1}=\frac{1}{2}\cdot\frac{2}{3}\cdot\frac{4}{5}\cdot\frac{16}{17}\cdot\frac{256}{257}\cdot\frac{65536}{65537}\approx 0.2500.$
Since $P(q)>1-10^{-19}$ for such large $q$, the limiting density $c(q)$ is about $0.25$. The exact finite-$n$ density $c_{n}(q)$ is then also essentially $0.25$, so about one in four random $n\times n$ matrices has a primitive-root determinant, and the security reduction loses a factor of about $4$.
7 Implications for PRIM-LWE
For a fixed prime modulus $p$, [1, Theorem 10] shows that decision-prim-lwe is at least as hard as decision-LWE, up to an $O(n^{2})$ factor. Separately, [1, Proposition 5] isolates a positive lower bound, uniform in the matrix dimension, for the probability that a random secret matrix has primitive-root determinant. In our notation, that dimension-uniform sampling constant is
$c=\inf_{m\geq 1}c_{m}(p).$
Since
$c_{m}(p)=\frac{\varphi(p-1)}{p-1}\prod_{j=1}^{m}\Bigl(1-\frac{1}{p^{j}}\Bigr)$
is decreasing in $m$, it follows that
$\inf_{m\geq 1}c_{m}(p)=\lim_{m\to\infty}c_{m}(p)=c(p).$
Thus $c(p)$ is the dimension-uniform sampling term entering the reduction. The present note concerns this quantity $c(p)$; it does not alter the separate $O(n^{2})$ factor from [1, Theorem 10].
Theorem 3 shows that no uniform lower bound on $c(p)$ holds across all primes, meaning the reduction tightness degrades for primes $p$ whose predecessor p1p-1 is divisible by many small primes. More precisely, let $p^{*}(k)$ denote the least prime satisfying $p^{*}(k)\equiv 1\pmod{N_{k}}$, as in the proof of Proposition 1. Then, for the absolute constants $C_{0},L$ from Theorem 4,
$N_{k}<p^{*}(k)\leq C_{0}N_{k}^{\,L}.$
The left inequality holds because $p^{*}(k)=1+jN_{k}$ with $j\geq 1$, while the right inequality is exactly Linnik’s theorem. Therefore
$\log p^{*}(k)\asymp\log N_{k},\qquad\log N_{k}\sim p_{k}$
by Lemma 2, and hence
$\log\log p^{*}(k)\asymp\log p_{k}.$
On the other hand, the proof of Proposition 1 gives
$c\bigl(p^{*}(k)\bigr)\leq\prod_{i=1}^{k}\Bigl(1-\frac{1}{p_{i}}\Bigr)\sim\frac{e^{-\gamma}}{\log p_{k}}.$
Thus
$c\bigl(p^{*}(k)\bigr)\ll\frac{1}{\log\log p^{*}(k)},$
so the construction yields an infinite sequence of primes $p=p^{*}(k)$ for which
$c(p)=O\!\left(\frac{1}{\log\log p}\right).$
The dimension-uniform upper bound on the rejection-sampling overhead, namely $1/c(p)$, can therefore grow like
$\Omega(\log\log p).$
In fixed dimension $n$, the exact expected number of trials is $1/c_{n}(p)\leq 1/c(p)$. By [1, Theorem 10], the reduction from decision-LWE to decision-prim-lwe introduces an advantage loss of $O(n^{2})\cdot(1/c)$, where $c=\inf_{m\geq 1}c_{m}(p)=c(p)$ is the dimension-uniform sampling constant. Our results therefore show that the $1/c(p)$ factor in this bound can grow like $\Omega(\log\log p)$ along a suitable subsequence of primes.
However, this is a modulus-dependent nuance rather than a structural obstruction. For any fixed prime $p$, the constant $c(p)>0$ and the reduction remains valid. Moreover, the explicit bounds of Section 6 show that standard choices of LWE moduli yield $c(q)$ concretely bounded away from zero. For safe primes $q=2r+1$ with $r$ an odd prime (i.e. $q\geq 7$), one has
$c(q)=\frac{r-1}{2r}\prod_{j\geq 1}\left(1-q^{-j}\right).$
For a safe prime $q=2r+1$, the prefactor (r1)/(2r)=1/21/(q1)(r-1)/(2r)=1/2-1/(q-1) is already close to $1/2$ when $q$ is large, and the Euler product $\prod_{j\geq 1}(1-q^{-j})$ tends to $1$ as $q\to\infty$. Hence $c(q)$ is close to $1/2$ for large safe primes. Note, however, that the conditional safe-prime and Fermat-prime observations are now superseded by Corollary 3, which gives $\limsup c(p)=1/2$ unconditionally. On the other hand, if
$P(t):=\prod_{j\geq 1}\Bigl(1-t^{-j}\Bigr)\qquad(t>1),$
then $P(t)$ is increasing in $t$, so for every safe prime $q\geq 7$ we have $P(q)\geq P(3)=P_{3}$. Therefore
$c(q)\geq P_{3}\,\frac{r-1}{2r}.$
More generally, Corollary 4 shows that for every fixed $\tau\in(0,1/2)$, both sets of primes
$\{p:c(p)\leq\tau\}\qquad\text{and}\qquad\{p:c(p)>\tau\}$
have positive relative density, namely $G(\tau)$ and $1-G(\tau)$. Moreover, $G(\tau)\to 0$ as $\tau\downarrow 0$. Thus moduli yielding a very loose constant-factor sampling term are rare only in the threshold-asymptotic sense $\tau\downarrow 0$, not absent for any fixed threshold $\tau<1/2$.
We note that the primes typically used in lattice-based cryptography are chosen for NTT compatibility (often $q\equiv 1\pmod{2^{t}}$ for moderate $t$). Mathematically, $q\equiv 1\pmod{2^{t}}$ controls only the 2-adic part of q1q-1; it does not by itself prevent q1q-1 from having many odd prime factors, and NTT-friendliness alone does not imply a good lower bound on $c(q)$. What our theorems actually use is the controlled factorization structure of q1q-1, quantified by $\omega(q-1)$ or by the exact factorization. The NTT condition is harmless with respect to the 2-part, since the large power of $2$ contributes only the single Euler factor $(1-\tfrac{1}{2})$ to $\varphi(q-1)/(q-1)$. In practice, many NTT primes of cryptographic interest additionally have few small odd prime factors dividing q1q-1, and it is this latter property—not NTT-friendliness per se—that yields the good lower bounds on $c(q)$ recorded in Table 1 and Remark 7. Thus our results do not by themselves indicate a practical weakness; they quantify how modulus choice affects the reduction loss in the prim-lwe framework.
Affine determinant programs (ADPs) [26] provide a natural external determinant-based connection to prim-lwe. In the ADP obfuscation framework, a Kilian-style randomization step left- and right-multiplies by invertible matrices $R,S$ with $\det(R)\det(S)=1$ to preserve the determinant polynomial. Here the ambient sampling space is $\mathrm{GL}_{n}(\mathbb{F}_{q})$, not all of $M_{n}(\mathbb{F}_{q})$. Since the determinant map
$\det:\mathrm{GL}_{n}(\mathbb{F}_{q})\longrightarrow\mathbb{F}_{q}^{\times}$
is a surjective homomorphism, all fibers have the same cardinality, and therefore the exact fraction of invertible matrices with primitive-root determinant is
$\frac{\varphi(q-1)}{q-1}.$
Hence, if $\omega(q-1)\leq 10$, then
$\frac{\varphi(q-1)}{q-1}\;\geq\;\prod_{i=1}^{10}\Bigl(1-\frac{1}{p_{i}}\Bigr)\approx 0.15795,$
so rejection sampling inside $\mathrm{GL}_{n}(\mathbb{F}_{q})$ succeeds in fewer than $6.4$ trials on average. If instead one samples from all matrices $M_{n}(\mathbb{F}_{q})$, then the exact finite-$n$ primitive-root density is $c_{n}(q)$, whose limit as $n\to\infty$ is $c(q)$. We do not know whether such a restriction improves security. We only note, for context, that the ADP framework of [26] includes left- and right-randomization by invertible matrices $R,S$ satisfying $\det(R)\det(S)=1$, while the cryptanalysis of [27] targets a weakness in one of the candidate randomization steps.
8 Concluding Remarks
We have resolved the open question posed in [1, Remark 3(ii)] in the affirmative: $\inf_{p\text{ prime}}c(p)=0$, and in fact
$\liminf_{\begin{subarray}{c}p\to\infty\\ p\text{ prime}\end{subarray}}c(p)=0.$
The proof is short and unconditional, requiring only Dirichlet’s theorem on primes in arithmetic progressions and Mertens’ product formula; the primorial prime hypothesis is entirely unnecessary. We have further established several results that illuminate the finer structure of the density $c(p)$:
  1. 1.
    The minimum of $c(p)$ up to $x$ has the sharp order
    $\min_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}c(p)\asymp\frac{1}{\log\log x}.$
    In particular, the decay to zero is extremely slow and is completely quantified.
  2. 2.
    The quantity $c(p)$ possesses a continuous purely singular limiting distribution over the primes, this law coincides with the classical shifted-prime distribution of $\varphi(p-1)/(p-1)$, and its support is exactly $0,\tfrac{1}{2}]$. Equivalently, the distribution function is strictly increasing on $0,\tfrac{1}{2}]$.
  3. 3.
    For many practical moduli of cryptographic interest—including the current NIST-standardized moduli and other NTT-friendly primes—the density $c(q)$ admits an explicit positive lower bound depending only on $\omega(q-1)$. Accordingly, the prim-lwe reduction has a finite, explicitly computable constant-factor sampling term for every fixed prime modulus, and this term is small in the concrete parameter regimes treated here.
On the cryptographic side, our results show that the rejection-sampling overhead $1/c(p)$ can grow like $\Omega(\log\log p)$ along a suitable subsequence of primes, but this represents a quantitative sensitivity to modulus selection, not a qualitative failure of the reduction. For any fixed prime $p$, the constant $c(p)>0$ and the reduction remains valid. The primes typically chosen for lattice-based cryptography are NTT-friendly, meaning $q\equiv 1\pmod{2^{t}}$ for moderate $t$. While NTT-friendliness alone does not imply a good bound on $c(q)$, the NTT condition is harmless with respect to the 2-part of q1q-1 (the large power of $2$ contributes only the single Euler factor $(1-\tfrac{1}{2})$), and many practical NTT primes additionally have few small odd prime factors dividing q1q-1; it is this controlled factorization structure that yields the good concrete bounds on $c(q)$. For the two NIST-standardized moduli $q=3329$ (ML-KEM) and $q=8380417$ (ML-DSA), the dimension-uniform expected rejection-sampling overheads $1/c(q)$ are at most $2.17$ and $3.42$, respectively; in fixed dimension $n$, the exact expected overheads are $1/c_{n}(q)\leq 1/c(q)$. As a simple conservative explicit estimate, for any prime $q>2^{30}$ one has $1/c(q)\leq 1.79\log q$. The sharper universal pointwise bound is Corollary 7, which gives $1/c(q)=O(\log\log q)$; equivalently, the worst-case overhead among primes $p\leq x$ is of order $\Theta(\log\log x)$. Thus our results do not by themselves indicate a practical weakness. They quantify how modulus choice affects the reduction loss in the prim-lwe framework.
A natural remaining question is to characterize the limiting law $G$ more precisely—for instance, to understand its moments, its modulus of continuity, or its finer local regularity inside $0,\tfrac{1}{2}]$. This should not be confused with the classical support question for the distribution of $\varphi(n)/n$ over all integers, which remains open.
Appendix A Positive Density Near $1/2$
The following lemma gives an independent, Siegel–Walfisz-based proof that intervals near the endpoint $1/2$ carry positive limiting measure for $\varphi(p-1)/(p-1)$.
Lemma 5 (Positive density of primes with totient ratio near 1/2).
For every $\varepsilon>0$,
$\liminf_{x\to\infty}\frac{1}{\pi(x)}\#\left\{p\leq x:p\text{ prime and }\frac{\varphi(p-1)}{p-1}>\frac{1}{2}-\varepsilon\right\}>0.$
Proof.
If $\varepsilon\geq 1$, then $\tfrac{1}{2}-\varepsilon\leq 0$, so the displayed set contains every prime and the claim is trivial. Hence assume throughout that $0<\varepsilon<1$.
Step 1: Sieving out small odd prime divisors. Let $R\geq 3$ be a fixed integer, to be chosen later in terms of $\varepsilon$, and define
$M_{R}:=2\!\!\prod_{\begin{subarray}{c}3\leq r\leq R\\ r\text{ prime}\end{subarray}}r,\qquad\nu_{R}:=\prod_{\begin{subarray}{c}3\leq r\leq R\\ r\text{ prime}\end{subarray}}(r-2),\qquad\delta(R):=\frac{\nu_{R}}{\varphi(M_{R})}=\prod_{\begin{subarray}{c}3\leq r\leq R\\ r\text{ prime}\end{subarray}}\frac{r-2}{r-1}.$
Also set
[…]
The finitely many odd primes $p\leq R$ contribute only $O_{R}(1)$ to $|\mathcal{A}_{R}(x)|$. Hence, by Theorem 5,
$|\mathcal{A}_{R}(x)|=\frac{\nu_{R}}{\varphi(M_{R})}\,\operatorname{Li}(x)+O_{R}\!\bigl(xe^{-c_{R}\sqrt{\log x}}\bigr)=\delta(R)\,\operatorname{Li}(x)+O_{R}\!\bigl(xe^{-c_{R}\sqrt{\log x}}\bigr).$
Applying Theorem 5 also with modulus $1$, we have
$\pi(x)=\operatorname{Li}(x)+O\!\bigl(xe^{-c\sqrt{\log x}}\bigr),$
and therefore
$\frac{|\mathcal{A}_{R}(x)|}{\pi(x)}\longrightarrow\delta(R)>0.$
In particular, $\mathcal{A}_{R}(x)$ has positive relative density $\delta(R)$ among the primes.
Step 2: Reduction to a mean-value estimate. For $p\in\mathcal{A}_{R}(x)$, the only prime $\leq R$ dividing p1p-1 is $2$, so
$\frac{\varphi(p-1)}{p-1}=\frac{1}{2}\prod_{\begin{subarray}{c}q\mid(p-1)\\ q>R\end{subarray}}\Bigl(1-\frac{1}{q}\Bigr).$
Set
$S(p):=\sum_{\begin{subarray}{c}q\mid(p-1)\\ q>R\end{subarray}}\frac{1}{q}.$
By the elementary inequality
$\prod_{i}(1-a_{i})\geq 1-\sum_{i}a_{i}\qquad(0\leq a_{i}\leq 1),$
any $p\in\mathcal{A}_{R}(x)$ with $S(p)\leq\varepsilon$ satisfies
$\frac{\varphi(p-1)}{p-1}\geq\frac{1}{2}\bigl(1-S(p)\bigr)\geq\frac{1}{2}(1-\varepsilon)>\frac{1}{2}-\varepsilon.$
It therefore suffices to show that a positive proportion of primes in $\mathcal{A}_{R}(x)$ satisfy $S(p)\leq\varepsilon$.
Step 3: Mean-value estimate for $S(p)$. Fix $Q=(\log x)^{2}$, and decompose $S(p)=S_{\leq Q}(p)+S_{>Q}(p)$, where
$S_{\leq Q}(p)=\sum_{\begin{subarray}{c}q\mid(p-1)\\ R<q\leq Q\end{subarray}}\frac{1}{q}.$
For the tail, each prime divisor $q>Q$ of p1p-1 contributes at most $1/Q$, and there are at most $\log p/\log Q$ such divisors. Hence, uniformly for $p\leq x$,
$S_{>Q}(p)\leq\frac{\log p}{Q\log Q}\leq\frac{\log x}{Q\log Q}=\frac{1}{2\log x\,\log\log x}=o(1).$
Now assume $x$ is large enough that $Q>R$. For each prime $q\in(R,Q]$, write
$B_{q}(x):=\#\{p\in\mathcal{A}_{R}(x):p\equiv 1\pmod{q}\}$.
Since $q>R$, we have $q\nmid M_{R}$. The conditions $p\in\mathcal{A}_{R}(x)$ and $p\equiv 1\pmod{q}$ therefore define exactly $\nu_{R}$ reduced residue classes modulo $M_{R}q$: modulo each odd prime $r\leq R$ there are r2r-2 admissible reduced classes, modulo $2$ there is one, and modulo $q$ the residue class is forced to be $1$. Because $\varphi(M_{R}q)=\varphi(M_{R})(q-1)$, we obtain from Theorem 5 that
$B_{q}(x)=\frac{\nu_{R}}{\varphi(M_{R}q)}\,\operatorname{Li}(x)+O_{R}\!\bigl(xe^{-c_{R}\sqrt{\log x}}\bigr)=\frac{\delta(R)}{q-1}\,\operatorname{Li}(x)+O_{R}\!\bigl(xe^{-c_{R}\sqrt{\log x}}\bigr),$
uniformly for $R<q\leq Q$. (Here we use that $M_{R}$ is fixed and $q\leq(\log x)^{2}$, so $M_{R}q\leq(\log x)^{A}$ for some fixed $A$ and all sufficiently large $x$.)
Dividing by the asymptotic formula for $|\mathcal{A}_{R}(x)|$ yields
$\frac{B_{q}(x)}{|\mathcal{A}_{R}(x)|}=\frac{1}{q-1}+O_{R}\!\bigl((\log x)e^{-c_{R}\sqrt{\log x}}\bigr)$
uniformly for $R<q\leq Q$. Therefore
$\frac{1}{|\mathcal{A}_{R}(x)|}\sum_{p\in\mathcal{A}_{R}(x)}S_{\leq Q}(p)$ $=\sum_{\begin{subarray}{c}R<q\leq Q\\ q\text{ prime}\end{subarray}}\frac{1}{q}\,\frac{B_{q}(x)}{|\mathcal{A}_{R}(x)|}$
$\leq\sum_{\begin{subarray}{c}R<q\leq Q\\ q\text{ prime}\end{subarray}}\frac{1}{q(q-1)}+O_{R}\!\Bigl((\log x)e^{-c_{R}\sqrt{\log x}}\sum_{q\leq Q}\frac{1}{q}\Bigr)$
$=\sum_{\begin{subarray}{c}R<q\leq Q\\ q\text{ prime}\end{subarray}}\frac{1}{q(q-1)}+o(1),$
since $\sum_{q\leq Q}1/q\ll\log\log Q\ll\log\log\log x$.
Combining this with the uniform bound on $S_{>Q}(p)$, we get
$\frac{1}{|\mathcal{A}_{R}(x)|}\sum_{p\in\mathcal{A}_{R}(x)}S(p)\leq\eta(R)+o(1),\qquad\eta(R):=\sum_{\begin{subarray}{c}q>R\\ q\text{ prime}\end{subarray}}\frac{1}{q(q-1)}.$
Since $\sum_{q}1/(q(q-1))$ converges, we have $\eta(R)\to 0$ as $R\to\infty$.
Step 4: Markov’s inequality and conclusion. Choose $R$ large enough that $\eta(R)<\varepsilon/2$. By Markov’s inequality, the fraction of primes in $\mathcal{A}_{R}(x)$ with $S(p)>\varepsilon$ is at most
$\frac{\eta(R)+o(1)}{\varepsilon}<\frac{1}{2}$
for all sufficiently large $x$. Hence at least half the primes in $\mathcal{A}_{R}(x)$ satisfy $S(p)\leq\varepsilon$, and therefore
$\frac{\varphi(p-1)}{p-1}>\frac{1}{2}-\varepsilon.$
It follows that for all sufficiently large $x$,
$\left\{p\leq x:p\text{ prime and }\frac{\varphi(p-1)}{p-1}>\frac{1}{2}-\varepsilon\right\}\geq\frac{1}{2}\,|\mathcal{A}_{R}(x)|.$
Dividing by $\pi(x)$ and taking $\liminf$, we conclude that
$\liminf_{x\to\infty}\frac{1}{\pi(x)}\#\left\{p\leq x:p\text{ prime and }\frac{\varphi(p-1)}{p-1}>\frac{1}{2}-\varepsilon\right\}$ $\geq\frac{1}{2}\lim_{x\to\infty}\frac{|\mathcal{A}_{R}(x)|}{\pi(x)}$
$=\frac{\delta(R)}{2}>0.$
This proves the lemma. ∎

### 18 Mar 2026 v2

@@
 The prim-lwe problem, introduced by Sehrawat, Yeo, and Desmedt (Theoretical Computer Science, 886 (2021)), is a variant of the Learning with Errors problem in which the secret matrix is required to have a primitive-root determinant.
-The dimension-uniform reduction constant is $c(p)=\inf_{m\geq 1}c_{m}(p)$, where $c_{n}(p)$ is the exact density of $n\times n$ matrices over $\mathbb{F}_{p}$ with primitive-root determinant.
+The dimension-uniform reduction constant is $c(p)=\inf_{n\geq 1}c_{n}(p)$, where $c_{n}(p)$ is the exact density of $n\times n$ matrices over $\mathbb{F}_{p}$ with primitive-root determinant.
 Sehrawat, Yeo, and Desmedt asked whether $\inf_{p\text{ prime}}c(p)=0$, observing that an affirmative answer would follow from the conjectural infinitude of primorial primes.
-We resolve this question unconditionally using only Dirichlet’s theorem and Mertens’ product formula, entirely bypassing the primorial-prime hypothesis. We further establish the sharp order
+We resolve this question unconditionally using only Dirichlet’s theorem and Mertens’ product formula, entirely bypassing the primorial-prime hypothesis.
+We further establish the sharp order
 $\min_{p\leq x}c(p)\asymp\frac{1}{\log\log x}\qquad(x\to\infty),$
-prove that $c(p)$ possesses a continuous purely singular limiting distribution over the primes with support exactly $[0,1/2]$, and derive explicit lower bounds on $c(q)$ for primes of cryptographic interest, parameterized solely by the number $\omega(q-1)$ of distinct prime factors of q1q-1.
-These bounds apply to every prime $q$ whose predecessor q1q-1 has controlled factorization structure, as measured by the number $\omega(q-1)$ of distinct prime factors; this includes many common NTT-friendly moduli, though NTT-friendliness alone does not imply the needed bound.
-For the NIST-standardized moduli $q=3329$ and $q=8380417$, the dimension-uniform expected rejection-sampling overhead $1/c(q)$ in the prim-lwe reduction is at most $2.17$ and $3.42$, respectively.
-As a simple conservative explicit bound, for any prime $q>2^{30}$ the dimension-uniform expected rejection-sampling overhead $1/c(q)$ is at most $1.79\log q$.
+and show that the limiting distribution of $c(p)$ over the primes has support exactly $[0,1/2]$.
+We have not found this full-support statement in the literature.
+The law coincides with the classical shifted-prime distribution of $\varphi(p-1)/(p-1)$ via a transport lemma and is moreover continuous and purely singular.
+We also derive explicit lower bounds on $c(q)$ for primes of cryptographic interest, parameterized solely by the number of distinct prime factors of q1q-1.
+As a simple conservative explicit bound, for any prime $q>2^{30}$ the expected overhead $1/c(q)$ is at most $1.79\log q$.
 On the other hand, our results show that the worst-case overhead among primes $p\leq x$ is of order $\Theta(\log\log x)$, and in particular $1/c(q)=O(\log\log q)$ pointwise.
@@
 Following [1], a matrix $M\in M_{n}(\mathbb{F}_{p})$ has primitive-root determinant if $\det(M)$ generates $\mathbb{F}_{p}^{*}$.
-The fraction of such matrices is
-$c_{n}(p)\;=\;\frac{\varphi(p-1)}{p-1}\;\prod_{j=1}^{n}\Bigl(1-\frac{1}{p^{j}}\Bigr),$
-and since $c_{m}(p)$ is decreasing in $m$, the dimension-uniform constant governing the reduction from decision-LWE to decision-prim-lwe is
-$c(p)\;:=\;\inf_{m\geq 1}c_{m}(p)\;=\;\frac{\varphi(p-1)}{p-1}\;\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{p^{j}}\Bigr).$
-Sehrawat, Yeo, and Desmedt [1, Remark 3(ii)] observed that $\inf_{p}c(p)=0$ would follow from the conjectural infinitude of primorial primes and asked whether this can be proved unconditionally.
+The fraction of such matrices is denoted $c_{n}(p)$; since it is decreasing in $n$, the dimension-uniform constant governing the reduction from decision-LWE to decision-prim-lwe is $c(p):=\inf_{n\geq 1}c_{n}(p)=\lim_{n\to\infty}c_{n}(p)$.
+The precise formulas are recorded in Section 2.
+Sehrawat, Yeo, and Desmedt [1, Remark 3(ii)] asked whether $\inf_{p}c(p)=0$ can be proved unconditionally; see Section 2 for the precise statement.
 Main results
@@
 Unconditional vanishing (Theorem 3).
-We prove $\inf_{p\text{ prime}}c(p)=0$ using only Dirichlet’s theorem on primes in arithmetic progressions and Mertens’ third theorem.
-The key observation is that for each primorial $N_{k}=\prod_{i=1}^{k}p_{i}$, Dirichlet’s theorem supplies infinitely many primes $p\equiv 1\pmod{N_{k}}$, and for any such prime the totient ratio $\varphi(p-1)/(p-1)$ is bounded by the Mertens product $\prod_{i=1}^{k}(1-1/p_{i})\sim e^{-\gamma}/\!\log p_{k}\to 0$.
-This bypasses the primorial-prime hypothesis entirely.
+We prove $\inf_{p\text{ prime}}c(p)=0$ using only Dirichlet’s theorem on primes in arithmetic progressions and Mertens’ third theorem, without assuming the primorial-prime hypothesis; see Section 3.
 2.
@@
 Hence the decay to zero is completely quantified and extremely slow.
-To put this rate in perspective, $\log\log(2^{128})\approx 4.5$ and $\log\log(2^{4096})\approx 7.95$, so the asymptotic lower bound $\min_{p\leq x}c(p)\gg 1/\!\log\log x$ already guarantees that the minimum cannot drop faster than the reciprocal of a slowly growing function.
-Nonetheless, the implied constant is modest (roughly $P_{3}e^{-\gamma}\approx 0.31$), so primes $p$ whose predecessor p1p-1 is divisible by many small primes can still have $c(p)<0.1$ well within practical ranges.
-The concrete impact on prim-lwe therefore depends on the factorization structure of the chosen modulus; the explicit bounds in Section 6 show that standard cryptographic moduli with controlled factorization incur only a small constant-factor overhead.
-These implied constants should in principle be capable of being made explicit by combining explicit estimates for the Mertens product and for $\vartheta(x)$ with an explicit least-prime-in-an-arithmetic-progression bound; see Remark 3.
-We do not pursue that bookkeeping here.
+The implied constant is $\approx 0.31$, so primes $p$ whose predecessor p1p-1 is divisible by many small primes can still have $c(p)<0.1$, well within practical ranges.
+The explicit bounds in Section 6 show that standard cryptographic moduli with controlled factorization incur only a small constant-factor overhead; see Remark 1 for effectivity.
 3.
 Limiting distribution (Theorems 6 and 7, Corollary 2).
-The quantity $c(p)$ possesses a continuous purely singular limiting distribution over the primes, with support exactly $0,\tfrac{1}{2}]$.
-If $G$ denotes the distribution function of this law, then
+The quantity $c(p)$ possesses a continuous limiting distribution over the primes whose support is exactly $0,\tfrac{1}{2}]$; the law is moreover purely singular.
+If $G$ denotes the distribution function, then
 $0<G(\tau)<1\qquad(0<\tau<1/2),$
 the set of primes with $c(p)\leq\tau$ has relative density $G(\tau)$, and $G(\tau)\to 0$ as $\tau\downarrow 0$.
-The existence and continuity of $G$ follow from the shifted-prime Erdős–Wintner framework for additive functions evaluated at p1p-1; for the relevant classical background see Tenenbaum [2] and Deshouillers–Hassani [3].
-The full-support conclusion on $0,\tfrac{1}{2}]$, which to our knowledge has not appeared explicitly in the literature, combines the left non-differentiability result of [3, Theorem 1.1] with the density of $\{\varphi(m)/m:m\text{ even}\}$ in $[0,1/2]$ (Lemma 3).
-Our novel contribution is the transfer of this classical law to $c(p)$ via a transport lemma (Lemma 4), which establishes that the dimension-uniform prim-lwe reduction constant shares the same limiting distribution.
+The existence and continuity of $G$ follow from the shifted-prime Erdős–Wintner framework for additive functions evaluated at p1p-1 [2]; for the relevant classical background see Tenenbaum [3] and Deshouillers–Hassani [4].
+Pure singularity is classical in the integer case (Erdős [5]) and carries over to shifted primes by the same method, as noted by Tenenbaum [3, p.
+263]; we include a self-contained proof for the reader’s convenience.
+Our new contributions on the distributional side are twofold.
+First, we establish that the support of the limiting law is exactly $0,\tfrac{1}{2}]$ by combining the left non-differentiability result of [4, Theorem 1.1] with the density of $\{\varphi(m)/m:m\text{ even}\}$ in $[0,1/2]$ (Lemma 4).
+Second, we transfer the classical shifted-prime law for $\varphi(p-1)/(p-1)$ to the prim-lwe density $c(p)$ via a transport lemma (Lemma 5).
 4.
@@
 For each fixed dimension $n$, the exact expected overhead is $1/c_{n}(q)\leq 1/c(q)$.
-The general $\omega(q-1)$-based bound is immediate from Theorem 8 and gives factors at most $5.36$ and $7.82$ for the NIST-standardized moduli $q=3329$ used in ML-KEM [4] and $q=8380417$ used in ML-DSA [5], respectively.
-The sharper NIST figures use the exact factorizations recorded in Section 6.
-Remark 1.
-The constant $P_{3}$ may also be written in standard $q$-series notation as
-$P_{3}=\left(\frac{1}{3};\frac{1}{3}\right)_{\infty},\qquad(a;q)_{\infty}:=\prod_{m=0}^{\infty}(1-aq^{m}).$
+The general $\omega(q-1)$-based bound is immediate from Theorem 8; the sharper NIST figures use the exact factorizations recorded in Section 6.
 Related work
@@
 Their setting is different from ours: we work over prime fields and with the primitive-root determinant density arising in prim-lwe.
-On the distributional side, Deshouillers and Hassani [3, Proposition 2.1 and equation (2.1)] proved limiting-distribution laws for $\varphi(p-1)/(p-1)$ in the residue classes $p\equiv 1\pmod{m}$, including continuity, and [3, Theorem 1.1] established left non-differentiability of the distribution function at every value $\varphi(m)/m$ with $m$ even.
+On the distributional side, Deshouillers and Hassani [4, Proposition 2.1 and equation (2.1)] proved limiting-distribution laws for $\varphi(p-1)/(p-1)$ in the residue classes $p\equiv 1\pmod{m}$, including continuity, and [4, Theorem 1.1] established left non-differentiability of the distribution function at every value $\varphi(m)/m$ with $m$ even.
 The analogous support question for the classical distribution of $\varphi(n)/n$ over the integers remains open; see [4].
@@
 The quantity $c_{n}(p)$ is the exact acceptance probability in dimension $n$.
-Since $c_{m}(p)$ is decreasing in $m$, the dimension-uniform constant that enters the prim-lwe reduction is
-$\inf_{m\geq 1}c_{m}(p)=\lim_{m\to\infty}c_{m}(p)=c(p).$
+Since $c_{n}(p)$ is decreasing in $n$, the dimension-uniform constant that enters the prim-lwe reduction is
+$\inf_{n\geq 1}c_{n}(p)=\lim_{n\to\infty}c_{n}(p)=c(p).$
 Thus $c(p)$ is not the exact finite-dimensional density but the limiting, dimension-uniform reduction constant.
@@
 $\frac{\varphi(p-1)}{p-1}=\prod_{i=1}^{k}\frac{p_{i}-1}{p_{i}}\longrightarrow 0$
-by Mertens’ theorem, which would imply $\inf_{p}c(p)=0$. They left open whether this can be proved unconditionally.
+by Mertens’ theorem, which would imply $\inf_{p}c(p)=0$. They left open whether this can be proved unconditionally; see [8] for computational results on primorial primes.
 3 Unconditional Proof that $\inf_{p}c(p)=0$
@@
 and the infimum statement follows immediately. ∎
-Remark 2.
-The key advantage of this argument over the approach suggested in [1] is that we require only a prime $p\equiv 1\pmod{N_{k}}$—an infinite supply of which is guaranteed by Dirichlet’s theorem—rather than a primorial prime $p=N_{k}+1$, whose infinitude remains a longstanding open problem; see [10] for computational results on primorial primes.
 4 Quantitative Refinements
@@
 We now turn to the distributional theory of $c(p)$ over the primes.
-A related density problem in $\mathrm{GL}_{n}(q)$, with $q$ ranging over prime powers, was studied by Aivazidis and Sofos [6], who proved a distribution law for the density of maximal-order elements.
-Our setting is different: we work over prime fields and with the prim-lwe density $c(p)$.
-Recall that
-$f(n):=\log\frac{n}{\varphi(n)}=\sum_{\ell\mid n}\log\frac{\ell}{\ell-1},\qquad\frac{\varphi(p-1)}{p-1}=e^{-f(p-1)},$
-and hence the problem sits inside the shifted-prime Erdős–Wintner framework [13] for additive functions on $p+a$ with fixed $a\neq 0$—for this viewpoint, see Tenenbaum [2].
-Its singularity is also essentially classical, ultimately resting on the Jessen–Wintner theory of infinite convolutions [14]: in the integer setting, pure singularity under the condition
-$\sum_{\ell\ \text{prime}}\frac{f(\ell)^{2}}{\ell}<\infty$
-is due to Erdős [15]; see also [16, Exercise 256].
-Tenenbaum [2, p.
-258] observes that the same conclusion carries over to the shifted-prime law by adapting the same argument to the shifted-prime characteristic function.
-The full-support conclusion on $0,\tfrac{1}{2}]$ then follows from Deshouillers and Hassani [3, Theorem 1.1] together with Lemma 3.
-Our novel contribution in this section is the transport of this classical law to $c(p)$ via Lemma 4, which establishes that the dimension-uniform prim-lwe reduction constant shares the same limiting distribution.
+The classical existence and continuity of the limiting distribution of $\varphi(p-1)/(p-1)$ over primes follow from the shifted-prime Erdős–Wintner framework; for the shift a=1a=-1 used here, direct references are Tenenbaum [3, pp.
+257–258] and Deshouillers–Hassani [4].
+Our contributions in this section are:
+  1. (i)
+    The full-support conclusion: the limiting law has support exactly $0,\tfrac{1}{2}]$ (Theorem 6), combining the left non-differentiability result of [4, Theorem 1.1] with the density of $\{\varphi(m)/m:m\text{ even}\}$ in $[0,1/2]$ (Lemma 4).
+  2. (ii)
+    The transport of the classical shifted-prime law to the prim-lwe density $c(p)$ via Lemma 5, establishing that $c(p)$ shares the same limiting distribution as $\varphi(p-1)/(p-1)$.
+We supply a self-contained proof of the shifted-prime pure singularity in Lemma 9.
 5.1 The distribution of $\varphi(p-1)/(p-1)$
-Historically, the distribution of additive functions on shifted primes $p+a$, with fixed $a\neq 0$, is governed by the shifted-prime analogue of the Erdős–Wintner theorem; see [17, 18, 2] for a concise overview and references.
-For the specific function needed here, however, one can bypass that general machinery and use the following direct p1p-1-specific input from Deshouillers and Hassani.
-Proposition 2 (Deshouillers–Hassani; special case of [3, Proposition 2.1 and equation (2.1)]).
+One can bypass the general Erdős–Wintner machinery and use the following direct p1p-1-specific input from Deshouillers and Hassani.
+Lemma 3 (Erdős–Wintner hypotheses for the totient function).
+Let $h$ be the strongly additive function $h(n)=\log(n/\varphi(n))=-\log(\varphi(n)/n)$. For every prime $\ell$,
+$0<h(\ell)=\log\frac{\ell}{\ell-1}\leq\log 2<1,\qquad h(\ell)=\frac{1}{\ell}+O\!\left(\frac{1}{\ell^{2}}\right).$
+Consequently the third Erdős–Wintner series $\sum_{|h(\ell)|>1}1/\ell$ is vacuous, while $\sum_{\ell}h(\ell)/\ell$ and $\sum_{\ell}h(\ell)^{2}/\ell$ both converge.
+Since $h(\ell)\neq 0$ for every prime $\ell$ and $\sum_{\ell}1/\ell=\infty$, the continuity criterion of Lemma 6 is satisfied.
+The same conclusions hold for $g(n)=\varphi(n)/n$ via $\log g(\ell)=-h(\ell)$.
+Proposition 2 (Deshouillers–Hassani; special case of [4, Proposition 2.1 and equation (2.1)]).
 For each integer $m\geq 1$, there exists a continuous distribution function $G_{m}\colon\mathbb{R}\to[0,1]$ such that for every $\alpha\in\mathbb{R}$,
@@
 Proof.
-Apply Deshouillers and Hassani [3, Proposition 2.1] to
+Apply Deshouillers and Hassani [4, Proposition 2.1] to
 $g(n):=\frac{\varphi(n)}{n}.$
-For a prime $\ell$,
-$\log g(\ell)=\log\!\left(1-\frac{1}{\ell}\right)=-\frac{1}{\ell}+O\!\left(\frac{1}{\ell^{2}}\right),$
-so, writing $h(\ell):=\log g(\ell)$ for each prime $\ell$, we have $|h(\ell)|=|\log(1-1/\ell)|\leq\log 2<1$ for every prime $\ell$.
-Therefore the third Erdős–Wintner series $\sum_{|h(\ell)|>1}1/\ell$ is vacuous, while
-$\frac{h(\ell)}{\ell}=-\frac{1}{\ell^{2}}+O\!\left(\frac{1}{\ell^{3}}\right),\qquad\frac{h(\ell)^{2}}{\ell}=\frac{1}{\ell^{3}}+O\!\left(\frac{1}{\ell^{4}}\right).$
-Hence
-$\sum_{\ell}\frac{h(\ell)}{\ell}\quad\text{converges absolutely, and}\quad\sum_{\ell}\frac{h(\ell)^{2}}{\ell}\quad\text{converges}.$
-Thus the three convergence hypotheses of [3, Proposition 2.1] are satisfied, and there exists a distribution function $G_{m}$ such that the displayed limit holds at every continuity point of $G_{m}$.
+By Lemma 3 applied to $\log g(\ell)=-\log(\ell/(\ell-1))$, the three convergence hypotheses of [4, Proposition 2.1] are satisfied, and there exists a distribution function $G_{m}$ such that the displayed limit holds at every continuity point of $G_{m}$.
 By the continuity criterion in the same proposition, $G_{m}$ is continuous if and only if
@@
 Proof.
-Apply Proposition 2 with $m=2$. Since
+Apply Proposition 2 with $m=2$. Since every odd prime satisfies $p\equiv 1\pmod{2}$, we have
 $\pi(x;2,1)=\pi(x)-1\sim\pi(x),$
-there exists a continuous distribution function $G\colon\mathbb{R}\to[0,1]$ such that for every $\alpha\in\mathbb{R}$,
+so the distribution function $G_{2}$ furnished by that proposition is the unconditional limiting distribution of $\varphi(p-1)/(p-1)$ over all primes.
+In particular, $G_{2}$ coincides with the distribution function denoted $F$ in [4, Theorem 1.1].
+Writing $G:=G_{2}$, there exists a continuous distribution function $G\colon\mathbb{R}\to[0,1]$ such that for every $\alpha\in\mathbb{R}$,
 $\lim_{x\to\infty}\frac{1}{\pi(x)}\#\bigl\{p\leq x:p\text{ prime},\;\varphi(p-1)/(p-1)\leq\alpha\bigr\}=G(\alpha).$
@@
 $x_{m}:=\frac{\varphi(m)}{m}.$
-By Deshouillers and Hassani [3, Theorem 1.1], the distribution function $G$ is not left-differentiable at $x_{m}$.
+By Deshouillers and Hassani [4, Theorem 1.1], the distribution function $G$ is not left-differentiable at $x_{m}$.
+Deshouillers and Hassani denote this distribution function by $F$; we write it as $G$.
 If $x_{m}\notin\operatorname{supp}(\mu_{\varphi})$, then some open interval $I$ containing $x_{m}$ has $\mu_{\varphi}(I)=0$.
@@
 This does not settle the classical support question for the distribution of $\varphi(n)/n$ over the integers, which remains open; see [4].
-The residue-class laws in [3, Proposition 2.1 and equation (2.1)] are consistent with this picture, since for primes $p\equiv 1\pmod{m}$ the corresponding limiting measure is supported in $0,\varphi(m)/m]$.
+The residue-class laws in [4, Proposition 2.1 and equation (2.1)] are consistent with this picture, since for primes $p\equiv 1\pmod{m}$ the corresponding limiting measure is supported in $0,\varphi(m)/m]$.
 5.2 From $\varphi(p-1)/(p-1)$ to $c(p)$
+The following transport lemma is the mechanism for contribution (ii) listed at the beginning of this section.
 Lemma 5 (Transport lemma).
@@
 By Theorem 6, $X(p)$ has continuous limiting distribution function $G$. Therefore Lemma 5 applies with $a(p)=P(p)$, and shows that $c(p)=a(p)X(p)$ has the same limiting distribution $G$. ∎
+Lemma 6 (Shifted-prime pure-type and continuity criterion).
+Let $f$ be a real strongly additive arithmetical function satisfying the three convergence conditions of the shifted-prime Erdős–Wintner theorem, so that f(p1)f(p-1) possesses a limiting distribution $\mu_{f}$ over the primes.
+Then $\mu_{f}$ is of pure type: it is either purely discrete, purely singular, or absolutely continuous with respect to Lebesgue measure.
+Moreover, $\mu_{f}$ is continuous (i.e.
+has no atoms) if and only if $\sum_{f(\ell)\neq 0}1/\ell=\infty$.
+Proof.
+The relevant shifted-prime Erdős–Wintner framework is due to Kátai [14] and Hildebrand [15] for the shift $a=1$; the methods extend to any fixed nonzero shift via the Bombieri–Vinogradov theorem, and for the shift a=1a=-1 used here, the explicit characteristic function is recorded in Tenenbaum [3, formula (3), p.
+258].
+The same conclusions hold for general (not necessarily strongly) additive functions under the Erdős–Wintner hypotheses, by the identical Jessen–Wintner argument applied to the full local factors; we state the strongly additive case since it is all that we require.
+By Tenenbaum [3, formula (3), p. 258], the characteristic function of $\mu_{f}$ is given by
+$\widehat{\mu_{f}}(\tau)\;=\;\prod_{p}\!\left(1-\frac{1}{p-1}+\sum_{\nu\geq 1}\frac{e^{i\tau f(p^{\nu})}}{p^{\nu}}\right)$
+The shift is p1p-1, corresponding to a=1a=-1 in the notation of [3, formula (3)].
+Since no prime divides 1-1, the product runs over all primes.
+For the shift a=1a=-1, the local factor at $\ell=2$ in the characteristic function reduces, when $f$ is strongly additive, to $1-\frac{1}{2-1}+\sum_{\nu\geq 1}\frac{e^{i\tau f(2)}}{2^{\nu}}=e^{i\tau f(2)}$, which is a deterministic translation by $f(2)$.
+This factor carries no randomness and does not affect pure type or continuity; it merely shifts the law.
+It therefore suffices to apply the Jessen–Wintner theorem [16] (see also Tenenbaum [13, Theorem 2.7(b)]) to the infinite convolution over the odd primes.
+Each odd-prime factor is the characteristic function of a probability measure supported on a countable set, so the Jessen–Wintner theorem gives that the resulting infinite convolution is of pure type.
+The continuity criterion follows from the same theorem: the convolution is continuous if and only if $\sum(1-\sigma_{\ell})=\infty$, where $\sigma_{\ell}$ is the maximal atom of the factor at $\ell$.
+Since $f$ is strongly additive, the local factor at an odd prime $\ell$ places mass $(\ell-2)/(\ell-1)$ at the origin and mass $1/(\ell-1)$ at $f(\ell)$.
+When $f(\ell)\neq 0$, these two points are distinct, so $\sigma_{\ell}=(\ell-2)/(\ell-1)$ and $1-\sigma_{\ell}=1/(\ell-1)\asymp 1/\ell$.
+It follows that  $\sum_{\ell\text{ odd}}(1-\sigma_{\ell})$ diverges if and only if $\sum_{f(\ell)\neq 0}1/\ell=\infty$, since the prime $\ell=2$ does not enter the odd-prime convolution.
+Since a deterministic translation preserves pure type and continuity, the same conclusions hold for $\mu_{f}$.
+∎
+Lemma 7 (Prime density for a prescribed smooth kernel).
+Let $y\geq 3$, and for each prime $p$ define $a_{y}(p-1):=\prod_{\begin{subarray}{c}\ell\leq y,\;\ell\textup{ prime}\\ \ell\mid(p-1)\end{subarray}}\ell$.
+Let $m$ be an even squarefree positive integer with $P^{+}(m)\leq y$.
+Then the set $\{p\textup{ prime}:a_{y}(p-1)=m\}$ has relative density
+$\delta_{y}(m)\;=\;\frac{1}{\varphi(m)}\prod_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\nmid m\end{subarray}}\frac{\ell-2}{\ell-1}$
+among the primes, in the sense that
+$\lim_{x\to\infty}\frac{1}{\pi(x)}\#\bigl\{p\leq x:p\textup{ prime},\;a_{y}(p-1)=m\bigr\}\;=\;\delta_{y}(m).$
+Moreover, writing $C_{y}:=\prod_{3\leq\ell\leq y}(\ell-2)/(\ell-1)\asymp 1/\!\log y$, we have $\delta_{y}(m)\geq C_{y}/\varphi(m)$.
+Proof.
+For each odd prime $\ell\leq y$ with $\ell\mid m$, the condition $\ell\mid(p-1)$ forces $p\equiv 1\pmod{\ell}$.
+This selects one residue class out of $\ell-1$ reduced classes modulo $\ell$.
+For each odd prime $\ell\leq y$ with $\ell\nmid m$, the condition $\ell\nmid(p-1)$ excludes the class $p\equiv 1\pmod{\ell}$, leaving $\ell-2$ admissible classes out of $\ell-1$.
+The prime $\ell=2$ imposes no constraint since $2\mid(p-1)$ for every odd prime $p$ and $2\mid m$ by hypothesis.
+By the Chinese Remainder Theorem, these local conditions combine to specify a union of reduced residue classes modulo the primorial $M_{y}:=\prod_{\ell\leq y,\;\ell\text{ prime}}\ell$.
+The finitely many primes $p\leq y$ need not satisfy these residue-class conditions, but they contribute $O_{y}(1)$ to the count and do not affect the relative density.
+Since $M_{y}$ is fixed for each $y$, the Siegel–Walfisz theorem (Theorem 5) gives the stated density.
+Since each $m$ is even and squarefree, $\varphi(m)=\prod_{\begin{subarray}{c}\ell\mid m\\ \ell>2\end{subarray}}(\ell-1)$, so the first product in the expression for $\delta_{y}(m)$ equals $1/\varphi(m)$.
+The lower bound follows because each factor $(\ell-2)/(\ell-1)\in(0,1)$.
+To justify the asymptotic $C_{y}\asymp 1/\!\log y$, write
+$\log C_{y}=\sum_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\text{ prime}\end{subarray}}\log\frac{\ell-1}{\ell-2}=\sum_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\text{ prime}\end{subarray}}\frac{1}{\ell}+O(1),$
+since $\log((\ell-1)/(\ell-2))=1/\ell+O(1/\ell^{2})$ and $\sum_{\ell}1/\ell^{2}<\infty$.
+By Mertens’ second theorem the sum over primes equals $\log\log y+O(1)$, so $C_{y}=e^{-\log\log y+O(1)}\asymp 1/\!\log y$.
+∎
+Lemma 8 (Tail-mean estimate for f(p1)f(p-1)).
+Let $f(n):=\log(n/\varphi(n))$, and for each prime $p$ and parameter $y\geq 3$ define
+$T_{y}(p):=\sum_{\begin{subarray}{c}\ell\mid(p-1)\\ \ell>y,\;\ell\textup{ prime}\end{subarray}}\log\frac{\ell}{\ell-1}.$
+Then, for each fixed $y\geq 3$,
+$\lim_{x\to\infty}\frac{1}{\pi(x)}\sum_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}T_{y}(p)\;=\;\eta(y),$
+where $\eta(y):=\sum_{\ell>y,\;\ell\textup{ prime}}f(\ell)/(\ell-1)\ll 1/y$. In particular, $\eta(y)\to 0$ as $y\to\infty$.
+Proof.
+Since $f(\ell)=1/\ell+O(1/\ell^{2})$, we have $\eta(y)=\sum_{\ell>y}1/(\ell(\ell-1))+O(\sum_{\ell>y}1/\ell^{3})\ll 1/y$. Interchanging sums gives
+$\sum_{p\leq x}T_{y}(p)\;=\;\sum_{\begin{subarray}{c}y<\ell<x\\ \ell\text{ prime}\end{subarray}}f(\ell)\,\pi(x;\ell,1).$
+Fix $A:=2$ and split the sum at the threshold $L:=(\log x)^{A}$.
+Small primes ($y<\ell\leq L$). Each such $\ell$ satisfies $\ell\leq(\log x)^{A}$, so by the Siegel–Walfisz theorem (Theorem 5),
+$\pi(x;\ell,1)\;=\;\frac{\operatorname{Li}(x)}{\ell-1}\;+\;E(x;\ell,1),\qquad|E(x;\ell,1)|\;\ll_{A}\;xe^{-c_{A}\sqrt{\log x}}.$
+Since $f(\ell)\leq 1$ for every prime $\ell$ and $\{\text{primes }\ell\in(y,L]\}\leq L$, the total error is
+$\sum_{y<\ell\leq L}f(\ell)\,|E(x;\ell,1)|\;\leq\;L\cdot O_{A}\!\bigl(xe^{-c_{A}\sqrt{\log x}}\bigr)\;=\;o\bigl(\pi(x)\bigr).$
+The main term from this range is $\operatorname{Li}(x)\sum_{y<\ell\leq L}f(\ell)/(\ell-1)=\operatorname{Li}(x)\bigl(\eta(y)+O(1/L)\bigr)$, since the tail $\sum_{\ell>L}f(\ell)/(\ell-1)\ll 1/L$.
+Large primes ($\ell>L$). For every prime $\ell\leq x$, the trivial bound $\pi(x;\ell,1)\leq x/\ell+1$ and $f(\ell)=\log(\ell/(\ell-1))\leq 2/\ell$ give
+$\sum_{\ell>L}f(\ell)\,\pi(x;\ell,1)$ $\;\leq\;\sum_{L<\ell\leq x}\frac{2}{\ell}\left(\frac{x}{\ell}+1\right)\;\leq\;2x\sum_{\ell>L}\frac{1}{\ell^{2}}\;+\;2\sum_{\ell\leq x}\frac{1}{\ell}\;\ll\;\frac{x}{L}\;+\;\log\log x.$
+Dividing by $\pi(x)\sim x/\!\log x$, this contributes $O(\log x/L)+o(1)=o(1)$. Combining both ranges and dividing by $\pi(x)$ yields the claim. ∎
+Lemma 9 (Shifted-prime pure singularity for $\log(n/\varphi(n))$).
+Let
+$f(n)\;:=\;\log\frac{n}{\varphi(n)}\;=\;\sum_{\begin{subarray}{c}\ell\mid n\\ \ell\textup{ prime}\end{subarray}}\log\frac{\ell}{\ell-1}.$
+The shifted-prime limiting measure $\mu_{f}$ is purely singular with respect to Lebesgue measure.
+Here $\mu_{f}$ denotes the weak limit of the empirical measures $\pi(x)^{-1}\sum_{p\leq x}\delta_{f(p-1)}$.
+Proof.
+We adapt the integer-case argument of Erdős [5] (see also [13, Exercise 256] with solution in [17, Solution 2.4, pp.
+147–148]) to the shifted-prime setting, using the Siegel–Walfisz theorem (Theorem 5) in place of the integer equidistribution input.
+By the shifted-prime Erdős–Wintner theorem (Tenenbaum [3, pp.
+257–258] for the shift a=1a=-1; Kátai [14], Hildebrand [15] for the original $a=1$ case), the limiting measure $\mu_{f}$ exists.
+Indeed, by Lemma 3 the three convergence hypotheses are satisfied and the continuity criterion holds, so the limiting measure $\mu_{f}$ exists and is continuous.
+By Lemma 6, $\mu_{f}$ is of pure type.
+Since it is continuous (as verified above), it is therefore either purely singular or absolutely continuous.
+We show that $\mu_{f}$ is not absolutely continuous, which forces pure singularity.
+Step 1: Tail estimate.
+Throughout the proof, the single prime $p=2$ is negligible in any asymptotic density statement.
+Fix a parameter $\delta\in(0,\tfrac{1}{2})$ and, for $y\geq 3$, set $R:=\lfloor y^{\delta}\rfloor$.
+Since $f$ is strongly additive, for each prime $p$ we decompose $f(p-1)=f(a_{y}(p-1))+T_{y}(p)$, where $a_{y}(p-1)$ and $T_{y}(p)$ are as in Lemma 8.
+By that lemma and Markov’s inequality,
+$\limsup_{x\to\infty}\;\frac{1}{\pi(x)}\#\bigl\{p\leq x:T_{y}(p)>1/R^{2}\bigr\}\;\leq\;R^{2}\,\eta(y)\;\ll\;y^{2\delta-1},$ (1)
+which tends to 0 as $y\to\infty$ (since $\delta<\tfrac{1}{2}$).
+Here and in Steps 2–3, the parameter $y$ is held fixed while $x\to\infty$; the conclusions are then strengthened by letting $y\to\infty$.
+Step 2: A set of small Lebesgue measure carrying positive relative prime density.
+Let $m_{1}<m_{2}<\cdots<m_{R}$ be the first $R=\lfloor y^{\delta}\rfloor$ even squarefree positive integers.
+Since the counting function of even squarefree integers satisfies $A(t):=\#\{m\leq t:m\text{ even squarefree}\}\sim(2/\pi^{2})\,t$, for any fixed $C>\pi^{2}/2$ we have $A(Cj)\geq j$ for all sufficiently large $j$.
+After enlarging $C$ if necessary, we may assume $m_{j}\leq Cj$ for all $j\geq 1$.
+In particular $m_{R}\leq CR$.
+Therefore $P^{+}(m_{j})\leq m_{j}\leq Cy^{\delta}<y$ for all sufficiently large $y$ (since $\delta<1$).
+Moreover,
+$\sum_{j=1}^{R}\frac{1}{\varphi(m_{j})}\;\geq\;\sum_{j=1}^{R}\frac{1}{m_{j}}\;\geq\;\frac{1}{C}\sum_{j=1}^{R}\frac{1}{j}\;\gg\;\log R\;=\;\delta\log y+O(1).$ (2)
+By Lemma 7, the relative density of $\{p:a_{y}(p-1)=m_{j}\}$ among the primes is $\delta_{y}(m_{j})\geq C_{y}/\varphi(m_{j})$, where $C_{y}\asymp 1/\!\log y$. Combining with (2),
+$\sum_{j=1}^{R}\delta_{y}(m_{j})\;\geq\;C_{y}\sum_{j=1}^{R}\frac{1}{\varphi(m_{j})}\;\gg\;\frac{\delta\log y}{\log y}\;=\;\delta\;>\;0.$ (3)
+Since the $m_{j}$ are distinct, the sets $\{p:a_{y}(p-1)=m_{j}\}$ are pairwise disjoint, so the relative density of their union $\mathcal{A}_{y}:=\{p:a_{y}(p-1)\in\{m_{1},\ldots,m_{R}\}\}$ is the sum $\sum_{j}\delta_{y}(m_{j})$.
+In particular, there exists an absolute constant $c_{0}>0$ (depending only on $\delta$) such that for all sufficiently large $y$, the relative density of $\mathcal{A}_{y}$ among the primes is at least $c_{0}$.
+Define the closed set
+$\bar{I}_{y}\;:=\;\bigcup_{j=1}^{R}\bigl[\,f(m_{j}),\;f(m_{j})+1/R^{2}\,\bigr].$
+By subadditivity of Lebesgue measure, $\lambda(\bar{I}_{y})\leq R\cdot(1/R^{2})=1/R=1/\lfloor y^{\delta}\rfloor\to 0$ as $y\to\infty$.
+For any prime $p$ with $a_{y}(p-1)=m_{j}$ and $T_{y}(p)\leq 1/R^{2}$, the identity $f(p-1)=f(a_{y}(p-1))+T_{y}(p)$ and the nonnegativity of $T_{y}(p)$ give $f(p-1)\in[f(m_{j}),\,f(m_{j})+1/R^{2}]\subset\bar{I}_{y}$.
+Step 3: Contradiction with absolute continuity. Recall that $y$ (and hence $R$) are still fixed. By (1),
+$\limsup_{x\to\infty}\;\frac{1}{\pi(x)}\#\bigl\{p\leq x:T_{y}(p)>1/R^{2}\bigr\}\;=\;o_{y}(1).$
+Hence, for all sufficiently large $y$,
+$\liminf_{x\to\infty}\frac{1}{\pi(x)}\#\bigl\{p\leq x:f(p-1)\in\bar{I}_{y}\bigr\}\;\geq\;c_{0}-o_{y}(1)\;\geq\;\frac{c_{0}}{2}\;>\;0.$
+[…]
+$\mu_{f}(\bar{I}_{y})\;=\;\lim_{x\to\infty}\frac{1}{\pi(x)}\#\bigl\{p\leq x:f(p-1)\in\bar{I}_{y}\bigr\}\;\geq\;\frac{c_{0}}{2}\;>\;0.$
+If $\mu_{f}$ were absolutely continuous with density $h\in L^{1}(\mathbb{R})$, then by the absolute continuity of the Lebesgue integral, $\mu_{f}(\bar{I}_{y})=\int_{\bar{I}_{y}}h\,\mathrm{d}t\to 0$ as $\lambda(\bar{I}_{y})\to 0$, contradicting $\mu_{f}(\bar{I}_{y})\geq c_{0}/2$ for all large $y$.
+Therefore $\mu_{f}$ is not absolutely continuous, and being of pure type and continuous, it is purely singular.
+∎
 Proposition 3 (Shifted-prime limiting law for $f$).
@@
 $f(n):=\log\frac{n}{\varphi(n)}=\sum_{\begin{subarray}{c}\ell\mid n\\ \ell\textup{ prime}\end{subarray}}\log\frac{\ell}{\ell-1}.$
-Then the values f(p1)f(p-1), as $p$ ranges over the primes, possess a continuous purely singular limiting distribution. More precisely, the empirical measures
+Then the empirical measures
 $\mu_{x}:=\frac{1}{\pi(x)}\sum_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}\delta_{f(p-1)}$
@@
 Proof.
-The shifted-prime analogue of the Erdős–Wintner theorem, due independently to Kátai [17] and Hildebrand [18], shows that a strongly additive function $h$ possesses a limiting distribution on the shifted primes p1p-1 if and only if the three classical Erdős–Wintner series
-$\sum_{\begin{subarray}{c}\ell\ \textup{prime}\\ |h(\ell)|\leq 1\end{subarray}}\frac{h(\ell)}{\ell},\qquad\sum_{\begin{subarray}{c}\ell\ \textup{prime}\\ |h(\ell)|\leq 1\end{subarray}}\frac{h(\ell)^{2}}{\ell},\qquad\sum_{\begin{subarray}{c}\ell\ \textup{prime}\\ |h(\ell)|>1\end{subarray}}\frac{1}{\ell}$
-converge; see Tenenbaum [2, pp. 257–258]. In the same reference it is noted that the shifted-prime limiting law is of pure type and is continuous if and only if
-$\sum_{\begin{subarray}{c}\ell\ \textup{prime}\\ h(\ell)\neq 0\end{subarray}}\frac{1}{\ell}=\infty.$
-We apply this to the strongly additive function
-$f(n)=\sum_{\begin{subarray}{c}\ell\mid n\\ \ell\textup{ prime}\end{subarray}}\log\frac{\ell}{\ell-1}.$
-For every prime $\ell$,
-$0<f(\ell)=\log\frac{\ell}{\ell-1}\leq\log 2<1,\qquad f(\ell)=\frac{1}{\ell}+O\!\left(\frac{1}{\ell^{2}}\right).$
-Hence
-$\sum_{\ell}\frac{f(\ell)}{\ell}<\infty,\qquad\sum_{\ell}\frac{f(\ell)^{2}}{\ell}<\infty,\qquad\sum_{|f(\ell)|>1}\frac{1}{\ell}=0.$
-Therefore the values f(p1)f(p-1), as $p$ ranges over the primes, possess a limiting distribution $\mu_{f}$.
-Since $f(\ell)\neq 0$ for every prime $\ell$ and $\sum_{\ell}1/\ell=\infty$, the limiting law $\mu_{f}$ is continuous.
-Invoking the shifted-prime extension of Erdős’ singularity criterion (Tenenbaum [2, p.
-258]; see also [3, p.
-77]): for a strongly additive function $h$, once the limiting law of h(p1)h(p-1) exists, the condition
-$\sum_{\ell\ \textup{prime}}\frac{h(\ell)^{2}}{\ell}<\infty$
-forces that limiting law to be singular.
-Applying this with $h=f$, and using the already verified convergence of $\sum_{\ell}f(\ell)^{2}/\ell$, we conclude that $\mu_{f}$ is singular.
-Since $\mu_{f}$ is also continuous, it is purely singular.
-∎
+By Lemma 3, the three convergence conditions of the shifted-prime Erdős–Wintner theorem (Kátai [14], Hildebrand [15] for $p+1$; Tenenbaum [3, pp.
+257–258] for the general shift including p1p-1) are satisfied, and the continuity criterion holds.
+Therefore f(p1)f(p-1) possesses a continuous limiting distribution $\mu_{f}$, which is of pure type by Lemma 6.
+It remains to rule out absolute continuity. By Lemma 9, $\mu_{f}$ is purely singular. ∎
 Corollary 2 (Structure of the limiting PRIM-LWE law).
@@
 Therefore $T_{\#}\mu_{f}$ is supported on the null set $T(E)$ and is therefore purely singular.
-By Theorem 7, the same is true for $c(p)$.
-Since the limiting law has support containing more than one point, it is not a point mass, and therefore $c(p)$ does not converge to a constant along the primes.
-∎
+By Theorem 7, the same is true for $c(p)$. ∎
 Note that $c(p)<1/2$ for every fixed prime $p$. Indeed, for $p=2$ one has
@@
 Proof.
-The lower-bound argument is the same primorial/Mertens argument used in the proof of Corollary 5, but now with Theorem 8 replacing the crude bound $P_{n}(p)\geq\tfrac{1}{2}$.
-Let
-$m(x):=\max\{k\geq 1:N_{k}\leq x\}.$
-Since $c(2)>0$, it suffices for large $x$ to consider odd primes $p\leq x$. For such $p$,
-$N_{\omega(p-1)}\leq\prod_{r\mid(p-1)}r\leq p-1\leq x,$
-so $\omega(p-1)\leq m(x)$. Therefore Theorem 8 gives
-$c(p)\geq P_{3}\prod_{i=1}^{\omega(p-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr)\geq P_{3}\prod_{i=1}^{m(x)}\Bigl(1-\frac{1}{p_{i}}\Bigr).$
-As in the proof of Corollary 5, the relation
-$N_{m(x)}\leq x<N_{m(x)+1}=N_{m(x)}\,p_{m(x)+1}$
-implies $p_{m(x)}\sim\log x$. Hence, by Mertens’ theorem,
-$\prod_{i=1}^{m(x)}\Bigl(1-\frac{1}{p_{i}}\Bigr)\sim\frac{e^{-\gamma}}{\log p_{m(x)}}\asymp\frac{1}{\log\log x}.$
-Therefore
-$\min_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}c(p)\gg\frac{1}{\log\log x}.$
-The matching upper bound is Proposition 1. ∎
-The lower bound in Corollary 6 may be viewed as the prime-shifted specialization n=p1n=p-1 of the classical universal estimate
-$\frac{\varphi(n)}{n}\gg\frac{1}{\log\log n},$
-(see, e.g., Rosser and Schoenfeld [19]) together with the uniform bound $\prod_{j\geq 1}(1-p^{-j})\geq\tfrac{1}{2}$.
-The proof above keeps the primorial structure explicit, which makes it parallel to the upper-bound construction of Proposition 1.
+The lower bound follows by the same primorial/Mertens argument used in the proof of Corollary 5: for every odd prime $p\leq x$, $\omega(p-1)\leq m(x)$, and Lemma 1 together with Mertens’ theorem gives $\min_{p\leq x,\,p\textup{ prime}}c(p)\gg 1/\!\log\log x$.
+Theorem 8 sharpens the constant from $\tfrac{1}{2}$ to $P_{3}$ but does not affect the asymptotic order.
+The matching upper bound is Proposition 1.
+∎
 Corollary 7 (Pointwise universal lower bound).
@@
 Remark 1 (On effectivity).
-The comparison
-$\min_{p\leq x}c(p)\asymp\frac{1}{\log\log x}$
-should in principle be capable of being made fully explicit by combining explicit estimates for the Mertens product and for $\vartheta(x)$ (see, e.g., Rosser–Schoenfeld [19] and Dusart [20]) with an explicit least-prime-in-an-arithmetic-progression bound of Linnik type; see also Xylouris [21] for an admissible Linnik exponent.
-We do not carry out that bookkeeping here, since the present note only needs the qualitative asymptotic order.
+The asymptotic $\min_{p\leq x}c(p)\asymp 1/\!\log\log x$ should be capable of being made fully explicit by combining explicit Mertens product and $\vartheta(x)$ estimates (e.g., Rosser–Schoenfeld [18], Dusart [19]) with an explicit Linnik-type bound (see Xylouris [20]).
+We do not pursue the bookkeeping here.
 Remark 2.
-The bound $1/2$ for the infinite product can be sharpened for specific primes.
-For instance, $\prod_{j\geq 1}(1-q^{-j})\geq 1-1/(q-1)$, and this lower bound exceeds $2/3$ for $q\geq 5$ and is at least $0.99$ for $q\geq 101$.
-For large implementation-oriented NTT primes (for example $q>2^{30}$), this factor is numerically indistinguishable from $1$.
-The current NIST-standardized lattice moduli $q=3329$ and $q=8380417$ are much smaller [4, 5], but even there the same elementary bound already gives $\prod_{j\geq 1}(1-q^{-j})>0.9996$ and $\prod_{j\geq 1}(1-q^{-j})>0.9999998$, respectively.
+The bound $1/2$ for the infinite product can be sharpened for specific primes: $\prod_{j\geq 1}(1-q^{-j})\geq 1-1/(q-1)$, which exceeds $2/3$ for $q\geq 5$ and is at least $0.99$ for $q\geq 101$.
+For large NTT primes with $q>2^{30}$, this factor is numerically indistinguishable from 11.
+Even for the smaller NIST moduli $q=3329$ and $q=8380417$ [21, 22], the elementary bound gives $\prod_{j\geq 1}(1-q^{-j})>0.9996$ and $\prod_{j\geq 1}(1-q^{-j})>0.9999998$, respectively.
 We now instantiate the general $\omega(q-1)$-based bound in several parameter regimes relevant to cryptographic applications, including many NTT-friendly moduli.
@@
 Remark 3 (On the case $\omega(q-1)=1$).
-If $\omega(q-1)=1$ and $q$ is an odd prime, then q1q-1 is an even prime power, hence $q-1=2^{m}$ for some $m\geq 1$; equivalently, $q$ is a Fermat prime. In this case
+If $\omega(q-1)=1$ and $q$ is an odd prime, then q1q-1 is an even prime power, hence $q-1=2^{m}$ for some $m\geq 1$. Equivalently, $q$ is a Fermat prime. In this case
 $\frac{\varphi(q-1)}{q-1}=\frac{1}{2},\qquad c(q)=\frac{1}{2}\,P(q),\qquad P(q):=\prod_{j\geq 1}\Bigl(1-\frac{1}{q^{j}}\Bigr).$
@@
 where the inequality holds because the right-hand side includes additional factors in $(0,1)$.
-By Dusart’s explicit product estimate [20, Theorem 5.9], which we use conservatively with threshold $x\geq 2\,278\,383$,
+By Dusart’s explicit product estimate [19, Theorem 5.9], which we use conservatively with threshold $x\geq 2\,278\,383$,
 $\prod_{\begin{subarray}{c}r\leq x\\ r\text{ prime}\end{subarray}}\Bigl(1-\frac{1}{r}\Bigr)\;=\;\frac{e^{-\gamma}}{\log x}\,\Bigl(1+O^{*}\!\Bigl(\frac{0.2}{\log^{3}x}\Bigr)\Bigr).$
-Since $q>2^{30}$, we have q1>2 278 383q-1>2\,278\,383 and $\log(q-1)>20$, so the parenthetical factor lies in $(1-2.5\times 10^{-5},\;1+2.5\times 10^{-5})$.
+Since $q>2^{30}$, we have q1>2 278 383q-1>2\,278\,383 and $\log(q-1)>20$, so the correction factor lies in $(1-2.5\times 10^{-5},\;1+2.5\times 10^{-5})$.
 Also,
@@
 using $e^{-\gamma}>0.5614$, $\log(q-1)<\log q$, and $0.2/\log^{3}(q-1)<2.5\times 10^{-5}$ for $q>2^{30}$. ∎
-The bound $c(q)>0.56/\!\log q$ of Corollary 8 is a simple, fully explicit estimate obtained by taking the Mertens product over all primes up to q1q-1. However, the sharper universal pointwise bound
-$c(q)\;\geq\;P_{3}\prod_{i=1}^{\omega(q-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr)\;\geq\;P_{3}\prod_{i=1}^{m(q-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr)\;\gg\;\frac{1}{\log\log q}$
-follows directly from Theorem 8 together with $\omega(q-1)\leq m(q-1)$ and Mertens’ theorem, where $m(n):=\max\{k:N_{k}\leq n\}$.
-This matches the sharp order established in Corollary 6 and shows that $1/\!\log\log q$, not $1/\!\log q$, is the correct worst-case scale for the rejection-sampling overhead.
+The bound $c(q)>0.56/\!\log q$ of Corollary 8 is a simple, fully explicit estimate obtained by taking the Mertens product over all primes up to q1q-1.
+The sharper universal pointwise bound $c(q)\gg 1/\!\log\log q$ of Corollary 7 shows that $1/\!\log\log q$, not $1/\!\log q$, is the correct worst-case scale for the rejection-sampling overhead.
 Remark 4 (Typical size of $\omega(q-1)$).
 Let $\Phi$ denote the standard normal distribution function.
-The shifted-prime analogue of the Erdős–Kac theorem (due to Barban [22]; see also Halberstam [23] and Elliott [24, Ch.
+The shifted-prime analogue of the Erdős–Kac theorem (due to Barban [23]; see also Halberstam [24] and Elliott [25, Ch.
 12]) asserts that for every real $z$,
@@
 For the two prime moduli appearing, across all parameter sets, in the current NIST standards for ML-KEM and ML-DSA [21, 22], Table 1 records both the general $\omega(q-1)$-based bounds and the sharper exact-factorization bounds.
-The fourth column records the density lower bound $B(\omega(q-1))$, so the corresponding general rejection-sampling overhead upper bound is its reciprocal $1/B(\omega(q-1))$; for the two rows of Table 1 these are $1/B(2)\approx 5.36$ and $1/B(4)\approx 7.82$, respectively.
-The exact factorizations give
-$\frac{\varphi(3328)}{3328}=\frac{6}{13},\qquad\frac{\varphi(8380416)}{8380416}=\frac{100}{341},$
-so together with $P(q)\geq 1-\frac{1}{q-1}$ one obtains
+The exact factorizations give $\varphi(3328)/3328=6/13$ and $\varphi(8380416)/8380416=100/341$.
+Together with $P(q)\geq 1-\frac{1}{q-1}$, one obtains
 $c(3329)\geq\frac{6}{13}\Bigl(1-\frac{1}{3328}\Bigr)\approx 0.4614,\qquad c(8380417)\geq\frac{100}{341}\Bigl(1-\frac{1}{8380416}\Bigr)\approx 0.2933.$
-Direct evaluation of the defining product shows that these lower bounds are already essentially exact:
-$c(3329)=\frac{6}{13}\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{3329^{j}}\Bigr)\approx 0.46139978,\qquad c(8380417)=\frac{100}{341}\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{8380417^{j}}\Bigr)\approx 0.29325510.$
-Equivalently, the corresponding dimension-uniform expected rejection-sampling overheads $1/c(q)$ are approximately $2.1674$ and $3.4101$.
-For each fixed dimension $n$, the exact overheads $1/c_{n}(q)$ are no larger.
+A short tail-bound argument (taking $J=2$ partial products of $P(q)$ and using the estimate $P_{J}(q)-P(q)\leq q^{-J}/(q-1)$) confirms that these lower bounds are tight to many decimal places:
+$c(3329)=0.4613997781\ldots,\qquad c(8380417)=0.2932550969\ldots.$
+The corresponding dimension-uniform expected rejection-sampling overheads $1/c(q)$ are approximately $2.17$ and $3.41$. For each fixed dimension $n$, the exact overheads $1/c_{n}(q)$ are no larger.
+Context $q$ factorization of q1q-1 $\omega(q-1)$ $1/c(q)$
+ML-KEM (FIPS 203) $3329$ $^{8}\cdot 13 22$ $2.17$
+ML-DSA (FIPS 204) $8380417$ $2^{13}\cdot 3\cdot 11\cdot 31$ 44 $3.41$
+NewHope / FALCON [^1] FALCON is being standardized by NIST as FN-DSA (FIPS 206). [26, 27] $12289$ $^{12}\cdot 3 22$ $3.00$
+BFV/BGV FHE (small) [28, 29, 30] $786433$ $^{18}\cdot 3 22$ $3.00$
+SEAL BFV default [31] $132120577$ $2^{21}\cdot 3^{2}\cdot 7$ 33 $3.50$
+BFV/BGV FHE (medium) [28, 29, 30] $2013265921$ $2^{27}\cdot 3\cdot 5$ 33 $3.75$
+NTTRU / BLISS [32, 33] $7681$ $2^{9}\cdot 3\cdot 5$ 33 $3.75$
+64-bit NTT prime [34] $2^{64}-2^{32}+1$ $2^{32}\cdot 3\cdot 5\cdot 17\cdot 257\cdot 65537$ 66 $4.00$
+Fermat prime ($2^{16}+1$) $65537$ $2^{16}$ 11 $2.00$
+Table 2: Representative NTT-friendly primes from lattice-based cryptography and FHE implementations, together with $\omega(q-1)$ and the dimension-uniform expected rejection-sampling overhead $1/c(q)$.
+All overheads are at most $4.00$.
+The first two rows reproduce the NIST entries from Table 1.
+The SEAL entry is the default coefficient modulus for poly_modulus_degree = 1024 at the 128-bit security level.
+Table 2 collects nine representative moduli drawn from current lattice-based and FHE implementations.
+Every modulus in the table satisfies $\omega(q-1)\leq 6$ and incurs a dimension-uniform rejection-sampling overhead $1/c(q)\leq 4.00$.
+The table substantiates the claim that the primes typically encountered in practice have controlled factorization structure, which yields a good lower bound on $c(q)$.
 Remark 5 (Concrete example).
-Consider $q=2^{64}-2^{32}+1=18\,446\,744\,069\,414\,584\,321$, a well-known NTT-friendly prime.
-This modulus is not the native modulus of the standardized NIST-PQC schemes, but it is used as a 64-bit NTT-friendly ambient field in implementation-oriented work on lattice cryptography, so it is a natural cryptographic test case [25].
-One verifies that $q-1=2^{32}\cdot 3\cdot 5\cdot 17\cdot 257\cdot 65537$, so $\omega(q-1)=6$ and
-$\frac{\varphi(q-1)}{q-1}=\frac{1}{2}\cdot\frac{2}{3}\cdot\frac{4}{5}\cdot\frac{16}{17}\cdot\frac{256}{257}\cdot\frac{65536}{65537}\approx 0.2500.$
-Since $P(q)>1-10^{-19}$ for such large $q$, the limiting density $c(q)$ is about $0.25$.
-The exact finite-$n$ density $c_{n}(q)$ is then also essentially $0.25$, so about one in four random $n\times n$ matrices has a primitive-root determinant, and the security reduction loses a factor of about $4$.
+For the 64-bit NTT prime $q=2^{64}-2^{32}+1$ (sometimes called the “Goldilocks prime” in the zero-knowledge community, not to be confused with the Ed448-Goldilocks prime $2^{448}-2^{224}-1$ [35]), used as an NTT-friendly ambient field in implementation-oriented lattice cryptography [34], one has $\omega(q-1)=6$ and $\varphi(q-1)/(q-1)\approx 0.2500$.
+Since $P(q)>1-10^{-19}$, one obtains $c(q)\approx 0.25$, so roughly a quarter of random $n\times n$ matrices have a primitive-root determinant and the security reduction loses a factor of about $4$.
 7 Implications for PRIM-LWE
@@
 In our notation, that dimension-uniform sampling constant is
-$c=\inf_{m\geq 1}c_{m}(p).$
-Since
-$c_{m}(p)=\frac{\varphi(p-1)}{p-1}\prod_{j=1}^{m}\Bigl(1-\frac{1}{p^{j}}\Bigr)$
-is decreasing in $m$, it follows that
-$\inf_{m\geq 1}c_{m}(p)=\lim_{m\to\infty}c_{m}(p)=c(p).$
-Thus $c(p)$ is the dimension-uniform sampling term entering the reduction. The present note concerns this quantity $c(p)$; it does not alter the separate $O(n^{2})$ factor from [1, Theorem 10].
+$c=\inf_{n\geq 1}c_{n}(p).$
+As shown in Section 2, $c_{n}(p)$ is decreasing in $n$ and $\inf_{n\geq 1}c_{n}(p)=c(p)$, so $c(p)$ is the dimension-uniform sampling term entering the reduction.
+The present note concerns this quantity $c(p)$; it does not alter the separate $O(n^{2})$ factor from [1, Theorem 10].
 Theorem 3 shows that no uniform lower bound on $c(p)$ holds across all primes, meaning the reduction tightness degrades for primes $p$ whose predecessor p1p-1 is divisible by many small primes.
@@
 In fixed dimension $n$, the exact expected number of trials is $1/c_{n}(p)\leq 1/c(p)$.
-By [1, Theorem 10], the reduction from decision-LWE to decision-prim-lwe introduces an advantage loss of $O(n^{2})\cdot(1/c)$, where $c=\inf_{m\geq 1}c_{m}(p)=c(p)$ is the dimension-uniform sampling constant.
+By [1, Theorem 10], the reduction from decision-LWE to decision-prim-lwe introduces an advantage loss of $O(n^{2})\cdot(1/c)$, where $c=\inf_{n\geq 1}c_{n}(p)=c(p)$ is the dimension-uniform sampling constant.
 Our results therefore show that the $1/c(p)$ factor in this bound can grow like $\Omega(\log\log p)$ along a suitable subsequence of primes.
@@
 For any fixed prime $p$, the constant $c(p)>0$ and the reduction remains valid.
-Moreover, the explicit bounds of Section 6 show that standard choices of LWE moduli yield $c(q)$ concretely bounded away from zero.
-For safe primes $q=2r+1$ with $r$ an odd prime (i.e.
-$q\geq 7$), one has
-$c(q)=\frac{r-1}{2r}\prod_{j\geq 1}\left(1-q^{-j}\right).$
-For a safe prime $q=2r+1$, the prefactor (r1)/(2r)=1/21/(q1)(r-1)/(2r)=1/2-1/(q-1) is already close to $1/2$ when $q$ is large, and the Euler product $\prod_{j\geq 1}(1-q^{-j})$ tends to $1$ as $q\to\infty$.
-Hence $c(q)$ is close to $1/2$ for large safe primes.
-Note, however, that the conditional safe-prime and Fermat-prime observations are now superseded by Corollary 3, which gives $\limsup c(p)=1/2$ unconditionally.
-On the other hand, if
-$P(t):=\prod_{j\geq 1}\Bigl(1-t^{-j}\Bigr)\qquad(t>1),$
-then $P(t)$ is increasing in $t$, so for every safe prime $q\geq 7$ we have $P(q)\geq P(3)=P_{3}$. Therefore
-$c(q)\geq P_{3}\,\frac{r-1}{2r}.$
+Moreover, the explicit bounds of Section 6 show that the specific moduli surveyed in Table 2—which cover the NIST standards and several widely deployed FHE primes—all have $c(q)$ concretely bounded away from zero.
+Corollary 3 gives $\limsup c(p)=1/2$ unconditionally, so there is no shortage of primes for which $c(p)$ is close to its maximal value.
+As a concrete illustration, for safe primes $q=2r+1$ with $r$ an odd prime and $q\geq 7$, one has
+$c(q)=\frac{r-1}{2r}\prod_{j\geq 1}\left(1-q^{-j}\right)\geq P_{3}\,\frac{r-1}{2r}.$
+The prefactor (r1)/(2r)=1/21/(q1)(r-1)/(2r)=1/2-1/(q-1) is already close to $1/2$ when $q$ is large, and the Euler product $\prod_{j\geq 1}(1-q^{-j})$ tends to $1$ as $q\to\infty$, so $c(q)$ approaches $1/2$ for large safe primes.
 More generally, Corollary 4 shows that for every fixed $\tau\in(0,1/2)$, both sets of primes
@@
 Thus moduli yielding a very loose constant-factor sampling term are rare only in the threshold-asymptotic sense $\tau\downarrow 0$, not absent for any fixed threshold $\tau<1/2$.
-We note that the primes typically used in lattice-based cryptography are chosen for NTT compatibility (often $q\equiv 1\pmod{2^{t}}$ for moderate $t$).
-Mathematically, $q\equiv 1\pmod{2^{t}}$ controls only the 2-adic part of q1q-1; it does not by itself prevent q1q-1 from having many odd prime factors, and NTT-friendliness alone does not imply a good lower bound on $c(q)$.
+We note that the primes typically used in lattice-based cryptography are chosen for NTT compatibility.
+Concretely, this typically means $q\equiv 1\pmod{2^{t}}$ for moderate $t$.
+Mathematically, this condition controls only the 2-adic part of q1q-1.
+It does not by itself prevent q1q-1 from having many odd prime factors, and NTT-friendliness alone does not imply a good lower bound on $c(q)$.
 What our theorems actually use is the controlled factorization structure of q1q-1, quantified by $\omega(q-1)$ or by the exact factorization.
 The NTT condition is harmless with respect to the 2-part, since the large power of $2$ contributes only the single Euler factor $(1-\tfrac{1}{2})$ to $\varphi(q-1)/(q-1)$.
-In practice, many NTT primes of cryptographic interest additionally have few small odd prime factors dividing q1q-1, and it is this latter property—not NTT-friendliness per se—that yields the good lower bounds on $c(q)$ recorded in Table 1 and Remark 7.
+Among the implementation moduli surveyed in Table 2, every prime satisfies $\omega(q-1)\leq 6$.
+It is this controlled factorization that yields the good lower bounds on $c(q)$ recorded in Remark 5, and Tables 1 and 2.
 Thus our results do not by themselves indicate a practical weakness; they quantify how modulus choice affects the reduction loss in the prim-lwe framework.
@@
 In the ADP obfuscation framework, a Kilian-style randomization step left- and right-multiplies by invertible matrices $R,S$ with $\det(R)\det(S)=1$ to preserve the determinant polynomial.
-Here the ambient sampling space is $\mathrm{GL}_{n}(\mathbb{F}_{q})$, not all of $M_{n}(\mathbb{F}_{q})$.
+Because the ambient sampling space is $\mathrm{GL}_{n}(\mathbb{F}_{q})$ rather than all of $M_{n}(\mathbb{F}_{q})$, the relevant primitive-root fraction reduces to $\varphi(q-1)/(q-1)$: the matrix-count Euler product $\prod_{j=1}^{n}(1-q^{-j})$ is already absorbed into the $\mathrm{GL}_{n}$ normalization.
 Since the determinant map
@@
 If instead one samples from all matrices $M_{n}(\mathbb{F}_{q})$, then the exact finite-$n$ primitive-root density is $c_{n}(q)$, whose limit as $n\to\infty$ is $c(q)$.
-We do not know whether such a restriction improves security.
-We only note, for context, that the ADP framework of [26] includes left- and right-randomization by invertible matrices $R,S$ satisfying $\det(R)\det(S)=1$, while the cryptanalysis of [27] targets a weakness in one of the candidate randomization steps.
+While our results do not establish a security advantage for imposing primitive-root determinant constraints, they confirm that the associated sampling overhead is mild whenever q1q-1 has controlled factorization.
 8 Concluding Remarks
@@
 $\liminf_{\begin{subarray}{c}p\to\infty\\ p\text{ prime}\end{subarray}}c(p)=0.$
-The proof is short and unconditional, requiring only Dirichlet’s theorem on primes in arithmetic progressions and Mertens’ product formula; the primorial prime hypothesis is entirely unnecessary.
+The proof is short and unconditional, requiring only Dirichlet’s theorem and Mertens’ product formula.
 We have further established several results that illuminate the finer structure of the density $c(p)$:
@@
   2. 2.
-    The quantity $c(p)$ possesses a continuous purely singular limiting distribution over the primes, this law coincides with the classical shifted-prime distribution of $\varphi(p-1)/(p-1)$, and its support is exactly $0,\tfrac{1}{2}]$.
-Equivalently, the distribution function is strictly increasing on $0,\tfrac{1}{2}]$.
+    The quantity $c(p)$ has a continuous limiting distribution over the primes whose support is exactly $0,\tfrac{1}{2}]$; equivalently, the distribution function is strictly increasing on $0,\tfrac{1}{2}]$.
+This full-support conclusion appears to be new.
+The law coincides with the classical shifted-prime distribution of $\varphi(p-1)/(p-1)$, as shown by a transport lemma, and is purely singular.
+Pure singularity was already known in the integer case from Erdős [5]; Tenenbaum [3, p.
+263] noted that it carries over to shifted primes.
   3. 3.
-    For many practical moduli of cryptographic interest—including the current NIST-standardized moduli and other NTT-friendly primes—the density $c(q)$ admits an explicit positive lower bound depending only on $\omega(q-1)$.
+    For the implementation moduli surveyed in Table 2—including the current NIST-standardized moduli and several widely used FHE primes—the density $c(q)$ admits an explicit positive lower bound depending only on $\omega(q-1)$.
 Accordingly, the prim-lwe reduction has a finite, explicitly computable constant-factor sampling term for every fixed prime modulus, and this term is small in the concrete parameter regimes treated here.
-On the cryptographic side, our results show that the rejection-sampling overhead $1/c(p)$ can grow like $\Omega(\log\log p)$ along a suitable subsequence of primes, but this represents a quantitative sensitivity to modulus selection, not a qualitative failure of the reduction.
-For any fixed prime $p$, the constant $c(p)>0$ and the reduction remains valid.
-The primes typically chosen for lattice-based cryptography are NTT-friendly, meaning $q\equiv 1\pmod{2^{t}}$ for moderate $t$.
-While NTT-friendliness alone does not imply a good bound on $c(q)$, the NTT condition is harmless with respect to the 2-part of q1q-1 (the large power of $2$ contributes only the single Euler factor $(1-\tfrac{1}{2})$), and many practical NTT primes additionally have few small odd prime factors dividing q1q-1; it is this controlled factorization structure that yields the good concrete bounds on $c(q)$.
-For the two NIST-standardized moduli $q=3329$ (ML-KEM) and $q=8380417$ (ML-DSA), the dimension-uniform expected rejection-sampling overheads $1/c(q)$ are at most $2.17$ and $3.42$, respectively; in fixed dimension $n$, the exact expected overheads are $1/c_{n}(q)\leq 1/c(q)$.
-As a simple conservative explicit estimate, for any prime $q>2^{30}$ one has $1/c(q)\leq 1.79\log q$.
-The sharper universal pointwise bound is Corollary 7, which gives $1/c(q)=O(\log\log q)$; equivalently, the worst-case overhead among primes $p\leq x$ is of order $\Theta(\log\log x)$.
-Thus our results do not by themselves indicate a practical weakness.
-They quantify how modulus choice affects the reduction loss in the prim-lwe framework.
+On the cryptographic side, the rejection-sampling overhead $1/c(p)$ can grow like $\Omega(\log\log p)$ along a suitable subsequence of primes, but as discussed in Section 7, this represents a quantitative sensitivity to modulus selection rather than a structural obstruction.
+Table 2 confirms that every implementation modulus we examined has $\omega(q-1)\leq 6$ and $1/c(q)\leq 4.00$, and the sharper universal pointwise bound of Corollary 7 gives $1/c(q)=O(\log\log q)$.
 A natural remaining question is to characterize the limiting law $G$ more precisely—for instance, to understand its moments, its modulus of continuity, or its finer local regularity inside $0,\tfrac{1}{2}]$.
@@
 It therefore suffices to show that a positive proportion of primes in $\mathcal{A}_{R}(x)$ satisfy $S(p)\leq\varepsilon$.
-Step 3: Mean-value estimate for $S(p)$. Fix $Q=(\log x)^{2}$, and decompose $S(p)=S_{\leq Q}(p)+S_{>Q}(p)$, where
+Step 3: Mean-value estimate for $S(p)$.
+Throughout Steps 3 and 4, the parameter $R$ (and hence $\varepsilon$) is fixed; all limits are as $x\to\infty$.
+Fix $Q=(\log x)^{2}$, and decompose $S(p)=S_{\leq Q}(p)+S_{>Q}(p)$, where
 $S_{\leq Q}(p)=\sum_{\begin{subarray}{c}q\mid(p-1)\\ R<q\leq Q\end{subarray}}\frac{1}{q}.$
@@
 Since $q>R$, we have $q\nmid M_{R}$.
-The conditions $p\in\mathcal{A}_{R}(x)$ and $p\equiv 1\pmod{q}$ therefore define exactly $\nu_{R}$ reduced residue classes modulo $M_{R}q$: modulo each odd prime $r\leq R$ there are r2r-2 admissible reduced classes, modulo $2$ there is one, and modulo $q$ the residue class is forced to be $1$.
+The conditions $p\in\mathcal{A}_{R}(x)$ and $p\equiv 1\pmod{q}$ therefore define exactly $\nu_{R}$ reduced residue classes modulo $M_{R}q$: modulo each odd prime $r\leq R$ there are r2r-2 admissible reduced classes, modulo $2$ there is exactly a single admissible class, and modulo $q$ the residue class is forced to be $1$.
 Because $\varphi(M_{R}q)=\varphi(M_{R})(q-1)$, we obtain from Theorem 5 that
 $B_{q}(x)=\frac{\nu_{R}}{\varphi(M_{R}q)}\,\operatorname{Li}(x)+O_{R}\!\bigl(xe^{-c_{R}\sqrt{\log x}}\bigr)=\frac{\delta(R)}{q-1}\,\operatorname{Li}(x)+O_{R}\!\bigl(xe^{-c_{R}\sqrt{\log x}}\bigr),$
-uniformly for $R<q\leq Q$. (Here we use that $M_{R}$ is fixed and $q\leq(\log x)^{2}$, so $M_{R}q\leq(\log x)^{A}$ for some fixed $A$ and all sufficiently large $x$.)
+uniformly for $R<q\leq Q$. Here we use that $M_{R}$ is fixed and $q\leq(\log x)^{2}$, so $M_{R}q\leq(\log x)^{A}$ for some fixed $A$ and all sufficiently large $x$.
 Dividing by the asymptotic formula for $|\mathcal{A}_{R}(x)|$ yields

### 30 Mar 2026 v3

@@
 Sehrawat, Yeo, and Desmedt asked whether $\inf_{p\text{ prime}}c(p)=0$, observing that an affirmative answer would follow from the conjectural infinitude of primorial primes.
-We resolve this question unconditionally using only Dirichlet’s theorem and Mertens’ product formula, entirely bypassing the primorial-prime hypothesis.
+We resolve this question unconditionally using only Dirichlet’s theorem and Mertens’ product formula, bypassing the primorial-prime hypothesis.
 We further establish the sharp order
@@
 and show that the limiting distribution of $c(p)$ over the primes has support exactly $[0,1/2]$.
-We have not found this full-support statement in the literature.
-The law coincides with the classical shifted-prime distribution of $\varphi(p-1)/(p-1)$ via a transport lemma and is moreover continuous and purely singular.
+This full-support property appears to be new.
+The law coincides with the classical shifted-prime distribution of $\varphi(p-1)/(p-1)$ via a transport lemma, is moreover continuous and purely singular, and the limiting measure has Hausdorff dimension 0.
+The moments of the limiting distribution are given by convergent Euler products that uniquely determine it, and the associated Mellin transform extends analytically to the right half-plane.
 We also derive explicit lower bounds on $c(q)$ for primes of cryptographic interest, parameterized solely by the number of distinct prime factors of q1q-1.
-As a simple conservative explicit bound, for any prime $q>2^{30}$ the expected overhead $1/c(q)$ is at most $1.79\log q$.
+As a simple conservative explicit bound, for any prime $q>2^{30}$ the dimension-uniform expected overhead $1/c(q)$ is at most $1.79\log q$.
 On the other hand, our results show that the worst-case overhead among primes $p\leq x$ is of order $\Theta(\log\log x)$, and in particular $1/c(q)=O(\log\log q)$ pointwise.
@@
 The matching lower bound follows from two observations: if $p\leq x$ is an odd prime, then $\omega(p-1)\leq m(x):=\max\{k:N_{k}\leq x\}$, and Mertens’ theorem converts this into a $1/\!\log\log x$ lower bound for $\varphi(p-1)/(p-1)$; the additional factor $\prod_{j\geq 1}(1-p^{-j})\geq P_{3}$ is handled by Theorem 11.
-Hence the decay to zero is completely quantified and extremely slow.
-The implied constant is $\approx 0.31$, so primes $p$ whose predecessor p1p-1 is divisible by many small primes can still have $c(p)<0.1$, well within practical ranges.
+Hence the decay to zero is quantified precisely.
+The implied constant is $\approx 0.31$, so primes $p$ whose predecessor p1p-1 is divisible by many small primes can still have $c(p)<0.1$.
 The explicit bounds in Section 6 show that standard cryptographic moduli with controlled factorization incur only a small constant-factor overhead; see Remark 1 for effectivity.
 3.
-Limiting distribution (Theorems 6 and 7, Corollary 2).
-The quantity $c(p)$ possesses a continuous limiting distribution over the primes whose support is exactly $0,\tfrac{1}{2}]$; the law is moreover purely singular.
+Limiting distribution (Theorems 6 and 7, Corollary 2, Proposition 4).
+The quantity $c(p)$ possesses a continuous limiting distribution over the primes whose support is exactly $0,\tfrac{1}{2}]$; the law is moreover purely singular, and Proposition 4 shows that the limiting measure has Hausdorff dimension 0.
 If $G$ denotes the distribution function, then
@@
 263]; we include a self-contained proof for the reader’s convenience.
-Our new contributions on the distributional side are twofold.
-First, we establish that the support of the limiting law is exactly $0,\tfrac{1}{2}]$ by combining the left non-differentiability result of [4, Theorem 1.1] with the density of $\{\varphi(m)/m:m\text{ even}\}$ in $[0,1/2]$ (Lemma 4).
-Second, we transfer the classical shifted-prime law for $\varphi(p-1)/(p-1)$ to the prim-lwe density $c(p)$ via a transport lemma (Lemma 5).
+Our new contribution on the distributional side is the full-support conclusion: the limiting law has support exactly $0,\tfrac{1}{2}]$, established by combining the left non-differentiability result of [4, Theorem 1.1] with the density of $\{\varphi(m)/m:m\text{ even}\}$ in $[0,1/2]$ (Lemma 4).
+Since the Euler product $P(p)\to 1$, the prim-lwe density $c(p)=\frac{\varphi(p-1)}{p-1}P(p)$ shares the same limiting distribution (Lemma 5).
+We also derive the moments of $G$ as convergent Euler products (Theorem 8) and establish the analytic continuation of the Mellin transform $\mathbb{E}[X^{s}]$ to the right half-plane (Theorem 10).
+At the right endpoint, $1-G(\tfrac{1}{2}-\varepsilon)\sim\kappa/\!\log(1/\varepsilon)$ as $\varepsilon\downarrow 0$ for an explicit constant $\kappa$ (Proposition 7), so $G$ is not Hölder of any positive order at $1/2$.
 4.
@@
 We give unconditional lower bounds on $c(q)$ for primes $q$ parameterized by $\omega(q-1)$, the number of distinct prime factors of q1q-1, yielding concrete bounds for the NIST-standardized primes and other moduli whose predecessors have controlled factorization structure.
-The following consequence singles out the cryptographic implication.
+The following consequence singles out the cryptographic implication: in the reduction from decision-LWE to decision-prim-lwe over $\mathbb{F}_{q}$, the dimension-uniform expected rejection-sampling overhead $1/c(q)$ admits an explicit upper bound depending only on $\omega(q-1)$.
 Corollary 1 (Cryptographic consequence).
@@
 $P_{3}:=\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{3^{j}}\Bigr)\approx 0.5601260779.$
-Let $q\geq 3$ be a prime modulus with $\omega(q-1)\leq K$.
-Then, in the reduction from decision-LWE to decision-prim-lwe over $\mathbb{F}_{q}$, the dimension-uniform expected rejection-sampling overhead $1/c(q)$ satisfies
+Let $q\geq 3$ be a prime modulus with $\omega(q-1)\leq K$. Then
 $\frac{1}{c(q)}\;\leq\;\frac{1}{P_{3}}\prod_{i=1}^{K}\frac{p_{i}}{p_{i}-1}.$
-For each fixed dimension $n$, the exact expected overhead is $1/c_{n}(q)\leq 1/c(q)$.
 The general $\omega(q-1)$-based bound is immediate from Theorem 11; the sharper NIST figures use the exact factorizations recorded in Section 6.
@@
 Section 4 proves the upper bound $\min_{p\leq x}c(p)\ll 1/\log\log x$.
-Section 5 develops the limiting distribution theory.
+Section 5 develops the limiting distribution theory, including the Hausdorff-dimension-zero result and the moment Euler product (Section 5.3).
 Section 6 derives explicit lower bounds for primes of cryptographic interest and, in particular, supplies the matching lower bound $\min_{p\leq x}c(p)\gg 1/\log\log x$.
@@
 We write $\omega(n)$ for the number of distinct prime factors of an integer $n$, $\pi(x)$ for the number of primes up to $x$, and adopt the convention that $p_{1}=2<p_{2}=3<p_{3}=5<\cdots$ denotes the sequence of all primes in increasing order.
+For a prime $p$, we write
+$P(p):=\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{p^{j}}\Bigr)$
+for the matrix-count Euler product; thus $c(p)=\frac{\varphi(p-1)}{p-1}\,P(p)$.
 All limits are taken as $x\to\infty$ unless stated otherwise.
@@
 Thus $c(p)$ is not the exact finite-dimensional density but the limiting, dimension-uniform reduction constant.
-Recall that one may equivalently bundle $n$ standard LWE secret vectors into a matrix and view decision-LWE in matrix form: given $\mathbf{A}\in\mathbb{Z}_{q}^{w\times n}$, one must distinguish
+In the matrix form of decision-LWE, given $\mathbf{A}\in\mathbb{Z}_{q}^{w\times n}$, one must distinguish
 $(\mathbf{A},\,\mathbf{A}\mathbf{S}+\mathbf{E})$
 from uniform over $\mathbb{Z}_{q}^{w\times n}\times\mathbb{Z}_{q}^{w\times n}$, where $\mathbf{S}\in\mathbb{Z}_{q}^{n\times n}$ is the secret matrix and $\mathbf{E}\in\mathbb{Z}_{q}^{w\times n}$ is the error matrix [7].
-Via the identification $\mathbb{Z}_{q}\cong\mathbb{F}_{q}$ for prime $q$, decision-prim-lwe [1] is the variant in which the secret matrix is additionally required to lie in $M_{n}^{\mathrm{prim}}(\mathbb{F}_{q})$.
+Via the identification $\mathbb{Z}/q\mathbb{Z}\cong\mathbb{F}_{q}$ for prime $q$, decision-prim-lwe [1] is the variant in which the secret matrix is additionally required to lie in $M_{n}^{\mathrm{prim}}(\mathbb{F}_{q})$.
 In [1, Remark 3(ii)], Sehrawat, Yeo, and Desmedt observed that if there exist infinitely many primorial primes
@@
 where $\gamma\approx 0.5772$ is the Euler–Mascheroni constant.
-We also record the following simple observation.
+We also record the following observation.
 Lemma 1 (Lower bound for the matrix-count Euler product).
@@
 $p(a,q)\leq C_{0}q^{L}.$
+Let $N_{k}:=\prod_{i=1}^{k}p_{i}$ denote the $k$-th primorial, $\vartheta(x):=\sum_{\begin{subarray}{c}p\leq x,\;p\text{ prime}\end{subarray}}\log p$ the Chebyshev function, and $m(x):=\max\{k\geq 1:N_{k}\leq x\}$.
 Lemma 2 (Prime-number-theorem consequences for primorials; cf. [12]).
-Let
-$N_{k}=\prod_{i=1}^{k}p_{i},\qquad\vartheta(x):=\sum_{\begin{subarray}{c}p\leq x\\ p\text{ prime}\end{subarray}}\log p.$
-Then
-$\log N_{k}=\vartheta(p_{k})\sim p_{k},\qquad p_{k+1}\sim p_{k}\qquad(k\to\infty).$
-Consequently, if
-$m(x):=\max\{k\geq 1:N_{k}\leq x\},$
-then
-$\log N_{m(x)}\sim\log x,\qquad p_{m(x)}\sim\log x\qquad(x\to\infty).$
+As $k\to\infty$,
+$\log N_{k}=\vartheta(p_{k})\sim p_{k},\qquad p_{k+1}\sim p_{k}.$
+As $x\to\infty$,
+$\log N_{m(x)}\sim\log x,\qquad p_{m(x)}\sim\log x.$
 Proof.
@@
 The classical existence and continuity of the limiting distribution of $\varphi(p-1)/(p-1)$ over primes follow from the shifted-prime Erdős–Wintner framework; for the shift a=1a=-1 used here, direct references are Tenenbaum [3, pp.
-257–258] and Deshouillers–Hassani [4].
-Our contributions in this section are:
-  1. (i)
-    The full-support conclusion: the limiting law has support exactly $0,\tfrac{1}{2}]$ (Theorem 6), combining the left non-differentiability result of [4, Theorem 1.1] with the density of $\{\varphi(m)/m:m\text{ even}\}$ in $[0,1/2]$ (Lemma 4).
-  2. (ii)
-    The transport of the classical shifted-prime law to the prim-lwe density $c(p)$ via Lemma 5, establishing that $c(p)$ shares the same limiting distribution as $\varphi(p-1)/(p-1)$.
-We supply a self-contained proof of the shifted-prime pure singularity in Lemma 9.
+262–263] and Deshouillers–Hassani [4].
+Throughout this section we write
+$f(n)\;:=\;\log\frac{n}{\varphi(n)}\;=\;\sum_{\begin{subarray}{c}\ell\mid n\\ \ell\textup{ prime}\end{subarray}}\log\frac{\ell}{\ell-1}$ (1)
+for the strongly additive function that records $\log(n/\varphi(n))$.
+Our contribution in this section is the full-support conclusion: the limiting law has support exactly $0,\tfrac{1}{2}]$ (Theorem 6), combining the left non-differentiability result of [4, Theorem 1.1] with the density of $\{\varphi(m)/m:m\text{ even}\}$ in $[0,1/2]$ (Lemma 4).
+Since $P(p)\to 1$, the prim-lwe density $c(p)$ shares the same limiting distribution (Lemma 5).
+We supply a self-contained proof of the shifted-prime pure singularity in Lemma 9; in Section 5.3 we show that the limiting measure has Hausdorff dimension 0 and derive the moments of $G$ as convergent Euler products.
 5.1 The distribution of $\varphi(p-1)/(p-1)$
@@
 Lemma 3 (Erdős–Wintner hypotheses for the totient function).
-Let $h$ be the strongly additive function $h(n)=\log(n/\varphi(n))=-\log(\varphi(n)/n)$. For every prime $\ell$,
-$0<h(\ell)=\log\frac{\ell}{\ell-1}\leq\log 2<1,\qquad h(\ell)=\frac{1}{\ell}+O\!\left(\frac{1}{\ell^{2}}\right).$
-Consequently the third Erdős–Wintner series $\sum_{|h(\ell)|>1}1/\ell$ is vacuous, while $\sum_{\ell}h(\ell)/\ell$ and $\sum_{\ell}h(\ell)^{2}/\ell$ both converge.
-Since $h(\ell)\neq 0$ for every prime $\ell$ and $\sum_{\ell}1/\ell=\infty$, the continuity criterion of Lemma 6 is satisfied.
-The same conclusions hold for $g(n)=\varphi(n)/n$ via $\log g(\ell)=-h(\ell)$.
+The function $f$ defined in (1) is strongly additive, with $f(n)=\log(n/\varphi(n))=-\log(\varphi(n)/n)$. For every prime $\ell$,
+$0<f(\ell)=\log\frac{\ell}{\ell-1}\leq\log 2<1,\qquad f(\ell)=\frac{1}{\ell}+O\!\left(\frac{1}{\ell^{2}}\right).$
+The three Erdős–Wintner convergence conditions are satisfied, as is the continuity criterion of Lemma 6. The same conclusions hold for $g(n)=\varphi(n)/n$.
+Proof.
+Since $|f(\ell)|\leq\log 2<1$ for every prime $\ell$, the series $\sum_{|f(\ell)|>1}1/\ell$ is vacuous.
+The estimate $f(\ell)=1/\ell+O(1/\ell^{2})$ gives $\sum_{\ell}f(\ell)/\ell<\infty$ and $\sum_{\ell}f(\ell)^{2}/\ell<\infty$.
+Since $f(\ell)\neq 0$ for every prime $\ell$ and $\sum_{\ell}1/\ell=\infty$, the continuity criterion of Lemma 6 is satisfied.
+The case $g(n)=\varphi(n)/n$ follows from $\log g(\ell)=-f(\ell)$.
+∎
 Proposition 2 (Deshouillers–Hassani; special case of [4, Proposition 2.1 and equation (2.1)]).
@@
 5.2 From $\varphi(p-1)/(p-1)$ to $c(p)$
-The following transport lemma is the mechanism for contribution (ii) listed at the beginning of this section.
+Since the Euler product $P(p)\to 1$, the density $c(p)=\frac{\varphi(p-1)}{p-1}P(p)$ inherits the limiting distribution of $\varphi(p-1)/(p-1)$. The following lemma records this fact.
 Lemma 5 (Transport lemma).
@@
 The relevant shifted-prime Erdős–Wintner framework is due to Kátai [14] and Hildebrand [15] for the shift $a=1$; the methods extend to any fixed nonzero shift via the Bombieri–Vinogradov theorem, and for the shift a=1a=-1 used here, the explicit characteristic function is recorded in Tenenbaum [3, formula (3), p.
-258].
+262].
 The same conclusions hold for general (not necessarily strongly) additive functions under the Erdős–Wintner hypotheses, by the identical Jessen–Wintner argument applied to the full local factors; we state the strongly additive case since it is all that we require.
-By Tenenbaum [3, formula (3), p. 258], the characteristic function of $\mu_{f}$ is given by
+By Tenenbaum [3, formula (3), p. 262], the characteristic function of $\mu_{f}$ is given by
 $\widehat{\mu_{f}}(\tau)\;=\;\prod_{p}\!\left(1-\frac{1}{p-1}+\sum_{\nu\geq 1}\frac{e^{i\tau f(p^{\nu})}}{p^{\nu}}\right)$
@@
 ∎
+For $y\geq 3$ and each prime $p$, define $a_{y}(p-1):=\prod_{\begin{subarray}{c}\ell\leq y,\;\ell\textup{ prime},\;\ell\mid(p-1)\end{subarray}}\ell$, and write $P^{+}(n)$ for the largest prime factor of $n$.
 Lemma 7 (Prime density for a prescribed smooth kernel).
-Let $y\geq 3$, and for each prime $p$ define $a_{y}(p-1):=\prod_{\begin{subarray}{c}\ell\leq y,\;\ell\textup{ prime}\\ \ell\mid(p-1)\end{subarray}}\ell$.
-Let $m$ be an even squarefree positive integer with $P^{+}(m)\leq y$.
-Then the set $\{p\textup{ prime}:a_{y}(p-1)=m\}$ has relative density
+Let $m$ be an even squarefree positive integer with $P^{+}(m)\leq y$. The set $\{p\textup{ prime}:a_{y}(p-1)=m\}$ has relative density
 $\delta_{y}(m)\;=\;\frac{1}{\varphi(m)}\prod_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\nmid m\end{subarray}}\frac{\ell-2}{\ell-1}$
@@
 Lemma 8 (Tail-mean estimate for f(p1)f(p-1)).
-Let $f(n):=\log(n/\varphi(n))$, and for each prime $p$ and parameter $y\geq 3$ define
+Let $f$ be as in (1), and for each prime $p$ and parameter $y\geq 3$ define
 $T_{y}(p):=\sum_{\begin{subarray}{c}\ell\mid(p-1)\\ \ell>y,\;\ell\textup{ prime}\end{subarray}}\log\frac{\ell}{\ell-1}.$
@@
 $\lim_{x\to\infty}\frac{1}{\pi(x)}\sum_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}T_{y}(p)\;=\;\eta(y),$
-where $\eta(y):=\sum_{\ell>y,\;\ell\textup{ prime}}f(\ell)/(\ell-1)\ll 1/y$. In particular, $\eta(y)\to 0$ as $y\to\infty$.
+where $\eta(y):=\sum_{\ell>y,\;\ell\textup{ prime}}f(\ell)/(\ell-1)\ll 1/y$.
 Proof.
@@
 Lemma 9 (Shifted-prime pure singularity for $\log(n/\varphi(n))$).
-Let
-$f(n)\;:=\;\log\frac{n}{\varphi(n)}\;=\;\sum_{\begin{subarray}{c}\ell\mid n\\ \ell\textup{ prime}\end{subarray}}\log\frac{\ell}{\ell-1}.$
+Let $f$ be the strongly additive function defined in (1).
 The shifted-prime limiting measure $\mu_{f}$ is purely singular with respect to Lebesgue measure.
@@
 By the shifted-prime Erdős–Wintner theorem (Tenenbaum [3, pp.
-257–258] for the shift a=1a=-1; Kátai [14], Hildebrand [15] for the original $a=1$ case), the limiting measure $\mu_{f}$ exists.
+262–263] for the shift a=1a=-1; Kátai [14], Hildebrand [15] for the original $a=1$ case), the limiting measure $\mu_{f}$ exists.
 Indeed, by Lemma 3 the three convergence hypotheses are satisfied and the continuity criterion holds, so the limiting measure $\mu_{f}$ exists and is continuous.
@@
 Proposition 3 (Shifted-prime limiting law for $f$).
-Let
-$f(n):=\log\frac{n}{\varphi(n)}=\sum_{\begin{subarray}{c}\ell\mid n\\ \ell\textup{ prime}\end{subarray}}\log\frac{\ell}{\ell-1}.$
-Then the empirical measures
+Let $f$ be as in (1). Then the empirical measures
 $\mu_{x}:=\frac{1}{\pi(x)}\sum_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}\delta_{f(p-1)}$
@@
 By Lemma 3, the three convergence conditions of the shifted-prime Erdős–Wintner theorem (Kátai [14], Hildebrand [15] for $p+1$; Tenenbaum [3, pp.
-257–258] for the general shift including p1p-1) are satisfied, and the continuity criterion holds.
+262–263] for the general shift including p1p-1) are satisfied, and the continuity criterion holds.
 Therefore f(p1)f(p-1) possesses a continuous limiting distribution $\mu_{f}$, which is of pure type by Lemma 6.
@@
 It remains only to justify pure singularity.
-Set
-$f(n):=\log\frac{n}{\varphi(n)}=\sum_{\ell\mid n}\log\frac{\ell}{\ell-1}.$
-Then $f$ is strongly additive and
+Let $f$ be as in (1). Then $f$ is strongly additive and
 $f(\ell)=\log\frac{\ell}{\ell-1}\sim\frac{1}{\ell}\qquad(\ell\to\infty,\ \ell\ \text{prime}),$
@@
 Y(p)=T(f(p1))Y(p)=T(f(p-1))
-has limiting law $T_{\#}\mu_{f}$.
+has limiting law $T_{\#}\mu_{f}$ (the push-forward of $\mu_{f}$ by $T$, defined by $T_{\#}\mu_{f}(B):=\mu_{f}(T^{-1}(B))$).
 Choose a Borel set $E\subset[0,\infty)$ with Lebesgue measure zero and $\mu_{f}(E)=1$.
@@
 so
-$c(2)=\prod_{j\geq 1}\Bigl(1-2^{-j}\Bigr)=\left(\frac{1}{2};\frac{1}{2}\right)_{\infty}\approx 0.288788<\frac{1}{2}.$
+$c(2)=\prod_{j\geq 1}\Bigl(1-2^{-j}\Bigr)\approx 0.288788<\frac{1}{2}.$
 For odd $p$, one has $\varphi(p-1)/(p-1)\leq 1/2$ together with $\prod_{j\geq 1}(1-p^{-j})<1$. Thus $1/2$ is only a boundary point of the support of the limiting law for $c(p)$, not an attained value.
@@
 which proves the claim. ∎
+5.3 Hausdorff dimension, moments, and Fourier structure of $G$
+We derive a sharper geometric description of the limiting distribution and then its moment and Fourier structure.
+Throughout, $f$ is the strongly additive function defined in (1), $\mu_{f}$ is the shifted-prime limiting measure of f(p1)f(p-1), and $X$ is a random variable with distribution $G$.
+By Corollary 2, $X\overset{d}{=}e^{-Y}$ where $Y\sim\mu_{f}$.
+Lemma 10 (Factor structure of $\mu_{f}$).
+The measure $\mu_{f}$ is the infinite convolution
+$\mu_{f}\;=\;\delta_{\log 2}\;*\;\mathop{{\ast}}_{\begin{subarray}{c}\ell\geq 3\\ \ell\textup{ prime}\end{subarray}}\nu_{\ell},$
+where $\delta_{\log 2}$ is the point mass at $\log 2$ and, for each odd prime $\ell$,
+$\nu_{\ell}\;:=\;\frac{\ell-2}{\ell-1}\,\delta_{0}\;+\;\frac{1}{\ell-1}\,\delta_{f(\ell)},\qquad f(\ell)=\log\frac{\ell}{\ell-1}.$
+Equivalently, $Y\overset{d}{=}\log 2+\sum_{\ell\geq 3,\;\ell\textup{ prime}}Z_{\ell}$ with the $Z_{\ell}$ independent, $\mathbb{P}(Z_{\ell}=0)=(\ell-2)/(\ell-1)$, $\mathbb{P}(Z_{\ell}=f(\ell))=1/(\ell-1)$.
+Proof.
+By Tenenbaum [3, formula (3), p. 262] for the shift a=1a=-1, the characteristic function of $\mu_{f}$ is
+$\widehat{\mu_{f}}(\tau)=\prod_{\ell\text{ prime}}\Bigl(1-\frac{1}{\ell-1}+\sum_{\nu\geq 1}\frac{e^{i\tau f(\ell^{\nu})}}{\ell^{\nu}}\Bigr).$
+Since $f$ is strongly additive, $f(\ell^{\nu})=f(\ell)$ for all $\nu\geq 1$, whence $\sum_{\nu\geq 1}e^{i\tau f(\ell^{\nu})}\ell^{-\nu}=e^{i\tau f(\ell)}/(\ell-1)$.
+At each odd prime $\ell$ the local factor equals $\frac{\ell-2}{\ell-1}+\frac{e^{i\tau f(\ell)}}{\ell-1}=\widehat{\nu_{\ell}}(\tau)$; at $\ell=2$ it equals $e^{i\tau\log 2}=\widehat{\delta_{\log 2}}(\tau)$.
+Since the characteristic function determines a probability measure uniquely, the convolution decomposition follows.
+∎
+For a Borel probability measure $\mu$ on $\mathbb{R}$, recall the Hausdorff dimension of $\mu$:
+$\dim_{H}(\mu):=\inf\{\dim_{H}(E):E\subset\mathbb{R}\text{ is Borel and }\mu(E)=1\}.$
+Proposition 4 (Hausdorff dimension zero of the limiting law).
+$\dim_{H}(\mu_{f})=0$. Consequently, if $\mu_{G}$ denotes the limiting measure on $0,\tfrac{1}{2}]$ with distribution function $G$, then $\dim_{H}(\mu_{G})=0$.
+Proof.
+By Lemma 10,
+$Y\overset{d}{=}\log 2+\sum_{\begin{subarray}{c}\ell\geq 3\\ \ell\text{ prime}\end{subarray}}B_{\ell}w_{\ell},\qquad w_{\ell}:=\log\frac{\ell}{\ell-1},$
+where the $B_{\ell}$ are independent Bernoulli random variables with
+$\mathbb{P}(B_{\ell}=1)=\frac{1}{\ell-1},\qquad\mathbb{P}(B_{\ell}=0)=\frac{\ell-2}{\ell-1}.$
+Fix $s>0$. We shall show that $\mu_{f}$ gives full mass to a Borel set of vanishing $s$-dimensional Hausdorff measure.
+For $y\geq 3$, let
+$I(y):=\sum_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\text{ prime}\end{subarray}}\left(B_{\ell}\log(\ell-1)+(1-B_{\ell})\log\frac{\ell-1}{\ell-2}\right).$
+If $\omega=(b_{\ell})_{3\leq\ell\leq y}$ is a prefix configuration, then
+$\mathbb{P}\bigl((B_{\ell})_{3\leq\ell\leq y}=\omega\bigr)=e^{-I_{y}(\omega)},$
+where $I_{y}(\omega)$ denotes the value of $I(y)$ at that configuration. Moreover,
+$\mathbb{E}I(y)\ll\sum_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\text{ prime}\end{subarray}}\frac{\log\ell}{\ell}\ll\log y$
+and
+$\operatorname{Var}(I(y))\leq\sum_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\text{ prime}\end{subarray}}\mathbb{E}\!\left[\left(B_{\ell}\log(\ell-1)+(1-B_{\ell})\log\frac{\ell-1}{\ell-2}\right)^{2}\right]\ll\sum_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\text{ prime}\end{subarray}}\frac{(\log\ell)^{2}}{\ell}\ll(\log y)^{2}.$
+Hence there exists an absolute constant $C_{0}>0$ such that
+$\mathbb{E}I(y)\leq C_{0}\log y,\qquad\operatorname{Var}(I(y))\leq C_{0}(\log y)^{2}\qquad(y\geq 3).$
+Choose $C:=4C_{0}$ (so that $Cn-a_{n}s/2\leq-Cn$ in the covering estimate below). Next choose an increasing sequence $y_{n}\to\infty$ so rapidly that, with
+$a_{n}:=\left\lceil\frac{4Cn}{s}\right\rceil\qquad(\text{recall }s>0\text{ is fixed}),\qquad z_{n}:=y_{n}^{a_{n}},\qquad r_{n}:=z_{n}^{-1/2},$
+the intervals $(y_{n},z_{n}]$ are pairwise disjoint, the sequence $(z_{n})$ is strictly increasing, and
+$\sum_{n\geq 1}r_{n}<\infty,\qquad\sum_{n\geq 1}y_{n}^{-Cn}<\infty.$
+For each $n$, set $I_{n}:=I(y_{n})$ and
+$A_{n}:=\{I_{n}\leq Cn\log y_{n}\}.$
+By Chebyshev’s inequality,
+$\mathbb{P}(A_{n}^{c})\leq\frac{\operatorname{Var}(I_{n})}{(Cn\log y_{n}-\mathbb{E}I_{n})^{2}}\ll\frac{1}{n^{2}}.$
+Let $T_{n}$ be the set of prefix configurations $\omega=(b_{\ell})_{3\leq\ell\leq y_{n}}$ satisfying $I_{n}(\omega)\leq Cn\log y_{n}$. Then
+$\mathbb{P}\bigl((B_{\ell})_{3\leq\ell\leq y_{n}}\in T_{n}\bigr)=1-O(n^{-2}),$
+and every atom in $T_{n}$ has probability at least $y_{n}^{-Cn}$. Since the total mass is at most $1$, it follows that
+$|T_{n}|\leq y_{n}^{Cn}.$
+Now define the empty-block event
+$E_{n}:=\{B_{\ell}=0\text{ for every prime }\ell\in(y_{n},z_{n}]\}.$
+Because the blocks $(y_{n},z_{n}]$ are disjoint, the events $E_{n}$ are independent. Moreover,
+$\mathbb{P}(E_{n})=\prod_{\begin{subarray}{c}y_{n}<\ell\leq z_{n}\\ \ell\text{ prime}\end{subarray}}\left(1-\frac{1}{\ell-1}\right).$
+Since
+$1-\frac{1}{\ell-1}=\left(1-\frac{1}{\ell}\right)\left(1-\frac{1}{(\ell-1)^{2}}\right),$
+Mertens’ theorem and the convergence of $\prod_{\ell\geq 3}(1-(\ell-1)^{-2})$ imply
+$\mathbb{P}(E_{n})\asymp\frac{\log y_{n}}{\log z_{n}}=\frac{1}{a_{n}}\asymp\frac{1}{n}.$
+Therefore $\sum_{n}\mathbb{P}(E_{n})=\infty$, and the second Borel–Cantelli lemma gives
+$\mathbb{P}(E_{n}\text{ i.o.})=1.$
+Let
+$R_{n}:=\sum_{\begin{subarray}{c}\ell>z_{n}\\ \ell\text{ prime}\end{subarray}}B_{\ell}w_{\ell},\qquad F_{n}:=\{R_{n}\leq r_{n}\}.$
+Since $w_{\ell}=\frac{1}{\ell}+O(\ell^{-2})$, we have
+$\mathbb{E}R_{n}=\sum_{\begin{subarray}{c}\ell>z_{n}\\ \ell\text{ prime}\end{subarray}}\frac{w_{\ell}}{\ell-1}\ll\sum_{\begin{subarray}{c}\ell>z_{n}\\ \ell\text{ prime}\end{subarray}}\frac{1}{\ell^{2}}\ll\frac{1}{z_{n}}.$
+By Markov’s inequality,
+$\mathbb{P}(F_{n}^{c})\leq\frac{\mathbb{E}R_{n}}{r_{n}}\ll z_{n}^{-1/2}=r_{n}.$
+Thus $\sum_{n}\mathbb{P}(F_{n}^{c})<\infty$, so
+$\mathbb{P}(F_{n}^{c}\text{ i.o.})=0.$
+For each $\omega=(b_{\ell})_{3\leq\ell\leq y_{n}}\in T_{n}$, define
+$c(\omega):=\log 2+\sum_{\begin{subarray}{c}3\leq\ell\leq y_{n}\\ \ell\text{ prime}\end{subarray}}b_{\ell}w_{\ell},$
+and let
+$U_{n}:=\bigcup_{\omega\in T_{n}}[c(\omega),\,c(\omega)+r_{n}].$
+If $A_{n}\cap E_{n}\cap F_{n}$ occurs, then the block $(y_{n},z_{n}]$ contributes nothing and the tail beyond $z_{n}$ contributes at most $r_{n}$, so $Y\in U_{n}$. Since
+$\sum_{n}\mathbb{P}(A_{n}^{c})<\infty,\qquad\sum_{n}\mathbb{P}(F_{n}^{c})<\infty,\qquad\mathbb{P}(E_{n}\text{ i.o.})=1,$
+it follows that
+$\mu_{f}\!\left(\limsup_{n\to\infty}U_{n}\right)=1.$
+On the other hand,
+$\sum_{n\geq 1}|T_{n}|\,r_{n}^{s}\ll\sum_{n\geq 1}y_{n}^{Cn}z_{n}^{-s/2}=\sum_{n\geq 1}y_{n}^{Cn-a_{n}s/2}\leq\sum_{n\geq 1}y_{n}^{-Cn}<\infty.$
+Because $r_{n}\downarrow 0$, for every $N\geq 1$ the set $\limsup_{n\to\infty}U_{n}$ is covered by the intervals comprising $U_{n}$ for $n\geq N$, each of diameter at most $r_{N}$. Hence
+$\mathcal{H}^{s}_{r_{N}}\!\left(\limsup_{n\to\infty}U_{n}\right)\leq\sum_{n\geq N}|T_{n}|\,r_{n}^{s}.$
+Letting $N\to\infty$ gives
+$\mathcal{H}^{s}\!\left(\limsup_{n\to\infty}U_{n}\right)=0.$
+Thus $\mu_{f}$ assigns full mass to a Borel set of zero $s$-dimensional Hausdorff measure, and therefore $\dim_{H}(\mu_{f})\leq s$. Since $s>0$ was arbitrary, $\dim_{H}(\mu_{f})=0$.
+Finally, $\mu_{G}$ is the push-forward of $\mu_{f}$ under the map $T(u):=e^{-u}$.
+Since $T$ is Lipschitz on $\log 2,\infty)$, the image of any set of Hausdorff dimension 0 again has Hausdorff dimension 0.
+Therefore $\dim_{H}(\mu_{G})=0$.
+∎
+Theorem 8 (Moment Euler product).
+For every real $k>0$,
+$\mathbb{E}[X^{k}]\;=\;2^{-k}\!\prod_{\begin{subarray}{c}\ell\geq 3\\ \ell\textup{ prime}\end{subarray}}\!\left(\frac{\ell-2}{\ell-1}+\frac{1}{\ell-1}\Bigl(\frac{\ell-1}{\ell}\Bigr)^{\!k}\right).$ (5)
+The product converges absolutely, with the $\ell$-th factor equal to $1+O(k/\ell^{2})$.
+Proof.
+Since $Y=\log 2+\sum_{\ell}Z_{\ell}$ with the $Z_{\ell}\geq 0$ independent (Lemma 10), the partial sums $S_{L}:=\sum_{3\leq\ell\leq L}Z_{\ell}\nearrow S:=\sum_{\ell}Z_{\ell}$ give $e^{-kS_{L}}\searrow e^{-kS}$.
+Since $0\leq e^{-kS_{L}}\leq 1$ for all $L$, by dominated convergence,
+$\mathbb{E}[e^{-kS}]=\prod_{\ell\geq 3}\mathbb{E}[e^{-kZ_{\ell}}]=\prod_{\ell\geq 3}\Bigl(\frac{\ell-2}{\ell-1}+\frac{1}{\ell-1}\Bigl(\frac{\ell-1}{\ell}\Bigr)^{\!k}\Bigr).$
+Since $\mathbb{E}[X^{k}]=2^{-k}\,\mathbb{E}[e^{-kS}]$, this gives (5).
+Writing the $\ell$-th factor as $1+a_{\ell}$ with $a_{\ell}=\bigl((\frac{\ell-1}{\ell})^{k}-1\bigr)/(\ell-1)=(e^{-kw_{\ell}}-1)/(\ell-1)$, where $w_{\ell}:=\log(\ell/(\ell-1))=1/\ell+O(1/\ell^{2})$.
+The bound $|e^{u}-1|\leq|u|\,e^{|u|}$ with $u=-kw_{\ell}$ gives $|a_{\ell}|\leq kw_{\ell}\,e^{kw_{\ell}}/(\ell-1)$.
+Since $w_{\ell}=1/\ell+O(1/\ell^{2})$, we obtain $|a_{\ell}|\leq C_{k}/\ell^{2}$ for all $\ell\geq 3$, where $C_{k}>0$ depends only on $k$.
+Since $\sum_{\ell}\ell^{-2}<\infty$, absolute convergence follows.
+∎
+Corollary 6 (First and second moments).
+  1. (i)
+    $\mathbb{E}[X]=\frac{1}{2}\prod_{\begin{subarray}{c}\ell\geq 3\\ \ell\textup{ prime}\end{subarray}}\frac{\ell^{2}-\ell-1}{\ell(\ell-1)}.$
+  2. (ii)
+    $\mathbb{E}[X^{2}]=\frac{1}{4}\prod_{\begin{subarray}{c}\ell\geq 3\\ \ell\textup{ prime}\end{subarray}}\frac{\ell^{3}-\ell^{2}-2\ell+1}{\ell^{2}(\ell-1)}.$
+  3. (iii)
+    $\operatorname{Var}(X)=\mathbb{E}[X^{2}]-(\mathbb{E}[X])^{2}$.
+Proof.
+(i) Set $k=1$ in (5): $\frac{\ell-2}{\ell-1}+\frac{1}{\ell}=\frac{\ell(\ell-2)+(\ell-1)}{\ell(\ell-1)}=\frac{\ell^{2}-\ell-1}{\ell(\ell-1)}$.
+(ii) Set $k=2$: $\frac{\ell-2}{\ell-1}+\frac{(\ell-1)}{\ell^{2}}=\frac{(\ell-2)\ell^{2}+(\ell-1)^{2}}{\ell^{2}(\ell-1)}=\frac{\ell^{3}-\ell^{2}-2\ell+1}{\ell^{2}(\ell-1)}$.
+(Verification at $\ell=3$: $\frac{13}{18}=\frac{1}{2}+\frac{2}{9}$.
+) ∎
+Theorem 9 (Moment determinacy).
+The moment sequence $\{\mathbb{E}[X^{k}]\}_{k=1}^{\infty}$ uniquely determines $G$.
+Proof.
+Since $\operatorname{supp}(G)=[0,\tfrac{1}{2}]$ (Theorem 6), the random variable $X$ is supported on the compact interval $0,\tfrac{1}{2}]$.
+By the Weierstrass approximation theorem, polynomials are uniformly dense in $C([0,\tfrac{1}{2}])$, so the moment sequence $\{\mathbb{E}[X^{k}]\}_{k=1}^{\infty}$ determines the distribution of $X$ uniquely.
+∎
+Theorem 10 (Analytic continuation).
+The function
+$M(s)\;:=\;\mathbb{E}[X^{s}]\;=\;2^{-s}\!\prod_{\begin{subarray}{c}\ell\geq 3\\ \ell\textup{ prime}\end{subarray}}F_{s}(\ell),\qquad F_{s}(\ell):=\frac{\ell-2}{\ell-1}+\frac{1}{\ell-1}\Bigl(\frac{\ell-1}{\ell}\Bigr)^{\!s},$
+extends to an analytic non-vanishing function on $\{s\in\mathbb{C}:\operatorname{Re}(s)>0\}$, with the Euler product converging uniformly on compact subsets.
+Proof.
+We first establish the product formula for complex $s$ with $\sigma:=\operatorname{Re}(s)>0$.
+By Lemma 10, $Y=\log 2+\sum_{\ell\geq 3}Z_{\ell}$ with independent $Z_{\ell}\geq 0$, and the series converges almost surely since $\sum_{\ell}\mathbb{E}[Z_{\ell}]<\infty$.
+Set $S_{L}:=\sum_{3\leq\ell\leq L}Z_{\ell}$.
+Since $Z_{\ell}\geq 0$ and $\sigma>0$,
+$|e^{-sS_{L}}|=e^{-\sigma S_{L}}\leq 1$
+for every $L$.
+As $S_{L}\to S:=\sum_{\ell}Z_{\ell}$ a.s., dominated convergence gives $\mathbb{E}[e^{-sS_{L}}]\to\mathbb{E}[e^{-sS}]$.
+Independence factors the finite partial products: $\mathbb{E}[e^{-sS_{L}}]=\prod_{3\leq\ell\leq L}\mathbb{E}[e^{-sZ_{\ell}}]=\prod_{3\leq\ell\leq L}F_{s}(\ell)$.
+Therefore $\mathbb{E}[X^{s}]=2^{-s}\,\mathbb{E}[e^{-sS}]=2^{-s}\prod_{\ell\geq 3}F_{s}(\ell)$, extending the real-variable identity of Theorem 8 to the half-plane $\operatorname{Re}(s)>0$.
+It remains to verify analyticity and non-vanishing.
+Write $w_{\ell}:=\log(\ell/(\ell-1))$ and $F_{s}(\ell)=1+a_{\ell}(s)$ with $a_{\ell}(s)=(e^{-sw_{\ell}}-1)/(\ell-1)$.
+The bound $|e^{z}-1|\leq|z|e^{|z|}$ gives $|a_{\ell}(s)|\leq|s|\,w_{\ell}\,e^{|s|w_{\ell}}/(\ell-1)$.
+Since $w_{\ell}=1/\ell+O(1/\ell^{2})$, on any compact $K\subset\{\operatorname{Re}(s)>0\}$ with $|s|\leq R$ we get $|a_{\ell}(s)|\leq C_{K}/\ell^{2}$ for all large $\ell$.
+Since $\sum\ell^{-2}<\infty$, Weierstrass’s theorem on products of analytic functions (see e.g.
+[18, Ch.
+VII]) gives that the product converges uniformly on compact subsets to an analytic function.
+As $2^{-s}$ is entire, $M(s)$ is analytic on $\{\operatorname{Re}(s)>0\}$.
+For non-vanishing with $\sigma:=\operatorname{Re}(s)>0$: each factor $F_{s}(\ell)\neq 0$ since $|F_{s}(\ell)|\geq\frac{\ell-2}{\ell-1}-\frac{e^{-\sigma w_{\ell}}}{\ell-1}$ and $e^{-\sigma w_{\ell}}<1$: at $\ell=3$, $|F_{s}(3)|\geq(1-(2/3)^{\sigma})/2>0$; at $\ell\geq 5$, $|F_{s}(\ell)|\geq(\ell-3)/(\ell-1)>0$.
+To conclude that the infinite product is nonzero, fix a compact $K\subset\{\operatorname{Re}(s)>0\}$.
+Choose $L$ so that $|a_{\ell}(s)|<1/2$ for all $\ell\geq L$ and $s\in K$ (possible since $|a_{\ell}(s)|\leq C_{K}/\ell^{2}$).
+For $\ell\geq L$ the principal branch $\log(1+a_{\ell}(s))$ is analytic, and $|\!\log(1+a_{\ell}(s))|\leq 2|a_{\ell}(s)|\leq 2C_{K}/\ell^{2}$, so $\sum_{\ell\geq L}\log(1+a_{\ell}(s))$ converges uniformly on $K$ and the tail product equals $\exp\!\bigl(\sum_{\ell\geq L}\log(1+a_{\ell}(s))\bigr)\neq 0$.
+The finitely many factors with $\ell<L$ are each nonzero, so $M(s)\neq 0$.
+∎
+Proposition 5 (Squared modulus of the characteristic function).
+For every $\tau\in\mathbb{R}$,
+$\bigl|\widehat{\mu_{f}}(\tau)\bigr|^{2}\;=\;\prod_{\begin{subarray}{c}\ell\geq 3\\ \ell\textup{ prime}\end{subarray}}\!\left(1-\frac{2(\ell-2)}{(\ell-1)^{2}}\bigl(1-\cos(\tau\log\tfrac{\ell}{\ell-1})\bigr)\right).$ (6)
+Proof.
+By Lemma 10, $|\widehat{\mu_{f}}(\tau)|^{2}=\prod_{\ell\geq 3}|\widehat{\nu_{\ell}}(\tau)|^{2}$.
+With $\alpha=\tau\log(\ell/(\ell-1))$ and $\widehat{\nu_{\ell}}(\tau)=\frac{\ell-2}{\ell-1}+\frac{e^{i\alpha}}{\ell-1}$,
+$|\widehat{\nu_{\ell}}(\tau)|^{2}=\frac{(\ell-2)^{2}+2(\ell-2)\cos\alpha+1}{(\ell-1)^{2}}=1-\frac{2(\ell-2)}{(\ell-1)^{2}}(1-\cos\alpha),$
+using $(\ell-2)^{2}+2(\ell-2)+1=(\ell-1)^{2}$. ∎
+Proposition 6 (Fourier non-decay implies singularity).
+If $\limsup_{|\tau|\to\infty}|\widehat{\mu_{f}}(\tau)|>0$, then $\mu_{f}$ is purely singular, and consequently $G$ is purely singular.
+Proof.
+By Lemma 6, $\mu_{f}$ is continuous and of pure type: either absolutely continuous or purely singular.
+The Riemann–Lebesgue lemma gives: if $\mu_{f}$ is absolutely continuous ($\mu_{f}=h\,\mathrm{d}t$, $h\in L^{1}$), then $\widehat{\mu_{f}}(\tau)=\widehat{h}(\tau)\to 0$.
+Contrapositively, $\limsup|\widehat{\mu_{f}}(\tau)|>0$ implies $\mu_{f}$ is not absolutely continuous, hence purely singular.
+For the conclusion about $G$: by Corollary 2, $X\overset{d}{=}e^{-Y}$ where $Y\sim\mu_{f}$, and $G$ is the distribution function of $X$.
+The map $T(u)=e^{-u}$ is locally Lipschitz on $0,\infty)$, so the push-forward $T_{\#}\mu_{f}$ preserves Lebesgue-null sets.
+Hence pure singularity of $\mu_{f}$ implies pure singularity of $G$ (cf.
+the proof of Corollary 2).
+∎
+5.4 Endpoint asymptotics at $1/2$
+Proposition 7 (Endpoint asymptotic at $1/2$).
+Let $G$ be the limiting distribution function from Theorem 6, and let $X$ be a random variable with distribution $G$. Define
+$\kappa:=2e^{-\gamma}\prod_{\begin{subarray}{c}\ell\geq 3\\ \ell\textup{ prime}\end{subarray}}\Bigl(1-\frac{1}{(\ell-1)^{2}}\Bigr).$
+Then, as $\varepsilon\downarrow 0$,
+$1-G\!\left(\frac{1}{2}-\varepsilon\right)=\mathbb{P}\!\left(X>\frac{1}{2}-\varepsilon\right)\sim\frac{\kappa}{\log(1/\varepsilon)}.$
+Proof.
+By Lemma 10,
+$X\overset{d}{=}\frac{1}{2}e^{-S},\qquad S:=\sum_{\begin{subarray}{c}\ell\geq 3\\ \ell\textup{ prime}\end{subarray}}Z_{\ell},$
+where the $Z_{\ell}$ are independent and satisfy
+$\mathbb{P}(Z_{\ell}=0)=\frac{\ell-2}{\ell-1},\qquad\mathbb{P}\!\left(Z_{\ell}=\log\frac{\ell}{\ell-1}\right)=\frac{1}{\ell-1}.$
+Set
+$M(y):=\prod_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\textup{ prime}\end{subarray}}\frac{\ell-2}{\ell-1}.$
+For $u>0$ one has
+$\frac{u}{1+u}<\log(1+u)\leq u.$
+Applying this with $u=1/(\ell-1)$ gives
+$\frac{1}{\ell}<\log\frac{\ell}{\ell-1}\leq\frac{1}{\ell-1}.$
+Hence, if $\ell\leq(2t)^{-1}$, then
+$Z_{\ell}\neq 0\;\Longrightarrow\;Z_{\ell}>\frac{1}{\ell}\geq 2t>t,$
+so the event $\{S\leq t\}$ forces $Z_{\ell}=0$ for every odd prime $\ell\leq(2t)^{-1}$. Therefore
+$\mathbb{P}(S\leq t)\leq M\!\bigl((2t)^{-1}\bigr).$ (7)
+For the lower bound, set
+$z_{t}:=\frac{\log(1/t)}{t}\qquad(0<t<e^{-1}).$
+By independence,
+$\mathbb{P}(S\leq t)\geq M(z_{t})\,\mathbb{P}\!\left(\sum_{\begin{subarray}{c}\ell>z_{t}\\ \ell\textup{ prime}\end{subarray}}Z_{\ell}\leq t\right).$ (8)
+Since $Z_{\ell}\leq 1/(\ell-1)$ almost surely,
+$\mathbb{E}\!\left[\sum_{\begin{subarray}{c}\ell>z_{t}\\ \ell\textup{ prime}\end{subarray}}Z_{\ell}\right]=\sum_{\begin{subarray}{c}\ell>z_{t}\\ \ell\textup{ prime}\end{subarray}}\frac{1}{\ell-1}\log\frac{\ell}{\ell-1}\leq\sum_{\begin{subarray}{c}\ell>z_{t}\\ \ell\textup{ prime}\end{subarray}}\frac{1}{(\ell-1)^{2}}\ll\frac{1}{z_{t}}=\frac{t}{\log(1/t)}.$
+Thus Markov’s inequality gives
+$\mathbb{P}\!\left(\sum_{\begin{subarray}{c}\ell>z_{t}\\ \ell\textup{ prime}\end{subarray}}Z_{\ell}>t\right)\ll\frac{1}{\log(1/t)}=o(1),$
+and (8) becomes
+$\mathbb{P}(S\leq t)\geq M(z_{t})\,(1-o(1)).$ (9)
+Now
+$M(y)=\prod_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\textup{ prime}\end{subarray}}\left(1-\frac{1}{\ell}\right)\prod_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\textup{ prime}\end{subarray}}\left(1-\frac{1}{(\ell-1)^{2}}\right).$
+The second product converges to $\prod_{\ell\geq 3}(1-(\ell-1)^{-2})$, while Theorem 2 gives
+$\prod_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\textup{ prime}\end{subarray}}\left(1-\frac{1}{\ell}\right)\sim\frac{2e^{-\gamma}}{\log y}.$
+Hence
+$M(y)\sim\frac{\kappa}{\log y}\qquad(y\to\infty).$
+Since $\log((2t)^{-1})\sim\log z_{t}\sim\log(1/t)$ as $t\downarrow 0$, (7) and (9) imply
+$\mathbb{P}(S\leq t)\sim\frac{\kappa}{\log(1/t)}\qquad(t\downarrow 0).$
+Since $\mu_{f}$ is continuous by Proposition 3, the law of $S=Y-\log 2$ is also continuous, so $\mathbb{P}(S<t)=\mathbb{P}(S\leq t)$. Therefore
+$1-G\!\left(\frac{1}{2}-\varepsilon\right)=\mathbb{P}\!\left(X>\frac{1}{2}-\varepsilon\right)=\mathbb{P}\!\left(S<-\log(1-2\varepsilon)\right).$
+As $\varepsilon\downarrow 0$,
+$\log(1-2\varepsilon)=2\varepsilon+O(\varepsilon^{2}),$
+so $\log(1/(-\log(1-2\varepsilon)))\sim\log(1/\varepsilon)$. The stated asymptotic follows. ∎
+Corollary 7 (Endpoint modulus at $1/2$).
+As $\varepsilon\downarrow 0$,
+$G\!\left(\frac{1}{2}\right)-G\!\left(\frac{1}{2}-\varepsilon\right)=1-G\!\left(\frac{1}{2}-\varepsilon\right)\asymp\frac{1}{\log(1/\varepsilon)}.$
+In particular, for every $\alpha>0$, one does not have
+$G\!\left(\frac{1}{2}\right)-G\!\left(\frac{1}{2}-\varepsilon\right)=O(\varepsilon^{\alpha})\qquad(\varepsilon\downarrow 0),$
+so $G$ is not Hölder of any positive order at the endpoint $1/2$.
+Proof.
+The two-sided order follows immediately from Proposition 7. If for some $\alpha>0$ one had
+$G\!\left(\frac{1}{2}\right)-G\!\left(\frac{1}{2}-\varepsilon\right)=O(\varepsilon^{\alpha}),$
+then Proposition 7 would force
+$\frac{1}{\log(1/\varepsilon)}=O(\varepsilon^{\alpha}),$
+which is impossible because
+$\varepsilon^{\alpha}\log(1/\varepsilon)\to 0\qquad(\varepsilon\downarrow 0).$
 6 Explicit Lower Bounds for Primes of Cryptographic Interest
@@
 Remark 1 (On effectivity).
-The asymptotic $\min_{p\leq x}c(p)\asymp 1/\!\log\log x$ should be capable of being made fully explicit by combining explicit Mertens product and $\vartheta(x)$ estimates (e.g., Rosser–Schoenfeld [18], Dusart [19]) with an explicit Linnik-type bound (see Xylouris [20]).
-We do not pursue the bookkeeping here.
+The asymptotic $\min_{p\leq x}c(p)\asymp 1/\!\log\log x$ can be made fully explicit by combining explicit Mertens product and $\vartheta(x)$ estimates (e.g., Rosser–Schoenfeld [19], Dusart [20]) with an explicit Linnik-type bound (see Xylouris [21]).
+We do not carry out the explicit computation here.
 Remark 2.
 The bound $1/2$ for the infinite product can be sharpened for specific primes: $\prod_{j\geq 1}(1-q^{-j})\geq 1-1/(q-1)$, which exceeds $2/3$ for $q\geq 5$ and is at least $0.99$ for $q\geq 101$.
-For large NTT primes with $q>2^{30}$, this factor is numerically indistinguishable from 11.
+For large NTT primes with $q>2^{30}$, this factor differs from 11 by less than $10^{-9}$.
 Even for the smaller NIST moduli $q=3329$ and $q=8380417$ [22, 23], the elementary bound gives $\prod_{j\geq 1}(1-q^{-j})>0.9996$ and $\prod_{j\geq 1}(1-q^{-j})>0.9999998$, respectively.
 We now instantiate the general $\omega(q-1)$-based bound in several parameter regimes relevant to cryptographic applications, including many NTT-friendly moduli.
-Proposition 4 (Bounds for primes with bounded $\omega(q-1)$).
-Let $q\geq 3$ be a prime. Define
-$B(\omega)=P_{3}\prod_{i=1}^{\omega}\Bigl(1-\frac{1}{p_{i}}\Bigr).$
-Then $c(q)\geq B(\omega(q-1))$. Explicit values:
-  1. (a)
-    $\omega(q-1)=1$: $B(1)=P_{3}/2\approx 0.280063$.
-  2. (b)
-    $\omega(q-1)\leq 5$: $B(5)\approx 0.116390$.
-  3. (c)
-    $\omega(q-1)\leq 10$: $B(10)\approx 0.0884704$.
-  4. (d)
-    $\omega(q-1)\leq 15$: $B(15)\approx 0.0776918$.
-  5. (e)
-    $\omega(q-1)\leq 20$: $B(20)\approx 0.0715828$.
-Proof.
-The bound $c(q)\geq B(\omega(q-1))$ is immediate from Theorem 8. Direct evaluation gives
-$B(1)=\frac{P_{3}}{2}\approx 0.280063,\qquad B(5)=\frac{16P_{3}}{77}\approx 0.116390,$
-$B(10)\approx 0.0884704,\qquad B(15)\approx 0.0776918,\qquad B(20)\approx 0.0715828.$
-This yields the stated bounds. ∎
+Define $B(\omega):=P_{3}\prod_{i=1}^{\omega}(1-1/p_{i})$.
+By Theorem 11, every prime $q\geq 3$ satisfies $c(q)\geq B(\omega(q-1))$.
+Direct evaluation gives the following values:
+$\omega(q-1)\leq K$ $B(K)$
+11 $P_{3}/2\approx 0.280063$
+55 $\approx 0.116390$
+$10$ $\approx 0.0884704$
+$15$ $\approx 0.0776918$
+2020 $\approx 0.0715828$
 Remark 3 (On the case $\omega(q-1)=1$).
@@
 Proof.
-The claim under $\omega(q-1)\leq 10$ follows from Proposition 4(c), since
+The claim under $\omega(q-1)\leq 10$ follows from the bound $c(q)\geq B(\omega(q-1))$ (Theorem 11), since
 $B(10)=P_{3}\prod_{i=1}^{10}\Bigl(1-\frac{1}{p_{i}}\Bigr)\approx 0.08847>0.088.$
@@
 where the inequality holds because the right-hand side includes additional factors in $(0,1)$.
-By Dusart’s explicit product estimate [19, Theorem 5.9], which we use conservatively with threshold $x\geq 2\,278\,383$,
+By Dusart’s explicit product estimate [20, Theorem 5.9], which we use conservatively with threshold $x\geq 2\,278\,383$,
 $\prod_{\begin{subarray}{c}r\leq x\\ r\text{ prime}\end{subarray}}\Bigl(1-\frac{1}{r}\Bigr)\;=\;\frac{e^{-\gamma}}{\log x}\,\Bigl(1+O^{*}\!\Bigl(\frac{0.2}{\log^{3}x}\Bigr)\Bigr).$
@@
 The sharper universal pointwise bound $c(q)\gg 1/\!\log\log q$ of Corollary 9 shows that $1/\!\log\log q$, not $1/\!\log q$, is the correct worst-case scale for the rejection-sampling overhead.
+By the shifted-prime Erdős–Kac theorem (Barban [24]; see also Halberstam [25] and Elliott [26, Ch.
+12]), $(\omega(q-1)-\log\log x)/\sqrt{\log\log x}$ converges in distribution to the standard normal as $x\to\infty$ over primes $q\leq x$.
+In particular, for every fixed $\delta>0$, almost all primes $q\leq x$ satisfy $\omega(q-1)\leq(1+\delta)\log\log x$.
 Remark 4 (Typical size of $\omega(q-1)$).
-Let $\Phi$ denote the standard normal distribution function.
-The shifted-prime analogue of the Erdős–Kac theorem (due to Barban [23]; see also Halberstam [24] and Elliott [25, Ch.
-12]) asserts that for every real $z$,
-$\lim_{x\to\infty}\frac{1}{\pi(x)}\#\left\{q\leq x:q\text{ prime and }\frac{\omega(q-1)-\log\log x}{\sqrt{\log\log x}}\leq z\right\}=\Phi(z).$
-Consequently, for every fixed $\delta>0$,
-$\frac{1}{\pi(x)}\#\left\{q\leq x:q\text{ prime and }\omega(q-1)>(1+\delta)\log\log x\right\}\longrightarrow 0.$
-Thus, for large $x$, almost all primes $q\leq x$ satisfy
-$\omega(q-1)\leq(1+\delta)\log\log x,$
-which explains why bounds depending only on $\omega(q-1)$ are relevant in cryptographic parameter ranges.
+The shifted-prime Erdős–Kac theorem explains why bounds depending only on $\omega(q-1)$ are relevant in cryptographic parameter ranges: for $q\approx 2^{128}$, one has $\log\log q\approx 4.5$, so $\omega(q-1)\leq 10$ holds for the vast majority of such primes.
 Scheme $q$ factorization of q1q-1 general lower bound for $c(q)$ exact-factorization lower bound
@@
 The fourth column gives the general $\omega(q-1)$-based lower bound for $c(q)$ from Theorem 11, while the fifth gives the sharper lower bound obtained from the exact factorization of q1q-1 together with $P(q)\geq 1-\frac{1}{q-1}$.
-The exact-factorization bounds imply expected rejection-sampling overheads at most $2.17$ and $3.42$, respectively.
+The exact-factorization bounds imply dimension-uniform expected rejection-sampling overheads at most $2.17$ and $3.42$, respectively.
 For the two prime moduli appearing, across all parameter sets, in the current NIST standards for ML-KEM and ML-DSA [22, 23], Table 1 records both the general $\omega(q-1)$-based bounds and the sharper exact-factorization bounds.
@@
 ML-DSA (FIPS 204) $8380417$ $2^{13}\cdot 3\cdot 11\cdot 31$ 44 $3.41$
-NewHope / FALCON [^1] FALCON is being standardized by NIST as FN-DSA (FIPS 206). [26, 27] $12289$ $^{12}\cdot 3 22$ $3.00$
+NewHope / FALCON [27, 28] $12289$ $^{12}\cdot 3 22$ $3.00$
 BFV/BGV FHE (small) [29, 30, 31] $786433$ $^{18}\cdot 3 22$ $3.00$
@@
 Table 2: Representative NTT-friendly primes from lattice-based cryptography and FHE implementations, together with $\omega(q-1)$ and the dimension-uniform expected rejection-sampling overhead $1/c(q)$.
+The displayed overheads are rounded to two decimal places; the exact values differ by less than $10^{-3}$ due to the Euler product $P(q)$.
 All overheads are at most $4.00$.
 The first two rows reproduce the NIST entries from Table 1.
-The SEAL entry is the default coefficient modulus for poly_modulus_degree = 1024 at the 128-bit security level.
 Table 2 collects nine representative moduli drawn from current lattice-based and FHE implementations.
@@
 For the 64-bit NTT prime $q=2^{64}-2^{32}+1$ (sometimes called the “Goldilocks prime” in the zero-knowledge community, not to be confused with the Ed448-Goldilocks prime $2^{448}-2^{224}-1$ [36]), used as an NTT-friendly ambient field in implementation-oriented lattice cryptography [35], one has $\omega(q-1)=6$ and $\varphi(q-1)/(q-1)\approx 0.2500$.
-Since $P(q)>1-10^{-19}$, one obtains $c(q)\approx 0.25$, so roughly a quarter of random $n\times n$ matrices have a primitive-root determinant and the security reduction loses a factor of about $4$.
+Since $P(q)>1-10^{-19}$, one obtains $c(q)\approx 0.25$, so in every dimension $n$ at least a fraction $\approx 0.25$ of random $n\times n$ matrices have a primitive-root determinant, and the dimension-uniform security-reduction overhead is about $4$.
 7 Implications for PRIM-LWE
@@
 Our results therefore show that the $1/c(p)$ factor in this bound can grow like $\Omega(\log\log p)$ along a suitable subsequence of primes.
-However, this is a modulus-dependent nuance rather than a structural obstruction.
+However, this is a modulus-dependent artifact rather than a structural obstruction.
 For any fixed prime $p$, the constant $c(p)>0$ and the reduction remains valid.
@@
 We note that the primes typically used in lattice-based cryptography are chosen for NTT compatibility.
-Concretely, this typically means $q\equiv 1\pmod{2^{t}}$ for moderate $t$.
+Concretely, this means $q\equiv 1\pmod{2^{t}}$ for moderate $t$.
 Mathematically, this condition controls only the 2-adic part of q1q-1.
@@
 The proof is short and unconditional, requiring only Dirichlet’s theorem and Mertens’ product formula.
-We have further established several results that illuminate the finer structure of the density $c(p)$:
+We have further established several results that describe the finer structure of the density $c(p)$:
   1. 1.
@@
     $\min_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}c(p)\asymp\frac{1}{\log\log x}.$
-    In particular, the decay to zero is extremely slow and is completely quantified.
+    In particular, the decay to zero is slow and quantified precisely.
   2. 2.
@@
 This full-support conclusion appears to be new.
-The law coincides with the classical shifted-prime distribution of $\varphi(p-1)/(p-1)$, as shown by a transport lemma, and is purely singular.
+The law coincides with the classical shifted-prime distribution of $\varphi(p-1)/(p-1)$, is purely singular, and, by Proposition 4, the limiting measure has Hausdorff dimension 0.
+The moments of $G$ are given by convergent Euler products (Theorem 8), and the Mellin transform $\mathbb{E}[X^{s}]$ extends analytically to $\{\operatorname{Re}(s)>0\}$ (Theorem 10).
 Pure singularity was already known in the integer case from Erdős [5]; Tenenbaum [3, p.
@@
 Table 2 confirms that every implementation modulus we examined has $\omega(q-1)\leq 6$ and $1/c(q)\leq 4.00$, and the sharper universal pointwise bound of Corollary 9 gives $1/c(q)=O(\log\log q)$.
-A natural remaining question is to characterize the limiting law $G$ more precisely—for instance, to understand its moments, its modulus of continuity, or its finer local regularity inside $0,\tfrac{1}{2}]$.
+Proposition 7 and Corollary 7 determine the right-endpoint behavior of $G$:
+$1-G\!\left(\frac{1}{2}-\varepsilon\right)\sim\frac{\kappa}{\log(1/\varepsilon)}\qquad(\varepsilon\downarrow 0),$
+so $G$ is not Hölder of any positive order at $1/2$.
+Proposition 4 shows that the limiting measure has Hausdorff dimension 0.
+Equivalently, the limiting PRIM-LWE law is a continuous, purely singular, full-support measure on $0,\tfrac{1}{2}]$ that is concentrated on a Borel set of Hausdorff dimension 0.
+A natural remaining question is therefore to determine a global modulus of continuity of $G$.
+The squared modulus of the characteristic function admits an explicit Euler product (Proposition 5), but whether
+$|\widehat{\mu_{f}}(\tau)|\to 0\qquad\text{or}\qquad\limsup_{|\tau|\to\infty}|\widehat{\mu_{f}}(\tau)|>0$
+remains open; by Proposition 6, the latter would give an independent Fourier-analytic proof of pure singularity.
 This should not be confused with the classical support question for the distribution of $\varphi(n)/n$ over all integers, which remains open.

### 21 Apr 2026 v4

@@
 Abstract
-The prim-lwe problem, introduced by Sehrawat, Yeo, and Desmedt (Theoretical Computer Science, 886 (2021)), is a variant of the Learning with Errors problem in which the secret matrix is required to have a primitive-root determinant.
-The dimension-uniform reduction constant is $c(p)=\inf_{n\geq 1}c_{n}(p)$, where $c_{n}(p)$ is the exact density of $n\times n$ matrices over $\mathbb{F}_{p}$ with primitive-root determinant.
-Sehrawat, Yeo, and Desmedt asked whether $\inf_{p\text{ prime}}c(p)=0$, observing that an affirmative answer would follow from the conjectural infinitude of primorial primes.
-We resolve this question unconditionally using only Dirichlet’s theorem and Mertens’ product formula, bypassing the primorial-prime hypothesis.
-We further establish the sharp order
-$\min_{p\leq x}c(p)\asymp\frac{1}{\log\log x}\qquad(x\to\infty),$
-and show that the limiting distribution of $c(p)$ over the primes has support exactly $[0,1/2]$.
-This full-support property appears to be new.
-The law coincides with the classical shifted-prime distribution of $\varphi(p-1)/(p-1)$ via a transport lemma, is moreover continuous and purely singular, and the limiting measure has Hausdorff dimension 0.
-The moments of the limiting distribution are given by convergent Euler products that uniquely determine it, and the associated Mellin transform extends analytically to the right half-plane.
-We also derive explicit lower bounds on $c(q)$ for primes of cryptographic interest, parameterized solely by the number of distinct prime factors of q1q-1.
-As a simple conservative explicit bound, for any prime $q>2^{30}$ the dimension-uniform expected overhead $1/c(q)$ is at most $1.79\log q$.
-On the other hand, our results show that the worst-case overhead among primes $p\leq x$ is of order $\Theta(\log\log x)$, and in particular $1/c(q)=O(\log\log q)$ pointwise.
+For a prime $p$, let $c_{n}(p)$ denote the fraction of $n\times n$ matrices over $\mathbb{F}_{p}$ whose determinant is a primitive root modulo $p$, and let $c(p)=\lim_{n\to\infty}c_{n}(p)$ be the limiting density.
+This quantity governs the efficiency of PRIM-LWE, a variant of the learning with errors (LWE) problem in which the secret matrix is required to have primitive-root determinant.
+The value $1/c(p)$ equals the expected rejection-sampling overhead in the reduction from standard LWE to PRIM-LWE.
+We prove unconditionally that $\inf_{p\text{ prime}}c(p)=0$, resolving an open question from the PRIM-LWE literature, and establish the rate $\min_{p\leq x}c(p)\asymp 1/\log\log x$ as $x\to\infty$.
+We show that $c(p)$, viewed as a function on the primes, has a continuous, purely singular limiting distribution with support $[0,1/2]$ and Hausdorff dimension 0.
+The moments of this limiting distribution are given by convergent Euler products, and the Mellin transform $\mathbb{E}[X^{s}]$ extends analytically to $\operatorname{Re}(s)>0$.
+Using the number of distinct prime factors $\omega(p-1)$, we derive explicit lower bounds: for every prime $p>2^{30}$ the overhead satisfies $1/c(p)<1.79\log p$, and in general $1/c(p)=O(\log\log p)$.
+We also describe a certified algorithm for evaluating $c(p)$ and tabulate values for primes of cryptographic interest.
+Keywords: Primitive root, determinant density, PRIM-LWE, limiting distribution, Hausdorff dimension
+MSC 2020: 11A07, 11N60, 94A60, 15B33
 1 Introduction
@@
 Following [36], a matrix $M\in M_{n}(\mathbb{F}_{p})$ has primitive-root determinant if $\det(M)$ generates $\mathbb{F}_{p}^{*}$.
-The fraction of such matrices is denoted $c_{n}(p)$; since it is decreasing in $n$, the dimension-uniform constant governing the reduction from decision-LWE to decision-prim-lwe is $c(p):=\inf_{n\geq 1}c_{n}(p)=\lim_{n\to\infty}c_{n}(p)$.
-The precise formulas are recorded in Section 2.
-Sehrawat, Yeo, and Desmedt [1, Remark 3(ii)] asked whether $\inf_{p}c(p)=0$ can be proved unconditionally; see Section 2 for the precise statement.
+Write $c_{n}(p)$ for the fraction of such matrices, and $c(p):=\inf_{n\geq 1}c_{n}(p)=\lim_{n\to\infty}c_{n}(p)$ for the limiting density; the precise formulas are recorded in Section 2.
+Since $c_{n}(p)$ is decreasing in $n$, one has $c(p)\leq c_{n}(p)$ for every $n\geq 1$.
+Consequently, every lower bound on $c(p)$ is simultaneously a lower bound on every finite-dimensional density $c_{n}(p)$.
+The quantity $c(p)$ arises as a reduction constant in the PRIM-LWE variant [36] of the popular learning with errors (LWE) problem [33], introduced to construct verifiable secret sharing schemes with “free” verification.
+The reduction proceeds by rejection-sampling and succeeds with probability $c_{n}(p)$ in dimension $n$, so the dimension-uniform overhead is $1/c(p)$.
+This overhead represents a tightness loss in the security reduction: the hardness guarantee for PRIM-LWE degrades by a factor of $c(p)$ relative to standard LWE, making tight bounds on $c(p)$ essential for concrete security estimates.
+Sehrawat, Yeo, and Desmedt [36, Remark 3(ii)] asked whether $\inf_{p}c(p)=0$ can be proved unconditionally.
 Main results
-We resolve this question and establish several natural refinements.
 1.
@@
 We prove $\inf_{p\text{ prime}}c(p)=0$ using only Dirichlet’s theorem on primes in arithmetic progressions and Mertens’ third theorem, without assuming the primorial-prime hypothesis; see Section 3.
-2.
-Sharp order of the minimum (Proposition 1, Corollary 8).
-By additionally invoking Linnik’s theorem on the least prime in an arithmetic progression, together with standard prime-number-theorem consequences for primorials, we establish
+2. Sharp order of the minimum (Proposition 1, Corollary 7). We establish
 $\min_{p\leq x}c(p)\;\asymp\;\frac{1}{\log\log x}\qquad(x\to\infty).$
-The matching lower bound follows from two observations: if $p\leq x$ is an odd prime, then $\omega(p-1)\leq m(x):=\max\{k:N_{k}\leq x\}$, and Mertens’ theorem converts this into a $1/\!\log\log x$ lower bound for $\varphi(p-1)/(p-1)$; the additional factor $\prod_{j\geq 1}(1-p^{-j})\geq P_{3}$ is handled by Theorem 11.
-Hence the decay to zero is quantified precisely.
-The implied constant is $\approx 0.31$, so primes $p$ whose predecessor p1p-1 is divisible by many small primes can still have $c(p)<0.1$.
-The explicit bounds in Section 6 show that standard cryptographic moduli with controlled factorization incur only a small constant-factor overhead; see Remark 1 for effectivity.
+See Remark 4 for effectivity.
 3.
-Limiting distribution (Theorems 6 and 7, Corollary 2, Proposition 4).
-The quantity $c(p)$ possesses a continuous limiting distribution over the primes whose support is exactly $0,\tfrac{1}{2}]$; the law is moreover purely singular, and Proposition 4 shows that the limiting measure has Hausdorff dimension 0.
-If $G$ denotes the distribution function, then
-$0<G(\tau)<1\qquad(0<\tau<1/2),$
-the set of primes with $c(p)\leq\tau$ has relative density $G(\tau)$, and $G(\tau)\to 0$ as $\tau\downarrow 0$.
-The existence and continuity of $G$ follow from the shifted-prime Erdős–Wintner framework for additive functions evaluated at p1p-1 [2]; for the relevant classical background see Tenenbaum [3] and Deshouillers–Hassani [4].
-Pure singularity is classical in the integer case (Erdős [5]) and carries over to shifted primes by the same method, as noted by Tenenbaum [3, p.
-263]; we include a self-contained proof for the reader’s convenience.
-Our new contribution on the distributional side is the full-support conclusion: the limiting law has support exactly $0,\tfrac{1}{2}]$, established by combining the left non-differentiability result of [4, Theorem 1.1] with the density of $\{\varphi(m)/m:m\text{ even}\}$ in $[0,1/2]$ (Lemma 4).
-Since the Euler product $P(p)\to 1$, the prim-lwe density $c(p)=\frac{\varphi(p-1)}{p-1}P(p)$ shares the same limiting distribution (Lemma 5).
-We also derive the moments of $G$ as convergent Euler products (Theorem 8) and establish the analytic continuation of the Mellin transform $\mathbb{E}[X^{s}]$ to the right half-plane (Theorem 10).
-At the right endpoint, $1-G(\tfrac{1}{2}-\varepsilon)\sim\kappa/\!\log(1/\varepsilon)$ as $\varepsilon\downarrow 0$ for an explicit constant $\kappa$ (Proposition 7), so $G$ is not Hölder of any positive order at $1/2$.
+Limiting distribution (Theorems 4 and 5, Corollary 2, Proposition 4).
+The quantity $c(p)$ possesses a continuous, purely singular limiting distribution over the primes whose support is exactly $0,\tfrac{1}{2}]$; the limiting measure has Hausdorff dimension 0.
+The full-support conclusion is new; it combines the left non-differentiability result of Deshouillers–Hassani [9, Theorem 1.1] with the density of even totient ratios (Lemma 4).
+The moments of $G$ are given by convergent Euler products (Theorem 6), the Mellin transform $\mathbb{E}[X^{s}]$ extends analytically to $\{\operatorname{Re}(s)>0\}$ (Theorem 8), and $G$ is not Hölder of any positive order at $1/2$ (Proposition 7).
+Since the limiting law is purely singular and supported on a set of Hausdorff dimension 0, it cannot certify $c(p)$ at any individual prime; this motivates the pointwise bounds of item 4.
 4.
-Explicit bounds for cryptographic moduli (Theorem 11, Table 1).
-We give unconditional lower bounds on $c(q)$ for primes $q$ parameterized by $\omega(q-1)$, the number of distinct prime factors of q1q-1, yielding concrete bounds for the NIST-standardized primes and other moduli whose predecessors have controlled factorization structure.
-The following consequence singles out the cryptographic implication: in the reduction from decision-LWE to decision-prim-lwe over $\mathbb{F}_{q}$, the dimension-uniform expected rejection-sampling overhead $1/c(q)$ admits an explicit upper bound depending only on $\omega(q-1)$.
+Explicit bounds via distinct prime factors (Theorem 9).
+We give unconditional lower bounds on $c(p)$ for primes $p$ parameterized by $\omega(p-1)$, the number of distinct prime factors of p1p-1, yielding computable bounds for any prime $p$ such that the factorization of p1p-1 is known.
+For instance, for the NIST-standardized primes $p=3329$ and $p=8380417$, the exact factorizations of p1p-1 give $c(3329)\geq 0.4613$ and $c(8380417)\geq 0.2932$; see Section 7.
+In particular, the dimension-uniform expected rejection-sampling overhead $1/c(p)$ in the LWE-to-PRIM-LWE reduction [36, Theorem 10] admits an explicit upper bound depending only on $\omega(p-1)$.
+From the computational viewpoint, our main contributions are: explicit lower bounds on $c(p)$ valid for all primes $p>2^{30}$ (Corollary 9), a certified rational-arithmetic algorithm for evaluating $c(p)$ to arbitrary precision given the factorization of p1p-1 (Proposition 8), and tables of $c(p)$ for all primes appearing in current NIST post-quantum cryptographic standards and representative lattice-based implementations (Tables 2 and 3).
+The explicit bound takes the following compact form:
 Corollary 1 (Cryptographic consequence).
-Let
-$P_{3}:=\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{3^{j}}\Bigr)\approx 0.5601260779.$
-Let $q\geq 3$ be a prime modulus with $\omega(q-1)\leq K$. Then
-$\frac{1}{c(q)}\;\leq\;\frac{1}{P_{3}}\prod_{i=1}^{K}\frac{p_{i}}{p_{i}-1}.$
-The general $\omega(q-1)$-based bound is immediate from Theorem 11; the sharper NIST figures use the exact factorizations recorded in Section 6.
+Let $P_{3}:=\prod_{j=1}^{\infty}(1-1/3^{j})\approx 0.5601260779$. If $p\geq 3$ is a prime modulus with $\omega(p-1)\leq K$, then
+$\frac{1}{c(p)}\;\leq\;\frac{1}{P_{3}}\prod_{i=1}^{K}\frac{p_{i}}{p_{i}-1}.$
+The general $\omega(p-1)$-based bound is immediate from Theorem 9. Table 1 lists representative values of $B(K)$, and Tables 2 and 3 collect these bounds for representative cryptographic moduli.
 Related work
@@
 Their setting is different from ours: we work over prime fields and with the primitive-root determinant density arising in PRIM-LWE.
-On the distributional side, Deshouillers and Hassani [4, Proposition 2.1 and equation (2.1)] proved limiting-distribution laws for $\varphi(p-1)/(p-1)$ in the residue classes $p\equiv 1\pmod{m}$, including continuity, and [4, Theorem 1.1] established left non-differentiability of the distribution function at every value $\varphi(m)/m$ with $m$ even.
+On the distributional side, Deshouillers and Hassani [9, Proposition 2.1 and equation (2.1)] proved limiting-distribution laws for $\varphi(p-1)/(p-1)$ in the residue classes $p\equiv 1\pmod{m}$, including continuity, and [9, Theorem 1.1] established left non-differentiability of the distribution function at every value $\varphi(m)/m$ with $m$ even.
 The analogous support question for the classical distribution of $\varphi(n)/n$ over the integers remains open; see [9].
-Organization
-Section 2 records the precise definitions and the open question from [1].
-Section 3 gives the unconditional proof that $\inf_{p}c(p)=0$.
-Section 4 proves the upper bound $\min_{p\leq x}c(p)\ll 1/\log\log x$.
-Section 5 develops the limiting distribution theory, including the Hausdorff-dimension-zero result and the moment Euler product (Section 5.3).
-Section 6 derives explicit lower bounds for primes of cryptographic interest and, in particular, supplies the matching lower bound $\min_{p\leq x}c(p)\gg 1/\log\log x$.
-Section 7 discusses the implications for prim-lwe.
-Notation
-We write $\omega(n)$ for the number of distinct prime factors of an integer $n$, $\pi(x)$ for the number of primes up to $x$, and adopt the convention that $p_{1}=2<p_{2}=3<p_{3}=5<\cdots$ denotes the sequence of all primes in increasing order.
-For a prime $p$, we write
-$P(p):=\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{p^{j}}\Bigr)$
-for the matrix-count Euler product; thus $c(p)=\frac{\varphi(p-1)}{p-1}\,P(p)$.
-All limits are taken as $x\to\infty$ unless stated otherwise.
-We use the standard Vinogradov notation $A\ll B$ and $A\gg B$ synonymously with $A=O(B)$ and $B=O(A)$, respectively.
-$A\asymp B$ means both $A\ll B$ and $A\gg B$.
-We write $\operatorname{Li}(x)=\int_{2}^{x}\mathrm{d}t/\!\log t$ for the logarithmic integral, and throughout $\log$ denotes the natural logarithm.
-We use the notation $A=O^{*}(B)$ to mean $|A|\leq B$, i.e.
-an implied constant of $1$.
 2 Background and the Open Question
-We now record the precise setup from [1] and state the open question. Following [1], define
+Throughout, $\log$ denotes the natural logarithm. We write $\varphi$ for Euler’s totient function, $\omega(n)$ for the number of distinct prime factors of $n$, and $p_{k}$ for the $k$-th prime.
+Following [36], define
 $M_{n}^{\mathrm{prim}}(\mathbb{F}_{p})\;=\;\bigl\{\,M\in M_{n}(\mathbb{F}_{p}):\det(M)\text{ generates }\mathbb{F}_{p}^{*}\,\bigr\}.$
@@
 $\det:\mathrm{GL}_{n}(\mathbb{F}_{p})\longrightarrow\mathbb{F}_{p}^{\times}$
-is a surjective homomorphism, so all its fibres have the same cardinality. Therefore a fraction $\varphi(p-1)/(p-1)$ of invertible matrices have primitive-root determinant. Since
-$|\mathrm{GL}_{n}(\mathbb{F}_{p})|=p^{n^{2}}\prod_{j=1}^{n}\Bigl(1-\frac{1}{p^{j}}\Bigr),$
-it follows that
+is a surjective group homomorphism, so a fraction $\varphi(p-1)/(p-1)$ of invertible matrices have primitive-root determinant. Since
+$|\mathrm{GL}_{n}(\mathbb{F}_{p})|=p^{n^{2}}\prod_{j=1}^{n}\Bigl(1-\frac{1}{p^{j}}\Bigr)$
+(see, e.g., Mullen and Panario [30, Theorem 13.3.3]), it follows that
 $c_{n}(p)\;:=\;\frac{|M_{n}^{\mathrm{prim}}(\mathbb{F}_{p})|}{|M_{n}(\mathbb{F}_{p})|}\;=\;\frac{\varphi(p-1)}{p-1}\;\prod_{j=1}^{n}\Bigl(1-\frac{1}{p^{j}}\Bigr),$
-recovering [1, Corollary 4]. Consequently, the limit
+cf. [36, Corollary 4]. Consequently, the limit
 $c(p)\;:=\;\lim_{n\to\infty}c_{n}(p)\;=\;\frac{\varphi(p-1)}{p-1}\;\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{p^{j}}\Bigr)$
-exists and is strictly positive for each fixed $p$.
-The quantity $c_{n}(p)$ is the exact acceptance probability in dimension $n$.
-Since $c_{n}(p)$ is decreasing in $n$, the dimension-uniform constant that enters the prim-lwe reduction is
-$\inf_{n\geq 1}c_{n}(p)=\lim_{n\to\infty}c_{n}(p)=c(p).$
-Thus $c(p)$ is not the exact finite-dimensional density but the limiting, dimension-uniform reduction constant.
-In the matrix form of decision-LWE, given $\mathbf{A}\in\mathbb{Z}_{q}^{w\times n}$, one must distinguish
-$(\mathbf{A},\,\mathbf{A}\mathbf{S}+\mathbf{E})$
-from uniform over $\mathbb{Z}_{q}^{w\times n}\times\mathbb{Z}_{q}^{w\times n}$, where $\mathbf{S}\in\mathbb{Z}_{q}^{n\times n}$ is the secret matrix and $\mathbf{E}\in\mathbb{Z}_{q}^{w\times n}$ is the error matrix [7].
-Via the identification $\mathbb{Z}/q\mathbb{Z}\cong\mathbb{F}_{q}$ for prime $q$, decision-prim-lwe [1] is the variant in which the secret matrix is additionally required to lie in $M_{n}^{\mathrm{prim}}(\mathbb{F}_{q})$.
-In [1, Remark 3(ii)], Sehrawat, Yeo, and Desmedt observed that if there exist infinitely many primorial primes
-$p=\prod_{i=1}^{k}p_{i}+1,$
-then
-$\frac{\varphi(p-1)}{p-1}=\prod_{i=1}^{k}\frac{p_{i}-1}{p_{i}}\longrightarrow 0$
-by Mertens’ theorem, which would imply $\inf_{p}c(p)=0$. They left open whether this can be proved unconditionally; see [8] for computational results on primorial primes.
+exists and is strictly positive for each fixed $p$. Since $c_{n}(p)$ is decreasing in $n$, $\inf_{n\geq 1}c_{n}(p)=c(p)$.
+Writing $p-1=r_{1}^{e_{1}}\cdots r_{k}^{e_{k}}$ for the prime factorization of p1p-1, one has
+$c(p)\;=\;\prod_{i=1}^{k}\Bigl(1-\frac{1}{r_{i}}\Bigr)\;\cdot\;\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{p^{j}}\Bigr).$ (1)
+Since $P(p):=\prod_{j=1}^{\infty}(1-p^{-j})\to 1$ as $p\to\infty$, the behavior of $c(p)$ over the primes is governed by $\varphi(p-1)/(p-1)$, and the distributional theory in Section 5 reduces to the classical shifted-prime distribution of this ratio.
+In [36, Remark 3(ii)], Sehrawat, Yeo, and Desmedt observed that if there exist infinitely many primorial primes $p=\prod_{i=1}^{k}p_{i}+1$, then $\varphi(p-1)/(p-1)=\prod_{i=1}^{k}(p_{i}-1)/p_{i}\to 0$ by Mertens’ third theorem, which would imply $\inf_{p}c(p)=0$.
+They left open whether this can be proved unconditionally; see [7] for computational results on primorial primes.
 3 Unconditional Proof that $\inf_{p}c(p)=0$
-The argument requires only two classical results.
-Theorem 1 (Dirichlet [9], 1837).
-For coprime integers $a$ and $q$, there are infinitely many primes $p\equiv a\pmod{q}$.
-Theorem 2 (Mertens’ Third Theorem [10], 1874).
+The argument requires only two classical results: Dirichlet’s theorem on primes in arithmetic progressions [10] and the following.
+Theorem 1 (Mertens’ Third Theorem [28]; see [29, Theorem 2.7(e)], 1874).
 As $x\to\infty$,
-$\prod_{\begin{subarray}{c}q\leq x\\ q\text{ prime}\end{subarray}}\Bigl(1-\frac{1}{q}\Bigr)\;\sim\;\frac{e^{-\gamma}}{\log x},$
+$\prod_{\begin{subarray}{c}\ell\leq x\\ \ell\textup{ prime}\end{subarray}}\Bigl(1-\frac{1}{\ell}\Bigr)\;\sim\;\frac{e^{-\gamma}}{\log x},$
 where $\gamma\approx 0.5772$ is the Euler–Mascheroni constant.
-We also record the following observation.
 Lemma 1 (Lower bound for the matrix-count Euler product).
@@
 Proof.
-For each $J\geq 1$, set $a_{j}:=p^{-j}$. Then $0<a_{j}\leq 1$ for all $j$, so the finite-product inequality
-$\prod_{j=1}^{J}(1-a_{j})\geq 1-\sum_{j=1}^{J}a_{j}$
-gives
-$\prod_{j=1}^{J}\Bigl(1-\frac{1}{p^{j}}\Bigr)\;\geq\;1-\sum_{j=1}^{J}\frac{1}{p^{j}}.$
-Letting $J\to\infty$ yields
-$\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{p^{j}}\Bigr)\;\geq\;1-\sum_{j=1}^{\infty}\frac{1}{p^{j}}=1-\frac{1}{p-1}\geq\frac{1}{2}\qquad(p\geq 3).$
-Theorem 3 (Unconditional vanishing of the infimal PRIM-LWE density).
+By the finite-product inequality $\prod(1-a_{j})\geq 1-\sum a_{j}$ (immediate by induction when $0\leq a_{j}\leq 1$) and the geometric series $\sum_{j\geq 1}p^{-j}=1/(p-1)\leq 1/2$. ∎
+Theorem 2 (Unconditional vanishing of $\inf c(p)$).
+The following identities hold:
 $\inf_{p\textup{ prime}}c(p)=0,\qquad\liminf_{\begin{subarray}{c}p\to\infty\\ p\textup{ prime}\end{subarray}}c(p)=0.$
@@
 $N_{k}\;=\;\prod_{i=1}^{k}p_{i}.$
-Since $\gcd(1,N_{k})=1$, Theorem 1 guarantees infinitely many primes in the arithmetic progression $\{1+j\,N_{k}:j\geq 1\}$. Fix any such prime $p\equiv 1\pmod{N_{k}}$, so that $N_{k}\mid(p-1)$.
-Because every prime $q\leq p_{k}$ divides p1p-1, the multiplicative formula for the totient gives
-$\frac{\varphi(p-1)}{p-1}\;=\;\prod_{q\mid(p-1)}\Bigl(1-\frac{1}{q}\Bigr)\;\leq\;\prod_{i=1}^{k}\Bigl(1-\frac{1}{p_{i}}\Bigr),$
-where the inequality holds because each factor $\bigl(1-\frac{1}{q}\bigr)$ lies in $(0,1)$, so the full product over all prime divisors $q\mid(p-1)$ is no larger than the partial product over the subset $p_{1},\dots,p_{k}$.
+Since $\gcd(1,N_{k})=1$, Dirichlet’s theorem guarantees infinitely many primes in the arithmetic progression $\{1+j\,N_{k}:j\geq 1\}$.
+Fix any such prime $p\equiv 1\pmod{N_{k}}$, so that $N_{k}\mid(p-1)$.
+Because every prime $\ell\leq p_{k}$ divides p1p-1, the multiplicative formula for the totient gives
+$\frac{\varphi(p-1)}{p-1}\;=\;\prod_{\ell\mid(p-1)}\Bigl(1-\frac{1}{\ell}\Bigr)\;\leq\;\prod_{i=1}^{k}\Bigl(1-\frac{1}{p_{i}}\Bigr).$
 By Mertens’ third theorem (Theorem 1),
@@
 4 Quantitative Refinements
-We first record the form of Linnik’s theorem used in the proof.
-Theorem 4 (Least prime in an arithmetic progression; Linnik [11]).
-There exist absolute constants $C_{0},L>0$ such that for every modulus $q\geq 1$ and every residue class $a\pmod{q}$ with $\gcd(a,q)=1$, the least prime $p(a,q)\equiv a\pmod{q}$ satisfies
-$p(a,q)\leq C_{0}q^{L}.$
-Let $N_{k}:=\prod_{i=1}^{k}p_{i}$ denote the $k$-th primorial, $\vartheta(x):=\sum_{\begin{subarray}{c}p\leq x,\;p\text{ prime}\end{subarray}}\log p$ the Chebyshev function, and $m(x):=\max\{k\geq 1:N_{k}\leq x\}$.
-Lemma 2 (Prime-number-theorem consequences for primorials; cf. [12]).
+By Linnik’s theorem [26], there exist absolute constants $C_{0},L>0$ such that the least prime in any reduced residue class modulo $d$ is at most $C_{0}d^{L}$.
+Let $N_{k}$ denote the $k$-th primorial, $\vartheta$ the Chebyshev function, and $m(x)$ the primorial-counting function:
+$N_{k}:=\prod_{i=1}^{k}p_{i},\qquad\vartheta(x):=\sum_{\begin{subarray}{c}p\leq x\\ p\text{ prime}\end{subarray}}\log p,\qquad m(x):=\max\{k\geq 1:N_{k}\leq x\}.$
+Lemma 2 (Prime-number-theorem consequences for primorials; cf. [29, Corollary 2.5]).
 As $k\to\infty$,
@@
 Proof.
-The first two assertions are standard consequences of the prime number theorem. Since
-$N_{m(x)}\leq x<N_{m(x)+1}=N_{m(x)}\,p_{m(x)+1},$
-we have
-$\log N_{m(x)}\leq\log x<\log N_{m(x)}+\log p_{m(x)+1}.$
-Now
-$\log p_{m(x)+1}=o\!\bigl(p_{m(x)}\bigr)=o\!\bigl(\log N_{m(x)}\bigr),$
-because $\log N_{m(x)}=\vartheta(p_{m(x)})\sim p_{m(x)}$ and $p_{m(x)+1}\sim p_{m(x)}$. Hence $\log x\sim\log N_{m(x)}$, and therefore $p_{m(x)}\sim\log x$. ∎
+The first two assertions follow from $\vartheta(x)\sim x$. For the last two, $N_{m(x)}\leq x<N_{m(x)}p_{m(x)+1}$ gives $\log x\sim\log N_{m(x)}\sim p_{m(x)}$, since $\log p_{m(x)+1}=o(p_{m(x)})$. ∎
 Proposition 1 (Upper bound for the minimum density up to $x$).
@@
 Proof.
-By Theorem 4, fix absolute constants $C_{0},L>0$ such that the least prime in any reduced residue class modulo $q$ is at most $C_{0}q^{L}$.
+Fix absolute constants $C_{0},L>0$ from Linnik’s theorem such that the least prime in any reduced residue class modulo $d$ is at most $C_{0}d^{L}$.
 For each $k\geq 1$, let $p^{*}(k)$ denote the least prime satisfying $p^{*}(k)\equiv 1\pmod{N_{k}}$.
@@
 as claimed. ∎
+The matching lower bound $\min_{p\leq x,\,p\text{ prime}}c(p)\gg 1/\!\log\log x$, which completes the proof that the order is sharp, is established in Corollary 7 below; its proof requires only the primorial counting argument and Lemma 1.
 5 The Limiting Distribution of $c(p)$
-We now turn to the distributional theory of $c(p)$ over the primes.
-The classical existence and continuity of the limiting distribution of $\varphi(p-1)/(p-1)$ over primes follow from the shifted-prime Erdős–Wintner framework; for the shift a=1a=-1 used here, direct references are Tenenbaum [3, pp.
-262–263] and Deshouillers–Hassani [4].
+The classical existence and continuity of the limiting distribution of $\varphi(p-1)/(p-1)$ over primes follow from the shifted-prime Erdős–Wintner framework [16] (Tenenbaum [38, pp.
+262–263], Deshouillers–Hassani [9]).
 Throughout this section we write
 $f(n)\;:=\;\log\frac{n}{\varphi(n)}\;=\;\sum_{\begin{subarray}{c}\ell\mid n\\ \ell\textup{ prime}\end{subarray}}\log\frac{\ell}{\ell-1}$ (2)
-for the strongly additive function that records $\log(n/\varphi(n))$.
-Our contribution in this section is the full-support conclusion: the limiting law has support exactly $0,\tfrac{1}{2}]$ (Theorem 6), combining the left non-differentiability result of [4, Theorem 1.1] with the density of $\{\varphi(m)/m:m\text{ even}\}$ in $[0,1/2]$ (Lemma 4).
-Since $P(p)\to 1$, the prim-lwe density $c(p)$ shares the same limiting distribution (Lemma 5).
-We supply a self-contained proof of the shifted-prime pure singularity in Lemma 9; in Section 5.3 we show that the limiting measure has Hausdorff dimension 0 and derive the moments of $G$ as convergent Euler products.
 5.1 The distribution of $\varphi(p-1)/(p-1)$
-One can bypass the general Erdős–Wintner machinery and use the following direct p1p-1-specific input from Deshouillers and Hassani.
-Lemma 3 (Erdős–Wintner hypotheses for the totient function).
+Lemma 3 (Verification of the Erdős–Wintner hypotheses).
 The function $f$ defined in (2) is strongly additive, with $f(n)=\log(n/\varphi(n))=-\log(\varphi(n)/n)$. For every prime $\ell$,
 $0<f(\ell)=\log\frac{\ell}{\ell-1}\leq\log 2<1,\qquad f(\ell)=\frac{1}{\ell}+O\!\left(\frac{1}{\ell^{2}}\right).$
-The three Erdős–Wintner convergence conditions are satisfied, as is the continuity criterion of Lemma 6. The same conclusions hold for $g(n)=\varphi(n)/n$.
+The three convergence conditions of the shifted-prime Erdős–Wintner theorem hold:
+$\sum_{|f(\ell)|>1}\frac{1}{\ell}<\infty,\qquad\sum_{\ell}\frac{f(\ell)}{\ell}<\infty,\qquad\sum_{\ell}\frac{f(\ell)^{2}}{\ell}<\infty,$
+as does the continuity criterion $\sum_{f(\ell)\neq 0}1/\ell=\infty$. The same conclusions hold for $g(n)=\varphi(n)/n$.
 Proof.
@@
 The estimate $f(\ell)=1/\ell+O(1/\ell^{2})$ gives $\sum_{\ell}f(\ell)/\ell<\infty$ and $\sum_{\ell}f(\ell)^{2}/\ell<\infty$.
-Since $f(\ell)\neq 0$ for every prime $\ell$ and $\sum_{\ell}1/\ell=\infty$, the continuity criterion of Lemma 6 is satisfied.
+Since $f(\ell)\neq 0$ for every prime $\ell$ and $\sum_{\ell}1/\ell=\infty$, the continuity criterion is satisfied.
 The case $g(n)=\varphi(n)/n$ follows from $\log g(\ell)=-f(\ell)$.
 ∎
-Proposition 2 (Deshouillers–Hassani; special case of [4, Proposition 2.1 and equation (2.1)]).
+Proposition 2 (Deshouillers–Hassani; special case of [9, Proposition 2.1 and equation (2.1)]).
 For each integer $m\geq 1$, there exists a continuous distribution function $G_{m}\colon\mathbb{R}\to[0,1]$ such that for every $\alpha\in\mathbb{R}$,
-$\lim_{x\to\infty}\frac{1}{\pi(x;m,1)}\#\left\{p\leq x:p\text{ prime},\ p\equiv 1\pmod{m},\;\frac{\varphi(p-1)}{p-1}\leq\alpha\right\}=G_{m}(\alpha).$
+$\lim_{x\to\infty}\frac{1}{\pi(x;m,1)}\#\left\{p\leq x:p\textup{ prime},\ p\equiv 1\pmod{m},\;\frac{\varphi(p-1)}{p-1}\leq\alpha\right\}=G_{m}(\alpha).$
 Moreover, the limiting measure is supported on $0,\varphi(m)/m]$.
 Proof.
-Apply Deshouillers and Hassani [4, Proposition 2.1] to
+Apply Deshouillers and Hassani [9, Proposition 2.1] to
 $g(n):=\frac{\varphi(n)}{n}.$
-By Lemma 3 applied to $\log g(\ell)=-\log(\ell/(\ell-1))$, the three convergence hypotheses of [4, Proposition 2.1] are satisfied, and there exists a distribution function $G_{m}$ such that the displayed limit holds at every continuity point of $G_{m}$.
+By Lemma 3 applied to $\log g(\ell)=-\log(\ell/(\ell-1))$, the three convergence hypotheses of [9, Proposition 2.1] are satisfied, and there exists a distribution function $G_{m}$ such that the displayed limit holds at every continuity point of $G_{m}$.
 By the continuity criterion in the same proposition, $G_{m}$ is continuous if and only if
@@
 $\sum_{\ell\equiv 1\pmod{m}}\frac{1}{\ell}=\infty$
-by the divergence of the prime harmonic series in every reduced residue class. Thus $G_{m}$ is continuous, so the displayed limit holds for every $\alpha\in\mathbb{R}$.
+by the divergence of the prime harmonic series in every reduced residue class (see, e.g., [29, Corollary 4.10]).
+Thus $G_{m}$ is continuous, so the displayed limit holds for every $\alpha\in\mathbb{R}$.
 If $p\equiv 1\pmod{m}$, then every prime divisor of $m$ also divides p1p-1, so
-$\frac{\varphi(p-1)}{p-1}=\prod_{q\mid(p-1)}\Bigl(1-\frac{1}{q}\Bigr)\leq\prod_{q\mid m}\Bigl(1-\frac{1}{q}\Bigr)=\frac{\varphi(m)}{m},$
+$\frac{\varphi(p-1)}{p-1}=\prod_{\ell\mid(p-1)}\Bigl(1-\frac{1}{\ell}\Bigr)\leq\prod_{\ell\mid m}\Bigl(1-\frac{1}{\ell}\Bigr)=\frac{\varphi(m)}{m},$
 which yields the support claim. ∎
-We shall also use the following standard Siegel–Walfisz (see [12, 13]) estimate in a range of polylogarithmic moduli.
+We shall also use the following standard Siegel–Walfisz (see [29, Corollary 11.21] or [39, Theorem II.8.17]) estimate in a range of polylogarithmic moduli.
 Theorem 3 (Siegel–Walfisz).
@@
 where
-$\pi(x;m,a)=\#\{p\leq x:p\text{ prime},\ p\equiv a\pmod{m}\}$.
-For an independent, self-contained proof (based on Siegel–Walfisz) that intervals near the endpoint $1/2$ carry positive limiting measure, see Lemma 11 in Appendix A.
-That result is not used in what follows.
+$\pi(x;m,a)=\#\{p\leq x:p\textup{ prime},\ p\equiv a\pmod{m}\}$.
 Lemma 4 (Density of even totient ratios).
@@
 Since $\varphi(m)/m$ depends only on the distinct prime divisors of $m$, we have
-$\left\{\frac{\varphi(m)}{m}:m\geq 1\text{ even}\right\}=\left\{\frac{1}{2}\prod_{p\in S}\left(1-\frac{1}{p}\right):S\subset\{\text{odd primes}\},\ S\text{ finite}\right\}.$
+$\left\{\frac{\varphi(m)}{m}:m\geq 1\text{ even}\right\}=\left\{\frac{1}{2}\prod_{r\in S}\left(1-\frac{1}{r}\right):S\subset\{\text{odd primes}\},\ S\text{ finite}\right\}.$
 Enumerate the odd primes as $r_{1}<r_{2}<\cdots$, and set
@@
 is dense in $(0,1/2]$. Since the finite subsums can be arbitrarily large, 0 is also a limit point. Hence the displayed set is dense in $[0,1/2]$, as claimed. ∎
-Theorem 6 (Continuous limiting distribution and full support of the totient ratio over primes).
+Theorem 4 (Limiting distribution and full support over primes).
 There exists a continuous distribution function $G\colon\mathbb{R}\to[0,1]$ such that for every $\alpha\in\mathbb{R}$,
-$\lim_{x\to\infty}\;\frac{1}{\pi(x)}\,\#\bigl\{p\leq x:p\text{ prime},\;\varphi(p-1)/(p-1)\leq\alpha\bigr\}\;=\;G(\alpha).$
+$\lim_{x\to\infty}\;\frac{1}{\pi(x)}\,\#\bigl\{p\leq x:p\textup{ prime},\;\varphi(p-1)/(p-1)\leq\alpha\bigr\}\;=\;G(\alpha).$
 The limiting law has support exactly $0,\tfrac{1}{2}]$, with $G(\alpha)=0$ for $\alpha\leq 0$ and $G(\alpha)=1$ for $\alpha\geq 1/2$. In particular, $G$ is strictly increasing on $0,\tfrac{1}{2}]$.
@@
 $\pi(x;2,1)=\pi(x)-1\sim\pi(x),$
-so the distribution function $G_{2}$ furnished by that proposition is the unconditional limiting distribution of $\varphi(p-1)/(p-1)$ over all primes.
-In particular, $G_{2}$ coincides with the distribution function denoted $F$ in [4, Theorem 1.1].
-Writing $G:=G_{2}$, there exists a continuous distribution function $G\colon\mathbb{R}\to[0,1]$ such that for every $\alpha\in\mathbb{R}$,
+so $G_{2}$ is the unconditional limiting distribution of $\varphi(p-1)/(p-1)$ over all primes.
+Writing $G:=G_{2}$ (denoted $F$ in [9, Theorem 1.1]), there exists a continuous distribution function $G\colon\mathbb{R}\to[0,1]$ such that for every $\alpha\in\mathbb{R}$,
 $\lim_{x\to\infty}\frac{1}{\pi(x)}\#\bigl\{p\leq x:p\text{ prime},\;\varphi(p-1)/(p-1)\leq\alpha\bigr\}=G(\alpha).$
@@
 $x_{m}:=\frac{\varphi(m)}{m}.$
-By Deshouillers and Hassani [4, Theorem 1.1], the distribution function $G$ is not left-differentiable at $x_{m}$.
+By Deshouillers and Hassani [9, Theorem 1.1], the distribution function $G$ is not left-differentiable at $x_{m}$.
 Deshouillers and Hassani denote this distribution function by $F$; we write it as $G$.
@@
 so $G$ is strictly increasing on $0,\tfrac{1}{2}]$. The displayed bounds $0<G(\tau)<1$ for $0<\tau<1/2$ follow immediately from this and the endpoint values $G(0)=0$, $G(1/2)=1$. ∎
-Theorem 6 completely determines the shifted-prime support: it is exactly $0,\tfrac{1}{2}]$.
-This does not settle the classical support question for the distribution of $\varphi(n)/n$ over the integers, which remains open; see [4].
-The residue-class laws in [4, Proposition 2.1 and equation (2.1)] are consistent with this picture, since for primes $p\equiv 1\pmod{m}$ the corresponding limiting measure is supported in $0,\varphi(m)/m]$.
+The support question for $\varphi(n)/n$ over the integers remains open; see Schoenberg [35] and [9].
 5.2 From $\varphi(p-1)/(p-1)$ to $c(p)$
-Since the Euler product $P(p)\to 1$, the density $c(p)=\frac{\varphi(p-1)}{p-1}P(p)$ inherits the limiting distribution of $\varphi(p-1)/(p-1)$. The following lemma records this fact.
 Lemma 5 (Transport lemma).
 Let $X(p)\geq 0$ and $a(p)>0$ be functions of the primes. Assume that $X(p)$ has a continuous limiting distribution function $G$ over the primes, i.e.
-$\lim_{x\to\infty}\frac{1}{\pi(x)}\#\bigl\{p\leq x:p\text{ prime},\ X(p)\leq\alpha\bigr\}=G(\alpha)\qquad(\alpha\in\mathbb{R}).$
+$\lim_{x\to\infty}\frac{1}{\pi(x)}\#\bigl\{p\leq x:p\textup{ prime},\ X(p)\leq\alpha\bigr\}=G(\alpha)\qquad(\alpha\in\mathbb{R}).$
 If $a(p)\to 1$ as $p\to\infty$ through the primes, then $a(p)X(p)$ has the same limiting distribution $G$.
 Proof.
-For $\alpha<0$ there is nothing to prove, since both distribution functions vanish. Fix $\alpha\geq 0$ and $0<\varepsilon<1$. Since $a(p)\to 1$, there exists $P_{0}=P_{0}(\varepsilon)$ such that
-$1-\varepsilon<a(p)<1+\varepsilon\qquad(p>P_{0},\ p\text{ prime}).$
-For such primes $p$,
-$X(p)\leq\frac{\alpha}{1+\varepsilon}\;\Longrightarrow\;a(p)X(p)\leq\alpha\;\Longrightarrow\;X(p)\leq\frac{\alpha}{1-\varepsilon}.$
-Hence
-$\frac{1}{\pi(x)}\#\left\{p\leq x:p\text{ prime},\ X(p)\leq\frac{\alpha}{1+\varepsilon}\right\}\leq\frac{1}{\pi(x)}\#\bigl\{p\leq x:p\text{ prime},\ a(p)X(p)\leq\alpha\bigr\}+\frac{\pi(P_{0})}{\pi(x)},$
-and
-$\frac{1}{\pi(x)}\#\bigl\{p\leq x:p\text{ prime},\ a(p)X(p)\leq\alpha\bigr\}\leq\frac{1}{\pi(x)}\#\left\{p\leq x:p\text{ prime},\ X(p)\leq\frac{\alpha}{1-\varepsilon}\right\}+\frac{\pi(P_{0})}{\pi(x)}.$
-Taking $\liminf$ and $\limsup$ as $x\to\infty$, and then letting $\varepsilon\downarrow 0$, continuity of $G$ gives the claim. ∎
-Theorem 7 (Transfer of the limiting distribution to the PRIM-LWE density).
+Fix $\alpha\geq 0$ and $0<\varepsilon<1$. For all sufficiently large primes, $1-\varepsilon<a(p)<1+\varepsilon$, so
+$\{X(p)\leq\alpha/(1+\varepsilon)\}\;\subseteq\;\{a(p)X(p)\leq\alpha\}\;\subseteq\;\{X(p)\leq\alpha/(1-\varepsilon)\}$
+up to finitely many exceptions. Taking densities as $x\to\infty$ and letting $\varepsilon\downarrow 0$, continuity of $G$ gives the claim. ∎
+Theorem 5 (Transfer to the PRIM-LWE density).
 The quantity $c(p)$ has the same limiting distribution as $\varphi(p-1)/(p-1)$. That is, for every $\alpha\in\mathbb{R}$,
-$\lim_{x\to\infty}\;\frac{1}{\pi(x)}\,\#\bigl\{p\leq x:p\text{ prime},\;c(p)\leq\alpha\bigr\}\;=\;G(\alpha),$
+$\lim_{x\to\infty}\;\frac{1}{\pi(x)}\,\#\bigl\{p\leq x:p\textup{ prime},\;c(p)\leq\alpha\bigr\}\;=\;G(\alpha),$
 where $G$ is the distribution function from Theorem 4.
@@
 By Theorem 4, $X(p)$ has continuous limiting distribution function $G$. Therefore Lemma 5 applies with $a(p)=P(p)$, and shows that $c(p)=a(p)X(p)$ has the same limiting distribution $G$. ∎
+5.3 Pure singularity
+Having established that $c(p)$ shares the limiting distribution of $\varphi(p-1)/(p-1)$, we now investigate the fine structure of this law.
+By the Jessen–Wintner theorem [24], any limiting distribution arising from an infinite convolution of discrete measures must be of pure type—purely discrete, purely singular, or absolutely continuous.
+Continuity was already verified in Lemma 3, so the remaining question is whether the law is absolutely continuous or purely singular.
+The latter turns out to hold: the values of $c(p)$ cluster on a fractal-like set of Lebesgue measure zero, as we prove by adapting a classical argument of Erdős to the shifted-prime setting.
 Lemma 6 (Shifted-prime pure-type and continuity criterion).
@@
 Proof.
-The relevant shifted-prime Erdős–Wintner framework is due to Kátai [14] and Hildebrand [15] for the shift $a=1$; the methods extend to any fixed nonzero shift via the Bombieri–Vinogradov theorem, and for the shift a=1a=-1 used here, the explicit characteristic function is recorded in Tenenbaum [3, formula (3), p.
+The relevant shifted-prime Erdős–Wintner framework is due to Kátai [25, Theorem 2] and Hildebrand [23, Theorem 5] for the shift $a=1$; the methods extend to any fixed nonzero shift via the Bombieri–Vinogradov theorem [4] (see also [39, Theorem II.8.34]): the only property of the shift used in these proofs is the Bombieri–Vinogradov estimate for primes in arithmetic progressions, which holds uniformly for any fixed nonzero $a$.
+Hildebrand [23, p.
+211] notes explicitly that the same methods apply to $\{p+a\}$ for arbitrary nonzero $a$.
+For the shift a=1a=-1 used here, the explicit characteristic function is recorded in Tenenbaum [38, formula (3), p.
 262].
 The same conclusions hold for general (not necessarily strongly) additive functions under the Erdős–Wintner hypotheses, by the identical Jessen–Wintner argument applied to the full local factors; we state the strongly additive case since it is all that we require.
-By Tenenbaum [3, formula (3), p. 262], the characteristic function of $\mu_{f}$ is given by
-$\widehat{\mu_{f}}(\tau)\;=\;\prod_{p}\!\left(1-\frac{1}{p-1}+\sum_{\nu\geq 1}\frac{e^{i\tau f(p^{\nu})}}{p^{\nu}}\right)$
-The shift is p1p-1, corresponding to a=1a=-1 in the notation of [3, formula (3)].
+By Tenenbaum [38, formula (3), p. 262], the characteristic function of $\mu_{f}$ is given by
+$\widehat{\mu_{f}}(\tau)\;=\;\prod_{\ell\textup{ prime}}\!\left(1-\frac{1}{\ell-1}+\sum_{\nu\geq 1}\frac{e^{i\tau f(\ell^{\nu})}}{\ell^{\nu}}\right).$
+The shift is p1p-1, corresponding to a=1a=-1 in the notation of [38, formula (3)].
 Since no prime divides 1-1, the product runs over all primes.
@@
 This factor carries no randomness and does not affect pure type or continuity; it merely shifts the law.
-It therefore suffices to apply the Jessen–Wintner theorem [16] (see also Tenenbaum [13, Theorem 2.7(b)]) to the infinite convolution over the odd primes.
-Each odd-prime factor is the characteristic function of a probability measure supported on a countable set, so the Jessen–Wintner theorem gives that the resulting infinite convolution is of pure type.
-The continuity criterion follows from the same theorem: the convolution is continuous if and only if $\sum(1-\sigma_{\ell})=\infty$, where $\sigma_{\ell}$ is the maximal atom of the factor at $\ell$.
-Since $f$ is strongly additive, the local factor at an odd prime $\ell$ places mass $(\ell-2)/(\ell-1)$ at the origin and mass $1/(\ell-1)$ at $f(\ell)$.
-When $f(\ell)\neq 0$, these two points are distinct, so $\sigma_{\ell}=(\ell-2)/(\ell-1)$ and $1-\sigma_{\ell}=1/(\ell-1)\asymp 1/\ell$.
+For each odd prime $\ell$, the remaining local factor is the characteristic function of an atomic probability measure $\nu_{\ell}$ placing mass $(\ell-2)/(\ell-1)$ at the origin and mass $1/(\ell-1)$ at $f(\ell)$.
+The product formula therefore represents the characteristic function of the infinite convolution $\delta_{f(2)}\ast\mathop{{\ast}}_{\ell\geq 3}\nu_{\ell}$.
+By Lévy’s uniqueness theorem (see, e.g., [39, Section III.2]), $\mu_{f}$ equals this infinite convolution.
+It therefore suffices to apply the Jessen–Wintner theorem (see also Tenenbaum [39, Theorem III.2.7(b)]) to the convolution over the odd primes.
+Each factor $\nu_{\ell}$ is supported on a countable (indeed two-element) set, so the Jessen–Wintner theorem gives that the resulting infinite convolution—and hence $\mu_{f}$ itself—is of pure type.
+The continuity criterion follows from the same theorem: the convolution is continuous if and only if $\sum(1-\sigma_{\ell})=\infty$, where $\sigma_{\ell}$ is the maximal atom of $\nu_{\ell}$.
+Since $\nu_{\ell}$ places mass $(\ell-2)/(\ell-1)$ at the origin and mass $1/(\ell-1)$ at $f(\ell)$, when $f(\ell)\neq 0$ these two points are distinct, so $\sigma_{\ell}=(\ell-2)/(\ell-1)$ and $1-\sigma_{\ell}=1/(\ell-1)\asymp 1/\ell$.
 It follows that  $\sum_{\ell\text{ odd}}(1-\sigma_{\ell})$ diverges if and only if $\sum_{f(\ell)\neq 0}1/\ell=\infty$, since the prime $\ell=2$ does not enter the odd-prime convolution.
@@
 ∎
-For $y\geq 3$ and each prime $p$, define $a_{y}(p-1):=\prod_{\begin{subarray}{c}\ell\leq y,\;\ell\textup{ prime},\;\ell\mid(p-1)\end{subarray}}\ell$, and write $P^{+}(n)$ for the largest prime factor of $n$.
+For $y\geq 3$ and each prime $p$, define
+$a_{y}(p-1):=\prod_{\begin{subarray}{c}\ell\leq y\\ \ell\textup{ prime}\\ \ell\mid(p-1)\end{subarray}}\ell,$
+and write $P^{+}(n)$ for the largest prime factor of $n$.
 Lemma 7 (Prime density for a prescribed smooth kernel).
@@
 $\lim_{x\to\infty}\frac{1}{\pi(x)}\#\bigl\{p\leq x:p\textup{ prime},\;a_{y}(p-1)=m\bigr\}\;=\;\delta_{y}(m).$
-Moreover, writing $C_{y}:=\prod_{3\leq\ell\leq y}(\ell-2)/(\ell-1)\asymp 1/\!\log y$, we have $\delta_{y}(m)\geq C_{y}/\varphi(m)$.
+Moreover, writing
+$C_{y}:=\prod_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\textup{ prime}\end{subarray}}\frac{\ell-2}{\ell-1}\;\asymp\;\frac{1}{\log y},$
+we have $\delta_{y}(m)\geq C_{y}/\varphi(m)$.
 Proof.
@@
 since $\log((\ell-1)/(\ell-2))=1/\ell+O(1/\ell^{2})$ and $\sum_{\ell}1/\ell^{2}<\infty$.
-By Mertens’ second theorem the sum over primes equals $\log\log y+O(1)$, so $C_{y}=e^{-\log\log y+O(1)}\asymp 1/\!\log y$.
+By Mertens’ second theorem [29, Theorem 2.7(d)] the sum over primes equals $\log\log y+O(1)$, so $C_{y}=e^{-\log\log y+O(1)}\asymp 1/\!\log y$.
 ∎
@@
 Proof.
-We adapt the integer-case argument of Erdős [5] (see also [13, Exercise 256] with solution in [17, Solution 2.4, pp.
+We adapt the integer-case argument of Erdős [14, pp.
+722–725] (see also [39, Exercise 256] with solution in [40, Solution 2.4, pp.
 147–148]) to the shifted-prime setting, using the Siegel–Walfisz theorem (Theorem 3) in place of the integer equidistribution input.
-By the shifted-prime Erdős–Wintner theorem (Tenenbaum [3, pp.
-262–263] for the shift a=1a=-1; Kátai [14], Hildebrand [15] for the original $a=1$ case), the limiting measure $\mu_{f}$ exists.
+By the shifted-prime Erdős–Wintner theorem (Lemma 6), the limiting measure $\mu_{f}$ exists.
 Indeed, by Lemma 3 the three convergence hypotheses are satisfied and the continuity criterion holds, so the limiting measure $\mu_{f}$ exists and is continuous.
@@
 Let $m_{1}<m_{2}<\cdots<m_{R}$ be the first $R=\lfloor y^{\delta}\rfloor$ even squarefree positive integers.
-Since the counting function of even squarefree integers satisfies $A(t):=\#\{m\leq t:m\text{ even squarefree}\}\sim(2/\pi^{2})\,t$, for any fixed $C>\pi^{2}/2$ we have $A(Cj)\geq j$ for all sufficiently large $j$.
+Since the counting function of even squarefree integers satisfies $A(t):=\#\{m\leq t:m\text{ even squarefree}\}\sim(2/\pi^{2})\,t$ (see, e.g., [39, Theorem I.3.10]), for any fixed $C>\pi^{2}/2$ we have $A(Cj)\geq j$ for all sufficiently large $j$.
 After enlarging $C$ if necessary, we may assume $m_{j}\leq Cj$ for all $j\geq 1$.
@@
 By Lemma 7, the relative density of $\{p:a_{y}(p-1)=m_{j}\}$ among the primes is $\delta_{y}(m_{j})\geq C_{y}/\varphi(m_{j})$, where $C_{y}\asymp 1/\!\log y$. Combining with (4),
-$\sum_{j=1}^{R}\delta_{y}(m_{j})\;\geq\;C_{y}\sum_{j=1}^{R}\frac{1}{\varphi(m_{j})}\;\gg\;\frac{\delta\log y}{\log y}\;=\;\delta\;>\;0.$ (4)
+$\sum_{j=1}^{R}\delta_{y}(m_{j})\;\geq\;C_{y}\sum_{j=1}^{R}\frac{1}{\varphi(m_{j})}\;\gg\;\frac{\delta\log y}{\log y}\;=\;\delta\;>\;0.$
 Since the $m_{j}$ are distinct, the sets $\{p:a_{y}(p-1)=m_{j}\}$ are pairwise disjoint, so the relative density of their union $\mathcal{A}_{y}:=\{p:a_{y}(p-1)\in\{m_{1},\ldots,m_{R}\}\}$ is the sum $\sum_{j}\delta_{y}(m_{j})$.
@@
 Proof.
-By Lemma 3, the three convergence conditions of the shifted-prime Erdős–Wintner theorem (Kátai [14], Hildebrand [15] for $p+1$; Tenenbaum [3, pp.
-262–263] for the general shift including p1p-1) are satisfied, and the continuity criterion holds.
+By Lemma 3, the three convergence conditions of the shifted-prime Erdős–Wintner theorem (Lemma 6) are satisfied, and the continuity criterion holds.
 Therefore f(p1)f(p-1) possesses a continuous limiting distribution $\mu_{f}$, which is of pure type by Lemma 6.
@@
 Proof.
-Write
-$Y(p):=\frac{\varphi(p-1)}{p-1}.$
-By Theorem 7, $c(p)$ has the same limiting law as $Y(p)$.
-By Theorem 6, this common law is continuous, has support exactly $0,\tfrac{1}{2}]$, and its distribution function is strictly increasing on $0,\tfrac{1}{2}]$.
-It remains only to justify pure singularity.
-Let $f$ be as in (1). Then $f$ is strongly additive and
-$f(\ell)=\log\frac{\ell}{\ell-1}\sim\frac{1}{\ell}\qquad(\ell\to\infty,\ \ell\ \text{prime}),$
-while
-$Y(p)=e^{-f(p-1)}.$
-By Proposition 3, the empirical measures of f(p1)f(p-1) converge weakly to a continuous purely singular law $\mu_{f}$.
-Since $T(u):=e^{-u}$ is continuous on $0,\infty)$, the continuous mapping theorem implies that
-Y(p)=T(f(p1))Y(p)=T(f(p-1))
-has limiting law $T_{\#}\mu_{f}$ (the push-forward of $\mu_{f}$ by $T$, defined by $T_{\#}\mu_{f}(B):=\mu_{f}(T^{-1}(B))$).
-Choose a Borel set $E\subset[0,\infty)$ with Lebesgue measure zero and $\mu_{f}(E)=1$.
-The map $T$ is strictly monotone and Lipschitz on each compact interval $[m,m+1]$, $m\geq 0$.
-Writing $E=\bigcup_{m\geq 0}(E\cap[m,m+1])$, each image $T(E\cap[m,m+1])$ has Lebesgue measure zero (since a Lipschitz map on a compact interval preserves null sets), and therefore $T(E)$ has Lebesgue measure zero by countable subadditivity.
-Therefore $T_{\#}\mu_{f}$ is supported on the null set $T(E)$ and is therefore purely singular.
-By Theorem 7, the same is true for $c(p)$. ∎
-Note that $c(p)<1/2$ for every fixed prime $p$. Indeed, for $p=2$ one has
-$\frac{\varphi(p-1)}{p-1}=\frac{\varphi(1)}{1}=1,$
-so
-$c(2)=\prod_{j\geq 1}\Bigl(1-2^{-j}\Bigr)\approx 0.288788<\frac{1}{2}.$
-For odd $p$, one has $\varphi(p-1)/(p-1)\leq 1/2$ together with $\prod_{j\geq 1}(1-p^{-j})<1$. Thus $1/2$ is only a boundary point of the support of the limiting law for $c(p)$, not an attained value.
+Combine Theorems 4 and 5: $c(p)$ has a continuous limiting law with support $0,\tfrac{1}{2}]$ whose distribution function is strictly increasing. It remains to justify pure singularity.
+By Proposition 3, the shifted-prime law $\mu_{f}$ of f(p1)f(p-1) (with $f$ as in (2)) is continuous and purely singular.
+Since $\varphi(p-1)/(p-1)=e^{-f(p-1)}$ and $T(u):=e^{-u}$ is locally Lipschitz on $0,\infty)$, the push-forward $T_{\#}\mu_{f}$ preserves Lebesgue-null sets and is therefore purely singular.
+By Theorem 5, the same holds for $c(p)$.
+∎
+Figure 1 compares the empirical distribution of $c(p)$ for primes $p\leq 10^{6}$ with the limiting law $G$ of Theorem 5; the close agreement already visible at this scale provides a useful numerical check on the distributional theory.
+Figure 1: Empirical CDF of $c(p)$ over the 78,49778{,}497 primes $3\leq p\leq 10^{6}$, overlaid with the theoretical limiting distribution function $G$ of Theorem 5, estimated by $10^{7}$ Monte Carlo samples of the infinite convolution of Lemma 10 truncated at the 10,00010{,}000th odd prime.
+The empirical values of $c(p)$ were computed via the truncated Euler product of Proposition 8 with $J=20$ terms, contributing error below $10^{-9}$ for every prime $p\geq 3$.
+Remark 1.
+The value $1/2$ is not attained: $c(p)<1/2$ for every prime $p$.
+Indeed, for $p=2$ one has $\varphi(p-1)/(p-1)=1$, so $c(2)=\prod_{j\geq 1}(1-2^{-j})\approx 0.288788<1/2$.
+For odd $p$, one has $\varphi(p-1)/(p-1)\leq 1/2$ together with $\prod_{j\geq 1}(1-p^{-j})<1$.
+Thus $1/2$ is only a boundary point of the support of the limiting law for $c(p)$, not an attained value.
 Corollary 3 (Unconditional limsup).
+We have
 $\limsup_{\begin{subarray}{c}p\to\infty\\ p\textup{ prime}\end{subarray}}c(p)=\frac{1}{2}.$
@@
 Proof.
-Since $c(p)<1/2$ for every prime, $\limsup c(p)\leq 1/2$.
+Since $c(p)<1/2$ for every prime (Remark 1), $\limsup c(p)\leq 1/2$.
 For the reverse inequality, fix $\varepsilon>0$.
@@
 In particular, infinitely many primes satisfy $c(p)>1/2-\varepsilon$, so $\limsup c(p)\geq 1/2-\varepsilon$. Letting $\varepsilon\downarrow 0$ yields the claim. ∎
-Corollary 4 (Small-density moduli).
-For every fixed $\tau\in(0,1/2)$, both sets $\{p:c(p)\leq\tau\}$ and $\{p:c(p)>\tau\}$ have positive relative density among the primes, equal to $G(\tau)$ and $1-G(\tau)$, respectively.
-Proof.
-By Theorem 7 with $\alpha=\tau$, the limiting density of $\{p:c(p)\leq\tau\}$ is $G(\tau)$.
-Since $G$ is strictly increasing on $0,\tfrac{1}{2}]$ by Theorem 6, we have $0<G(\tau)<1$, giving positive density for both sets.
-∎
 Corollary 4 (Fixed-dimension analogue).
@@
 Proof.
-Write
-$c_{n}(p)=\frac{\varphi(p-1)}{p-1}\,P_{n}(p),\qquad P_{n}(p):=\prod_{j=1}^{n}\Bigl(1-\frac{1}{p^{j}}\Bigr).$
-Since $P_{n}(p)\to 1$ as $p\to\infty$ through the primes, Lemma 5 shows that $c_{n}(p)$ has the same limiting law as $\varphi(p-1)/(p-1)$.
-Since $c_{n}(2)>0$, it suffices for large $x$ to consider odd primes $p\leq x$. Let
-$m(x):=\max\{k\geq 1:N_{k}\leq x\}.$
-For each odd prime $p\leq x$, we have
-$N_{\omega(p-1)}\leq\prod_{r\mid(p-1)}r\leq p-1\leq x,$
-so $\omega(p-1)\leq m(x)$. Let
-$r_{1}<r_{2}<\cdots<r_{d}$
-be the distinct prime divisors of p1p-1, where $d=\omega(p-1)$. Then $d\leq m(x)$, and $r_{i}\geq p_{i}$ for $1\leq i\leq d$. Hence
-$\frac{\varphi(p-1)}{p-1}=\prod_{i=1}^{d}\Bigl(1-\frac{1}{r_{i}}\Bigr)\geq\prod_{i=1}^{d}\Bigl(1-\frac{1}{p_{i}}\Bigr)\geq\prod_{i=1}^{m(x)}\Bigl(1-\frac{1}{p_{i}}\Bigr),$
-because each factor $1-\frac{1}{p_{i}}$ lies in $(0,1)$. Also, by the same elementary estimate used in Lemma 1,
-$P_{n}(p)\geq 1-\sum_{j=1}^{n}\frac{1}{p^{j}}\geq 1-\frac{1}{p-1}\geq\frac{1}{2}.$
-Therefore
-$c_{n}(p)\geq\frac{1}{2}\prod_{i=1}^{m(x)}\Bigl(1-\frac{1}{p_{i}}\Bigr).$
-Now $N_{m(x)}\leq x<N_{m(x)+1}=N_{m(x)}p_{m(x)+1}$, so $\log x\sim\log N_{m(x)}$. Since
-$\log N_{m(x)}=\sum_{i=1}^{m(x)}\log p_{i}\sim p_{m(x)},$
-we get $p_{m(x)}\sim\log x$. By Mertens’ theorem,
-$\prod_{i=1}^{m(x)}\Bigl(1-\frac{1}{p_{i}}\Bigr)\sim\frac{e^{-\gamma}}{\log p_{m(x)}}\asymp\frac{1}{\log\log x},$
-and therefore
-$\min_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}c_{n}(p)\gg\frac{1}{\log\log x}.$
-For the upper bound, the argument of Proposition 1 applies verbatim, with the additional observation that $P_{n}(p)\leq 1$ for every prime $p$.
-If $p^{*}(k)\leq x$ is the least prime congruent to $1\pmod{N_{k}}$ arising there, then
-$c_{n}\bigl(p^{*}(k)\bigr)=\frac{\varphi(p^{*}(k)-1)}{p^{*}(k)-1}\,P_{n}\bigl(p^{*}(k)\bigr)\leq\frac{\varphi(p^{*}(k)-1)}{p^{*}(k)-1}\leq\prod_{i=1}^{k}\Bigl(1-\frac{1}{p_{i}}\Bigr)\ll\frac{1}{\log\log x}.$
-Hence
-$\min_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}c_{n}(p)\ll\frac{1}{\log\log x},$
-which proves the claim. ∎
-5.3 Hausdorff dimension, moments, and Fourier structure of $G$
-We derive a sharper geometric description of the limiting distribution and then its moment and Fourier structure.
+Write $c_{n}(p)=\frac{\varphi(p-1)}{p-1}P_{n}(p)$ with $P_{n}(p):=\prod_{j=1}^{n}(1-p^{-j})$.
+Since $P_{n}(p)\to 1$, Lemma 5 gives the distributional claim.
+For the lower bound on $\min_{p\leq x}c_{n}(p)$: every odd prime $p\leq x$ satisfies $\omega(p-1)\leq m(x)$, so
+$\frac{\varphi(p-1)}{p-1}\;\geq\;\prod_{i=1}^{m(x)}\Bigl(1-\frac{1}{p_{i}}\Bigr),$
+and $P_{n}(p)\geq 1/2$ by Lemma 1. By Lemma 2 and Mertens’ third theorem,
+$\prod_{i=1}^{m(x)}\Bigl(1-\frac{1}{p_{i}}\Bigr)\;\asymp\;\frac{1}{\log\log x}.$
+The upper bound is identical to Proposition 1, since $P_{n}(p)\leq 1$. ∎
+6 Hausdorff Dimension, Moments, and Fourier Structure of $G$
+The purely singular limiting law established in Section 5 raises a natural question: how thin is the set carrying full mass of the limiting measure?
+Because $\mu_{f}$ decomposes as an infinite convolution of discrete measures (one for each odd prime), one can exploit the resulting independence structure to build efficient coverings and bound the Hausdorff dimension.
+We show that $\dim_{H}(\mu_{f})=0$, meaning that the mass concentrates on a Borel set of Hausdorff dimension zero even though the topological support is the full interval $0,\tfrac{1}{2}]$; in particular, the fractal-like clustering visible in Figure 2 is not an artifact of finite sampling.
+We then derive an explicit Euler product for every moment of $G$, prove that these moments determine the distribution uniquely, and establish the analytic continuation of the associated Mellin transform.
+The section concludes with a precise asymptotic for the behavior of $G$ near the endpoint $1/2$.
 Throughout, $f$ is the strongly additive function defined in (2), $\mu_{f}$ is the shifted-prime limiting measure of f(p1)f(p-1), and $X$ is a random variable with distribution $G$.
@@
 Proof.
-By Tenenbaum [3, formula (3), p. 262] for the shift a=1a=-1, the characteristic function of $\mu_{f}$ is
+By Tenenbaum [38, formula (3), p. 262] for the shift a=1a=-1, the characteristic function of $\mu_{f}$ is
 $\widehat{\mu_{f}}(\tau)=\prod_{\ell\text{ prime}}\Bigl(1-\frac{1}{\ell-1}+\sum_{\nu\geq 1}\frac{e^{i\tau f(\ell^{\nu})}}{\ell^{\nu}}\Bigr).$
@@
 ∎
-For a Borel probability measure $\mu$ on $\mathbb{R}$, recall the Hausdorff dimension of $\mu$:
+For a Borel probability measure $\mu$ on $\mathbb{R}$, recall the Hausdorff dimension of $\mu$ (see, e.g., Falconer [17, Chapter 2]):
 $\dim_{H}(\mu):=\inf\{\dim_{H}(E):E\subset\mathbb{R}\text{ is Borel and }\mu(E)=1\}.$
@@
 $I(y):=\sum_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\text{ prime}\end{subarray}}\left(B_{\ell}\log(\ell-1)+(1-B_{\ell})\log\frac{\ell-1}{\ell-2}\right).$
-If $\omega=(b_{\ell})_{3\leq\ell\leq y}$ is a prefix configuration, then
-$\mathbb{P}\bigl((B_{\ell})_{3\leq\ell\leq y}=\omega\bigr)=e^{-I_{y}(\omega)},$
-where $I_{y}(\omega)$ denotes the value of $I(y)$ at that configuration. Moreover,
+If $\boldsymbol{b}=(b_{\ell})_{3\leq\ell\leq y}$ is a prefix configuration, then
+$\mathbb{P}\bigl((B_{\ell})_{3\leq\ell\leq y}=\boldsymbol{b}\bigr)=e^{-I_{y}(\boldsymbol{b})},$
+where $I_{y}(\boldsymbol{b})$ denotes the value of $I(y)$ at that configuration. Moreover,
 $\mathbb{E}I(y)\ll\sum_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\text{ prime}\end{subarray}}\frac{\log\ell}{\ell}\ll\log y$
@@
 $\mathbb{E}I(y)\leq C_{0}\log y,\qquad\operatorname{Var}(I(y))\leq C_{0}(\log y)^{2}\qquad(y\geq 3).$
-Choose $C:=4C_{0}$ (so that $Cn-a_{n}s/2\leq-Cn$ in the covering estimate below). Next choose an increasing sequence $y_{n}\to\infty$ so rapidly that, with
-$a_{n}:=\left\lceil\frac{4Cn}{s}\right\rceil\qquad(\text{recall }s>0\text{ is fixed}),\qquad z_{n}:=y_{n}^{a_{n}},\qquad r_{n}:=z_{n}^{-1/2},$
-the intervals $(y_{n},z_{n}]$ are pairwise disjoint, the sequence $(z_{n})$ is strictly increasing, and
-$\sum_{n\geq 1}r_{n}<\infty,\qquad\sum_{n\geq 1}y_{n}^{-Cn}<\infty.$
+Choose $C:=4C_{0}$ (so that $Cn-a_{n}s/2\leq-Cn$ in the covering estimate below).
+$a_{n}:=\left\lceil\frac{4Cn}{s}\right\rceil\qquad(\text{recall }s>0\text{ is fixed}),\qquad z_{n}:=y_{n}^{a_{n}},\qquad r_{n}:=z_{n}^{-1/2}.$
+We construct the sequence $(y_{n})$ recursively.
+Set $y_{1}:=3$.
+Given $y_{n}$, define $a_{n}$, $z_{n}$, $r_{n}$ as above and put $y_{n+1}:=z_{n}+1$.
+Then the intervals $(y_{n},z_{n}]$ are pairwise disjoint by construction, $z_{n}=y_{n}^{a_{n}}$ grows super-exponentially, and the summability conditions $\sum r_{n}<\infty$ and $\sum y_{n}^{-Cn}<\infty$ are immediate since $r_{n}=z_{n}^{-1/2}$ and $y_{n}^{-Cn}\leq y_{n}^{-C}$ with $y_{n}\to\infty$ super-exponentially.
 For each $n$, set $I_{n}:=I(y_{n})$ and
@@
 $\mathbb{P}(A_{n}^{c})\leq\frac{\operatorname{Var}(I_{n})}{(Cn\log y_{n}-\mathbb{E}I_{n})^{2}}\ll\frac{1}{n^{2}}.$
-Let $T_{n}$ be the set of prefix configurations $\omega=(b_{\ell})_{3\leq\ell\leq y_{n}}$ satisfying $I_{n}(\omega)\leq Cn\log y_{n}$. Then
+Let $T_{n}$ be the set of prefix configurations $\boldsymbol{b}=(b_{\ell})_{3\leq\ell\leq y_{n}}$ satisfying $I_{n}(\boldsymbol{b})\leq Cn\log y_{n}$. Then
 $\mathbb{P}\bigl((B_{\ell})_{3\leq\ell\leq y_{n}}\in T_{n}\bigr)=1-O(n^{-2}),$
@@
 $1-\frac{1}{\ell-1}=\left(1-\frac{1}{\ell}\right)\left(1-\frac{1}{(\ell-1)^{2}}\right),$
-Mertens’ theorem and the convergence of $\prod_{\ell\geq 3}(1-(\ell-1)^{-2})$ imply
+Mertens’ third theorem and the convergence of $\prod_{\ell\geq 3}(1-(\ell-1)^{-2})$ imply
 $\mathbb{P}(E_{n})\asymp\frac{\log y_{n}}{\log z_{n}}=\frac{1}{a_{n}}\asymp\frac{1}{n}.$
-Therefore $\sum_{n}\mathbb{P}(E_{n})=\infty$, and the second Borel–Cantelli lemma gives
+Therefore $\sum_{n}\mathbb{P}(E_{n})=\infty$.
+Since the random variables $B_{\ell}$ are mutually independent (Lemma 10) and the blocks $(y_{n},z_{n}]$ are pairwise disjoint, the events $E_{n}$ are mutually independent.
+The second Borel–Cantelli lemma therefore gives
 $\mathbb{P}(E_{n}\text{ i.o.})=1.$
@@
 $\mathbb{P}(F_{n}^{c}\text{ i.o.})=0.$
-For each $\omega=(b_{\ell})_{3\leq\ell\leq y_{n}}\in T_{n}$, define
-$c(\omega):=\log 2+\sum_{\begin{subarray}{c}3\leq\ell\leq y_{n}\\ \ell\text{ prime}\end{subarray}}b_{\ell}w_{\ell},$
+For each $\boldsymbol{b}=(b_{\ell})_{3\leq\ell\leq y_{n}}\in T_{n}$, define
+$v(\boldsymbol{b}):=\log 2+\sum_{\begin{subarray}{c}3\leq\ell\leq y_{n}\\ \ell\text{ prime}\end{subarray}}b_{\ell}w_{\ell},$
 and let
-$U_{n}:=\bigcup_{\omega\in T_{n}}[c(\omega),\,c(\omega)+r_{n}].$
+$U_{n}:=\bigcup_{\boldsymbol{b}\in T_{n}}[v(\boldsymbol{b}),\,v(\boldsymbol{b})+r_{n}].$
 If $A_{n}\cap E_{n}\cap F_{n}$ occurs, then the block $(y_{n},z_{n}]$ contributes nothing and the tail beyond $z_{n}$ contributes at most $r_{n}$, so $Y\in U_{n}$. Since
@@
 Finally, $\mu_{G}$ is the push-forward of $\mu_{f}$ under the map $T(u):=e^{-u}$.
-Since $T$ is Lipschitz on $\log 2,\infty)$, the image of any set of Hausdorff dimension 0 again has Hausdorff dimension 0.
+Since $T$ is Lipschitz on $\log 2,\infty)$, the image of any set of Hausdorff dimension 0 again has Hausdorff dimension 0 (see, e.g., [17, Corollary 2.4]).
 Therefore $\dim_{H}(\mu_{G})=0$.
 ∎
+The vanishing Hausdorff dimension of the limiting measure explains the clustering visible in Figure 2: the purely singular law forces the values of $c(p)$ to concentrate near arithmetic anchor points of the form $\frac{1}{2}\prod_{r\in S}(1-1/r)$, determined by the small prime divisors of p1p-1, rather than spreading uniformly over $0,\frac{1}{2}]$.
+Figure 2: Empirical bin masses of $c(p)$ over the 78,49778{,}497 primes $3\leq p\leq 10^{6}$, computed via Proposition 8 with truncation depth $J=20$, using $100$ equal-width bins covering $[0,0.5]$.
+The dashed bars mark selected anchor values $\varphi(m)/m$ for small even squarefree $m$.
+The concentration of mass near these arithmetic anchors reflects the dependence of $c(p)$ on the small prime divisors of p1p-1 and illustrates the clustering predicted by the purely singular limiting law of Corollary 2.
 Theorem 6 (Moment Euler product).
@@
 Proof.
-(i) Set $k=1$ in (5): $\frac{\ell-2}{\ell-1}+\frac{1}{\ell}=\frac{\ell(\ell-2)+(\ell-1)}{\ell(\ell-1)}=\frac{\ell^{2}-\ell-1}{\ell(\ell-1)}$.
-(ii) Set $k=2$: $\frac{\ell-2}{\ell-1}+\frac{(\ell-1)}{\ell^{2}}=\frac{(\ell-2)\ell^{2}+(\ell-1)^{2}}{\ell^{2}(\ell-1)}=\frac{\ell^{3}-\ell^{2}-2\ell+1}{\ell^{2}(\ell-1)}$.
-(Verification at $\ell=3$: $\frac{13}{18}=\frac{1}{2}+\frac{2}{9}$.
-) ∎
+Immediate from (5) with $k=1$ and $k=2$. ∎
 Theorem 7 (Moment determinacy).
@@
 Since $\sum\ell^{-2}<\infty$, Weierstrass’s theorem on products of analytic functions (see e.g.
-[18, Ch.
+[8, Ch.
 VII]) gives that the product converges uniformly on compact subsets to an analytic function.
@@
 For every $\tau\in\mathbb{R}$,
-$\bigl|\widehat{\mu_{f}}(\tau)\bigr|^{2}\;=\;\prod_{\begin{subarray}{c}\ell\geq 3\\ \ell\textup{ prime}\end{subarray}}\!\left(1-\frac{2(\ell-2)}{(\ell-1)^{2}}\bigl(1-\cos(\tau\log\tfrac{\ell}{\ell-1})\bigr)\right).$ (6)
+$\bigl|\widehat{\mu_{f}}(\tau)\bigr|^{2}\;=\;\prod_{\begin{subarray}{c}\ell\geq 3\\ \ell\textup{ prime}\end{subarray}}\!\left(1-\frac{2(\ell-2)}{(\ell-1)^{2}}\bigl(1-\cos(\tau\log\tfrac{\ell}{\ell-1})\bigr)\right).$
 Proof.
@@
 ∎
+Figure 3: Squared modulus $|\widehat{\mu_{f}}(\tau)|^{2}$ computed via the Euler product of Proposition 5 using all primes $\ell\leq 10^{7}$.
+The top panel uses a linear scale and the bottom panel a log-log scale.
+The function oscillates but its local averages decrease across successive decades of $\tau$.
+Remark 2 (Numerical evidence on Fourier decay).
+We evaluated $|\widehat{\mu_{f}}(\tau)|^{2}$ via the Euler product of Proposition 5 with all primes $\ell\leq 10^{7}$.
+Figure 3 displays the result for $\tau\in[1,10^{6}]$.
+The function oscillates wildly, with local maxima near $0.14$ persisting at moderately large $\tau$.
+However, the mean value within each decade decreases steadily: approximately $0.06$ for $\tau\in[10,10^{2}]$, $0.03$ for $10^{2},10^{3}]$, $0.02$ for $10^{3},10^{4}]$, $0.01$ for $10^{4},10^{5}]$, and $0.008$ for $10^{5},10^{6}]$.
+A power-law fit to the envelope of local maxima yields $|\widehat{\mu_{f}}(\tau)|^{2}\approx\tau^{-0.19}$.
+This is consistent with decay to 0, which would mean $\mu_{f}$ is a Rajchman measure; establishing this rigorously remains open (Section 8).
 6.1 Endpoint asymptotics at $1/2$
+The value $1/2$ is the supremum of the support of $G$ but is never attained (Remark 1).
+The rate at which $G$ approaches its endpoint governs, in particular, the density of primes $p$ for which the rejection-sampling overhead $1/c(p)$ is close to its minimum value of $2$.
+We now determine this rate precisely.
 Proposition 7 (Endpoint asymptotic at $1/2$).
 Let $G$ be the limiting distribution function from Theorem 4, and let $X$ be a random variable with distribution $G$. Define
-$\kappa:=2e^{-\gamma}\prod_{\begin{subarray}{c}\ell\geq 3\\ \ell\textup{ prime}\end{subarray}}\Bigl(1-\frac{1}{(\ell-1)^{2}}\Bigr).$
+$\kappa:=2e^{-\gamma}\prod_{\begin{subarray}{c}\ell\geq 3\\ \ell\textup{ prime}\end{subarray}}\Bigl(1-\frac{1}{(\ell-1)^{2}}\Bigr)\approx 0.7413.$
 Then, as $\varepsilon\downarrow 0$,
@@
 so $\log(1/(-\log(1-2\varepsilon)))\sim\log(1/\varepsilon)$. The stated asymptotic follows. ∎
+Remark 3 (Twin prime constant).
+The constant $\kappa$ equals $C_{2}\,e^{-\gamma}$, where $C_{2}=2\prod_{p\geq 3}(1-(p-1)^{-2})\approx 1.3203$ is the twin prime constant of Hardy and Littlewood [22] (denote $2C_{2}$ therein).
+This follows from the factorisation $(\ell-2)/(\ell-1)=(1-1/\ell)(1-(\ell-1)^{-2})$: the first factor produces the Mertens asymptotic $2e^{-\gamma}/\!\log y$, while the product of the second converges to $C_{2}/2$.
 Corollary 6 (Endpoint modulus at $1/2$).
@@
 $\varepsilon^{\alpha}\log(1/\varepsilon)\to 0\qquad(\varepsilon\downarrow 0).$
-6 Explicit Lower Bounds for Primes of Cryptographic Interest
-We now give explicit lower bounds on $c(q)$ for primes $q$ whose predecessor q1q-1 has few distinct prime factors—a condition satisfied by many primes commonly used in lattice-based cryptography, especially Number Theoretic Transform (NTT)-friendly ones.
-Theorem 11 (Explicit lower bound on $c(q)$ via the number of distinct prime factors).
-Let $q\geq 3$ be a prime, and let $\omega=\omega(q-1)$ denote the number of distinct prime factors of q1q-1. Then
-$c(q)\;\geq\;P_{3}\prod_{i=1}^{\omega}\Bigl(1-\frac{1}{p_{i}}\Bigr),\qquad P_{3}:=\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{3^{j}}\Bigr)\approx 0.5601260779,$
+7 Explicit Lower Bounds via the Number of Distinct Prime Factors
+Theorem 9 (Explicit lower bound on $c(p)$ via the number of distinct prime factors).
+Let $p\geq 3$ be a prime, and let $\omega=\omega(p-1)$ denote the number of distinct prime factors of p1p-1. Then
+$c(p)\;\geq\;P_{3}\prod_{i=1}^{\omega}\Bigl(1-\frac{1}{p_{i}}\Bigr),\qquad P_{3}:=\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{3^{j}}\Bigr)\approx 0.5601260779,$
 where $p_{1}<p_{2}<\cdots<p_{\omega}$ are the first $\omega$ primes.
 Proof.
-Write the distinct prime factors of q1q-1 as $r_{1}<r_{2}<\cdots<r_{\omega}$. Then
-$\frac{\varphi(q-1)}{q-1}=\prod_{i=1}^{\omega}\Bigl(1-\frac{1}{r_{i}}\Bigr)\geq\prod_{i=1}^{\omega}\Bigl(1-\frac{1}{p_{i}}\Bigr),$
-since the product is minimized by choosing the smallest possible distinct primes.
-Now define
-$P(t):=\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{t^{j}}\Bigr)\qquad(t>1).$
-Each factor $1-t^{-j}$ is increasing in $t$, so $P(t)$ is increasing on $(1,\infty)$. Since $q\geq 3$ is prime, we have $P(q)\geq P(3)=P_{3}$. Therefore
-$c(q)=\frac{\varphi(q-1)}{q-1}\,P(q)\geq P_{3}\prod_{i=1}^{\omega}\Bigl(1-\frac{1}{p_{i}}\Bigr),$
+Write the distinct prime factors of p1p-1 as $r_{1}<r_{2}<\cdots<r_{\omega}$. Then
+$\frac{\varphi(p-1)}{p-1}=\prod_{i=1}^{\omega}\Bigl(1-\frac{1}{r_{i}}\Bigr)\geq\prod_{i=1}^{\omega}\Bigl(1-\frac{1}{p_{i}}\Bigr).$
+Since $P(t):=\prod_{j=1}^{\infty}(1-t^{-j})$ is increasing in $t>1$, we have $P(p)\geq P(3)=P_{3}$. Therefore, by (1),
+$c(p)=\frac{\varphi(p-1)}{p-1}\,P(p)\geq P_{3}\prod_{i=1}^{\omega}\Bigl(1-\frac{1}{p_{i}}\Bigr),$
 as claimed. ∎
@@
 Proof.
-The lower bound follows by the same primorial/Mertens argument used in the proof of Corollary 5: for every odd prime $p\leq x$, $\omega(p-1)\leq m(x)$, and Lemma 1 together with Mertens’ theorem gives $\min_{p\leq x,\,p\textup{ prime}}c(p)\gg 1/\!\log\log x$.
+The lower bound follows by the same primorial/Mertens argument used in the proof of Corollary 4: for every odd prime $p\leq x$, $\omega(p-1)\leq m(x)$, and Lemma 1 together with Mertens’ third theorem gives $\min_{p\leq x,\,p\textup{ prime}}c(p)\gg 1/\!\log\log x$.
 Theorem 9 sharpens the constant from $\tfrac{1}{2}$ to $P_{3}$ but does not affect the asymptotic order.
@@
 $m(n):=\max\{k\geq 1:N_{k}\leq n\}.$
-Then for every prime $q\geq 3$,
-$c(q)\;\geq\;P_{3}\prod_{i=1}^{m(q-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr).$
-In particular, there exists an absolute constant $c_{0}>0$ such that, for all sufficiently large primes $q$,
-$c(q)\geq\frac{c_{0}}{\log\log q}.$
+Then for every prime $p\geq 3$,
+$c(p)\;\geq\;P_{3}\prod_{i=1}^{m(p-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr).$
+In particular, there exists an absolute constant $c_{0}>0$ such that, for all sufficiently large primes $p$,
+$c(p)\geq\frac{c_{0}}{\log\log p}.$
 Proof.
 By Theorem 9,
-$c(q)\geq P_{3}\prod_{i=1}^{\omega(q-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr).$
+$c(p)\geq P_{3}\prod_{i=1}^{\omega(p-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr).$
 Also,
-$N_{\omega(q-1)}\leq\prod_{r\mid(q-1)}r\leq q-1,$
-so $\omega(q-1)\leq m(q-1)$. Since each factor $1-\frac{1}{p_{i}}$ lies in $(0,1)$,
-$\prod_{i=1}^{\omega(q-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr)\geq\prod_{i=1}^{m(q-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr),$
+$N_{\omega(p-1)}\leq\prod_{r\mid(p-1)}r\leq p-1,$
+so $\omega(p-1)\leq m(p-1)$. Since each factor $1-\frac{1}{p_{i}}$ lies in $(0,1)$,
+$\prod_{i=1}^{\omega(p-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr)\geq\prod_{i=1}^{m(p-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr),$
 which proves the first claim.
-Finally, by Mertens’ theorem and Lemma 2,
-$\prod_{i=1}^{m(q-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr)\asymp\frac{1}{\log p_{m(q-1)}}\asymp\frac{1}{\log\log q},$
-as $q\to\infty$ through the primes. This gives the second claim; equivalently, $1/c(q)\ll\log\log q$. ∎
+Finally, by Mertens’ third theorem and Lemma 2,
+$\prod_{i=1}^{m(p-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr)\asymp\frac{1}{\log p_{m(p-1)}}\asymp\frac{1}{\log\log p},$
+as $p\to\infty$ through the primes. This gives the second claim; equivalently, $1/c(p)\ll\log\log p$. ∎
 Remark 4 (On effectivity).
-The asymptotic $\min_{p\leq x}c(p)\asymp 1/\!\log\log x$ can be made fully explicit by combining explicit Mertens product and $\vartheta(x)$ estimates (e.g., Rosser–Schoenfeld [19], Dusart [20]) with an explicit Linnik-type bound (see Xylouris [21]).
-We do not carry out the explicit computation here.
-Remark 2.
-The bound $1/2$ for the infinite product can be sharpened for specific primes: $\prod_{j\geq 1}(1-q^{-j})\geq 1-1/(q-1)$, which exceeds $2/3$ for $q\geq 5$ and is at least $0.99$ for $q\geq 101$.
-For large NTT primes with $q>2^{30}$, this factor differs from 11 by less than $10^{-9}$.
-Even for the smaller NIST moduli $q=3329$ and $q=8380417$ [22, 23], the elementary bound gives $\prod_{j\geq 1}(1-q^{-j})>0.9996$ and $\prod_{j\geq 1}(1-q^{-j})>0.9999998$, respectively.
-We now instantiate the general $\omega(q-1)$-based bound in several parameter regimes relevant to cryptographic applications, including many NTT-friendly moduli.
-Define $B(\omega):=P_{3}\prod_{i=1}^{\omega}(1-1/p_{i})$.
-By Theorem 11, every prime $q\geq 3$ satisfies $c(q)\geq B(\omega(q-1))$.
-Direct evaluation gives the following values:
-$\omega(q-1)\leq K$ $B(K)$
+The asymptotic $\min_{p\leq x}c(p)\asymp 1/\!\log\log x$ can be made fully effective by combining explicit Mertens estimates (Rosser–Schoenfeld [34, Theorems 4 and 7], Dusart [12, Theorem 5.9]) with the Linnik-type bound $L\leq 5.18$ of Xylouris [41, Theorem 1.1].
+Remark 5 (Sharper bounds for specific primes).
+Direct computation of the Euler product gives $P(5)\approx 0.7603$, $P(101)\approx 0.9900$, and $P(p)>1-10^{-9}$ for $p>2^{30}$.
+For the NIST moduli $p=3329$ and $p=8380417$ [31, 32], $\prod_{j\geq 1}(1-p^{-j})>0.9996$ and >0.9999998>0.9999998, respectively.
+Define $B(\omega):=P_{3}\prod_{i=1}^{\omega}(1-1/p_{i})$; by Theorem 9, $c(p)\geq B(\omega(p-1))$ for every prime $p\geq 3$. Table 1 lists representative values of $B(K)$ for small values of $K$.
+$\omega(p-1)\leq K$ $B(K)$
 11 $P_{3}/2\approx 0.280063$
@@
 2020 $\approx 0.0715828$
-Remark 3 (On the case $\omega(q-1)=1$).
-If $\omega(q-1)=1$ and $q$ is an odd prime, then q1q-1 is an even prime power, hence $q-1=2^{m}$ for some $m\geq 1$. Equivalently, $q$ is a Fermat prime. In this case
-$\frac{\varphi(q-1)}{q-1}=\frac{1}{2},\qquad c(q)=\frac{1}{2}\,P(q),\qquad P(q):=\prod_{j\geq 1}\Bigl(1-\frac{1}{q^{j}}\Bigr).$
-Hence $c(q)$ is close to $1/2$ for large Fermat primes; so if infinitely many Fermat primes exist, then $c(q)\to 1/2$ along that subsequence.
-Thus the universal bound $B(1)=P_{3}/2\approx 0.280063$ is still not tight in this special case.
-Corollary 10 (Practical bounds for cryptographic primes).
-Let $q$ be a prime with $q>2^{30}$. Then
-$c(q)>\frac{0.56}{\log q}.$
-Moreover, if $\omega(q-1)\leq 10$, then
-$c(q)>0.088.$
-In particular, this applies to cryptographic-size primes $q\in[2^{64},2^{128}]$ satisfying $\omega(q-1)\leq 10$.
+Table 1: Representative values of the explicit lower bound $B(K)=P_{3}\prod_{i=1}^{K}(1-1/p_{i})$ appearing in Theorem 9. Thus every prime $p\geq 3$ with $\omega(p-1)\leq K$ satisfies $c(p)\geq B(K)$.
+Remark 6 (On the case $\omega(p-1)=1$).
+If $\omega(p-1)=1$ and $p$ is an odd prime, then p1p-1 is an even prime power, hence $p-1=2^{m}$ for some $m\geq 1$. Equivalently, $p$ is a Fermat prime. In this case
+$\frac{\varphi(p-1)}{p-1}=\frac{1}{2},\qquad c(p)=\frac{1}{2}\,P(p),\qquad P(p):=\prod_{j\geq 1}\Bigl(1-\frac{1}{p^{j}}\Bigr).$
+Hence $c(p)\to 1/2$ along any infinite subsequence of Fermat primes, so $B(1)$ is not tight for Fermat primes.
+Corollary 9 (Explicit lower bounds for large primes).
+Let $p$ be a prime with $p>2^{30}$. Then
+$c(p)>\frac{0.56}{\log p}.$
+Moreover, if $\omega(p-1)\leq 10$, then
+$c(p)>0.088.$
 Proof.
-The claim under $\omega(q-1)\leq 10$ follows from the bound $c(q)\geq B(\omega(q-1))$ (Theorem 11), since
+The claim under $\omega(p-1)\leq 10$ follows from the bound $c(p)\geq B(\omega(p-1))$ (Theorem 9), since
 $B(10)=P_{3}\prod_{i=1}^{10}\Bigl(1-\frac{1}{p_{i}}\Bigr)\approx 0.08847>0.088.$
-For the general bound, since every prime divisor of q1q-1 is at most q1q-1,
-$\frac{\varphi(q-1)}{q-1}=\prod_{r\mid(q-1)}\Bigl(1-\frac{1}{r}\Bigr)\;\geq\;\prod_{\begin{subarray}{c}r\leq q-1\\ r\text{ prime}\end{subarray}}\Bigl(1-\frac{1}{r}\Bigr),$
-where the inequality holds because the right-hand side includes additional factors in $(0,1)$.
-By Dusart’s explicit product estimate [20, Theorem 5.9], which we use conservatively with threshold $x\geq 2\,278\,383$,
+For the general bound, since every prime divisor of p1p-1 is at most p1p-1,
+$\frac{\varphi(p-1)}{p-1}=\prod_{r\mid(p-1)}\Bigl(1-\frac{1}{r}\Bigr)\;\geq\;\prod_{\begin{subarray}{c}r\leq p-1\\ r\text{ prime}\end{subarray}}\Bigl(1-\frac{1}{r}\Bigr).$
+By Dusart’s explicit product estimate [12, Theorem 5.9], which we use with threshold $x\geq 2\,278\,382$ (noting that $2\,278\,382<2^{22}\ll 2^{30}$, so the threshold is satisfied for all primes $p>2^{30}$),
 $\prod_{\begin{subarray}{c}r\leq x\\ r\text{ prime}\end{subarray}}\Bigl(1-\frac{1}{r}\Bigr)\;=\;\frac{e^{-\gamma}}{\log x}\,\Bigl(1+O^{*}\!\Bigl(\frac{0.2}{\log^{3}x}\Bigr)\Bigr).$
-Since $q>2^{30}$, we have q1>2 278 383q-1>2\,278\,383 and $\log(q-1)>20$, so the correction factor lies in $(1-2.5\times 10^{-5},\;1+2.5\times 10^{-5})$.
+Since $p>2^{30}$, we have p1>2 278 382p-1>2\,278\,382 and $\log(p-1)>20$, so the correction factor lies in $(1-2.5\times 10^{-5},\;1+2.5\times 10^{-5})$.
 Also,
-$\prod_{j\geq 1}\Bigl(1-\frac{1}{q^{j}}\Bigr)\geq 1-\sum_{j\geq 1}\frac{1}{q^{j}}=1-\frac{1}{q-1}>0.9999.$
+$\prod_{j\geq 1}\Bigl(1-\frac{1}{p^{j}}\Bigr)\geq 1-\sum_{j\geq 1}\frac{1}{p^{j}}=1-\frac{1}{p-1}>0.9999.$
 Therefore
-$c(q)=\frac{\varphi(q-1)}{q-1}\prod_{j\geq 1}\Bigl(1-\frac{1}{q^{j}}\Bigr)\;>\;\Bigl(1-\frac{0.2}{\log^{3}(q-1)}\Bigr)\cdot 0.9999\cdot\frac{e^{-\gamma}}{\log(q-1)}\;>\;\frac{0.56}{\log q},$
-using $e^{-\gamma}>0.5614$, $\log(q-1)<\log q$, and $0.2/\log^{3}(q-1)<2.5\times 10^{-5}$ for $q>2^{30}$. ∎
-The bound $c(q)>0.56/\!\log q$ of Corollary 10 is a simple, fully explicit estimate obtained by taking the Mertens product over all primes up to q1q-1.
-The sharper universal pointwise bound $c(q)\gg 1/\!\log\log q$ of Corollary 9 shows that $1/\!\log\log q$, not $1/\!\log q$, is the correct worst-case scale for the rejection-sampling overhead.
-By the shifted-prime Erdős–Kac theorem (Barban [24]; see also Halberstam [25] and Elliott [26, Ch.
-12]), $(\omega(q-1)-\log\log x)/\sqrt{\log\log x}$ converges in distribution to the standard normal as $x\to\infty$ over primes $q\leq x$.
-In particular, for every fixed $\delta>0$, almost all primes $q\leq x$ satisfy $\omega(q-1)\leq(1+\delta)\log\log x$.
-Remark 4 (Typical size of $\omega(q-1)$).
-The shifted-prime Erdős–Kac theorem explains why bounds depending only on $\omega(q-1)$ are relevant in cryptographic parameter ranges: for $q\approx 2^{128}$, one has $\log\log q\approx 4.5$, so $\omega(q-1)\leq 10$ holds for the vast majority of such primes.
-Scheme $q$ factorization of q1q-1 general lower bound for $c(q)$ exact-factorization lower bound
-ML-KEM $3329$ $2^{8}\cdot 13$ $B(2)=P_{3}/3\approx 0.1867$ $\dfrac{6}{13}\Bigl(1-\dfrac{1}{3328}\Bigr)\approx 0.4614$
-ML-DSA $8380417$ $2^{13}\cdot 3\cdot 11\cdot 31$ $B(4)=8P_{3}/35\approx 0.1280$ $\dfrac{100}{341}\Bigl(1-\dfrac{1}{8380416}\Bigr)\approx 0.2933$
-Table 1: The prime moduli appearing in the current NIST standards for ML-KEM [22] and ML-DSA [23].
-The fourth column gives the general $\omega(q-1)$-based lower bound for $c(q)$ from Theorem 11, while the fifth gives the sharper lower bound obtained from the exact factorization of q1q-1 together with $P(q)\geq 1-\frac{1}{q-1}$.
-The exact-factorization bounds imply dimension-uniform expected rejection-sampling overheads at most $2.17$ and $3.42$, respectively.
-For the two prime moduli appearing, across all parameter sets, in the current NIST standards for ML-KEM and ML-DSA [22, 23], Table 1 records both the general $\omega(q-1)$-based bounds and the sharper exact-factorization bounds.
-The exact factorizations give $\varphi(3328)/3328=6/13$ and $\varphi(8380416)/8380416=100/341$.
-Together with $P(q)\geq 1-\frac{1}{q-1}$, one obtains
-$c(3329)\geq\frac{6}{13}\Bigl(1-\frac{1}{3328}\Bigr)\approx 0.4614,\qquad c(8380417)\geq\frac{100}{341}\Bigl(1-\frac{1}{8380416}\Bigr)\approx 0.2933.$
-A short tail-bound argument (taking $J=2$ partial products of $P(q)$ and using the estimate $P_{J}(q)-P(q)\leq q^{-J}/(q-1)$) confirms that these lower bounds are tight to many decimal places:
-$c(3329)=0.4613997781\ldots,\qquad c(8380417)=0.2932550969\ldots.$
-The corresponding dimension-uniform expected rejection-sampling overheads $1/c(q)$ are approximately $2.17$ and $3.41$. For each fixed dimension $n$, the exact overheads $1/c_{n}(q)$ are no larger.
-Context $q$ factorization of q1q-1 $\omega(q-1)$ $1/c(q)$
-ML-KEM (FIPS 203) $3329$ $^{8}\cdot 13 22$ $2.17$
-ML-DSA (FIPS 204) $8380417$ $2^{13}\cdot 3\cdot 11\cdot 31$ 44 $3.41$
-NewHope / FALCON [27, 28] $12289$ $^{12}\cdot 3 22$ $3.00$
-BFV/BGV FHE (small) [29, 30, 31] $786433$ $^{18}\cdot 3 22$ $3.00$
-SEAL BFV default [32] $132120577$ $2^{21}\cdot 3^{2}\cdot 7$ 33 $3.50$
-BFV/BGV FHE (medium) [29, 30, 31] $2013265921$ $2^{27}\cdot 3\cdot 5$ 33 $3.75$
-NTTRU / BLISS [33, 34] $7681$ $2^{9}\cdot 3\cdot 5$ 33 $3.75$
-64-bit NTT prime [35] $2^{64}-2^{32}+1$ $2^{32}\cdot 3\cdot 5\cdot 17\cdot 257\cdot 65537$ 66 $4.00$
-Fermat prime ($2^{16}+1$) $65537$ $2^{16}$ 11 $2.00$
-Table 2: Representative NTT-friendly primes from lattice-based cryptography and FHE implementations, together with $\omega(q-1)$ and the dimension-uniform expected rejection-sampling overhead $1/c(q)$.
-The displayed overheads are rounded to two decimal places; the exact values differ by less than $10^{-3}$ due to the Euler product $P(q)$.
-All overheads are at most $4.00$.
-The first two rows reproduce the NIST entries from Table 1.
-Table 2 collects nine representative moduli drawn from current lattice-based and FHE implementations.
-Every modulus in the table satisfies $\omega(q-1)\leq 6$ and incurs a dimension-uniform rejection-sampling overhead $1/c(q)\leq 4.00$.
-The table substantiates the claim that the primes typically encountered in practice have controlled factorization structure, which yields a good lower bound on $c(q)$.
+$c(p)=\frac{\varphi(p-1)}{p-1}\prod_{j\geq 1}\Bigl(1-\frac{1}{p^{j}}\Bigr)\;>\;\Bigl(1-\frac{0.2}{\log^{3}(p-1)}\Bigr)\cdot 0.9999\cdot\frac{e^{-\gamma}}{\log(p-1)}\;>\;\frac{0.56}{\log p},$
+using $e^{-\gamma}>0.5614$, $\log(p-1)<\log p$, and $0.2/\log^{3}(p-1)<2.5\times 10^{-5}$ for $p>2^{30}$. Equivalently, for $p>2^{30}$,
+$\frac{1}{c(p)}\;<\;\frac{\log p}{0.56}\;<\;1.79\log p.$
+Remark 7.
+By the shifted-prime Erdős–Kac theorem (Barban [3]; see also Halberstam [20] and Elliott [13, Ch.
+12]), extending the classical result of Erdős and Kac [15], $(\omega(p-1)-\log\log x)/\sqrt{\log\log x}$ converges in distribution to the standard normal over primes $p\leq x$.
+For $p\approx 2^{128}$, $\log\log p\approx 4.5$, so $\omega(p-1)\leq 10$ holds for the vast majority of such primes.
+Standard $p$ generic $B(\omega(p-1))$ bound sharper factorization bound
+ML-KEM $3329$ $B(2)=P_{3}/3\approx 0.1867$ $\approx 0.4614$
+ML-DSA $8380417$ $B(4)=8P_{3}/35\approx 0.1280$ $\approx 0.2933$
+Table 2: The prime moduli appearing in FIPS 203 and FIPS 204 [31, 32].
+The third column gives the generic bound $B(\omega(p-1))$ from Theorem 9, depending only on the number of distinct prime factors of p1p-1, while the fourth gives the sharper lower bound obtained from the exact factorization of p1p-1 together with $P(p)\geq 1-\frac{1}{p-1}$.
+For the NIST-standardized moduli [31, 32], the exact factorizations $\varphi(3328)/3328=6/13$ and $\varphi(8380416)/8380416=100/341$ together with $P(p)\geq 1-1/(p-1)$ give $c(3329)\geq 0.4613$ and $c(8380417)\geq 0.2932$.
+The certified evaluation of Proposition 8 yields $c(3329)=0.461399\ldots$ and $c(8380417)=0.293255\ldots$, so the sharper factorization bounds agree with the exact values to $10$ and $21$ decimal places, respectively; the gap is $O(p^{-3})$, since replacing $P(p)$ by 11/(p1)1-1/(p-1) incurs an error of order $p^{-1}\cdot p^{-2}$ from the leading cross-term in the Euler product expansion.
+Context $p$ $\omega(p-1)$ $1/c(p)$
+ML-KEM (FIPS 203) $3329$ 22 $2.17$
+ML-DSA (FIPS 204) $8380417$ 44 $3.41$
+NewHope / FALCON [2, 19] $12289$ 22 $3.00$
+BFV/BGV FHE (small) [5, 18, 6] $786433$ 22 $3.00$
+BFV/BGV FHE (medium) [5, 18, 6] $2013265921$ 33 $3.75$
+NTTRU / BLISS [27, 11] $7681$ 33 $3.75$
+64-bit NTT prime [37] $2^{64}-2^{32}+1$ 66 $4.00$
+Fermat prime ($2^{16}+1$) $65537$ 11 $2.00$
+Table 3: Representative NTT-friendly primes from lattice-based cryptography and FHE implementations, together with $\omega(p-1)$ and the dimension-uniform expected rejection-sampling overhead $1/c(p)$.
+The displayed overheads are rounded to two decimal places; the exact values differ by less than $3\times 10^{-3}$.
 Remark 8 (Concrete example).
-For the 64-bit NTT prime $q=2^{64}-2^{32}+1$ (sometimes called the “Goldilocks prime” in the zero-knowledge community, not to be confused with the Ed448-Goldilocks prime $2^{448}-2^{224}-1$ [36]), used as an NTT-friendly ambient field in implementation-oriented lattice cryptography [35], one has $\omega(q-1)=6$ and $\varphi(q-1)/(q-1)\approx 0.2500$.
-Since $P(q)>1-10^{-19}$, one obtains $c(q)\approx 0.25$, so in every dimension $n$ at least a fraction $\approx 0.25$ of random $n\times n$ matrices have a primitive-root determinant, and the dimension-uniform security-reduction overhead is about $4$.
-7 Implications for PRIM-LWE
-For a fixed prime modulus $p$, [1, Theorem 10] shows that decision-prim-lwe is at least as hard as decision-LWE, up to an $O(n^{2})$ factor.
-Separately, [1, Proposition 5] isolates a positive lower bound, uniform in the matrix dimension, for the probability that a random secret matrix has primitive-root determinant.
-In our notation, that dimension-uniform sampling constant is
-$c=\inf_{n\geq 1}c_{n}(p).$
-As shown in Section 2, $c_{n}(p)$ is decreasing in $n$ and $\inf_{n\geq 1}c_{n}(p)=c(p)$, so $c(p)$ is the dimension-uniform sampling term entering the reduction.
-The present note concerns this quantity $c(p)$; it does not alter the separate $O(n^{2})$ factor from [1, Theorem 10].
-Theorem 3 shows that no uniform lower bound on $c(p)$ holds across all primes, meaning the reduction tightness degrades for primes $p$ whose predecessor p1p-1 is divisible by many small primes.
-More precisely, let $p^{*}(k)$ denote the least prime satisfying $p^{*}(k)\equiv 1\pmod{N_{k}}$, as in the proof of Proposition 1.
-Then, for the absolute constants $C_{0},L$ from Theorem 4,
-$N_{k}<p^{*}(k)\leq C_{0}N_{k}^{\,L}.$
-The left inequality holds because $p^{*}(k)=1+jN_{k}$ with $j\geq 1$, while the right inequality is exactly Linnik’s theorem. Therefore
-$\log p^{*}(k)\asymp\log N_{k},\qquad\log N_{k}\sim p_{k}$
-by Lemma 2, and hence
-$\log\log p^{*}(k)\asymp\log p_{k}.$
-On the other hand, the proof of Proposition 1 gives
-$c\bigl(p^{*}(k)\bigr)\leq\prod_{i=1}^{k}\Bigl(1-\frac{1}{p_{i}}\Bigr)\sim\frac{e^{-\gamma}}{\log p_{k}}.$
-Thus
-$c\bigl(p^{*}(k)\bigr)\ll\frac{1}{\log\log p^{*}(k)},$
-so the construction yields an infinite sequence of primes $p=p^{*}(k)$ for which
-$c(p)=O\!\left(\frac{1}{\log\log p}\right).$
-The dimension-uniform upper bound on the rejection-sampling overhead, namely $1/c(p)$, can therefore grow like
-$\Omega(\log\log p).$
-In fixed dimension $n$, the exact expected number of trials is $1/c_{n}(p)\leq 1/c(p)$.
-By [1, Theorem 10], the reduction from decision-LWE to decision-prim-lwe introduces an advantage loss of $O(n^{2})\cdot(1/c)$, where $c=\inf_{n\geq 1}c_{n}(p)=c(p)$ is the dimension-uniform sampling constant.
-Our results therefore show that the $1/c(p)$ factor in this bound can grow like $\Omega(\log\log p)$ along a suitable subsequence of primes.
-However, this is a modulus-dependent artifact rather than a structural obstruction.
-For any fixed prime $p$, the constant $c(p)>0$ and the reduction remains valid.
-Moreover, the explicit bounds of Section 6 show that the specific moduli surveyed in Table 2—which cover the NIST standards and several widely deployed FHE primes—all have $c(q)$ concretely bounded away from zero.
-Corollary 3 gives $\limsup c(p)=1/2$ unconditionally, so there is no shortage of primes for which $c(p)$ is close to its maximal value.
-As a concrete illustration, for safe primes $q=2r+1$ with $r$ an odd prime and $q\geq 7$, one has
-$c(q)=\frac{r-1}{2r}\prod_{j\geq 1}\left(1-q^{-j}\right)\geq P_{3}\,\frac{r-1}{2r}.$
-The prefactor (r1)/(2r)=1/21/(q1)(r-1)/(2r)=1/2-1/(q-1) is already close to $1/2$ when $q$ is large, and the Euler product $\prod_{j\geq 1}(1-q^{-j})$ tends to $1$ as $q\to\infty$, so $c(q)$ approaches $1/2$ for large safe primes.
-More generally, Corollary 4 shows that for every fixed $\tau\in(0,1/2)$, both sets of primes
-$\{p:c(p)\leq\tau\}\qquad\text{and}\qquad\{p:c(p)>\tau\}$
-have positive relative density, namely $G(\tau)$ and $1-G(\tau)$.
-Moreover, $G(\tau)\to 0$ as $\tau\downarrow 0$.
-Thus moduli yielding a very loose constant-factor sampling term are rare only in the threshold-asymptotic sense $\tau\downarrow 0$, not absent for any fixed threshold $\tau<1/2$.
-We note that the primes typically used in lattice-based cryptography are chosen for NTT compatibility.
-Concretely, this means $q\equiv 1\pmod{2^{t}}$ for moderate $t$.
-Mathematically, this condition controls only the 2-adic part of q1q-1.
-It does not by itself prevent q1q-1 from having many odd prime factors, and NTT-friendliness alone does not imply a good lower bound on $c(q)$.
-What our theorems actually use is the controlled factorization structure of q1q-1, quantified by $\omega(q-1)$ or by the exact factorization.
-The NTT condition is harmless with respect to the 2-part, since the large power of $2$ contributes only the single Euler factor $(1-\tfrac{1}{2})$ to $\varphi(q-1)/(q-1)$.
-Among the implementation moduli surveyed in Table 2, every prime satisfies $\omega(q-1)\leq 6$.
-It is this controlled factorization that yields the good lower bounds on $c(q)$ recorded in Remark 5, and Tables 1 and 2.
-Thus our results do not by themselves indicate a practical weakness; they quantify how modulus choice affects the reduction loss in the prim-lwe framework.
-Affine determinant programs (ADPs) [37] provide a natural external determinant-based connection to prim-lwe.
-In the ADP obfuscation framework, a Kilian-style randomization step left- and right-multiplies by invertible matrices $R,S$ with $\det(R)\det(S)=1$ to preserve the determinant polynomial.
-Because the ambient sampling space is $\mathrm{GL}_{n}(\mathbb{F}_{q})$ rather than all of $M_{n}(\mathbb{F}_{q})$, the relevant primitive-root fraction reduces to $\varphi(q-1)/(q-1)$: the matrix-count Euler product $\prod_{j=1}^{n}(1-q^{-j})$ is already absorbed into the $\mathrm{GL}_{n}$ normalization.
-Since the determinant map
-$\det:\mathrm{GL}_{n}(\mathbb{F}_{q})\longrightarrow\mathbb{F}_{q}^{\times}$
-is a surjective homomorphism, all fibers have the same cardinality, and therefore the exact fraction of invertible matrices with primitive-root determinant is
-$\frac{\varphi(q-1)}{q-1}.$
-Hence, if $\omega(q-1)\leq 10$, then
-$\frac{\varphi(q-1)}{q-1}\;\geq\;\prod_{i=1}^{10}\Bigl(1-\frac{1}{p_{i}}\Bigr)\approx 0.15795,$
-so rejection sampling inside $\mathrm{GL}_{n}(\mathbb{F}_{q})$ succeeds in fewer than $6.4$ trials on average.
-If instead one samples from all matrices $M_{n}(\mathbb{F}_{q})$, then the exact finite-$n$ primitive-root density is $c_{n}(q)$, whose limit as $n\to\infty$ is $c(q)$.
-While our results do not establish a security advantage for imposing primitive-root determinant constraints, they confirm that the associated sampling overhead is mild whenever q1q-1 has controlled factorization.
+For the 64-bit NTT prime $p=2^{64}-2^{32}+1$ (sometimes called the “Goldilocks prime”; not to be confused with the Ed448-Goldilocks prime [21]), used in lattice cryptography [37], one has $\omega(p-1)=6$, $\varphi(p-1)/(p-1)\approx 0.2500$, $P(p)>1-10^{-19}$, and $1/c(p)\approx 4$.
+Proposition 8 (Certified evaluation of $c(p)$).
+Given a prime $p\geq 3$ and the complete factorization $p-1=r_{1}^{e_{1}}\cdots r_{k}^{e_{k}}$, the following procedure computes $c(p)$ to $D$ correct decimal digits using $O(k+D\log 10/\!\log p)$ rational-arithmetic operations in the unit-cost rational-arithmetic model:
+  1. (i)
+    Compute the totient ratio $\varphi(p-1)/(p-1)=\prod_{i=1}^{k}(1-1/r_{i})$, which is exact as a rational number.
+  2. (ii)
+    Set the truncation depth $J:=\lceil D\log 10/\!\log p\rceil$ and compute the partial Euler product $P_{J}(p):=\prod_{j=1}^{J}(1-p^{-j})$, again exact in rational arithmetic.
+  3. (iii)
+    Return the certified interval
+    $\frac{\varphi(p-1)}{p-1}\!\left(P_{J}(p)-\frac{p^{-J}}{p-1}\right)\;\leq\;c(p)\;\leq\;\frac{\varphi(p-1)}{p-1}\,P_{J}(p).$
+The width of this certified interval is at most $p^{-J}/(2(p-1))$.
+Moreover, in step (ii) the exact rational $P_{J}(p)$ has denominator $p^{J(J+1)/2}$, so the intermediate numerators and denominators have $O(J^{2}\log p)$ bits.
+Proof.
+Steps (i) and (ii) use k1k-1 and J1J-1 multiplications respectively.
+The certified bound in (iii) follows from the identity $P(p)=P_{J}(p)\prod_{j>J}(1-p^{-j})$ together with $\prod_{j>J}(1-p^{-j})\geq 1-\sum_{j>J}p^{-j}=1-p^{-J}/(p-1)$ and $P(p)\leq P_{J}(p)$, giving $0\leq P_{J}(p)-P(p)\leq p^{-J}/(p-1)$.
+The interval width bound uses $\varphi(p-1)/(p-1)\leq 1/2$.
+∎
 8 Concluding Remarks
-We have resolved the open question posed in [1, Remark 3(ii)] in the affirmative: $\inf_{p\text{ prime}}c(p)=0$, and in fact
-$\liminf_{\begin{subarray}{c}p\to\infty\\ p\text{ prime}\end{subarray}}c(p)=0.$
-The proof is short and unconditional, requiring only Dirichlet’s theorem and Mertens’ product formula.
-We have further established several results that describe the finer structure of the density $c(p)$:
-  1. 1.
-    The minimum of $c(p)$ up to $x$ has the sharp order
-    $\min_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}c(p)\asymp\frac{1}{\log\log x}.$
-    In particular, the decay to zero is slow and quantified precisely.
-  2. 2.
-    The quantity $c(p)$ has a continuous limiting distribution over the primes whose support is exactly $0,\tfrac{1}{2}]$; equivalently, the distribution function is strictly increasing on $0,\tfrac{1}{2}]$.
-This full-support conclusion appears to be new.
-The law coincides with the classical shifted-prime distribution of $\varphi(p-1)/(p-1)$, is purely singular, and, by Proposition 4, the limiting measure has Hausdorff dimension 0.
-The moments of $G$ are given by convergent Euler products (Theorem 8), and the Mellin transform $\mathbb{E}[X^{s}]$ extends analytically to $\{\operatorname{Re}(s)>0\}$ (Theorem 10).
-Pure singularity was already known in the integer case from Erdős [5]; Tenenbaum [3, p.
-263] noted that it carries over to shifted primes.
-  3. 3.
-    For the implementation moduli surveyed in Table 2—including the current NIST-standardized moduli and several widely used FHE primes—the density $c(q)$ admits an explicit positive lower bound depending only on $\omega(q-1)$.
-Accordingly, the prim-lwe reduction has a finite, explicitly computable constant-factor sampling term for every fixed prime modulus, and this term is small in the concrete parameter regimes treated here.
-On the cryptographic side, the rejection-sampling overhead $1/c(p)$ can grow like $\Omega(\log\log p)$ along a suitable subsequence of primes, but as discussed in Section 7, this represents a quantitative sensitivity to modulus selection rather than a structural obstruction.
-Table 2 confirms that every implementation modulus we examined has $\omega(q-1)\leq 6$ and $1/c(q)\leq 4.00$, and the sharper universal pointwise bound of Corollary 9 gives $1/c(q)=O(\log\log q)$.
-Proposition 7 and Corollary 7 determine the right-endpoint behavior of $G$:
-$1-G\!\left(\frac{1}{2}-\varepsilon\right)\sim\frac{\kappa}{\log(1/\varepsilon)}\qquad(\varepsilon\downarrow 0),$
-so $G$ is not Hölder of any positive order at $1/2$.
-Proposition 4 shows that the limiting measure has Hausdorff dimension 0.
-Equivalently, the limiting PRIM-LWE law is a continuous, purely singular, full-support measure on $0,\tfrac{1}{2}]$ that is concentrated on a Borel set of Hausdorff dimension 0.
-A natural remaining question is therefore to determine a global modulus of continuity of $G$.
-The squared modulus of the characteristic function admits an explicit Euler product (Proposition 5), but whether
+This article studies the PRIM-LWE density constant $c(p)$ unconditionally from extremal, distributional, and computational viewpoints.
+We proved that $\inf_{p\text{ prime}}c(p)=0$ and that $\min_{p\leq x}c(p)\asymp 1/\log\log x$, so the worst-case rejection-sampling overhead grows only on a doubly logarithmic scale.
+We also showed that $c(p)$ admits a continuous, purely singular limiting distribution with support exactly $0,\tfrac{1}{2}]$ and Hausdorff dimension 0; in particular, the limiting measure is supported on a metrically negligible subset of $0,\tfrac{1}{2}]$ despite having full topological support.
+Finally, the explicit $\omega(p-1)$-based bounds and the certified evaluation algorithm of Proposition 8 make these asymptotic results effective for concrete cryptographic moduli.
+Several questions remain open. The squared modulus of the characteristic function admits an explicit Euler product (Proposition 5), but whether
 $|\widehat{\mu_{f}}(\tau)|\to 0\qquad\text{or}\qquad\limsup_{|\tau|\to\infty}|\widehat{\mu_{f}}(\tau)|>0$
 remains open; by Proposition 6, the latter would give an independent Fourier-analytic proof of pure singularity.
-This should not be confused with the classical support question for the distribution of $\varphi(n)/n$ over all integers, which remains open.
-Appendix A Positive Density Near $1/2$
-The following lemma gives an independent, Siegel–Walfisz-based proof that intervals near the endpoint $1/2$ carry positive limiting measure for $\varphi(p-1)/(p-1)$.
-Lemma 11 (Positive density of primes with totient ratio near 1/2).
-For every $\varepsilon>0$,
-$\liminf_{x\to\infty}\frac{1}{\pi(x)}\#\left\{p\leq x:p\text{ prime and }\frac{\varphi(p-1)}{p-1}>\frac{1}{2}-\varepsilon\right\}>0.$
-Proof.
-If $\varepsilon\geq 1$, then $\tfrac{1}{2}-\varepsilon\leq 0$, so the displayed set contains every prime and the claim is trivial. Hence assume throughout that $0<\varepsilon<1$.
-Step 1: Sieving out small odd prime divisors. Let $R\geq 3$ be a fixed integer, to be chosen later in terms of $\varepsilon$, and define
-$M_{R}:=2\!\!\prod_{\begin{subarray}{c}3\leq r\leq R\\ r\text{ prime}\end{subarray}}r,\qquad\nu_{R}:=\prod_{\begin{subarray}{c}3\leq r\leq R\\ r\text{ prime}\end{subarray}}(r-2),\qquad\delta(R):=\frac{\nu_{R}}{\varphi(M_{R})}=\prod_{\begin{subarray}{c}3\leq r\leq R\\ r\text{ prime}\end{subarray}}\frac{r-2}{r-1}.$
-Also set
-[…]
-[…]
-The finitely many odd primes $p\leq R$ contribute only $O_{R}(1)$ to $|\mathcal{A}_{R}(x)|$. Hence, by Theorem 5,
-$|\mathcal{A}_{R}(x)|=\frac{\nu_{R}}{\varphi(M_{R})}\,\operatorname{Li}(x)+O_{R}\!\bigl(xe^{-c_{R}\sqrt{\log x}}\bigr)=\delta(R)\,\operatorname{Li}(x)+O_{R}\!\bigl(xe^{-c_{R}\sqrt{\log x}}\bigr).$
-Applying Theorem 5 also with modulus $1$, we have
-$\pi(x)=\operatorname{Li}(x)+O\!\bigl(xe^{-c\sqrt{\log x}}\bigr),$
-and therefore
-$\frac{|\mathcal{A}_{R}(x)|}{\pi(x)}\longrightarrow\delta(R)>0.$
-In particular, $\mathcal{A}_{R}(x)$ has positive relative density $\delta(R)$ among the primes.
-Step 2: Reduction to a mean-value estimate. For $p\in\mathcal{A}_{R}(x)$, the only prime $\leq R$ dividing p1p-1 is $2$, so
-$\frac{\varphi(p-1)}{p-1}=\frac{1}{2}\prod_{\begin{subarray}{c}q\mid(p-1)\\ q>R\end{subarray}}\Bigl(1-\frac{1}{q}\Bigr).$
-Set
-$S(p):=\sum_{\begin{subarray}{c}q\mid(p-1)\\ q>R\end{subarray}}\frac{1}{q}.$
-By the elementary inequality
-$\prod_{i}(1-a_{i})\geq 1-\sum_{i}a_{i}\qquad(0\leq a_{i}\leq 1),$
-any $p\in\mathcal{A}_{R}(x)$ with $S(p)\leq\varepsilon$ satisfies
-$\frac{\varphi(p-1)}{p-1}\geq\frac{1}{2}\bigl(1-S(p)\bigr)\geq\frac{1}{2}(1-\varepsilon)>\frac{1}{2}-\varepsilon.$
-It therefore suffices to show that a positive proportion of primes in $\mathcal{A}_{R}(x)$ satisfy $S(p)\leq\varepsilon$.
-Step 3: Mean-value estimate for $S(p)$.
-Throughout Steps 3 and 4, the parameter $R$ (and hence $\varepsilon$) is fixed; all limits are as $x\to\infty$.
-Fix $Q=(\log x)^{2}$, and decompose $S(p)=S_{\leq Q}(p)+S_{>Q}(p)$, where
-$S_{\leq Q}(p)=\sum_{\begin{subarray}{c}q\mid(p-1)\\ R<q\leq Q\end{subarray}}\frac{1}{q}.$
-For the tail, each prime divisor $q>Q$ of p1p-1 contributes at most $1/Q$, and there are at most $\log p/\log Q$ such divisors. Hence, uniformly for $p\leq x$,
-$S_{>Q}(p)\leq\frac{\log p}{Q\log Q}\leq\frac{\log x}{Q\log Q}=\frac{1}{2\log x\,\log\log x}=o(1).$
-Now assume $x$ is large enough that $Q>R$. For each prime $q\in(R,Q]$, write
-$B_{q}(x):=\#\{p\in\mathcal{A}_{R}(x):p\equiv 1\pmod{q}\}$.
-Since $q>R$, we have $q\nmid M_{R}$.
-The conditions $p\in\mathcal{A}_{R}(x)$ and $p\equiv 1\pmod{q}$ therefore define exactly $\nu_{R}$ reduced residue classes modulo $M_{R}q$: modulo each odd prime $r\leq R$ there are r2r-2 admissible reduced classes, modulo $2$ there is exactly a single admissible class, and modulo $q$ the residue class is forced to be $1$.
-Because $\varphi(M_{R}q)=\varphi(M_{R})(q-1)$, we obtain from Theorem 5 that
-$B_{q}(x)=\frac{\nu_{R}}{\varphi(M_{R}q)}\,\operatorname{Li}(x)+O_{R}\!\bigl(xe^{-c_{R}\sqrt{\log x}}\bigr)=\frac{\delta(R)}{q-1}\,\operatorname{Li}(x)+O_{R}\!\bigl(xe^{-c_{R}\sqrt{\log x}}\bigr),$
-uniformly for $R<q\leq Q$. Here we use that $M_{R}$ is fixed and $q\leq(\log x)^{2}$, so $M_{R}q\leq(\log x)^{A}$ for some fixed $A$ and all sufficiently large $x$.
-Dividing by the asymptotic formula for $|\mathcal{A}_{R}(x)|$ yields
-$\frac{B_{q}(x)}{|\mathcal{A}_{R}(x)|}=\frac{1}{q-1}+O_{R}\!\bigl((\log x)e^{-c_{R}\sqrt{\log x}}\bigr)$
-uniformly for $R<q\leq Q$. Therefore
-$\frac{1}{|\mathcal{A}_{R}(x)|}\sum_{p\in\mathcal{A}_{R}(x)}S_{\leq Q}(p)$ $=\sum_{\begin{subarray}{c}R<q\leq Q\\ q\text{ prime}\end{subarray}}\frac{1}{q}\,\frac{B_{q}(x)}{|\mathcal{A}_{R}(x)|}$
-$\leq\sum_{\begin{subarray}{c}R<q\leq Q\\ q\text{ prime}\end{subarray}}\frac{1}{q(q-1)}+O_{R}\!\Bigl((\log x)e^{-c_{R}\sqrt{\log x}}\sum_{q\leq Q}\frac{1}{q}\Bigr)$
-$=\sum_{\begin{subarray}{c}R<q\leq Q\\ q\text{ prime}\end{subarray}}\frac{1}{q(q-1)}+o(1),$
-since $\sum_{q\leq Q}1/q\ll\log\log Q\ll\log\log\log x$.
-Combining this with the uniform bound on $S_{>Q}(p)$, we get
-$\frac{1}{|\mathcal{A}_{R}(x)|}\sum_{p\in\mathcal{A}_{R}(x)}S(p)\leq\eta(R)+o(1),\qquad\eta(R):=\sum_{\begin{subarray}{c}q>R\\ q\text{ prime}\end{subarray}}\frac{1}{q(q-1)}.$
-Since $\sum_{q}1/(q(q-1))$ converges, we have $\eta(R)\to 0$ as $R\to\infty$.
-Step 4: Markov’s inequality and conclusion.
-Choose $R$ large enough that $\eta(R)<\varepsilon/2$.
-By Markov’s inequality, the fraction of primes in $\mathcal{A}_{R}(x)$ with $S(p)>\varepsilon$ is at most
-$\frac{\eta(R)+o(1)}{\varepsilon}<\frac{1}{2}$
-for all sufficiently large $x$. Hence at least half the primes in $\mathcal{A}_{R}(x)$ satisfy $S(p)\leq\varepsilon$, and therefore
-$\frac{\varphi(p-1)}{p-1}>\frac{1}{2}-\varepsilon.$
-It follows that for all sufficiently large $x$,
-$\left\{p\leq x:p\text{ prime and }\frac{\varphi(p-1)}{p-1}>\frac{1}{2}-\varepsilon\right\}\geq\frac{1}{2}\,|\mathcal{A}_{R}(x)|.$
-Dividing by $\pi(x)$ and taking $\liminf$, we conclude that
-$\liminf_{x\to\infty}\frac{1}{\pi(x)}\#\left\{p\leq x:p\text{ prime and }\frac{\varphi(p-1)}{p-1}>\frac{1}{2}-\varepsilon\right\}$ $\geq\frac{1}{2}\lim_{x\to\infty}\frac{|\mathcal{A}_{R}(x)|}{\pi(x)}$
-$=\frac{\delta(R)}{2}>0.$
-This proves the lemma. ∎
+The numerical evidence of Remark 2 and Figure 3 suggests slow decay, but the oscillatory structure makes rigorous conclusions difficult.
+Determining a global modulus of continuity of $G$ is also open.
+Table 3 shows $1/c(p)\leq 4$ for all surveyed moduli; Corollary 8 gives $1/c(p)=O(\log\log p)$ pointwise.
+Closing the gap between the empirical bound $4$ and the asymptotic order $\log\log p$ remains an interesting open problem.
+Declaration of competing interests
+The author declares that there are no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
+Funding
+This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors.
+Research Data Statement
+No external datasets were used; all numerical values are reproducible from the explicit formulas and Monte Carlo procedures described in the paper.
+All empirical values of $c(p)$ in Figures 12 and Table 3 were computed by evaluating the totient ratio $\varphi(p-1)/(p-1)$ from the factorization of p1p-1, followed by the truncated Euler product $P_{J}(p)$ of Proposition 8 with $J=20$; the resulting error is below $10^{-9}$ for every prime $p\geq 3$.
+The Monte Carlo approximation of $G$ in Figure 1 samples the product distribution of Lemma 10 truncated at the 10,00010{,}000th odd prime $\ell_{N}$.
+Write $T_{N}:=\sum_{\ell>\ell_{N}}Z_{\ell}$ for the omitted tail and $S_{N}:=\sum_{3\leq\ell\leq\ell_{N}}Z_{\ell}$ for the retained partial sum, so that $X_{N}:=\tfrac{1}{2}e^{-S_{N}}$ approximates $X=\tfrac{1}{2}e^{-S}$ with $S=S_{N}+T_{N}$.
+The tail mean satisfies $\mathbb{E}[T_{N}]=\sum_{\ell>\ell_{N}}\frac{1}{\ell-1}\log\frac{\ell}{\ell-1}<10^{-5}$.
+Since $T_{N}\geq 0$ and $X_{N}\leq\tfrac{1}{2}$, the pointwise coupling bound $0\leq X_{N}-X=X_{N}(1-e^{-T_{N}})\leq T_{N}/2$ holds almost surely, giving $W_{1}(G_{N},G)\leq\tfrac{1}{2}\,\mathbb{E}[T_{N}]<5\times 10^{-6}$ for the Wasserstein-$1$ distance between the truncated and full distribution functions.
+Markov’s inequality further gives $\mathbb{P}(T_{N}>\varepsilon)<10^{-5}/\varepsilon$ for every $\varepsilon>0$.
+Figure 3 was produced by evaluating the closed-form Euler product of Proposition 5 over all primes $\ell\leq 10^{7}$.
+Tables 1 and 2 were computed from the explicit formula $B(K)=P_{3}\prod_{i=1}^{K}(1-1/p_{i})$ of Theorem 9 and the exact factorizations of p1p-1 for the listed moduli, respectively, using only rational arithmetic.

### 06 May 2026 v5

@@
-Primitive-Root Determinant Densities over Prime Fields and Implications for PRIM-LWE
-Abstract
-For a prime $p$, let $c_{n}(p)$ denote the fraction of $n\times n$ matrices over $\mathbb{F}_{p}$ whose determinant is a primitive root modulo $p$, and let $c(p)=\lim_{n\to\infty}c_{n}(p)$ be the limiting density.
-This quantity governs the efficiency of PRIM-LWE, a variant of the learning with errors (LWE) problem in which the secret matrix is required to have primitive-root determinant.
-The value $1/c(p)$ equals the expected rejection-sampling overhead in the reduction from standard LWE to PRIM-LWE.
-We prove unconditionally that $\inf_{p\text{ prime}}c(p)=0$, resolving an open question from the PRIM-LWE literature, and establish the rate $\min_{p\leq x}c(p)\asymp 1/\log\log x$ as $x\to\infty$.
-We show that $c(p)$, viewed as a function on the primes, has a continuous, purely singular limiting distribution with support $[0,1/2]$ and Hausdorff dimension 0.
-The moments of this limiting distribution are given by convergent Euler products, and the Mellin transform $\mathbb{E}[X^{s}]$ extends analytically to $\operatorname{Re}(s)>0$.
-Using the number of distinct prime factors $\omega(p-1)$, we derive explicit lower bounds: for every prime $p>2^{30}$ the overhead satisfies $1/c(p)<1.79\log p$, and in general $1/c(p)=O(\log\log p)$.
-We also describe a certified algorithm for evaluating $c(p)$ and tabulate values for primes of cryptographic interest.
-Keywords: Primitive root, determinant density, PRIM-LWE, limiting distribution, Hausdorff dimension
-MSC 2020: 11A07, 11N60, 94A60, 15B33
-1 Introduction
-Let $p$ be a prime and $M_{n}(\mathbb{F}_{p})$ denote the ring of $n\times n$ matrices over $\mathbb{F}_{p}$.
-Following [36], a matrix $M\in M_{n}(\mathbb{F}_{p})$ has primitive-root determinant if $\det(M)$ generates $\mathbb{F}_{p}^{*}$.
-Write $c_{n}(p)$ for the fraction of such matrices, and $c(p):=\inf_{n\geq 1}c_{n}(p)=\lim_{n\to\infty}c_{n}(p)$ for the limiting density; the precise formulas are recorded in Section 2.
-Since $c_{n}(p)$ is decreasing in $n$, one has $c(p)\leq c_{n}(p)$ for every $n\geq 1$.
-Consequently, every lower bound on $c(p)$ is simultaneously a lower bound on every finite-dimensional density $c_{n}(p)$.
-The quantity $c(p)$ arises as a reduction constant in the PRIM-LWE variant [36] of the popular learning with errors (LWE) problem [33], introduced to construct verifiable secret sharing schemes with “free” verification.
-The reduction proceeds by rejection-sampling and succeeds with probability $c_{n}(p)$ in dimension $n$, so the dimension-uniform overhead is $1/c(p)$.
-This overhead represents a tightness loss in the security reduction: the hardness guarantee for PRIM-LWE degrades by a factor of $c(p)$ relative to standard LWE, making tight bounds on $c(p)$ essential for concrete security estimates.
-Sehrawat, Yeo, and Desmedt [36, Remark 3(ii)] asked whether $\inf_{p}c(p)=0$ can be proved unconditionally.
+Primitive-root ratio over prime fields: a shifted-prime distribution of Hausdorff dimension zero and Implications for PRIM-LWE
+Abstract.
+For a prime $p$, let $c(p)$ denote the limiting fraction of $n\times n$ matrices over $\mathbb{F}_{p}$ whose determinant is a primitive root modulo $p$.
+The quantity $c(p)$ is a natural multiplicative deformation of the totient ratio $\varphi(p-1)/(p-1)$ and inherits its distributional behaviour over the primes.
+Existence and continuity of the limiting law follow from the shifted-prime Erdős–Wintner–Hildebrand framework.
+We prove the following new results: unconditionally, $\inf_{p}c(p)=0$ and the sharp order is $\min_{p\leq x}c(p)\asymp 1/\log\log x$; the reciprocal satisfies $\limsup_{p\to\infty,\,p\textup{ prime}}1/(c(p)\log\log p)=e^{\gamma}$, and no smaller constant suffices.
+We give a complete proof, combining an adaptation of Erdős’s argument with the Jessen–Wintner pure-type dichotomy, that the limiting distribution is purely singular, and strengthen this to $\dim_{H}(\mu_{G})=0$, i.e.
+the limiting measure is carried by a Borel set of Hausdorff dimension zero.
+The distribution has full topological support $0,\tfrac{1}{2}]$ and admits a Bernoulli-product representation indexed by the odd primes.
+The moments are given by convergent Euler products, and the Mellin transform $\mathbb{E}[X^{s}]$ extends to an entire function of $s$, non-vanishing on $\operatorname{Re}(s)>0$.
+Near the right endpoint, $1-G(\tfrac{1}{2}-\varepsilon)\sim\kappa/\log(1/\varepsilon)$ with an explicit constant $\kappa$.
+The quantity $1/c(p)$ equals the dimension-uniform expected rejection-sampling overhead in the reduction from learning with errors (LWE) to PRIM-LWE in lattice-based cryptography.
+The explicit bounds yield concrete overhead estimates for all primes appearing in current NIST post-quantum standards and representative NTT-friendly moduli.
+2020 Mathematics Subject Classification:
+11N60, 11A07, 11K65, 28A78, 94A60
+1. Introduction
+The distribution of the Euler totient ratio $\varphi(n)/n$ over the positive integers was first investigated by Schoenberg [30], who showed that the sequence possesses a continuous asymptotic distribution function.
+The existence of such a distribution was placed on a rigorous probabilistic footing by the Erdős–Wintner theorem [11], which identifies the class of real additive functions admitting a limiting law.
+Erdős [10] subsequently proved that this distribution is purely singular with respect to Lebesgue measure, and the Jessen–Wintner zero–one law [18] supplies the underlying dichotomy.
+The shifted-prime analogue—the behaviour of $\varphi(p-1)/(p-1)$ as $p$ ranges over primes—was addressed by Kátai [20] and, in definitive form, by Hildebrand [16], who extended the Erdős–Wintner framework to additive functions evaluated on shifted primes.
+Deshouillers and Hassani [7] refined this picture by establishing continuity in arithmetic progressions and proving left non-differentiability of the distribution function at every value $\varphi(m)/m$ with $m$ even.
+The quantity
+$c(p)\;:=\;\frac{\varphi(p-1)}{p-1}\;\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{p^{j}}\Bigr)$
+arises as the limiting density of $n\times n$ matrices over $\mathbb{F}_{p}$ whose determinant is a primitive root modulo $p$; the precise derivation is recalled in Section 2.
+Writing $P(p):=\prod_{j=1}^{\infty}(1-p^{-j})$, one has $P(p)\to 1$ as $p\to\infty$, so the behaviour of $c(p)$ is governed by the classical totient ratio $\varphi(p-1)/(p-1)$, and the distributional theory of $c(p)$ reduces to the shifted-prime Erdős–Wintner–Hildebrand framework.
+In this sense $c(p)$ is a natural multiplicative deformation of $\varphi(p-1)/(p-1)$.
+The open question whether $\inf_{p}c(p)=0$ was raised in [31, Remark 3(ii)]; the cryptographic implications of the explicit bounds are developed in Section 8.
 Main results
-1.
-Unconditional vanishing (Theorem 2).
-We prove $\inf_{p\text{ prime}}c(p)=0$ using only Dirichlet’s theorem on primes in arithmetic progressions and Mertens’ third theorem, without assuming the primorial-prime hypothesis; see Section 3.
-2. Sharp order of the minimum (Proposition 1, Corollary 7). We establish
+We consolidate our principal findings into four theorems.
+[…]
 $\min_{p\leq x}c(p)\;\asymp\;\frac{1}{\log\log x}\qquad(x\to\infty).$
-See Remark 4 for effectivity.
-3.
-Limiting distribution (Theorems 4 and 5, Corollary 2, Proposition 4).
-The quantity $c(p)$ possesses a continuous, purely singular limiting distribution over the primes whose support is exactly $0,\tfrac{1}{2}]$; the limiting measure has Hausdorff dimension 0.
-The full-support conclusion is new; it combines the left non-differentiability result of Deshouillers–Hassani [9, Theorem 1.1] with the density of even totient ratios (Lemma 4).
-The moments of $G$ are given by convergent Euler products (Theorem 6), the Mellin transform $\mathbb{E}[X^{s}]$ extends analytically to $\{\operatorname{Re}(s)>0\}$ (Theorem 8), and $G$ is not Hölder of any positive order at $1/2$ (Proposition 7).
-Since the limiting law is purely singular and supported on a set of Hausdorff dimension 0, it cannot certify $c(p)$ at any individual prime; this motivates the pointwise bounds of item 4.
-4.
-Explicit bounds via distinct prime factors (Theorem 9).
-We give unconditional lower bounds on $c(p)$ for primes $p$ parameterized by $\omega(p-1)$, the number of distinct prime factors of p1p-1, yielding computable bounds for any prime $p$ such that the factorization of p1p-1 is known.
-For instance, for the NIST-standardized primes $p=3329$ and $p=8380417$, the exact factorizations of p1p-1 give $c(3329)\geq 0.4613$ and $c(8380417)\geq 0.2932$; see Section 7.
-In particular, the dimension-uniform expected rejection-sampling overhead $1/c(p)$ in the LWE-to-PRIM-LWE reduction [36, Theorem 10] admits an explicit upper bound depending only on $\omega(p-1)$.
-From the computational viewpoint, our main contributions are: explicit lower bounds on $c(p)$ valid for all primes $p>2^{30}$ (Corollary 9), a certified rational-arithmetic algorithm for evaluating $c(p)$ to arbitrary precision given the factorization of p1p-1 (Proposition 8), and tables of $c(p)$ for all primes appearing in current NIST post-quantum cryptographic standards and representative lattice-based implementations (Tables 2 and 3).
-The explicit bound takes the following compact form:
-Corollary 1 (Cryptographic consequence).
-Let $P_{3}:=\prod_{j=1}^{\infty}(1-1/3^{j})\approx 0.5601260779$. If $p\geq 3$ is a prime modulus with $\omega(p-1)\leq K$, then
-$\frac{1}{c(p)}\;\leq\;\frac{1}{P_{3}}\prod_{i=1}^{K}\frac{p_{i}}{p_{i}-1}.$
-The general $\omega(p-1)$-based bound is immediate from Theorem 9. Table 1 lists representative values of $B(K)$, and Tables 2 and 3 collect these bounds for representative cryptographic moduli.
-Related work
-A related distribution problem was studied by Aivazidis and Sofos [1], who considered the density of maximal-order elements in $\mathrm{GL}_{n}(q)$ as $q$ varies over prime powers and proved corresponding distribution laws.
-Their setting is different from ours: we work over prime fields and with the primitive-root determinant density arising in PRIM-LWE.
-On the distributional side, Deshouillers and Hassani [9, Proposition 2.1 and equation (2.1)] proved limiting-distribution laws for $\varphi(p-1)/(p-1)$ in the residue classes $p\equiv 1\pmod{m}$, including continuity, and [9, Theorem 1.1] established left non-differentiability of the distribution function at every value $\varphi(m)/m$ with $m$ even.
-The analogous support question for the classical distribution of $\varphi(n)/n$ over the integers remains open; see [9].
-2 Background and the Open Question
-Throughout, $\log$ denotes the natural logarithm. We write $\varphi$ for Euler’s totient function, $\omega(n)$ for the number of distinct prime factors of $n$, and $p_{k}$ for the $k$-th prime.
-Following [36], define
+The extremal order of the reciprocal is sharp:
+$\limsup_{p\to\infty,\,p\textup{ prime}}1/(c(p)\,\log\log p)=e^{\gamma}\approx 1.781.$
+In particular, $1/c(p)\leq(e^{\gamma}+o(1))\log\log p$ for every prime $p$, and no smaller constant suffices.
+The limsup is realised asymptotically along the least primes $p_{k}^{*}\equiv 1\pmod{N_{k}}$, whose existence is guaranteed by Dirichlet’s theorem and whose size is controlled by Linnik’s theorem.
+See Theorem 3.3, Proposition 4.2, Corollary 7.2, and Theorem 7.4. Remark 7.5 discusses effectivity.
+Theorem B (Limiting law).The quantity $c(p)$ possesses a continuous limiting distribution $G$ over the primes, coinciding with the distribution of $\varphi(p-1)/(p-1)$.
+The topological support of $G$ is the full interval $0,\tfrac{1}{2}]$, and $G$ admits a Bernoulli-convolution representation: if $B_{\ell}$ are independent Bernoulli random variables with $\mathbb{P}(B_{\ell}=1)=1/(\ell-1)$ for each odd prime $\ell$, then
+$X\;\stackrel{{\scriptstyle d}}{{=}}\;\tfrac{1}{2}\prod_{\ell\geq 3,\;\ell\textup{ prime}}\Bigl(1-\frac{1}{\ell}\Bigr)^{\!B_{\ell}}.$
+See Theorems 5.5 and 5.8, Corollary 5.14, and Lemma 6.1. Equivalently, the logarithmic transform of the distribution is an infinite convolution of two-point measures (Lemma 6.1).
+Theorem C (Hausdorff dimension zero and singularity).The Hausdorff dimension of the limiting measure $\mu_{G}$ is zero: $\dim_{H}(\mu_{G})=0$.
+In particular, $\mu_{G}$ is purely singular with respect to Lebesgue measure.
+See Theorem 6.2 and Section 5.3.
+Theorem D (Mellin transform and endpoint law).For every $s\in\mathbb{C}$, the Mellin transform $M(s)=\mathbb{E}[X^{s}]$ is given by a convergent Euler product; it extends to an entire function of $s$, non-vanishing on $\{\operatorname{Re}(s)>0\}$.
+Near the right endpoint, $1-G(\tfrac{1}{2}-\varepsilon)\sim\kappa/\log(1/\varepsilon)$ as $\varepsilon\downarrow 0$, where $\kappa=\mathfrak{S}_{2}\,e^{-\gamma}$ and $\mathfrak{S}_{2}$ is the Hardy–Littlewood twin-prime singular series.
+See Theorems 6.3 and 6.4, Proposition 6.9, and Corollary 6.10.
+The twin-prime singular series enters because the event that no small prime $\ell$ divides p1p-1 forces $\varphi(p-1)/(p-1)$ close to $1/2$.
+The resulting local factors $1-1/(\ell-1)^{2}$ are precisely those governing the density of twin primes in the Hardy–Littlewood heuristic [15].
+As a consequence, explicit lower bounds on $c(p)$ via $\omega(p-1)$ yield effective estimates for any prime whose factorisation of p1p-1 is known; see Section 7.
+Davenport [6] independently studied the distribution of $\sigma(n)/n$ over the integers, where $\sigma$ denotes the sum-of-divisors function; for $\varphi(n)/n$ the topological support is the full interval $[0,1]$ and the global modulus of continuity remains unknown (see Section 9).
+Hildebrand’s proof [16] of the shifted-prime Erdős–Wintner theorem builds on the equidistribution of primes in arithmetic progressions on average established by Bombieri [3]; see also Tenenbaum [32, 33].
+Tenenbaum [32] remarked that the pure singularity of the shifted-prime distribution can be obtained by adapting an exercise solution in [34, Exercise 2.4] (corresponding to [33, Exercise 256]), and Deshouillers–Hassani [7] cite the same exercise, but neither source carries out the argument.
+We give a complete proof in Section 5.3 below, combining an explicit adaptation of Erdős’s argument with the Jessen–Wintner pure-type dichotomy [18].
+We strengthen this to the metric result $\dim_{H}(\mu_{G})=0$ in Section 6.1.
+Wintner’s convergence theory [36]—the founding document for infinite convolutions of discrete measures—supplies the underlying framework.
+The Hausdorff dimension of Bernoulli convolutions has a substantial literature; see Kahane [19] for an early observation that Erdős-type arguments yield Hausdorff dimension estimates, and Peres, Schlag, and Solomyak [26] and Peres–Solomyak [27] for further dimensional results.
+The classical Bernoulli convolution involves a fixed contraction ratio.
+The infinite convolution arising here involves varying ratios $\log(\ell/(\ell-1))$ indexed by the odd primes, so the standard Pisot–Salem machinery does not apply directly.
+Billingsley [2] and Elliott [13] provide modern probabilistic perspectives.
+The extremal behaviour of $\varphi(n)/n$ dates to Mertens [23], whose third theorem gives $\prod_{\ell\leq x}(1-1/\ell)\sim e^{-\gamma}/\log x$; the systematic study of extremal values of arithmetic functions was initiated by Ramanujan [28].
+Over the primes, primes $p\equiv 1\pmod{N_{k}}$—for which p1p-1 has many small prime factors—realise the smallest values of the totient ratio, and the size of the least such prime is controlled by Linnik’s theorem [21].
+Aivazidis and Sofos [1] considered the density of maximal-order elements in $\mathrm{GL}_{n}(q)$ over prime powers.
+There is also a conceptual kinship with the Artin primitive-root conjecture (see Hooley [17] for a conditional proof under GRH): while Artin’s conjecture concerns the behaviour of a fixed base as $p$ varies, the present work concerns the aggregate density $c(p)$ in the complementary regime where the base varies uniformly.
+Both problems ultimately depend on the factorisation of p1p-1.
+Organisation
+Section 2 recalls the derivation of $c(p)$ and the open question from [31].
+Section 3 resolves this question unconditionally, and Section 4 gives quantitative refinements via Linnik’s theorem.
+Section 5 develops the distributional theory: existence and continuity of the limiting law (Section 5.1), transfer from $\varphi(p-1)/(p-1)$ to $c(p)$ (Section 5.2), and the complete proof of pure singularity (Section 5.3).
+Section 6 establishes the convolution structure and Hausdorff dimension zero (Section 6.1), the moment Euler product and Mellin transform (Section 6.2), the Fourier structure (Section 6.3), and the endpoint asymptotic (Section 6.4).
+Section 7 provides explicit lower bounds, Section 8 applies these bounds to the PRIM-LWE reduction in lattice-based cryptography, and Section 9 concludes by discussing two open questions.
+2. Background and the Open Question
+Throughout, $\log$ denotes the natural logarithm. We write $\omega(n)$ for the number of distinct prime factors of $n$ and $p_{k}$ for the $k$-th prime.
+Define (cf. [31])
 $M_{n}^{\mathrm{prim}}(\mathbb{F}_{p})\;=\;\bigl\{\,M\in M_{n}(\mathbb{F}_{p}):\det(M)\text{ generates }\mathbb{F}_{p}^{*}\,\bigr\}.$
@@
 $|\mathrm{GL}_{n}(\mathbb{F}_{p})|=p^{n^{2}}\prod_{j=1}^{n}\Bigl(1-\frac{1}{p^{j}}\Bigr)$
-(see, e.g., Mullen and Panario [30, Theorem 13.3.3]), it follows that
+(see, e.g., Mullen and Panario [25, Theorem 13.3.3]), it follows that
 $c_{n}(p)\;:=\;\frac{|M_{n}^{\mathrm{prim}}(\mathbb{F}_{p})|}{|M_{n}(\mathbb{F}_{p})|}\;=\;\frac{\varphi(p-1)}{p-1}\;\prod_{j=1}^{n}\Bigl(1-\frac{1}{p^{j}}\Bigr),$
-cf. [36, Corollary 4]. Consequently, the limit
+cf. [31, Corollary 4]. Consequently, the limit
 $c(p)\;:=\;\lim_{n\to\infty}c_{n}(p)\;=\;\frac{\varphi(p-1)}{p-1}\;\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{p^{j}}\Bigr)$
 exists and is strictly positive for each fixed $p$. Since $c_{n}(p)$ is decreasing in $n$, $\inf_{n\geq 1}c_{n}(p)=c(p)$.
-Writing $p-1=r_{1}^{e_{1}}\cdots r_{k}^{e_{k}}$ for the prime factorization of p1p-1, one has
-$c(p)\;=\;\prod_{i=1}^{k}\Bigl(1-\frac{1}{r_{i}}\Bigr)\;\cdot\;\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{p^{j}}\Bigr).$ (1)
-Since $P(p):=\prod_{j=1}^{\infty}(1-p^{-j})\to 1$ as $p\to\infty$, the behavior of $c(p)$ over the primes is governed by $\varphi(p-1)/(p-1)$, and the distributional theory in Section 5 reduces to the classical shifted-prime distribution of this ratio.
-In [36, Remark 3(ii)], Sehrawat, Yeo, and Desmedt observed that if there exist infinitely many primorial primes $p=\prod_{i=1}^{k}p_{i}+1$, then $\varphi(p-1)/(p-1)=\prod_{i=1}^{k}(p_{i}-1)/p_{i}\to 0$ by Mertens’ third theorem, which would imply $\inf_{p}c(p)=0$.
+Writing $p-1=r_{1}^{e_{1}}\cdots r_{k}^{e_{k}}$ for the prime factorisation of p1p-1, one has
+(2.1) $c(p)\;=\;\prod_{i=1}^{k}\Bigl(1-\frac{1}{r_{i}}\Bigr)\;\cdot\;\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{p^{j}}\Bigr).$
+Since $P(p):=\prod_{j=1}^{\infty}(1-p^{-j})\to 1$ as $p\to\infty$, the behaviour of $c(p)$ over the primes is governed by $\varphi(p-1)/(p-1)$, and the distributional theory in Section 5 reduces to the classical shifted-prime distribution of this ratio.
+In [31, Remark 3(ii)], Sehrawat, Yeo, and Desmedt observed that if there exist infinitely many primorial primes $p=\prod_{i=1}^{k}p_{i}+1$, then $\varphi(p-1)/(p-1)=\prod_{i=1}^{k}(p_{i}-1)/p_{i}\to 0$ by Mertens’ third theorem, which would imply $\inf_{p}c(p)=0$.
 They left open whether this can be proved unconditionally; see [4] for computational results on primorial primes.
-3 Unconditional Proof that $\inf_{p}c(p)=0$
+3. Unconditional Proof that $\inf_{p}c(p)=0$
 The argument requires only two classical results: Dirichlet’s theorem on primes in arithmetic progressions [8] and the following.
-Theorem 1 (Mertens’ Third Theorem [28]; see [29, Theorem 2.7(e)], 1874).
+Theorem 3.1 (Mertens’ Third Theorem [23]; see [24, Theorem 2.7(e)], 1874).
 As $x\to\infty$,
@@
 where $\gamma\approx 0.5772$ is the Euler–Mascheroni constant.
-Lemma 1 (Lower bound for the matrix-count Euler product).
+Lemma 3.2 (Lower bound for the matrix-count Euler product).
 For all primes $p\geq 3$,
@@
 By the finite-product inequality $\prod(1-a_{j})\geq 1-\sum a_{j}$ (immediate by induction when $0\leq a_{j}\leq 1$) and the geometric series $\sum_{j\geq 1}p^{-j}=1/(p-1)\leq 1/2$. ∎
-Theorem 2 (Unconditional vanishing of $\inf c(p)$).
+Theorem 3.3 (Unconditional vanishing of $\inf c(p)$).
 The following identities hold:
@@
 $\frac{\varphi(p-1)}{p-1}\;=\;\prod_{\ell\mid(p-1)}\Bigl(1-\frac{1}{\ell}\Bigr)\;\leq\;\prod_{i=1}^{k}\Bigl(1-\frac{1}{p_{i}}\Bigr).$
-By Mertens’ third theorem (Theorem 1),
+By Mertens’ third theorem (Theorem 3.1),
 $\prod_{i=1}^{k}\Bigl(1-\frac{1}{p_{i}}\Bigr)\;\sim\;\frac{e^{-\gamma}}{\log p_{k}}\;\xrightarrow{k\to\infty}\;0.$
@@
 and the infimum statement follows immediately. ∎
-4 Quantitative Refinements
+4. Quantitative Refinements
 By Linnik’s theorem [21], there exist absolute constants $C_{0},L>0$ such that the least prime in any reduced residue class modulo $d$ is at most $C_{0}d^{L}$.
@@
 $N_{k}:=\prod_{i=1}^{k}p_{i},\qquad\vartheta(x):=\sum_{\begin{subarray}{c}p\leq x\\ p\text{ prime}\end{subarray}}\log p,\qquad m(x):=\max\{k\geq 1:N_{k}\leq x\}.$
-Lemma 2 (Prime-number-theorem consequences for primorials; cf. [29, Corollary 2.5]).
+Lemma 4.1 (Prime-number-theorem consequences for primorials; cf. [24, Theorem 6.9 and Corollary 2.5]).
 As $k\to\infty$,
@@
 The first two assertions follow from $\vartheta(x)\sim x$. For the last two, $N_{m(x)}\leq x<N_{m(x)}p_{m(x)+1}$ gives $\log x\sim\log N_{m(x)}\sim p_{m(x)}$, since $\log p_{m(x)+1}=o(p_{m(x)})$. ∎
-Proposition 1 (Upper bound for the minimum density up to $x$).
+Proposition 4.2 (Upper bound for the minimum density up to $x$).
 There exists an absolute constant $C>0$ such that for all sufficiently large $x$,
@@
 Such a $k$ exists for all sufficiently large $x$, and $k\to\infty$ as $x\to\infty$.
-Since $\log N_{k}=\sum_{i=1}^{k}\log p_{i}\sim p_{k}$ by the prime number theorem, the inequality $N_{k}^{\,L}\leq x$ gives
-$L\,\log N_{k}\leq\log x,$
+Since $\log N_{k}=\sum_{i=1}^{k}\log p_{i}\sim p_{k}$ by the prime number theorem, the inequality $C_{0}N_{k}^{\,L}\leq x$ gives
+$L\,\log N_{k}\leq\log x-\log C_{0},$
 and hence $p_{k}\ll\log x$. Therefore $\log p_{k}=O(\log\log x)$.
@@
 $\log p_{k}\asymp\log\log x.$
-By construction the least prime $p^{*}(k)\equiv 1\pmod{N_{k}}$ satisfies $p^{*}(k)\leq x$, and the argument in the proof of Theorem 2 gives
+By construction the least prime $p^{*}(k)\equiv 1\pmod{N_{k}}$ satisfies $p^{*}(k)\leq x$, and the argument in the proof of Theorem 3.3 gives
 $c\bigl(p^{*}(k)\bigr)\;\leq\;\prod_{i=1}^{k}\Bigl(1-\frac{1}{p_{i}}\Bigr)\;\sim\;\frac{e^{-\gamma}}{\log p_{k}}\;\asymp\;\frac{1}{\log\log x}.$
@@
 as claimed. ∎
-The matching lower bound $\min_{p\leq x,\,p\text{ prime}}c(p)\gg 1/\!\log\log x$, which completes the proof that the order is sharp, is established in Corollary 7 below; its proof requires only the primorial counting argument and Lemma 1.
-5 The Limiting Distribution of $c(p)$
-The classical existence and continuity of the limiting distribution of $\varphi(p-1)/(p-1)$ over primes follow from the shifted-prime Erdős–Wintner framework [16] (Tenenbaum [38, pp.
-262–263], Deshouillers–Hassani [9]).
+The matching lower bound $\min_{p\leq x,\,p\text{ prime}}c(p)\gg 1/\!\log\log x$, which completes the proof that the order is sharp, is established in Corollary 7.2; its proof requires only the primorial counting argument and Lemma 3.2.
+In particular, the lower bound is independent of the distributional machinery in Sections 56: it uses the explicit inequality $c(p)\geq P_{3}\prod_{i=1}^{\omega(p-1)}(1-1/p_{i})$ of Theorem 7.1 together with the primorial comparison $\omega(p-1)\leq m(p-1)$ and Mertens’ third theorem.
+5. The Limiting Distribution of $c(p)$
+The classical existence and continuity of the limiting distribution of $\varphi(p-1)/(p-1)$ over primes follow from the shifted-prime Erdős–Wintner framework [11] (Tenenbaum [32], Deshouillers–Hassani [7]; see also [34]).
+We recall these results in Sections 5.1 and 5.2 for completeness.
+Pure singularity of the shifted-prime law was anticipated by a remark of Tenenbaum [32], who noted that the integer-case argument of Erdős [10] (with solution in [34, Exercise 2.4]) can be adapted.
+We carry out this adaptation in Section 5.3, providing a complete proof in the shifted-prime setting.
+The strictly stronger result $\dim_{H}(\mu_{G})=0$ is established in Section 6.1.
 Throughout this section we write
-$f(n)\;:=\;\log\frac{n}{\varphi(n)}\;=\;\sum_{\begin{subarray}{c}\ell\mid n\\ \ell\textup{ prime}\end{subarray}}\log\frac{\ell}{\ell-1}$ (2)
-5.1 The distribution of $\varphi(p-1)/(p-1)$
-Lemma 3 (Verification of the Erdős–Wintner hypotheses).
-The function $f$ defined in (2) is strongly additive, with $f(n)=\log(n/\varphi(n))=-\log(\varphi(n)/n)$. For every prime $\ell$,
+(5.1) $f(n)\;:=\;\log\frac{n}{\varphi(n)}\;=\;\sum_{\begin{subarray}{c}\ell\mid n\\ \ell\textup{ prime}\end{subarray}}\log\frac{\ell}{\ell-1}\,.$
+5.1. The distribution of $\varphi(p-1)/(p-1)$
+Lemma 5.1 (Verification of the Erdős–Wintner hypotheses).
+The function $f$ defined in (5.1) is strongly additive, with $f(n)=\log(n/\varphi(n))=-\log(\varphi(n)/n)$. For every prime $\ell$,
 $0<f(\ell)=\log\frac{\ell}{\ell-1}\leq\log 2<1,\qquad f(\ell)=\frac{1}{\ell}+O\!\left(\frac{1}{\ell^{2}}\right).$
@@
 $\sum_{|f(\ell)|>1}\frac{1}{\ell}<\infty,\qquad\sum_{\ell}\frac{f(\ell)}{\ell}<\infty,\qquad\sum_{\ell}\frac{f(\ell)^{2}}{\ell}<\infty,$
-as does the continuity criterion $\sum_{f(\ell)\neq 0}1/\ell=\infty$. The same conclusions hold for $g(n)=\varphi(n)/n$.
+as does the continuity criterion $\sum_{f(\ell)\neq 0}1/\ell=\infty$.
+The same convergence conditions and continuity criterion hold for f-f, and the limiting distribution of $g(p-1)=\varphi(p-1)/(p-1)$ is the push-forward of the limiting distribution of f(p1)f(p-1) under $u\mapsto e^{-u}$.
 Proof.
@@
 Since $f(\ell)\neq 0$ for every prime $\ell$ and $\sum_{\ell}1/\ell=\infty$, the continuity criterion is satisfied.
-The case $g(n)=\varphi(n)/n$ follows from $\log g(\ell)=-f(\ell)$.
+Since $\log(\varphi(n)/n)=-f(n)$ is also strongly additive with the same absolute values at primes, all three series conditions and the continuity criterion transfer to f-f.
+The distributional conclusion for $g(n)=\varphi(n)/n$ then follows by applying the continuous map $u\mapsto e^{-u}$.
 ∎
-Proposition 2 (Deshouillers–Hassani; special case of [9, Proposition 2.1 and equation (2.1)]).
+Proposition 5.2 (Deshouillers–Hassani; special case of [7, Proposition 2.1 and equation (2.1)]).
 For each integer $m\geq 1$, there exists a continuous distribution function $G_{m}\colon\mathbb{R}\to[0,1]$ such that for every $\alpha\in\mathbb{R}$,
-$\lim_{x\to\infty}\frac{1}{\pi(x;m,1)}\#\left\{p\leq x:p\textup{ prime},\ p\equiv 1\pmod{m},\;\frac{\varphi(p-1)}{p-1}\leq\alpha\right\}=G_{m}(\alpha).$
+$\lim_{x\to\infty}\frac{1}{\pi(x;m,1)}\#\Bigl\{p\leq x:p\textup{ prime},\ p\equiv 1\pmod{m},\;\frac{\varphi(p-1)}{p-1}\leq\alpha\Bigr\}\\ =G_{m}(\alpha).$
 Moreover, the limiting measure is supported on $0,\varphi(m)/m]$.
 Proof.
-Apply Deshouillers and Hassani [9, Proposition 2.1] to
+Apply Deshouillers and Hassani [7, Proposition 2.1] to
 $g(n):=\frac{\varphi(n)}{n}.$
-By Lemma 3 applied to $\log g(\ell)=-\log(\ell/(\ell-1))$, the three convergence hypotheses of [9, Proposition 2.1] are satisfied, and there exists a distribution function $G_{m}$ such that the displayed limit holds at every continuity point of $G_{m}$.
-By the continuity criterion in the same proposition, $G_{m}$ is continuous if and only if
+Since $\log g(\ell)=-\log(\ell/(\ell-1))=-f(\ell)$, Lemma 5.1 (which verifies the convergence conditions for both $f$ and f-f) shows that the three convergence hypotheses of [7, Proposition 2.1] are satisfied, and there exists a distribution function $G_{m}$ such that the displayed limit holds at every continuity point of $G_{m}$.
+By the Jessen–Wintner continuity criterion (applied in [7, Proposition 2.1, equation (2.1)]), $G_{m}$ is continuous if
 $\sum_{\begin{subarray}{c}\ell\equiv 1\pmod{m}\\ g(\ell)\neq 1\end{subarray}}\frac{1}{\ell}$
-diverges. Here $g(\ell)=1-\frac{1}{\ell}\neq 1$ for every prime $\ell$, and
+diverges.
+(The full criterion requires $\sum_{\ell\nmid m,\,g(\ell)\neq 1}1/\ell=\infty$, which holds a fortiori since $g(\ell)\neq 1$ for every prime.) Here $g(\ell)=1-\frac{1}{\ell}\neq 1$ for every prime $\ell$, and
 $\sum_{\ell\equiv 1\pmod{m}}\frac{1}{\ell}=\infty$
-by the divergence of the prime harmonic series in every reduced residue class (see, e.g., [29, Corollary 4.10]).
+by the divergence of the prime harmonic series in every reduced residue class (see, e.g., [24, Corollary 4.10]).
 Thus $G_{m}$ is continuous, so the displayed limit holds for every $\alpha\in\mathbb{R}$.
@@
 which yields the support claim. ∎
-We shall also use the following standard Siegel–Walfisz (see [29, Corollary 11.21] or [39, Theorem II.8.17]) estimate in a range of polylogarithmic moduli.
-Theorem 3 (Siegel–Walfisz).
+We shall also use the following standard Siegel–Walfisz (see [24, Corollary 11.21] or [33, Theorem II.8.17]) estimate in a range of polylogarithmic moduli.
+Theorem 5.3 (Siegel–Walfisz).
 For every fixed $A>0$ there exists $c_{A}>0$ such that uniformly for integers $m\leq(\log x)^{A}$ and reduced residue classes $a\pmod{m}$,
@@
 $\pi(x;m,a)=\#\{p\leq x:p\textup{ prime},\ p\equiv a\pmod{m}\}$.
-Lemma 4 (Density of even totient ratios).
+Lemma 5.4 (Density of even totient ratios).
 The set
@@
 is dense in $(0,1/2]$. Since the finite subsums can be arbitrarily large, 0 is also a limit point. Hence the displayed set is dense in $[0,1/2]$, as claimed. ∎
-Theorem 4 (Limiting distribution and full support over primes).
+Theorem 5.5 (Limiting distribution and full support over primes).
 There exists a continuous distribution function $G\colon\mathbb{R}\to[0,1]$ such that for every $\alpha\in\mathbb{R}$,
@@
 Proof.
-Apply Proposition 2 with $m=2$. Since every odd prime satisfies $p\equiv 1\pmod{2}$, we have
+Apply Proposition 5.2 with $m=2$. Since every odd prime satisfies $p\equiv 1\pmod{2}$, we have
 $\pi(x;2,1)=\pi(x)-1\sim\pi(x),$
 so $G_{2}$ is the unconditional limiting distribution of $\varphi(p-1)/(p-1)$ over all primes.
-Writing $G:=G_{2}$ (denoted $F$ in [9, Theorem 1.1]), there exists a continuous distribution function $G\colon\mathbb{R}\to[0,1]$ such that for every $\alpha\in\mathbb{R}$,
+Writing $G:=G_{2}$ (denoted $F$ in [7, Theorem 1.1]), there exists a continuous distribution function $G\colon\mathbb{R}\to[0,1]$ such that for every $\alpha\in\mathbb{R}$,
 $\lim_{x\to\infty}\frac{1}{\pi(x)}\#\bigl\{p\leq x:p\text{ prime},\;\varphi(p-1)/(p-1)\leq\alpha\bigr\}=G(\alpha).$
@@
 $G(\alpha)=0\qquad(\alpha\leq 0),\qquad G(\alpha)=1\qquad(\alpha\geq 1/2).$
-Let $\mu_{\varphi}$ denote the limiting measure with distribution function $G$. Fix an even integer $m\geq 2$, and write
+Let $\mu_{G}$ denote the limiting measure with distribution function $G$. Fix an even integer $m\geq 2$, and write
 $x_{m}:=\frac{\varphi(m)}{m}.$
-By Deshouillers and Hassani [9, Theorem 1.1], the distribution function $G$ is not left-differentiable at $x_{m}$.
+By Deshouillers and Hassani [7, Theorem 1.1], the distribution function $G$ is not left-differentiable at $x_{m}$.
 Deshouillers and Hassani denote this distribution function by $F$; we write it as $G$.
-If $x_{m}\notin\operatorname{supp}(\mu_{\varphi})$, then some open interval $I$ containing $x_{m}$ has $\mu_{\varphi}(I)=0$.
+If $x_{m}\notin\operatorname{supp}(\mu_{G})$, then some open interval $I$ containing $x_{m}$ has $\mu_{G}(I)=0$.
 Hence $G$ is locally constant on $I$, and in particular left-differentiable at $x_{m}$, a contradiction.
-Therefore $x_{m}\in\operatorname{supp}(\mu_{\varphi})$ for every even $m$.
-By Lemma 4, the support contains a dense subset of $0,\tfrac{1}{2}]$. Since the support is closed and already lies inside $0,\tfrac{1}{2}]$, it follows that
-$\operatorname{supp}(\mu_{\varphi})=[0,\tfrac{1}{2}].$
+Therefore $x_{m}\in\operatorname{supp}(\mu_{G})$ for every even $m$.
+By Lemma 5.4, the support contains a dense subset of $0,\tfrac{1}{2}]$. Since the support is closed and already lies inside $0,\tfrac{1}{2}]$, it follows that
+$\operatorname{supp}(\mu_{G})=[0,\tfrac{1}{2}].$
 Now let $0\leq a<b\leq 1/2$.
 Choose any $x\in(a,b)$.
-Since $x\in\operatorname{supp}(\mu_{\varphi})$, every neighbourhood of $x$ has positive $\mu_{\varphi}$-measure.
+Since $x\in\operatorname{supp}(\mu_{G})$, every neighbourhood of $x$ has positive $\mu_{G}$-measure.
 Taking $r>0$ so small that $(x-r,x+r)\subset(a,b)$, we obtain
-$\mu_{\varphi}((a,b])\geq\mu_{\varphi}((x-r,x+r))>0.$
+$\mu_{G}((a,b])\geq\mu_{G}((x-r,x+r))>0.$
 Hence
-$G(b)-G(a)=\mu_{\varphi}((a,b])>0,$
+$G(b)-G(a)=\mu_{G}((a,b])>0,$
 so $G$ is strictly increasing on $0,\tfrac{1}{2}]$. The displayed bounds $0<G(\tau)<1$ for $0<\tau<1/2$ follow immediately from this and the endpoint values $G(0)=0$, $G(1/2)=1$. ∎
-The support question for $\varphi(n)/n$ over the integers remains open; see Schoenberg [35] and [9].
-5.2 From $\varphi(p-1)/(p-1)$ to $c(p)$
-Lemma 5 (Transport lemma).
+Remark 5.6 (Alternative support argument).
+Full support also follows from the Bernoulli-convolution representation established in Section 6.1: the support of the infinite convolution $\mu_{f}$ of Lemma 6.1 is the closure of the set of partial sums $\log 2+\sum_{\ell\in S}\log(\ell/(\ell-1))$ over finite sets $S$ of odd primes, and the density argument of Lemma 5.4 shows this closure equals $\log 2,\infty)$.
+The proof above uses instead the non-differentiability result of Deshouillers–Hassani, which connects the support question to the local analytic behaviour of $G$.
+For the classical distribution of $\varphi(n)/n$ over the integers, the topological support is $[0,1]$: the argument of Lemma 5.4, applied over all primes and without the forced factor $1/2$, gives density of $\{\varphi(n)/n:n\geq 1\}$ in $[0,1]$, and the Erdős–Wintner theorem [11] (see also Schoenberg [30]) identifies this closure with the support.
+The global modulus of continuity remains open; see Section 9.
+5.2. From $\varphi(p-1)/(p-1)$ to $c(p)$
+Lemma 5.7 (Perturbation lemma).
 Let $X(p)\geq 0$ and $a(p)>0$ be functions of the primes. Assume that $X(p)$ has a continuous limiting distribution function $G$ over the primes, i.e.
@@
 Proof.
-Fix $\alpha\geq 0$ and $0<\varepsilon<1$. For all sufficiently large primes, $1-\varepsilon<a(p)<1+\varepsilon$, so
+For $\alpha<0$ both sides are zero since $X(p)\geq 0$ and $a(p)>0$, so the claim is immediate. We may therefore assume $\alpha\geq 0$.
+Fix $0<\varepsilon<1$. For all sufficiently large primes, $1-\varepsilon<a(p)<1+\varepsilon$, so
 $\{X(p)\leq\alpha/(1+\varepsilon)\}\;\subseteq\;\{a(p)X(p)\leq\alpha\}\;\subseteq\;\{X(p)\leq\alpha/(1-\varepsilon)\}$
 up to finitely many exceptions. Taking densities as $x\to\infty$ and letting $\varepsilon\downarrow 0$, continuity of $G$ gives the claim. ∎
-Theorem 5 (Transfer to the PRIM-LWE density).
+Theorem 5.8 (Transfer to the primitive-root determinant density).
 The quantity $c(p)$ has the same limiting distribution as $\varphi(p-1)/(p-1)$. That is, for every $\alpha\in\mathbb{R}$,
 $\lim_{x\to\infty}\;\frac{1}{\pi(x)}\,\#\bigl\{p\leq x:p\textup{ prime},\;c(p)\leq\alpha\bigr\}\;=\;G(\alpha),$
-where $G$ is the distribution function from Theorem 4.
+where $G$ is the distribution function from Theorem 5.5.
 Proof.
@@
 hence
-$P(p)=1+O\!\left(\frac{1}{p}\right)\qquad(p\to\infty\text{ through the primes}).$
-By Theorem 4, $X(p)$ has continuous limiting distribution function $G$. Therefore Lemma 5 applies with $a(p)=P(p)$, and shows that $c(p)=a(p)X(p)$ has the same limiting distribution $G$. ∎
-5.3 Pure singularity
+$P(p)=1-O\!\left(\frac{1}{p}\right)\qquad(p\to\infty\text{ through the primes}).$
+By Theorem 5.5, $X(p)$ has continuous limiting distribution function $G$. Therefore Lemma 5.7 applies with $a(p)=P(p)$, and shows that $c(p)=a(p)X(p)$ has the same limiting distribution $G$. ∎
+5.3. Pure singularity
 Having established that $c(p)$ shares the limiting distribution of $\varphi(p-1)/(p-1)$, we now investigate the fine structure of this law.
-By the Jessen–Wintner theorem [24], any limiting distribution arising from an infinite convolution of discrete measures must be of pure type—purely discrete, purely singular, or absolutely continuous.
-Continuity was already verified in Lemma 3, so the remaining question is whether the law is absolutely continuous or purely singular.
-The latter turns out to hold: the values of $c(p)$ cluster on a fractal-like set of Lebesgue measure zero, as we prove by adapting a classical argument of Erdős to the shifted-prime setting.
-Lemma 6 (Shifted-prime pure-type and continuity criterion).
-Let $f$ be a real strongly additive arithmetical function satisfying the three convergence conditions of the shifted-prime Erdős–Wintner theorem, so that f(p1)f(p-1) possesses a limiting distribution $\mu_{f}$ over the primes.
-Then $\mu_{f}$ is of pure type: it is either purely discrete, purely singular, or absolutely continuous with respect to Lebesgue measure.
-Moreover, $\mu_{f}$ is continuous (i.e.
-has no atoms) if and only if $\sum_{f(\ell)\neq 0}1/\ell=\infty$.
+Note.
+The strictly stronger result $\dim_{H}(\mu_{G})=0$—that is, the limiting measure is carried by a Borel set of Hausdorff dimension zero—is established independently in Section 6.1 by a different method (a Borel–Cantelli covering argument exploiting the independence structure of the Bernoulli convolution).
+That result implies pure singularity as an immediate corollary.
+We include the Erdős-style argument for its independent methodological interest: it provides the first complete adaptation of Erdős’s classical technique [10] to the shifted-prime setting.
+The shifted-prime Erdős–Wintner theorem (Theorem 5.9) gives the existence and characteristic function of the limit law for any additive function on $\{p-1\}$ satisfying three convergence conditions; for strongly additive functions it also yields that the law is of pure type and identifies a continuity criterion.
+Since Lemma 5.1 verifies the hypotheses for $f=\log(n/\varphi(n))$ and the continuity criterion holds, the remaining question is whether the law is absolutely continuous or purely singular.
+The latter turns out to hold: we prove pure singularity by adapting a classical argument of Erdős to the shifted-prime setting.
+Theorem 5.9 (Shifted-prime Erdős–Wintner theorem; Hildebrand [16], Jessen–Wintner [18]).
+Let $f$ be a real-valued additive arithmetical function. The following are equivalent:
+  1. (i)
+    The distribution functions
+    $\frac{1}{\pi(x)}\,\#\bigl\{p\leq x:p\textup{ prime},\;f(p-1)\leq z\bigr\}$
+    converge weakly, as $x\to\infty$, to a distribution function $G(z)$.
+  2. (ii)
+    The three series
+    (5.2)  $\sum_{\begin{subarray}{c}\ell\textup{ prime}\\ |f(\ell)|>1\end{subarray}}\frac{1}{\ell},\qquad\sum_{\ell\textup{ prime}}\frac{f^{*}(\ell)^{2}}{\ell},\qquad\sum_{\ell\textup{ prime}}\frac{f^{*}(\ell)}{\ell}$
+    all converge, where $f^{*}(n):=f(n)\mathbf{1}_{|f(n)|\leq 1}$.
+When these conditions hold, the characteristic function of the limit law is
+(5.3) $\widehat{G}(\tau)\;=\;\prod_{\ell\textup{ prime}}\biggl(1-\frac{1}{\ell-1}+\sum_{\nu\geq 1}\frac{e^{i\tau f(\ell^{\nu})}}{\ell^{\nu}}\biggr).$
+Furthermore, if $f$ is strongly additive (i.e. $f(\ell^{\nu})=f(\ell)$ for every prime $\ell$ and every $\nu\geq 1$):
+  1. (a)
+    (Pure type.) The limit law is of pure type: purely discrete, purely singular, or absolutely continuous with respect to Lebesgue measure.
+  2. (b)
+    (Continuity criterion.) The limit law is continuous (i.e. has no atoms) if and only if $\sum_{f(\ell)\neq 0}1/\ell=\infty$.
 Proof.
-The relevant shifted-prime Erdős–Wintner framework is due to Kátai [25, Theorem 2] and Hildebrand [23, Theorem 5] for the shift $a=1$; the methods extend to any fixed nonzero shift via the Bombieri–Vinogradov theorem [4] (see also [39, Theorem II.8.34]): the only property of the shift used in these proofs is the Bombieri–Vinogradov estimate for primes in arithmetic progressions, which holds uniformly for any fixed nonzero $a$.
-Hildebrand [23, p.
-211] notes explicitly that the same methods apply to $\{p+a\}$ for arbitrary nonzero $a$.
-For the shift a=1a=-1 used here, the explicit characteristic function is recorded in Tenenbaum [38, formula (3), p.
-262].
-The same conclusions hold for general (not necessarily strongly) additive functions under the Erdős–Wintner hypotheses, by the identical Jessen–Wintner argument applied to the full local factors; we state the strongly additive case since it is all that we require.
-By Tenenbaum [38, formula (3), p. 262], the characteristic function of $\mu_{f}$ is given by
-$\widehat{\mu_{f}}(\tau)\;=\;\prod_{\ell\textup{ prime}}\!\left(1-\frac{1}{\ell-1}+\sum_{\nu\geq 1}\frac{e^{i\tau f(\ell^{\nu})}}{\ell^{\nu}}\right).$
-The shift is p1p-1, corresponding to a=1a=-1 in the notation of [38, formula (3)].
-Since no prime divides 1-1, the product runs over all primes.
-For the shift a=1a=-1, the local factor at $\ell=2$ in the characteristic function reduces, when $f$ is strongly additive, to $1-\frac{1}{2-1}+\sum_{\nu\geq 1}\frac{e^{i\tau f(2)}}{2^{\nu}}=e^{i\tau f(2)}$, which is a deterministic translation by $f(2)$.
-This factor carries no randomness and does not affect pure type or continuity; it merely shifts the law.
-For each odd prime $\ell$, the remaining local factor is the characteristic function of an atomic probability measure $\nu_{\ell}$ placing mass $(\ell-2)/(\ell-1)$ at the origin and mass $1/(\ell-1)$ at $f(\ell)$.
-The product formula therefore represents the characteristic function of the infinite convolution $\delta_{f(2)}\ast\mathop{{\ast}}_{\ell\geq 3}\nu_{\ell}$.
-By Lévy’s uniqueness theorem (see, e.g., [39, Section III.2]), $\mu_{f}$ equals this infinite convolution.
-It therefore suffices to apply the Jessen–Wintner theorem (see also Tenenbaum [39, Theorem III.2.7(b)]) to the convolution over the odd primes.
-Each factor $\nu_{\ell}$ is supported on a countable (indeed two-element) set, so the Jessen–Wintner theorem gives that the resulting infinite convolution—and hence $\mu_{f}$ itself—is of pure type.
-The continuity criterion follows from the same theorem: the convolution is continuous if and only if $\sum(1-\sigma_{\ell})=\infty$, where $\sigma_{\ell}$ is the maximal atom of $\nu_{\ell}$.
-Since $\nu_{\ell}$ places mass $(\ell-2)/(\ell-1)$ at the origin and mass $1/(\ell-1)$ at $f(\ell)$, when $f(\ell)\neq 0$ these two points are distinct, so $\sigma_{\ell}=(\ell-2)/(\ell-1)$ and $1-\sigma_{\ell}=1/(\ell-1)\asymp 1/\ell$.
-It follows that  $\sum_{\ell\text{ odd}}(1-\sigma_{\ell})$ diverges if and only if $\sum_{f(\ell)\neq 0}1/\ell=\infty$, since the prime $\ell=2$ does not enter the odd-prime convolution.
-Since a deterministic translation preserves pure type and continuity, the same conclusions hold for $\mu_{f}$.
+[…]
+[…]
+Concretely, Hildebrand’s Theorem 5 [16] is stated for the distribution functions $\nu_{x}(p;\,f(p+1)\leq z):=(1/\pi(x))\#\{p\leq x:f(p+1)\leq z\}$.
+Replacing $p+1$ by p1p-1 throughout changes only the residue class in the Bombieri–Vinogradov input:
+$\pi(x;q,-1)\;\longrightarrow\;\pi(x;q,1),$
+for which the same level of distribution holds (see, e.g., [16]). The resulting theorem for $\{p-1\}$ is therefore identical in statement and proof, and no new estimates are required.
+The explicit characteristic function (5.3) is recorded by Tenenbaum [32, formula (3)] for a general nonzero shift $a$, confirming the Euler-product form directly.
+For a=1a=-1 no prime divides $a$, so the product in (5.3) runs over all primes $\ell$ without restriction.
+Pure type and continuity (strongly additive case).
+Assume now that $f$ is strongly additive; then $f(\ell^{\nu})=f(\ell)$ for all $\nu\geq 1$, so each local factor in (5.3) reduces to the characteristic function of an atomic probability measure supported on $\{0,f(\ell)\}$.
+(For $\ell=2$ the weight at 0 vanishes and the factor degenerates to the point mass $\delta_{f(2)}$; for odd $\ell$ it is a genuine two-point measure.) By Lévy’s uniqueness theorem (see, e.g., [33, Section III.2]), the limit law equals the infinite convolution of these local factors.
+The Jessen–Wintner theorem [18] (see also [33, Theorem III.2.7(b)]) states that every infinite convolution of discrete probability measures is of pure type, giving (a).
+For the continuity criterion (b), since $f$ is strongly additive each local factor reduces to $\frac{\ell-2}{\ell-1}\,\delta_{0}+\frac{1}{\ell-1}\,\delta_{f(\ell)}$, a two-point measure.
+Let $\sigma_{\ell}$ denote its maximal atom.
+The Jessen–Wintner theorem gives: the convolution is continuous if and only if $\sum_{\ell}(1-\sigma_{\ell})=\infty$.
+When $f(\ell)\neq 0$, the atoms at 0 and at $f(\ell)$ are distinct, so $\sigma_{\ell}=\max\!\bigl(\frac{\ell-2}{\ell-1},\,\frac{1}{\ell-1}\bigr)=\frac{\ell-2}{\ell-1}$ for $\ell\geq 3$, whence $1-\sigma_{\ell}=1/(\ell-1)\asymp 1/\ell$.
+When $f(\ell)=0$, the two-point measure collapses to $\delta_{0}$ and $1-\sigma_{\ell}=0$.
+Therefore $\sum_{\ell}(1-\sigma_{\ell})$ diverges if and only if $\sum_{f(\ell)\neq 0}1/\ell=\infty$.
 ∎
@@
 and write $P^{+}(n)$ for the largest prime factor of $n$.
-Lemma 7 (Prime density for a prescribed smooth kernel).
+Lemma 5.10 (Prime density for a prescribed smooth kernel).
 Let $m$ be an even squarefree positive integer with $P^{+}(m)\leq y$. The set $\{p\textup{ prime}:a_{y}(p-1)=m\}$ has relative density
@@
 Moreover, writing
-$C_{y}:=\prod_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\textup{ prime}\end{subarray}}\frac{\ell-2}{\ell-1}\;\asymp\;\frac{1}{\log y},$
-we have $\delta_{y}(m)\geq C_{y}/\varphi(m)$.
+$Q(y):=\prod_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\textup{ prime}\end{subarray}}\frac{\ell-2}{\ell-1}\;\asymp\;\frac{1}{\log y},$
+we have $\delta_{y}(m)\geq Q(y)/\varphi(m)$.
 Proof.
@@
 The finitely many primes $p\leq y$ need not satisfy these residue-class conditions, but they contribute $O_{y}(1)$ to the count and do not affect the relative density.
-Since $M_{y}$ is fixed for each $y$, the Siegel–Walfisz theorem (Theorem 3) gives the stated density.
+Since $M_{y}$ is fixed for each $y$, the Siegel–Walfisz theorem (Theorem 5.3) gives the stated density.
 Since each $m$ is even and squarefree, $\varphi(m)=\prod_{\begin{subarray}{c}\ell\mid m\\ \ell>2\end{subarray}}(\ell-1)$, so the first product in the expression for $\delta_{y}(m)$ equals $1/\varphi(m)$.
 The lower bound follows because each factor $(\ell-2)/(\ell-1)\in(0,1)$.
-To justify the asymptotic $C_{y}\asymp 1/\!\log y$, write
-$\log C_{y}=\sum_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\text{ prime}\end{subarray}}\log\frac{\ell-1}{\ell-2}=\sum_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\text{ prime}\end{subarray}}\frac{1}{\ell}+O(1),$
+To justify the asymptotic $Q(y)\asymp 1/\!\log y$, write
+$\log Q(y)=\sum_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\text{ prime}\end{subarray}}\log\frac{\ell-1}{\ell-2}=\sum_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\text{ prime}\end{subarray}}\frac{1}{\ell}+O(1),$
 since $\log((\ell-1)/(\ell-2))=1/\ell+O(1/\ell^{2})$ and $\sum_{\ell}1/\ell^{2}<\infty$.
-By Mertens’ second theorem [29, Theorem 2.7(d)] the sum over primes equals $\log\log y+O(1)$, so $C_{y}=e^{-\log\log y+O(1)}\asymp 1/\!\log y$.
+By Mertens’ second theorem [24, Theorem 2.7(d)] the sum over primes equals $\log\log y+O(1)$, so $Q(y)=e^{-\log\log y+O(1)}\asymp 1/\!\log y$.
 ∎
-Lemma 8 (Tail-mean estimate for f(p1)f(p-1)).
-Let $f$ be as in (2), and for each prime $p$ and parameter $y\geq 3$ define
+Lemma 5.11 (Tail-mean estimate for f(p1)f(p-1)).
+Let $f$ be as in (5.1), and for each prime $p$ and parameter $y\geq 3$ define
 $T_{y}(p):=\sum_{\begin{subarray}{c}\ell\mid(p-1)\\ \ell>y,\;\ell\textup{ prime}\end{subarray}}\log\frac{\ell}{\ell-1}.$
@@
 Fix $A:=2$ and split the sum at the threshold $L:=(\log x)^{A}$.
-Small primes ($y<\ell\leq L$). Each such $\ell$ satisfies $\ell\leq(\log x)^{A}$, so by the Siegel–Walfisz theorem (Theorem 3),
+Small primes ($y<\ell\leq L$). Each such $\ell$ satisfies $\ell\leq(\log x)^{A}$, so by the Siegel–Walfisz theorem (Theorem 5.3),
 $\pi(x;\ell,1)\;=\;\frac{\operatorname{Li}(x)}{\ell-1}\;+\;E(x;\ell,1),\qquad|E(x;\ell,1)|\;\ll_{A}\;xe^{-c_{A}\sqrt{\log x}}.$
@@
 Large primes ($\ell>L$). For every prime $\ell\leq x$, the trivial bound $\pi(x;\ell,1)\leq x/\ell+1$ and $f(\ell)=\log(\ell/(\ell-1))\leq 2/\ell$ give
-$\sum_{\ell>L}f(\ell)\,\pi(x;\ell,1)$ $\;\leq\;\sum_{L<\ell\leq x}\frac{2}{\ell}\left(\frac{x}{\ell}+1\right)\;\leq\;2x\sum_{\ell>L}\frac{1}{\ell^{2}}\;+\;2\sum_{\ell\leq x}\frac{1}{\ell}\;\ll\;\frac{x}{L}\;+\;\log\log x.$
+$\sum_{\ell>L}f(\ell)\,\pi(x;\ell,1)$ $\;\leq\;\sum_{L<\ell\leq x}\frac{2}{\ell}\left(\frac{x}{\ell}+1\right)\;\leq\;2x\sum_{\ell>L}\frac{1}{\ell^{2}}\;+\;2\sum_{\ell\leq x}\frac{1}{\ell}$
+$\;\ll\;\frac{x}{L}\;+\;\log\log x.$
 Dividing by $\pi(x)\sim x/\!\log x$, this contributes $O(\log x/L)+o(1)=o(1)$. Combining both ranges and dividing by $\pi(x)$ yields the claim. ∎
-Lemma 9 (Shifted-prime pure singularity for $\log(n/\varphi(n))$).
-Let $f$ be the strongly additive function defined in (2).
+Lemma 5.12 (Shifted-prime pure singularity for $\log(n/\varphi(n))$).
+Let $f$ be the strongly additive function defined in (5.1).
 The shifted-prime limiting measure $\mu_{f}$ is purely singular with respect to Lebesgue measure.
@@
 Proof.
-We adapt the integer-case argument of Erdős [14, pp.
-722–725] (see also [39, Exercise 256] with solution in [40, Solution 2.4, pp.
-147–148]) to the shifted-prime setting, using the Siegel–Walfisz theorem (Theorem 3) in place of the integer equidistribution input.
-By the shifted-prime Erdős–Wintner theorem (Lemma 6), the limiting measure $\mu_{f}$ exists.
-Indeed, by Lemma 3 the three convergence hypotheses are satisfied and the continuity criterion holds, so the limiting measure $\mu_{f}$ exists and is continuous.
-By Lemma 6, $\mu_{f}$ is of pure type.
-Since it is continuous (as verified above), it is therefore either purely singular or absolutely continuous.
-We show that $\mu_{f}$ is not absolutely continuous, which forces pure singularity.
-Step 1: Tail estimate.
-Throughout the proof, the single prime $p=2$ is negligible in any asymptotic density statement.
-Fix a parameter $\delta\in(0,\tfrac{1}{2})$ and, for $y\geq 3$, set $R:=\lfloor y^{\delta}\rfloor$.
-Since $f$ is strongly additive, for each prime $p$ we decompose $f(p-1)=f(a_{y}(p-1))+T_{y}(p)$, where $a_{y}(p-1)$ and $T_{y}(p)$ are as in Lemma 8.
-By that lemma and Markov’s inequality,
-$\limsup_{x\to\infty}\;\frac{1}{\pi(x)}\#\bigl\{p\leq x:T_{y}(p)>1/R^{2}\bigr\}\;\leq\;R^{2}\,\eta(y)\;\ll\;y^{2\delta-1},$ (3)
-which tends to 0 as $y\to\infty$ (since $\delta<\tfrac{1}{2}$).
-Here and in Steps 2–3, the parameter $y$ is held fixed while $x\to\infty$; the conclusions are then strengthened by letting $y\to\infty$.
-Step 2: A set of small Lebesgue measure carrying positive relative prime density.
-Let $m_{1}<m_{2}<\cdots<m_{R}$ be the first $R=\lfloor y^{\delta}\rfloor$ even squarefree positive integers.
-Since the counting function of even squarefree integers satisfies $A(t):=\#\{m\leq t:m\text{ even squarefree}\}\sim(2/\pi^{2})\,t$ (see, e.g., [39, Theorem I.3.10]), for any fixed $C>\pi^{2}/2$ we have $A(Cj)\geq j$ for all sufficiently large $j$.
-After enlarging $C$ if necessary, we may assume $m_{j}\leq Cj$ for all $j\geq 1$.
-In particular $m_{R}\leq CR$.
-Therefore $P^{+}(m_{j})\leq m_{j}\leq Cy^{\delta}<y$ for all sufficiently large $y$ (since $\delta<1$).
-Moreover,
-$\sum_{j=1}^{R}\frac{1}{\varphi(m_{j})}\;\geq\;\sum_{j=1}^{R}\frac{1}{m_{j}}\;\geq\;\frac{1}{C}\sum_{j=1}^{R}\frac{1}{j}\;\gg\;\log R\;=\;\delta\log y+O(1).$ (4)
-By Lemma 7, the relative density of $\{p:a_{y}(p-1)=m_{j}\}$ among the primes is $\delta_{y}(m_{j})\geq C_{y}/\varphi(m_{j})$, where $C_{y}\asymp 1/\!\log y$. Combining with (4),
-$\sum_{j=1}^{R}\delta_{y}(m_{j})\;\geq\;C_{y}\sum_{j=1}^{R}\frac{1}{\varphi(m_{j})}\;\gg\;\frac{\delta\log y}{\log y}\;=\;\delta\;>\;0.$
-Since the $m_{j}$ are distinct, the sets $\{p:a_{y}(p-1)=m_{j}\}$ are pairwise disjoint, so the relative density of their union $\mathcal{A}_{y}:=\{p:a_{y}(p-1)\in\{m_{1},\ldots,m_{R}\}\}$ is the sum $\sum_{j}\delta_{y}(m_{j})$.
-In particular, there exists an absolute constant $c_{0}>0$ (depending only on $\delta$) such that for all sufficiently large $y$, the relative density of $\mathcal{A}_{y}$ among the primes is at least $c_{0}$.
-Define the closed set
-$\bar{I}_{y}\;:=\;\bigcup_{j=1}^{R}\bigl[\,f(m_{j}),\;f(m_{j})+1/R^{2}\,\bigr].$
-By subadditivity of Lebesgue measure, $\lambda(\bar{I}_{y})\leq R\cdot(1/R^{2})=1/R=1/\lfloor y^{\delta}\rfloor\to 0$ as $y\to\infty$.
-For any prime $p$ with $a_{y}(p-1)=m_{j}$ and $T_{y}(p)\leq 1/R^{2}$, the identity $f(p-1)=f(a_{y}(p-1))+T_{y}(p)$ and the nonnegativity of $T_{y}(p)$ give $f(p-1)\in[f(m_{j}),\,f(m_{j})+1/R^{2}]\subset\bar{I}_{y}$.
-Step 3: Contradiction with absolute continuity. Recall that $y$ (and hence $R$) are still fixed. By (3),
-$\limsup_{x\to\infty}\;\frac{1}{\pi(x)}\#\bigl\{p\leq x:T_{y}(p)>1/R^{2}\bigr\}\;=\;o_{y}(1).$
-Hence, for all sufficiently large $y$,
-$\liminf_{x\to\infty}\frac{1}{\pi(x)}\#\bigl\{p\leq x:f(p-1)\in\bar{I}_{y}\bigr\}\;\geq\;c_{0}-o_{y}(1)\;\geq\;\frac{c_{0}}{2}\;>\;0.$
-[…]
-$\mu_{f}(\bar{I}_{y})\;=\;\lim_{x\to\infty}\frac{1}{\pi(x)}\#\bigl\{p\leq x:f(p-1)\in\bar{I}_{y}\bigr\}\;\geq\;\frac{c_{0}}{2}\;>\;0.$
-If $\mu_{f}$ were absolutely continuous with density $h\in L^{1}(\mathbb{R})$, then by the absolute continuity of the Lebesgue integral, $\mu_{f}(\bar{I}_{y})=\int_{\bar{I}_{y}}h\,\mathrm{d}t\to 0$ as $\lambda(\bar{I}_{y})\to 0$, contradicting $\mu_{f}(\bar{I}_{y})\geq c_{0}/2$ for all large $y$.
-Therefore $\mu_{f}$ is not absolutely continuous, and being of pure type and continuous, it is purely singular.
+We adapt the argument of Erdős [10, §3] (see also [34, Exercise 2.4]) to the shifted-prime setting, replacing the integer equidistribution input with the Siegel–Walfisz theorem (Theorem 5.3).
+Throughout, $y$ is fixed while $x\to\infty$; the moduli arising from the construction divide $M_{y}:=\prod_{\ell\leq y}\ell$ and hence lie in the Siegel–Walfisz range.
+Only after taking $x\to\infty$ do we let $y\to\infty$.
+By Lemma 5.1 and Theorem 5.9, $\mu_{f}$ exists, is continuous, and of pure type. It remains to rule out absolute continuity.
+Step 1: Tail estimate. Fix $\delta\in(0,\tfrac{1}{2})$ and set $R:=\lfloor y^{\delta}\rfloor$. Decompose $f(p-1)=f(a_{y}(p-1))+T_{y}(p)$. By Lemma 5.11 and Markov’s inequality,
+(5.4) $\limsup_{x\to\infty}\;\frac{1}{\pi(x)}\#\bigl\{p\leq x:T_{y}(p)>1/R^{2}\bigr\}\;\leq\;R^{2}\,\eta(y)\;\ll\;y^{2\delta-1}\;\xrightarrow{y\to\infty}\;0.$
+Step 2: Small-measure set with positive prime density.
+Let $m_{1}<\cdots<m_{R}$ be the smallest $R$ even squarefree positive integers.
+Since even squarefree integers have positive asymptotic density (equal to $2/\pi^{2}$), there is an absolute constant $C>0$ with $m_{j}\leq Cj$; in particular $m_{j}\leq CR=C\lfloor y^{\delta}\rfloor\ll y^{1/2}$ since $\delta<\tfrac{1}{2}$, so $P^{+}(m_{j})\leq m_{j}<y$ for all sufficiently large $y$.
+Because $\varphi(m_{j})\leq m_{j}\leq Cj$, we have $\sum_{j\leq R}1/\varphi(m_{j})\geq C^{-1}\sum_{j\leq R}1/j\gg\log R$.
+By Lemma 5.10, $\delta_{y}(m_{j})\geq Q(y)/\varphi(m_{j})$ with $Q(y)\asymp 1/\!\log y$, so
+(5.5) $\sum_{j=1}^{R}\delta_{y}(m_{j})\;\geq\;Q(y)\sum_{j=1}^{R}\frac{1}{\varphi(m_{j})}\;\gg\;\frac{\log R}{\log y}\;=\;\delta+o(1)\;\gg_{\delta}\;1$
+for all sufficiently large $y$, where the implied constants are absolute (from Mertens’ theorem and the density of even squarefree integers) and $\log R/\!\log y=\delta+o(1)$ because $R=\lfloor y^{\delta}\rfloor$.
+Define the closed set $\bar{I}_{y}:=\bigcup_{j=1}^{R}[f(m_{j}),\,f(m_{j})+1/R^{2}]$.
+Then $\lambda(\bar{I}_{y})\leq 1/R\to 0$ as $y\to\infty$, while for any $p$ with $a_{y}(p-1)=m_{j}$ and $T_{y}(p)\leq 1/R^{2}$ one has $f(p-1)\in\bar{I}_{y}$.
+Step 3: Contradiction.
+Since $a_{y}(p-1)$ takes a unique value for each prime $p$, the sets $\{p:a_{y}(p-1)=m_{j}\}$ are pairwise disjoint.
+Finite additivity of the natural density therefore gives $\lim_{x\to\infty}\frac{1}{\pi(x)}\#\{p\leq x:a_{y}(p-1)\in\{m_{1},\dots,m_{R}\}\}=\sum_{j\leq R}\delta_{y}(m_{j})$.
+Removing the primes for which $T_{y}(p)>1/R^{2}$ loses at most the quantity in (5.4), so combining with (5.5): there exists $c_{0}>0$ independent of $y$ such that for all large $y$,
+$\mu_{f}(\bar{I}_{y})\;=\;\lim_{x\to\infty}\frac{1}{\pi(x)}\#\bigl\{p\leq x:f(p-1)\in\bar{I}_{y}\bigr\}\;\geq\;c_{0}\;>\;0,$
+the limit being justified by the Portmanteau theorem (see, e.g., [2, Theorem 2.1]): $\bar{I}_{y}$ is a $\mu_{f}$-continuity set because its boundary consists of at most $2R$ points and $\mu_{f}$ is atom-free (Theorem 5.9(b)), so weak convergence of the empirical measures to $\mu_{f}$ gives $\mu_{x}(\bar{I}_{y})\to\mu_{f}(\bar{I}_{y})$.
+Since $\lambda(\bar{I}_{y})\to 0$ while $\mu_{f}(\bar{I}_{y})\geq c_{0}$, the measure $\mu_{f}$ cannot be absolutely continuous.
+Being continuous and of pure type, it is purely singular.
 ∎
-Proposition 3 (Shifted-prime limiting law for $f$).
-Let $f$ be as in (2). Then the empirical measures
+Proposition 5.13 (Shifted-prime limiting law for $f$).
+Let $f$ be as in (5.1). Then the empirical measures
 $\mu_{x}:=\frac{1}{\pi(x)}\sum_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}\delta_{f(p-1)}$
@@
 Proof.
-By Lemma 3, the three convergence conditions of the shifted-prime Erdős–Wintner theorem (Lemma 6) are satisfied, and the continuity criterion holds.
-Therefore f(p1)f(p-1) possesses a continuous limiting distribution $\mu_{f}$, which is of pure type by Lemma 6.
-It remains to rule out absolute continuity. By Lemma 9, $\mu_{f}$ is purely singular. ∎
-Corollary 2 (Structure of the limiting PRIM-LWE law).
-The density $c(p)$ has a continuous, purely singular limiting distribution with support exactly $0,\tfrac{1}{2}]$. Its distribution function $G$ is strictly increasing on $0,\tfrac{1}{2}]$.
+By Lemma 5.1, the three convergence conditions of the shifted-prime Erdős–Wintner theorem (Theorem 5.9) are satisfied, and the continuity criterion holds.
+Therefore f(p1)f(p-1) possesses a continuous limiting distribution $\mu_{f}$, which is of pure type by Theorem 5.9.
+It remains to rule out absolute continuity. By Lemma 5.12, $\mu_{f}$ is purely singular. ∎
+Corollary 5.14 (Structure of the limiting law for $c(p)$).
+The ratio $c(p)$ has a continuous, purely singular limiting distribution with support exactly $0,\tfrac{1}{2}]$. Its distribution function $G$ is strictly increasing on $0,\tfrac{1}{2}]$.
 Proof.
-Combine Theorems 4 and 5: $c(p)$ has a continuous limiting law with support $0,\tfrac{1}{2}]$ whose distribution function is strictly increasing. It remains to justify pure singularity.
-By Proposition 3, the shifted-prime law $\mu_{f}$ of f(p1)f(p-1) (with $f$ as in (2)) is continuous and purely singular.
-Since $\varphi(p-1)/(p-1)=e^{-f(p-1)}$ and $T(u):=e^{-u}$ is locally Lipschitz on $0,\infty)$, the push-forward $T_{\#}\mu_{f}$ preserves Lebesgue-null sets and is therefore purely singular.
-By Theorem 5, the same holds for $c(p)$.
+Combine Theorems 5.5 and 5.8: $c(p)$ has a continuous limiting law with support $0,\tfrac{1}{2}]$ whose distribution function is strictly increasing. It remains to justify pure singularity.
+By Proposition 5.13, the shifted-prime law $\mu_{f}$ of f(p1)f(p-1) (with $f$ as in (5.1)) is continuous and purely singular.
+Since $\varphi(p-1)/(p-1)=e^{-f(p-1)}$ and $T(u):=e^{-u}$ is Lipschitz on $0,\infty)$ (since $|T^{\prime}(u)|=e^{-u}\leq 1$), the push-forward $T_{\#}\mu_{f}$ preserves Lebesgue-null sets.
+Moreover, $T$ is injective on $\log 2,\infty)$ and $\mu_{f}$ concentrates on this half-line (since $Y\geq\log 2$ a.s.), so the carrying null set of $\mu_{f}$ maps bijectively to a null set carrying $T_{\#}\mu_{f}$.
+Hence $T_{\#}\mu_{f}$ is purely singular.
+Since Theorem 5.8 shows that $c(p)$ has the same limiting distribution $G$ as $\varphi(p-1)/(p-1)$, the limiting measure for $c(p)$ is exactly the same measure $\mu_{G}$, and pure singularity carries over.
 ∎
-Figure 1 compares the empirical distribution of $c(p)$ for primes $p\leq 10^{6}$ with the limiting law $G$ of Theorem 5; the close agreement already visible at this scale provides a useful numerical check on the distributional theory.
-Figure 1: Empirical CDF of $c(p)$ over the 78,49778{,}497 primes $3\leq p\leq 10^{6}$, overlaid with the theoretical limiting distribution function $G$ of Theorem 5, estimated by $10^{7}$ Monte Carlo samples of the infinite convolution of Lemma 10 truncated at the 10,00010{,}000th odd prime.
-The empirical values of $c(p)$ were computed via the truncated Euler product of Proposition 8 with $J=20$ terms, contributing error below $10^{-9}$ for every prime $p\geq 3$.
-Remark 1.
+The strictly stronger metric result $\dim_{H}(\mu_{G})=0$ is established in Theorem 6.2 by an independent method (a Borel–Cantelli covering argument exploiting the Bernoulli-convolution independence structure).
+Figure 1 compares the empirical distribution of $c(p)$ for primes $p\leq 10^{6}$ with the limiting law $G$ of Theorem 5.8; the close agreement already visible at this scale provides a useful numerical check on the distributional theory.
+Remark 5.15 (Truncation error for numerical approximation).
+Truncate the Bernoulli series of Lemma 6.1 at the first $N$ odd primes ($\ell\leq\ell_{N}$) and let $G_{N}$ denote the distribution function of the truncated law $X_{N}:=\tfrac{1}{2}\prod_{3\leq\ell\leq\ell_{N}}(1-1/\ell)^{B_{\ell}}$.
+Since $|X-X_{N}|\leq\tfrac{1}{2}T_{N}$ almost surely, where $T_{N}:=\sum_{\ell>\ell_{N},\,\ell\textup{ prime}}B_{\ell}\log(\ell/(\ell-1))$, the coupling characterisation of the Wasserstein-1 distance gives
+$W_{1}(G_{N},G)\;\leq\;\tfrac{1}{2}\,\mathbb{E}[T_{N}]\;=\;\tfrac{1}{2}\sum_{\begin{subarray}{c}\ell>\ell_{N}\\ \ell\textup{ prime}\end{subarray}}\frac{1}{\ell-1}\log\frac{\ell}{\ell-1}\,.$
+For N=10, [^0] N=10{,}000 (so $\ell_{N}=104{,}743$), $\mathbb{E}[T_{N}]<10^{-5}$ and therefore $W_{1}(G_{N},G)<5\times 10^{-6}$.
+Since $G_{N}$ has an atom of mass $Q(\ell_{N})\approx 0.064$ at $\tfrac{1}{2}$—where $Q(y)$ is the product defined in Lemma 5.10, equal to the probability that every Bernoulli indicator up to $\ell_{N}$ equals zero—while $G$ is continuous, the truncation error is controlled in $W_{1}$, not in the Kolmogorov metric.
+The plotted theoretical curve is a Monte Carlo approximation to $G_{N}$; the DKW inequality controls the sampling error (see Figure 1).
+Figure 1.
+Empirical CDF of $c(p)$ over the 78,49778{,}497 primes $3\leq p\leq 10^{6}$, overlaid with the theoretical CDF $G$ of Theorem 5.8, approximated by $10^{7}$ Monte Carlo samples of the truncated convolution $G_{N}$ with N=10, [^0] N=10{,}000 odd primes.
+See Remark 5.15 for truncation and sampling error bounds.
+Empirical $c(p)$ values use Remark 7.9 with $J=20$.
+Remark 5.16 (The value $1/2$: boundary but not attained).
 The value $1/2$ is not attained: $c(p)<1/2$ for every prime $p$.
-Indeed, for $p=2$ one has $\varphi(p-1)/(p-1)=1$, so $c(2)=\prod_{j\geq 1}(1-2^{-j})\approx 0.288788<1/2$.
+Indeed, for $p=2$ one has $\varphi(p-1)/(p-1)=1$, so $c(2)=\prod_{j\geq 1}(1-2^{-j})\approx 0.2888<1/2$.
 For odd $p$, one has $\varphi(p-1)/(p-1)\leq 1/2$ together with $\prod_{j\geq 1}(1-p^{-j})<1$.
-Thus $1/2$ is only a boundary point of the support of the limiting law for $c(p)$, not an attained value.
-Corollary 3 (Unconditional limsup).
-We have
-$\limsup_{\begin{subarray}{c}p\to\infty\\ p\textup{ prime}\end{subarray}}c(p)=\frac{1}{2}.$
-Moreover, for every $\varepsilon>0$, the set of primes with $c(p)>1/2-\varepsilon$ has relative density $1-G(1/2-\varepsilon)$ among the primes.
-Proof.
-Since $c(p)<1/2$ for every prime (Remark 1), $\limsup c(p)\leq 1/2$.
-For the reverse inequality, fix $\varepsilon>0$.
-By Corollary 2, the limiting distribution of $c(p)$ has support exactly $0,\tfrac{1}{2}]$, so the limiting measure assigns positive mass to the interval $(1/2-\varepsilon,\,1/2]$.
-Hence
-$\lim_{x\to\infty}\frac{1}{\pi(x)}\#\bigl\{p\leq x:c(p)>1/2-\varepsilon\bigr\}=1-G(1/2-\varepsilon)>0.$
-In particular, infinitely many primes satisfy $c(p)>1/2-\varepsilon$, so $\limsup c(p)\geq 1/2-\varepsilon$. Letting $\varepsilon\downarrow 0$ yields the claim. ∎
-Corollary 4 (Fixed-dimension analogue).
-For each fixed $n\geq 1$, the exact density $c_{n}(p)$ has the same continuous limiting distribution as $\varphi(p-1)/(p-1)$; in particular, its limiting law has support exactly $0,\tfrac{1}{2}]$.
-Moreover,
-$\min_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}c_{n}(p)\;\asymp\;\frac{1}{\log\log x}.$
-Proof.
-Write $c_{n}(p)=\frac{\varphi(p-1)}{p-1}P_{n}(p)$ with $P_{n}(p):=\prod_{j=1}^{n}(1-p^{-j})$.
-Since $P_{n}(p)\to 1$, Lemma 5 gives the distributional claim.
-For the lower bound on $\min_{p\leq x}c_{n}(p)$: every odd prime $p\leq x$ satisfies $\omega(p-1)\leq m(x)$, so
-$\frac{\varphi(p-1)}{p-1}\;\geq\;\prod_{i=1}^{m(x)}\Bigl(1-\frac{1}{p_{i}}\Bigr),$
-and $P_{n}(p)\geq 1/2$ by Lemma 1. By Lemma 2 and Mertens’ third theorem,
-$\prod_{i=1}^{m(x)}\Bigl(1-\frac{1}{p_{i}}\Bigr)\;\asymp\;\frac{1}{\log\log x}.$
-The upper bound is identical to Proposition 1, since $P_{n}(p)\leq 1$. ∎
-6 Hausdorff Dimension, Moments, and Fourier Structure of $G$
+Thus $1/2$ is only a boundary point of the support, not an attained value.
+Since the support of $G$ is exactly $0,\tfrac{1}{2}]$ (Corollary 5.14), it follows that $\limsup_{p\to\infty,\,p\textup{ prime}}c(p)=1/2$; for every $\varepsilon>0$, the set of primes with $c(p)>1/2-\varepsilon$ has relative density $1-G(1/2-\varepsilon)>0$ among the primes.
+Remark 5.17 (Fixed-dimension analogue).
+For each fixed $n\geq 1$, $c_{n}(p)=\frac{\varphi(p-1)}{p-1}P_{n}(p)$ with $P_{n}(p):=\prod_{j=1}^{n}(1-p^{-j})\to 1$, so Lemma 5.7 gives $c_{n}(p)$ the same limiting distribution as $\varphi(p-1)/(p-1)$, and the primorial/Mertens argument yields $\min_{p\leq x,\,p\textup{ prime}}c_{n}(p)\asymp 1/\!\log\log x$.
+6. Hausdorff Dimension, Moments, and Fourier Structure
 The purely singular limiting law established in Section 5 raises a natural question: how thin is the set carrying full mass of the limiting measure?
+Theorem 6.2 below provides the strictly stronger result $\dim_{H}(\mu_{f})=0$, giving in particular an independent proof of pure singularity.
+The Erdős-style singularity argument in Section 5.3 is included for its independent methodological interest.
 Because $\mu_{f}$ decomposes as an infinite convolution of discrete measures (one for each odd prime), one can exploit the resulting independence structure to build efficient coverings and bound the Hausdorff dimension.
-We show that $\dim_{H}(\mu_{f})=0$, meaning that the mass concentrates on a Borel set of Hausdorff dimension zero even though the topological support is the full interval $0,\tfrac{1}{2}]$; in particular, the fractal-like clustering visible in Figure 2 is not an artifact of finite sampling.
+We show that $\dim_{H}(\mu_{f})=0$, meaning that the mass concentrates on a Borel set of Hausdorff dimension zero even though the topological support of the push-forward $\mu_{G}$ is the full interval $0,\tfrac{1}{2}]$.
 We then derive an explicit Euler product for every moment of $G$, prove that these moments determine the distribution uniquely, and establish the analytic continuation of the associated Mellin transform.
-The section concludes with a precise asymptotic for the behavior of $G$ near the endpoint $1/2$.
-Throughout, $f$ is the strongly additive function defined in (2), $\mu_{f}$ is the shifted-prime limiting measure of f(p1)f(p-1), and $X$ is a random variable with distribution $G$.
-By Corollary 2, $X\overset{d}{=}e^{-Y}$ where $Y\sim\mu_{f}$.
-Lemma 10 (Factor structure of $\mu_{f}$).
+The section concludes with a precise asymptotic for the behaviour of $G$ near the endpoint $1/2$.
+Throughout, $f$ is the strongly additive function defined in (5.1), $\mu_{f}$ is the shifted-prime limiting measure of f(p1)f(p-1), and $X$ is a random variable with distribution $G$.
+By Corollary 5.14, $X\overset{d}{=}e^{-Y}$ where $Y\sim\mu_{f}$.
+The Bernoulli indicators $B_{\ell}$ and the scaled jumps $Z_{\ell}:=B_{\ell}\,w_{\ell}$ (with $w_{\ell}=\log(\ell/(\ell-1))$) are used interchangeably below; the former appears in the covering construction (Theorem 6.2), the latter in the convolution decomposition (Lemma 6.1).
+6.1. Convolution structure and Hausdorff dimension
+Lemma 6.1 (Factor structure of $\mu_{f}$).
 The measure $\mu_{f}$ is the infinite convolution
@@
 Proof.
-By Tenenbaum [38, formula (3), p. 262] for the shift a=1a=-1, the characteristic function of $\mu_{f}$ is
+By Theorem 5.9, the characteristic function of $\mu_{f}$ is given by (5.3):
 $\widehat{\mu_{f}}(\tau)=\prod_{\ell\text{ prime}}\Bigl(1-\frac{1}{\ell-1}+\sum_{\nu\geq 1}\frac{e^{i\tau f(\ell^{\nu})}}{\ell^{\nu}}\Bigr).$
@@
 ∎
-For a Borel probability measure $\mu$ on $\mathbb{R}$, recall the Hausdorff dimension of $\mu$ (see, e.g., Falconer [17, Chapter 2]):
+For a Borel probability measure $\mu$ on $\mathbb{R}$, recall the Hausdorff dimension of $\mu$ (see, e.g., Falconer [14, Chapter 2]):
 $\dim_{H}(\mu):=\inf\{\dim_{H}(E):E\subset\mathbb{R}\text{ is Borel and }\mu(E)=1\}.$
-Proposition 4 (Hausdorff dimension zero of the limiting law).
+Theorem 6.2 (Hausdorff dimension zero of the limiting law).
 $\dim_{H}(\mu_{f})=0$. Consequently, if $\mu_{G}$ denotes the limiting measure on $0,\tfrac{1}{2}]$ with distribution function $G$, then $\dim_{H}(\mu_{G})=0$.
 Proof.
-By Lemma 10,
+By Lemma 6.1,
 $Y\overset{d}{=}\log 2+\sum_{\begin{subarray}{c}\ell\geq 3\\ \ell\text{ prime}\end{subarray}}B_{\ell}w_{\ell},\qquad w_{\ell}:=\log\frac{\ell}{\ell-1},$
@@
 $\mathbb{P}(B_{\ell}=1)=\frac{1}{\ell-1},\qquad\mathbb{P}(B_{\ell}=0)=\frac{\ell-2}{\ell-1}.$
-Fix $s>0$. We shall show that $\mu_{f}$ gives full mass to a Borel set of vanishing $s$-dimensional Hausdorff measure.
-For $y\geq 3$, let
+Fix $s>0$.
+We shall show that $\mu_{f}$ gives full mass to a Borel set of vanishing $s$-dimensional Hausdorff measure.
+The argument proceeds in five steps: (1) construct recursively a sequence of super-exponentially growing covering scales; (2) bound the number of prefix configurations via Chebyshev’s inequality; (3) show that empty-block events occur infinitely often by the second Borel–Cantelli lemma; (4) control the tail contribution by Markov’s inequality; (5) assemble the covering and estimate its Hausdorff measure.
+Step 1: Recursive covering scales. For $y\geq 3$, let
 $I(y):=\sum_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\text{ prime}\end{subarray}}\left(B_{\ell}\log(\ell-1)+(1-B_{\ell})\log\frac{\ell-1}{\ell-2}\right).$
@@
 and
-$\operatorname{Var}(I(y))\leq\sum_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\text{ prime}\end{subarray}}\mathbb{E}\!\left[\left(B_{\ell}\log(\ell-1)+(1-B_{\ell})\log\frac{\ell-1}{\ell-2}\right)^{2}\right]\ll\sum_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\text{ prime}\end{subarray}}\frac{(\log\ell)^{2}}{\ell}\ll(\log y)^{2}.$
-Hence there exists an absolute constant $C_{0}>0$ such that
+$\operatorname{Var}(I(y))$ $\leq\sum_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\text{ prime}\end{subarray}}\mathbb{E}\!\left[\left(B_{\ell}\log(\ell-1)+(1-B_{\ell})\log\frac{\ell-1}{\ell-2}\right)^{\!2}\right]$
+$\ll\sum_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\text{ prime}\end{subarray}}\frac{(\log\ell)^{2}}{\ell}\;\ll\;(\log y)^{2},$
+where the final estimate follows by partial summation and the prime number theorem. Hence there exists an absolute constant $C_{0}>0$ such that
 $\mathbb{E}I(y)\leq C_{0}\log y,\qquad\operatorname{Var}(I(y))\leq C_{0}(\log y)^{2}\qquad(y\geq 3).$
-Choose $C:=4C_{0}$ (so that $Cn-a_{n}s/2\leq-Cn$ in the covering estimate below).
+Choose $C:=4C_{0}$ and define
 $a_{n}:=\left\lceil\frac{4Cn}{s}\right\rceil\qquad(\text{recall }s>0\text{ is fixed}),\qquad z_{n}:=y_{n}^{a_{n}},\qquad r_{n}:=z_{n}^{-1/2}.$
@@
 Then the intervals $(y_{n},z_{n}]$ are pairwise disjoint by construction, $z_{n}=y_{n}^{a_{n}}$ grows super-exponentially, and the summability conditions $\sum r_{n}<\infty$ and $\sum y_{n}^{-Cn}<\infty$ are immediate since $r_{n}=z_{n}^{-1/2}$ and $y_{n}^{-Cn}\leq y_{n}^{-C}$ with $y_{n}\to\infty$ super-exponentially.
-For each $n$, set $I_{n}:=I(y_{n})$ and
+Since $a_{n}\geq 4Cn/s$, it follows that
+[…]
+this inequality drives the convergence of the covering estimate below.
+Step 2: Concentration and cardinality of prefix configurations. For each $n$, set $I_{n}:=I(y_{n})$ and
 $A_{n}:=\{I_{n}\leq Cn\log y_{n}\}.$
@@
 $|T_{n}|\leq y_{n}^{Cn}.$
-Now define the empty-block event
+Step 3: Empty-block divergence (second Borel–Cantelli). Define the empty-block event
 $E_{n}:=\{B_{\ell}=0\text{ for every prime }\ell\in(y_{n},z_{n}]\}.$
@@
 $1-\frac{1}{\ell-1}=\left(1-\frac{1}{\ell}\right)\left(1-\frac{1}{(\ell-1)^{2}}\right),$
-Mertens’ third theorem and the convergence of $\prod_{\ell\geq 3}(1-(\ell-1)^{-2})$ imply
+Mertens’ third theorem and the convergence of $\prod_{\ell\geq 3}(1-(\ell-1)^{-2})$ imply that the product of $(1-1/\ell)$ over primes in $(y_{n},z_{n}]$ is asymptotic to $(\log y_{n})/(\log z_{n})=1/a_{n}$, since $\log z_{n}=a_{n}\log y_{n}$.
+Therefore
 $\mathbb{P}(E_{n})\asymp\frac{\log y_{n}}{\log z_{n}}=\frac{1}{a_{n}}\asymp\frac{1}{n}.$
 Therefore $\sum_{n}\mathbb{P}(E_{n})=\infty$.
-Since the random variables $B_{\ell}$ are mutually independent (Lemma 10) and the blocks $(y_{n},z_{n}]$ are pairwise disjoint, the events $E_{n}$ are mutually independent.
+Since the random variables $B_{\ell}$ are mutually independent (Lemma 6.1) and the blocks $(y_{n},z_{n}]$ are pairwise disjoint, the events $E_{n}$ are mutually independent.
 The second Borel–Cantelli lemma therefore gives
 $\mathbb{P}(E_{n}\text{ i.o.})=1.$
-Let
+Step 4: Tail control. Let
 $R_{n}:=\sum_{\begin{subarray}{c}\ell>z_{n}\\ \ell\text{ prime}\end{subarray}}B_{\ell}w_{\ell},\qquad F_{n}:=\{R_{n}\leq r_{n}\}.$
@@
 $\mathbb{P}(F_{n}^{c}\text{ i.o.})=0.$
-For each $\boldsymbol{b}=(b_{\ell})_{3\leq\ell\leq y_{n}}\in T_{n}$, define
+Step 5: Covering and Hausdorff measure estimate. For each $\boldsymbol{b}=(b_{\ell})_{3\leq\ell\leq y_{n}}\in T_{n}$, define
 $v(\boldsymbol{b}):=\log 2+\sum_{\begin{subarray}{c}3\leq\ell\leq y_{n}\\ \ell\text{ prime}\end{subarray}}b_{\ell}w_{\ell},$
@@
 $\sum_{n}\mathbb{P}(A_{n}^{c})<\infty,\qquad\sum_{n}\mathbb{P}(F_{n}^{c})<\infty,\qquad\mathbb{P}(E_{n}\text{ i.o.})=1,$
-it follows that
+the first Borel–Cantelli lemma implies that $A_{n}$ and $F_{n}$ each hold for all sufficiently large $n$ almost surely, while the second Borel–Cantelli lemma (applicable because the $E_{n}$ are independent) gives $E_{n}$ infinitely often almost surely.
+Hence $A_{n}\cap E_{n}\cap F_{n}$ occurs infinitely often almost surely, and on each such occurrence $Y\in U_{n}$.
+Therefore
 $\mu_{f}\!\left(\limsup_{n\to\infty}U_{n}\right)=1.$
@@
 $\sum_{n\geq 1}|T_{n}|\,r_{n}^{s}\ll\sum_{n\geq 1}y_{n}^{Cn}z_{n}^{-s/2}=\sum_{n\geq 1}y_{n}^{Cn-a_{n}s/2}\leq\sum_{n\geq 1}y_{n}^{-Cn}<\infty.$
-Because $r_{n}\downarrow 0$, for every $N\geq 1$ the set $\limsup_{n\to\infty}U_{n}$ is covered by the intervals comprising $U_{n}$ for $n\geq N$, each of diameter at most $r_{N}$. Hence
+Because $r_{n}\downarrow 0$, for every $N\geq 1$ the set $\limsup_{n\to\infty}U_{n}$ is covered by the intervals comprising $U_{n}$ for $n\geq N$, each of diameter at most $r_{N}$.
+Denoting by $\mathcal{H}^{s}_{\rho}(E):=\inf\bigl\{\sum_{i}(\operatorname{diam}E_{i})^{s}:E\subset\bigcup_{i}E_{i},\;\operatorname{diam}E_{i}\leq\rho\bigr\}$ the $s$-dimensional Hausdorff pre-measure at scale $\rho$ (see [14, Section 2.1]),
 $\mathcal{H}^{s}_{r_{N}}\!\left(\limsup_{n\to\infty}U_{n}\right)\leq\sum_{n\geq N}|T_{n}|\,r_{n}^{s}.$
@@
 Finally, $\mu_{G}$ is the push-forward of $\mu_{f}$ under the map $T(u):=e^{-u}$.
-Since $T$ is Lipschitz on $\log 2,\infty)$, the image of any set of Hausdorff dimension 0 again has Hausdorff dimension 0 (see, e.g., [17, Corollary 2.4]).
+Since every $Z_{\ell}\geq 0$, we have $Y\geq\log 2$ almost surely, so the carrying Borel set lies in $\log 2,\infty)$.
+Since $T$ is Lipschitz on $\log 2,\infty)$ (with $|T^{\prime}(u)|=e^{-u}\leq\tfrac{1}{2}$), the image of any set of Hausdorff dimension 0 again has Hausdorff dimension 0 (see, e.g., [14, Corollary 2.4]).
 Therefore $\dim_{H}(\mu_{G})=0$.
 ∎
-The vanishing Hausdorff dimension of the limiting measure explains the clustering visible in Figure 2: the purely singular law forces the values of $c(p)$ to concentrate near arithmetic anchor points of the form $\frac{1}{2}\prod_{r\in S}(1-1/r)$, determined by the small prime divisors of p1p-1, rather than spreading uniformly over $0,\frac{1}{2}]$.
-Figure 2: Empirical bin masses of $c(p)$ over the 78,49778{,}497 primes $3\leq p\leq 10^{6}$, computed via Proposition 8 with truncation depth $J=20$, using $100$ equal-width bins covering $[0,0.5]$.
+The vanishing Hausdorff dimension of the limiting measure explains why the values of $c(p)$ concentrate near arithmetic anchor points of the form $\frac{1}{2}\prod_{r\in S}(1-1/r)$, determined by the small prime divisors of p1p-1, rather than spreading uniformly over $0,\frac{1}{2}]$.
+Figure 2.
+Empirical bin masses of $c(p)$ over the 78,49778{,}497 primes $3\leq p\leq 10^{6}$, computed via Remark 7.9 with truncation depth $J=20$, using $100$ equal-width bins covering $[0,0.5]$.
 The dashed bars mark selected anchor values $\varphi(m)/m$ for small even squarefree $m$.
-The concentration of mass near these arithmetic anchors reflects the dependence of $c(p)$ on the small prime divisors of p1p-1 and illustrates the clustering predicted by the purely singular limiting law of Corollary 2.
-Theorem 6 (Moment Euler product).
+The concentration of mass near these arithmetic anchors reflects the dependence of $c(p)$ on the small prime divisors of p1p-1 and illustrates the clustering predicted by the purely singular limiting law of Corollary 5.14.
+6.2. Moments and Mellin transform
+Theorem 6.3 (Moment Euler product).
 For every real $k>0$,
-$\mathbb{E}[X^{k}]\;=\;2^{-k}\!\prod_{\begin{subarray}{c}\ell\geq 3\\ \ell\textup{ prime}\end{subarray}}\!\left(\frac{\ell-2}{\ell-1}+\frac{1}{\ell-1}\Bigl(\frac{\ell-1}{\ell}\Bigr)^{\!k}\right).$ (5)
+(6.1) $\mathbb{E}[X^{k}]\;=\;2^{-k}\!\prod_{\begin{subarray}{c}\ell\geq 3\\ \ell\textup{ prime}\end{subarray}}\!\left(\frac{\ell-2}{\ell-1}+\frac{1}{\ell-1}\Bigl(\frac{\ell-1}{\ell}\Bigr)^{\!k}\right).$
 The product converges absolutely, with the $\ell$-th factor equal to $1+O(k/\ell^{2})$.
 Proof.
-Since $Y=\log 2+\sum_{\ell}Z_{\ell}$ with the $Z_{\ell}\geq 0$ independent (Lemma 10), the partial sums $S_{L}:=\sum_{3\leq\ell\leq L}Z_{\ell}\nearrow S:=\sum_{\ell}Z_{\ell}$ give $e^{-kS_{L}}\searrow e^{-kS}$.
+Since $Y=\log 2+\sum_{\ell}Z_{\ell}$ with the $Z_{\ell}\geq 0$ independent (Lemma 6.1), the partial sums $S_{L}:=\sum_{3\leq\ell\leq L}Z_{\ell}\nearrow S:=\sum_{\ell}Z_{\ell}$ give $e^{-kS_{L}}\searrow e^{-kS}$.
 Since $0\leq e^{-kS_{L}}\leq 1$ for all $L$, by dominated convergence,
 $\mathbb{E}[e^{-kS}]=\prod_{\ell\geq 3}\mathbb{E}[e^{-kZ_{\ell}}]=\prod_{\ell\geq 3}\Bigl(\frac{\ell-2}{\ell-1}+\frac{1}{\ell-1}\Bigl(\frac{\ell-1}{\ell}\Bigr)^{\!k}\Bigr).$
-Since $\mathbb{E}[X^{k}]=2^{-k}\,\mathbb{E}[e^{-kS}]$, this gives (5).
+Since $\mathbb{E}[X^{k}]=2^{-k}\,\mathbb{E}[e^{-kS}]$, this gives (6.1).
 Writing the $\ell$-th factor as $1+a_{\ell}$ with $a_{\ell}=\bigl((\frac{\ell-1}{\ell})^{k}-1\bigr)/(\ell-1)=(e^{-kw_{\ell}}-1)/(\ell-1)$, where $w_{\ell}:=\log(\ell/(\ell-1))=1/\ell+O(1/\ell^{2})$.
@@
 ∎
-Corollary 5 (First and second moments).
-  1. (i)
-    $\mathbb{E}[X]=\frac{1}{2}\prod_{\begin{subarray}{c}\ell\geq 3\\ \ell\textup{ prime}\end{subarray}}\frac{\ell^{2}-\ell-1}{\ell(\ell-1)}.$
-  2. (ii)
-    $\mathbb{E}[X^{2}]=\frac{1}{4}\prod_{\begin{subarray}{c}\ell\geq 3\\ \ell\textup{ prime}\end{subarray}}\frac{\ell^{3}-\ell^{2}-2\ell+1}{\ell^{2}(\ell-1)}.$
-  3. (iii)
-    $\operatorname{Var}(X)=\mathbb{E}[X^{2}]-(\mathbb{E}[X])^{2}$.
+For example, substituting $k=1$ and $k=2$ into (6.1) gives
+$\mathbb{E}[X]=\frac{1}{2}\prod_{\begin{subarray}{c}\ell\geq 3\\ \ell\textup{ prime}\end{subarray}}\frac{\ell^{2}-\ell-1}{\ell(\ell-1)},\qquad\mathbb{E}[X^{2}]=\frac{1}{4}\prod_{\begin{subarray}{c}\ell\geq 3\\ \ell\textup{ prime}\end{subarray}}\frac{\ell^{3}-\ell^{2}-2\ell+1}{\ell^{2}(\ell-1)}.$
+Since $X$ is supported on the compact interval $0,\tfrac{1}{2}]$ (Theorem 5.5), the Weierstrass approximation theorem implies that the moment sequence $\{\mathbb{E}[X^{k}]\}_{k=1}^{\infty}$ determines $G$ uniquely.
+Theorem 6.4 (Analytic continuation).
+The function
+$M(s)\;:=\;2^{-s}\!\prod_{\begin{subarray}{c}\ell\geq 3\\ \ell\textup{ prime}\end{subarray}}F_{s}(\ell),\qquad F_{s}(\ell):=\frac{\ell-2}{\ell-1}+\frac{1}{\ell-1}\Bigl(\frac{\ell-1}{\ell}\Bigr)^{\!s},$
+is an entire function of $s$, with the Euler product converging uniformly on compact subsets of $\mathbb{C}$.
+Moreover, $M(s)=\mathbb{E}[X^{s}]$ for every $s\in\mathbb{C}$, and $M(s)\neq 0$ whenever $\operatorname{Re}(s)>0$.
 Proof.
-Immediate from (5) with $k=1$ and $k=2$. ∎
-Theorem 7 (Moment determinacy).
-The moment sequence $\{\mathbb{E}[X^{k}]\}_{k=1}^{\infty}$ uniquely determines $G$.
-Proof.
-Since $\operatorname{supp}(G)=[0,\tfrac{1}{2}]$ (Theorem 4), the random variable $X$ is supported on the compact interval $0,\tfrac{1}{2}]$.
-By the Weierstrass approximation theorem, polynomials are uniformly dense in $C([0,\tfrac{1}{2}])$, so the moment sequence $\{\mathbb{E}[X^{k}]\}_{k=1}^{\infty}$ determines the distribution of $X$ uniquely.
-∎
-Theorem 8 (Analytic continuation).
-The function
-$M(s)\;:=\;\mathbb{E}[X^{s}]\;=\;2^{-s}\!\prod_{\begin{subarray}{c}\ell\geq 3\\ \ell\textup{ prime}\end{subarray}}F_{s}(\ell),\qquad F_{s}(\ell):=\frac{\ell-2}{\ell-1}+\frac{1}{\ell-1}\Bigl(\frac{\ell-1}{\ell}\Bigr)^{\!s},$
-extends to an analytic non-vanishing function on $\{s\in\mathbb{C}:\operatorname{Re}(s)>0\}$, with the Euler product converging uniformly on compact subsets.
-Proof.
-We first establish the product formula for complex $s$ with $\sigma:=\operatorname{Re}(s)>0$.
-By Lemma 10, $Y=\log 2+\sum_{\ell\geq 3}Z_{\ell}$ with independent $Z_{\ell}\geq 0$, and the series converges almost surely since $\sum_{\ell}\mathbb{E}[Z_{\ell}]<\infty$.
-Set $S_{L}:=\sum_{3\leq\ell\leq L}Z_{\ell}$.
-Since $Z_{\ell}\geq 0$ and $\sigma>0$,
-$|e^{-sS_{L}}|=e^{-\sigma S_{L}}\leq 1$
+We first establish the product formula $\mathbb{E}[X^{s}]=M(s)$ for every $s\in\mathbb{C}$.
+Write $\sigma:=\operatorname{Re}(s)$.
+By Lemma 6.1, $Y=\log 2+\sum_{\ell\geq 3}Z_{\ell}$ with independent $Z_{\ell}\geq 0$, and the series converges almost surely since $\sum_{\ell}\mathbb{E}[Z_{\ell}]<\infty$.
+Set $S_{L}:=\sum_{3\leq\ell\leq L}Z_{\ell}$ and $S:=\sum_{\ell}Z_{\ell}$.
+Since $Z_{\ell}\geq 0$, the partial sums are monotone: $S_{L}\leq S$ for every $L$.
+Hence
+$|e^{-sS_{L}}|=e^{-\sigma S_{L}}\leq e^{|\sigma|S}$
 for every $L$.
-As $S_{L}\to S:=\sum_{\ell}Z_{\ell}$ a.s., dominated convergence gives $\mathbb{E}[e^{-sS_{L}}]\to\mathbb{E}[e^{-sS}]$.
+The dominator $e^{|\sigma|S}$ is integrable: independence gives $\mathbb{E}[e^{|\sigma|S}]=\prod_{\ell\geq 3}\mathbb{E}[e^{|\sigma|Z_{\ell}}]$, and each factor equals $\frac{\ell-2}{\ell-1}+\frac{1}{\ell-1}\bigl(\frac{\ell}{\ell-1}\bigr)^{|\sigma|}=1+O_{|\sigma|}(1/\ell^{2})$, so the product converges.
+As $S_{L}\to S$ a.s., dominated convergence gives $\mathbb{E}[e^{-sS_{L}}]\to\mathbb{E}[e^{-sS}]$.
 Independence factors the finite partial products: $\mathbb{E}[e^{-sS_{L}}]=\prod_{3\leq\ell\leq L}\mathbb{E}[e^{-sZ_{\ell}}]=\prod_{3\leq\ell\leq L}F_{s}(\ell)$.
-Therefore $\mathbb{E}[X^{s}]=2^{-s}\,\mathbb{E}[e^{-sS}]=2^{-s}\prod_{\ell\geq 3}F_{s}(\ell)$, extending the real-variable identity of Theorem 6 to the half-plane $\operatorname{Re}(s)>0$.
-It remains to verify analyticity and non-vanishing.
+Therefore $\mathbb{E}[X^{s}]=2^{-s}\,\mathbb{E}[e^{-sS}]=2^{-s}\prod_{\ell\geq 3}F_{s}(\ell)=M(s)$ for every $s\in\mathbb{C}$, extending the real-variable identity of Theorem 6.3.
+In particular, all negative moments $\mathbb{E}[X^{-\sigma}]$ ($\sigma>0$) are finite, a non-trivial consequence of the integrability of the dominator established above.
+It remains to verify that $M(s)$ is entire and non-vanishing on $\{\operatorname{Re}(s)>0\}$.
 Write $w_{\ell}:=\log(\ell/(\ell-1))$ and $F_{s}(\ell)=1+a_{\ell}(s)$ with $a_{\ell}(s)=(e^{-sw_{\ell}}-1)/(\ell-1)$.
 The bound $|e^{z}-1|\leq|z|e^{|z|}$ gives $|a_{\ell}(s)|\leq|s|\,w_{\ell}\,e^{|s|w_{\ell}}/(\ell-1)$.
-Since $w_{\ell}=1/\ell+O(1/\ell^{2})$, on any compact $K\subset\{\operatorname{Re}(s)>0\}$ with $|s|\leq R$ we get $|a_{\ell}(s)|\leq C_{K}/\ell^{2}$ for all large $\ell$.
+Since $w_{\ell}=1/\ell+O(1/\ell^{2})$, on any compact $K\subset\mathbb{C}$ with $|s|\leq R$ we get $|a_{\ell}(s)|\leq C_{R}/\ell^{2}$ for all large $\ell$.
 Since $\sum\ell^{-2}<\infty$, Weierstrass’s theorem on products of analytic functions (see e.g.
-[8, Ch.
-VII]) gives that the product converges uniformly on compact subsets to an analytic function.
-As $2^{-s}$ is entire, $M(s)$ is analytic on $\{\operatorname{Re}(s)>0\}$.
+[5, Ch.
+VII]) gives that the product converges uniformly on compact subsets of $\mathbb{C}$ to an analytic function.
+Each $F_{s}(\ell)$ is entire in $s$, and $2^{-s}$ is entire, so $M(s)$ is entire.
 For non-vanishing with $\sigma:=\operatorname{Re}(s)>0$: each factor $F_{s}(\ell)\neq 0$ since $|F_{s}(\ell)|\geq\frac{\ell-2}{\ell-1}-\frac{e^{-\sigma w_{\ell}}}{\ell-1}$ and $e^{-\sigma w_{\ell}}<1$: at $\ell=3$, $|F_{s}(3)|\geq(1-(2/3)^{\sigma})/2>0$; at $\ell\geq 5$, $|F_{s}(\ell)|\geq(\ell-3)/(\ell-1)>0$.
 To conclude that the infinite product is nonzero, fix a compact $K\subset\{\operatorname{Re}(s)>0\}$.
-Choose $L$ so that $|a_{\ell}(s)|<1/2$ for all $\ell\geq L$ and $s\in K$ (possible since $|a_{\ell}(s)|\leq C_{K}/\ell^{2}$).
-For $\ell\geq L$ the principal branch $\log(1+a_{\ell}(s))$ is analytic, and $|\!\log(1+a_{\ell}(s))|\leq 2|a_{\ell}(s)|\leq 2C_{K}/\ell^{2}$, so $\sum_{\ell\geq L}\log(1+a_{\ell}(s))$ converges uniformly on $K$ and the tail product equals $\exp\!\bigl(\sum_{\ell\geq L}\log(1+a_{\ell}(s))\bigr)\neq 0$.
+Choose $L$ so that $|a_{\ell}(s)|<1/2$ for all $\ell\geq L$ and $s\in K$ (possible since $|a_{\ell}(s)|\leq C_{R}/\ell^{2}$).
+For $\ell\geq L$ the principal branch $\log(1+a_{\ell}(s))$ is analytic, and $|\!\log(1+a_{\ell}(s))|\leq 2|a_{\ell}(s)|\leq 2C_{R}/\ell^{2}$, so $\sum_{\ell\geq L}\log(1+a_{\ell}(s))$ converges uniformly on $K$ and the tail product equals $\exp\!\bigl(\sum_{\ell\geq L}\log(1+a_{\ell}(s))\bigr)\neq 0$.
 The finitely many factors with $\ell<L$ are each nonzero, so $M(s)\neq 0$.
 ∎
-Proposition 5 (Squared modulus of the characteristic function).
+Remark 6.5 (Zeros on the imaginary axis).
+The non-vanishing assertion in Theorem 6.4 is sharp: $M(s)$ has zeros on the imaginary axis.
+Indeed, $F_{s}(3)=\tfrac{1}{2}+\tfrac{1}{2}(2/3)^{s}$ vanishes at $s=i\pi(2k+1)/\!\log(3/2)$ for every $k\in\mathbb{Z}$.
+For $\ell\geq 5$, one checks $|F_{it}(\ell)|\geq(\ell-3)/(\ell-1)>0$ for all $t\in\mathbb{R}$, so every zero of $M$ on the imaginary axis arises from the factor at $\ell=3$.
+6.3. Fourier structure
+Proposition 6.6 (Squared modulus of the characteristic function).
 For every $\tau\in\mathbb{R}$,
@@
 Proof.
-By Lemma 10, $|\widehat{\mu_{f}}(\tau)|^{2}=\prod_{\ell\geq 3}|\widehat{\nu_{\ell}}(\tau)|^{2}$.
+By Lemma 6.1, $|\widehat{\mu_{f}}(\tau)|^{2}=\prod_{\ell\geq 3}|\widehat{\nu_{\ell}}(\tau)|^{2}$.
 With $\alpha=\tau\log(\ell/(\ell-1))$ and $\widehat{\nu_{\ell}}(\tau)=\frac{\ell-2}{\ell-1}+\frac{e^{i\alpha}}{\ell-1}$,
 $|\widehat{\nu_{\ell}}(\tau)|^{2}=\frac{(\ell-2)^{2}+2(\ell-2)\cos\alpha+1}{(\ell-1)^{2}}=1-\frac{2(\ell-2)}{(\ell-1)^{2}}(1-\cos\alpha),$
-using $(\ell-2)^{2}+2(\ell-2)+1=(\ell-1)^{2}$. ∎
-Proposition 6 (Fourier non-decay implies singularity).
-If $\limsup_{|\tau|\to\infty}|\widehat{\mu_{f}}(\tau)|>0$, then $\mu_{f}$ is purely singular, and consequently $G$ is purely singular.
-Proof.
-By Lemma 6, $\mu_{f}$ is continuous and of pure type: either absolutely continuous or purely singular.
-The Riemann–Lebesgue lemma gives: if $\mu_{f}$ is absolutely continuous ($\mu_{f}=h\,\mathrm{d}t$, $h\in L^{1}$), then $\widehat{\mu_{f}}(\tau)=\widehat{h}(\tau)\to 0$.
-Contrapositively, $\limsup|\widehat{\mu_{f}}(\tau)|>0$ implies $\mu_{f}$ is not absolutely continuous, hence purely singular.
-For the conclusion about $G$: by Corollary 2, $X\overset{d}{=}e^{-Y}$ where $Y\sim\mu_{f}$, and $G$ is the distribution function of $X$.
-The map $T(u)=e^{-u}$ is locally Lipschitz on $0,\infty)$, so the push-forward $T_{\#}\mu_{f}$ preserves Lebesgue-null sets.
-Hence pure singularity of $\mu_{f}$ implies pure singularity of $G$ (cf.
-the proof of Corollary 2).
+using $(\ell-2)^{2}+2(\ell-2)+1=(\ell-1)^{2}$.
+Each factor lies in $[0,1]$, so the infinite product converges in $[0,1]$; the question of whether it tends to zero as $|\tau|\to\infty$ is addressed in Section 9.
 ∎
-Figure 3: Squared modulus $|\widehat{\mu_{f}}(\tau)|^{2}$ computed via the Euler product of Proposition 5 using all primes $\ell\leq 10^{7}$.
+Remark 6.7 (Fourier non-decay and singularity).
+If $\limsup_{|\tau|\to\infty}|\widehat{\mu_{f}}(\tau)|>0$, then $\mu_{f}$ is purely singular by the Riemann–Lebesgue lemma and the pure-type dichotomy (Theorem 5.9); the Lipschitz push-forward under $u\mapsto e^{-u}$ then gives singularity of $G$.
+Thus Fourier non-decay would provide an independent, Fourier-analytic proof of singularity.
+Remark 6.8 (Numerical evidence on Fourier decay).
+Numerical evaluation of $|\widehat{\mu_{f}}(\tau)|^{2}$ via the Euler product of Proposition 6.6, with the product truncated at all primes $\ell\leq 10^{7}$ (664,579664{,}579 factors), shows that the function oscillates, with local maxima near $0.14$ persisting at moderately large $\tau$.
+Convergence was verified by comparing partial products up to $10^{6}$, $5\times 10^{6}$, and $10^{7}$: the relative change from the second to the third cutoff is below $10^{-4}$ throughout the range tested.
+The arithmetic mean within each decade (over $10^{3}$ logarithmically spaced points in $1,10^{6}]$) decreases steadily: approximately $0.06$ for $\tau\in[10,10^{2}]$, declining to $0.008$ for $10^{5},10^{6}]$.
+This monotone decline is consistent with $|\widehat{\mu_{f}}(\tau)|^{2}\to 0$, which would mean $\mu_{f}$ is a Rajchman measure—a continuous measure whose Fourier transform vanishes at infinity (see Lyons [22]); establishing this rigorously remains open (Section 9).
+Figure 3.
+Squared modulus $|\widehat{\mu_{f}}(\tau)|^{2}$ computed via the Euler product of Proposition 6.6 using all primes $\ell\leq 10^{7}$.
 The top panel uses a linear scale and the bottom panel a log-log scale.
 The function oscillates but its local averages decrease across successive decades of $\tau$.
-Remark 2 (Numerical evidence on Fourier decay).
-We evaluated $|\widehat{\mu_{f}}(\tau)|^{2}$ via the Euler product of Proposition 5 with all primes $\ell\leq 10^{7}$.
-Figure 3 displays the result for $\tau\in[1,10^{6}]$.
-The function oscillates wildly, with local maxima near $0.14$ persisting at moderately large $\tau$.
-However, the mean value within each decade decreases steadily: approximately $0.06$ for $\tau\in[10,10^{2}]$, $0.03$ for $10^{2},10^{3}]$, $0.02$ for $10^{3},10^{4}]$, $0.01$ for $10^{4},10^{5}]$, and $0.008$ for $10^{5},10^{6}]$.
-A power-law fit to the envelope of local maxima yields $|\widehat{\mu_{f}}(\tau)|^{2}\approx\tau^{-0.19}$.
-This is consistent with decay to 0, which would mean $\mu_{f}$ is a Rajchman measure; establishing this rigorously remains open (Section 8).
-6.1 Endpoint asymptotics at $1/2$
-The value $1/2$ is the supremum of the support of $G$ but is never attained (Remark 1).
-The rate at which $G$ approaches its endpoint governs, in particular, the density of primes $p$ for which the rejection-sampling overhead $1/c(p)$ is close to its minimum value of $2$.
+6.4. Endpoint asymptotics at $1/2$
+The value $1/2$ is the right endpoint of the support of $G$ and carries zero mass; in particular, $c(p)<1/2$ for every prime $p$ (Remark 5.16), even though $\varphi(p-1)/(p-1)=1/2$ for Fermat primes.
+The rate at which $G$ approaches its endpoint governs, in particular, the density of primes $p$ for which the reciprocal $1/c(p)$ is close to its limiting infimum $2$.
 We now determine this rate precisely.
-Proposition 7 (Endpoint asymptotic at $1/2$).
-Let $G$ be the limiting distribution function from Theorem 4, and let $X$ be a random variable with distribution $G$. Define
+Proposition 6.9 (Endpoint asymptotic at $1/2$).
+Let $G$ be the limiting distribution function from Theorem 5.5, and let $X$ be a random variable with distribution $G$. Define
 $\kappa:=2e^{-\gamma}\prod_{\begin{subarray}{c}\ell\geq 3\\ \ell\textup{ prime}\end{subarray}}\Bigl(1-\frac{1}{(\ell-1)^{2}}\Bigr)\approx 0.7413.$
@@
 Proof.
-By Lemma 10,
+By Lemma 6.1,
 $X\overset{d}{=}\frac{1}{2}e^{-S},\qquad S:=\sum_{\begin{subarray}{c}\ell\geq 3\\ \ell\textup{ prime}\end{subarray}}Z_{\ell},$
@@
 $\mathbb{P}(Z_{\ell}=0)=\frac{\ell-2}{\ell-1},\qquad\mathbb{P}\!\left(Z_{\ell}=\log\frac{\ell}{\ell-1}\right)=\frac{1}{\ell-1}.$
-Set
-$M(y):=\prod_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\textup{ prime}\end{subarray}}\frac{\ell-2}{\ell-1}.$
-For $u>0$ one has
+Recall from Lemma 5.10 the notation $Q(y):=\prod_{3\leq\ell\leq y,\,\ell\textup{ prime}}(\ell-2)/(\ell-1)$. For $u>0$ one has
 $\frac{u}{1+u}<\log(1+u)\leq u.$
@@
 so the event $\{S\leq t\}$ forces $Z_{\ell}=0$ for every odd prime $\ell\leq(2t)^{-1}$. Therefore
-$\mathbb{P}(S\leq t)\leq M\!\bigl((2t)^{-1}\bigr).$ (6)
+(6.2) $\mathbb{P}(S\leq t)\leq Q\!\bigl((2t)^{-1}\bigr).$
 For the lower bound, set
@@
 By independence,
-$\mathbb{P}(S\leq t)\geq M(z_{t})\,\mathbb{P}\!\left(\sum_{\begin{subarray}{c}\ell>z_{t}\\ \ell\textup{ prime}\end{subarray}}Z_{\ell}\leq t\right).$ (7)
+(6.3) $\mathbb{P}(S\leq t)\geq Q(z_{t})\,\mathbb{P}\!\left(\sum_{\begin{subarray}{c}\ell>z_{t}\\ \ell\textup{ prime}\end{subarray}}Z_{\ell}\leq t\right).$
 Since $Z_{\ell}\leq 1/(\ell-1)$ almost surely,
@@
 $\mathbb{P}\!\left(\sum_{\begin{subarray}{c}\ell>z_{t}\\ \ell\textup{ prime}\end{subarray}}Z_{\ell}>t\right)\ll\frac{1}{\log(1/t)}=o(1),$
-and (7) becomes
-$\mathbb{P}(S\leq t)\geq M(z_{t})\,(1-o(1)).$ (8)
+and (6.3) becomes
+(6.4) $\mathbb{P}(S\leq t)\geq Q(z_{t})\,(1-o(1)).$
 Now
-$M(y)=\prod_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\textup{ prime}\end{subarray}}\left(1-\frac{1}{\ell}\right)\prod_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\textup{ prime}\end{subarray}}\left(1-\frac{1}{(\ell-1)^{2}}\right).$
-The second product converges to $\prod_{\ell\geq 3}(1-(\ell-1)^{-2})$, while Theorem 1 gives
+$Q(y)=\prod_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\textup{ prime}\end{subarray}}\left(1-\frac{1}{\ell}\right)\prod_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\textup{ prime}\end{subarray}}\left(1-\frac{1}{(\ell-1)^{2}}\right).$
+The second product converges to $\prod_{\ell\geq 3}(1-(\ell-1)^{-2})$, while Theorem 3.1 gives
 $\prod_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\textup{ prime}\end{subarray}}\left(1-\frac{1}{\ell}\right)\sim\frac{2e^{-\gamma}}{\log y}.$
 Hence
-$M(y)\sim\frac{\kappa}{\log y}\qquad(y\to\infty).$
-Since $\log((2t)^{-1})\sim\log z_{t}\sim\log(1/t)$ as $t\downarrow 0$, (6) and (8) imply
+$Q(y)\sim\frac{\kappa}{\log y}\qquad(y\to\infty).$
+Since $\log((2t)^{-1})\sim\log z_{t}\sim\log(1/t)$ as $t\downarrow 0$, (6.2) and (6.4) imply
 $\mathbb{P}(S\leq t)\sim\frac{\kappa}{\log(1/t)}\qquad(t\downarrow 0).$
-Since $\mu_{f}$ is continuous by Proposition 3, the law of $S=Y-\log 2$ is also continuous, so $\mathbb{P}(S<t)=\mathbb{P}(S\leq t)$. Therefore
+Since $\mu_{f}$ is continuous by Proposition 5.13, the law of $S=Y-\log 2$ is also continuous, so $\mathbb{P}(S<t)=\mathbb{P}(S\leq t)$. Therefore
 $1-G\!\left(\frac{1}{2}-\varepsilon\right)=\mathbb{P}\!\left(X>\frac{1}{2}-\varepsilon\right)=\mathbb{P}\!\left(S<-\log(1-2\varepsilon)\right).$
@@
 so $\log(1/(-\log(1-2\varepsilon)))\sim\log(1/\varepsilon)$. The stated asymptotic follows. ∎
-Remark 3 (Twin prime constant).
-The constant $\kappa$ equals $C_{2}\,e^{-\gamma}$, where $C_{2}=2\prod_{p\geq 3}(1-(p-1)^{-2})\approx 1.3203$ is the twin prime constant of Hardy and Littlewood [22] (denote $2C_{2}$ therein).
-This follows from the factorisation $(\ell-2)/(\ell-1)=(1-1/\ell)(1-(\ell-1)^{-2})$: the first factor produces the Mertens asymptotic $2e^{-\gamma}/\!\log y$, while the product of the second converges to $C_{2}/2$.
-Corollary 6 (Endpoint modulus at $1/2$).
+Corollary 6.10 (Twin-prime constant).
+The constant $\kappa$ equals $\mathfrak{S}_{2}\,e^{-\gamma}$, where
+$\mathfrak{S}_{2}\;:=\;2\prod_{p\geq 3}\Bigl(1-\frac{1}{(p-1)^{2}}\Bigr)\;\approx\;1.3203$
+is the singular series for twin primes in the sense of Hardy and Littlewood [15].
+Hardy and Littlewood write $2C_{2}$ for this quantity; we use $\mathfrak{S}_{2}$ to avoid collision with the convention that reserves $C_{2}\approx 0.6602$ for $\prod_{p\geq 3}p(p-2)/(p-1)^{2}$.
+Proof.
+The identity $\kappa=\mathfrak{S}_{2}\,e^{-\gamma}$ follows from the factorisation $(\ell-2)/(\ell-1)=(1-1/\ell)(1-(\ell-1)^{-2})$: the first factor produces the Mertens asymptotic $2e^{-\gamma}/\!\log y$, while the product of the second converges to $\mathfrak{S}_{2}/2$.
+∎
+Corollary 6.11 (Endpoint modulus at $1/2$).
 As $\varepsilon\downarrow 0$,
@@
 Proof.
-The two-sided order follows immediately from Proposition 7. If for some $\alpha>0$ one had
+The two-sided order follows immediately from Proposition 6.9. If for some $\alpha>0$ one had
 $G\!\left(\frac{1}{2}\right)-G\!\left(\frac{1}{2}-\varepsilon\right)=O(\varepsilon^{\alpha}),$
-then Proposition 7 would force
+then Proposition 6.9 would force
 $\frac{1}{\log(1/\varepsilon)}=O(\varepsilon^{\alpha}),$
@@
 $\varepsilon^{\alpha}\log(1/\varepsilon)\to 0\qquad(\varepsilon\downarrow 0).$
-7 Explicit Lower Bounds via the Number of Distinct Prime Factors
-Theorem 9 (Explicit lower bound on $c(p)$ via the number of distinct prime factors).
+7. Explicit Lower Bounds via the Number of Distinct Prime Factors
+The explicit bounds below are deterministic and hold for every prime $p\geq 3$; their strength depends on $\omega(p-1)$.
+By the shifted-prime analogue of the Erdős–Kac law [12] (proved in [16, Theorem 6]), $\omega(p-1)$ is typically of order $\log\log p$.
+Since the $K$-th prime satisfies $p_{K}\sim K\log K$ by the prime number theorem, Mertens’ third theorem gives $B(\omega(p-1))\asymp 1/\!\log\log\log p$ for a typical prime.
+The sharper scale $1/\!\log\log p$ arises in Corollary 7.2 and Corollary 7.3 from the primorial comparison $\omega(p-1)\leq m(p-1)$ together with $p_{m(p-1)}\sim\log p$.
+Actual divisibility of p1p-1 by all primes up to $p_{m(p-1)}$ occurs only in the Linnik/primorial-progression construction used for the matching upper bound and sharpness arguments (Proposition 4.2, Theorem 7.4).
+For primes with unusually many prime factors in p1p-1, the bound is correspondingly weaker.
+Theorem 7.1 (Explicit lower bound on $c(p)$ via the number of distinct prime factors).
 Let $p\geq 3$ be a prime, and let $\omega=\omega(p-1)$ denote the number of distinct prime factors of p1p-1. Then
-$c(p)\;\geq\;P_{3}\prod_{i=1}^{\omega}\Bigl(1-\frac{1}{p_{i}}\Bigr),\qquad P_{3}:=\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{3^{j}}\Bigr)\approx 0.5601260779,$
-where $p_{1}<p_{2}<\cdots<p_{\omega}$ are the first $\omega$ primes.
+$c(p)\geq P_{3}\prod_{i=1}^{\omega}\Bigl(1-\frac{1}{p_{i}}\Bigr),\>P_{3}:=\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{3^{j}}\Bigr)\approx 0.5601260779$
+(that is, $P_{3}=P(3)$), where $p_{1}<p_{2}<\cdots<p_{\omega}$ are the first $\omega$ primes.
 Proof.
@@
 $\frac{\varphi(p-1)}{p-1}=\prod_{i=1}^{\omega}\Bigl(1-\frac{1}{r_{i}}\Bigr)\geq\prod_{i=1}^{\omega}\Bigl(1-\frac{1}{p_{i}}\Bigr).$
-Since $P(t):=\prod_{j=1}^{\infty}(1-t^{-j})$ is increasing in $t>1$, we have $P(p)\geq P(3)=P_{3}$. Therefore, by (1),
+Since $P(t):=\prod_{j=1}^{\infty}(1-t^{-j})$ is increasing in $t>1$ (each factor $1-t^{-j}$ is positive and increasing for $t>1$, and the product converges), we have $P(p)\geq P(3)=P_{3}$.
+Therefore, by (2.1),
 $c(p)=\frac{\varphi(p-1)}{p-1}\,P(p)\geq P_{3}\prod_{i=1}^{\omega}\Bigl(1-\frac{1}{p_{i}}\Bigr),$
 as claimed. ∎
-Corollary 7 (Matching lower bound).
+Corollary 7.2 (Matching lower bound).
 There exists an absolute constant $c>0$ such that for all sufficiently large $x$,
@@
 Proof.
-The lower bound follows by the same primorial/Mertens argument used in the proof of Corollary 4: for every odd prime $p\leq x$, $\omega(p-1)\leq m(x)$, and Lemma 1 together with Mertens’ third theorem gives $\min_{p\leq x,\,p\textup{ prime}}c(p)\gg 1/\!\log\log x$.
-Theorem 9 sharpens the constant from $\tfrac{1}{2}$ to $P_{3}$ but does not affect the asymptotic order.
-The matching upper bound is Proposition 1.
-∎
-Corollary 8 (Pointwise universal lower bound).
-Let
+The prime $p=2$ satisfies $c(2)=\prod_{j\geq 1}(1-2^{-j})\approx 0.2888$, a fixed positive constant that exceeds $c/\!\log\log x$ for all large $x$.
+For every odd prime $p\leq x$, the primorial bound $N_{\omega(p-1)}\leq\prod_{r\mid(p-1)}r\leq p-1<x$ gives $\omega(p-1)\leq m(x)$.
+By Theorem 7.1 and the fact that each factor $(1-1/p_{i})\in(0,1)$,
+$c(p)\;\geq\;P_{3}\prod_{i=1}^{\omega(p-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr)\;\geq\;P_{3}\prod_{i=1}^{m(x)}\Bigl(1-\frac{1}{p_{i}}\Bigr).$
+Mertens’ third theorem and Lemma 4.1 ($p_{m(x)}\sim\log x$) now yield
+$\min_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}c(p)\;\geq\;P_{3}\prod_{i=1}^{m(x)}\Bigl(1-\frac{1}{p_{i}}\Bigr)\;\asymp\;\frac{1}{\log\log x}.$
+The matching upper bound is Proposition 4.2. ∎
+Corollary 7.3 (Pointwise universal lower bound).
+Recall the primorial-counting function from Section 4:
 $m(n):=\max\{k\geq 1:N_{k}\leq n\}.$
@@
 Proof.
-By Theorem 9,
+By Theorem 7.1,
 $c(p)\geq P_{3}\prod_{i=1}^{\omega(p-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr).$
@@
 which proves the first claim.
-Finally, by Mertens’ third theorem and Lemma 2,
+Finally, by Mertens’ third theorem and Lemma 4.1,
 $\prod_{i=1}^{m(p-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr)\asymp\frac{1}{\log p_{m(p-1)}}\asymp\frac{1}{\log\log p},$
 as $p\to\infty$ through the primes. This gives the second claim; equivalently, $1/c(p)\ll\log\log p$. ∎
-Remark 4 (On effectivity).
-The asymptotic $\min_{p\leq x}c(p)\asymp 1/\!\log\log x$ can be made fully effective by combining explicit Mertens estimates (Rosser–Schoenfeld [34, Theorems 4 and 7], Dusart [12, Theorem 5.9]) with the Linnik-type bound $L\leq 5.18$ of Xylouris [41, Theorem 1.1].
-Remark 5 (Sharper bounds for specific primes).
-Direct computation of the Euler product gives $P(5)\approx 0.7603$, $P(101)\approx 0.9900$, and $P(p)>1-10^{-9}$ for $p>2^{30}$.
-For the NIST moduli $p=3329$ and $p=8380417$ [31, 32], $\prod_{j\geq 1}(1-p^{-j})>0.9996$ and >0.9999998>0.9999998, respectively.
-Define $B(\omega):=P_{3}\prod_{i=1}^{\omega}(1-1/p_{i})$; by Theorem 9, $c(p)\geq B(\omega(p-1))$ for every prime $p\geq 3$. Table 1 lists representative values of $B(K)$ for small values of $K$.
+Theorem 7.4 (Sharp constant for the extremal order).
+$\limsup_{\begin{subarray}{c}p\to\infty\\ p\textup{ prime}\end{subarray}}\frac{1}{c(p)\,\log\log p}\;=\;e^{\gamma}.$
+In particular, $1/c(p)\leq(e^{\gamma}+o(1))\log\log p$ for every prime $p$, and this leading coefficient is best possible.
+Proof.
+Upper bound.
+Let $p\geq 3$ be prime, with $\omega:=\omega(p-1)$ distinct prime factors $r_{1}<\cdots<r_{\omega}$ of p1p-1.
+Since $t\mapsto t/(t-1)$ is strictly decreasing for $t>1$ and $r_{i}\geq p_{i}$ for each $i$,
+$\frac{p-1}{\varphi(p-1)}=\prod_{i=1}^{\omega}\frac{r_{i}}{r_{i}-1}\;\leq\;\prod_{i=1}^{\omega}\frac{p_{i}}{p_{i}-1}.$
+The primorial bound $N_{\omega}=\prod_{i=1}^{\omega}p_{i}\leq\prod_{r\mid(p-1)}r\leq p-1$ gives $\omega\leq m(p-1)$, so extending the product (each factor exceeds $1$):
+$\frac{p-1}{\varphi(p-1)}\;\leq\;\prod_{i=1}^{m(p-1)}\frac{p_{i}}{p_{i}-1}.$
+By Mertens’ third theorem (Theorem 3.1) and Lemma 4.1,
+$\prod_{i=1}^{m(p-1)}\frac{p_{i}}{p_{i}-1}\;\sim\;e^{\gamma}\log p_{m(p-1)}\;\sim\;e^{\gamma}\log\log p.$
+Since $1/P(p)=1+O(1/p)\to 1$, we obtain $1/c(p)\leq(e^{\gamma}+o(1))\log\log p$.
+Lower bound.
+For each $k\geq 1$, let $p^{*}(k)$ denote the least prime satisfying $p^{*}(k)\equiv 1\pmod{N_{k}}$, which exists by Dirichlet’s theorem.
+By Linnik’s theorem, $p^{*}(k)\leq C_{0}N_{k}^{\,L}$ for absolute constants $C_{0},L>0$.
+Since $N_{k}\mid(p^{*}(k)-1)$, every prime $p_{i}$ with $i\leq k$ divides $p^{*}(k)-1$, and additional prime factors of $p^{*}(k)-1$ beyond $p_{k}$ only increase the product. Hence
+$\frac{p^{*}(k)-1}{\varphi(p^{*}(k)-1)}\;\geq\;\prod_{i=1}^{k}\frac{p_{i}}{p_{i}-1}\;\sim\;e^{\gamma}\log p_{k}.$
+For the asymptotics of $\log\log p^{*}(k)$: by Linnik’s bound, $\log p^{*}(k)\leq\log C_{0}+L\log N_{k}\sim Lp_{k}$ (using $\log N_{k}=\vartheta(p_{k})\sim p_{k}$ from Lemma 4.1), so $\log\log p^{*}(k)\leq\log p_{k}+\log L+o(1)$.
+Since $p^{*}(k)>N_{k}$ and $\log N_{k}\sim p_{k}$, we also have $\log\log p^{*}(k)\geq\log p_{k}+o(1)$.
+Therefore $\log\log p^{*}(k)\sim\log p_{k}$.
+Combined with $P(p^{*}(k))\to 1$:
+$\frac{1}{c(p^{*}(k))\,\log\log p^{*}(k)}\;\geq\;\frac{(1+o(1))\,e^{\gamma}\log p_{k}}{(1+o(1))\,\log p_{k}}\;\to\;e^{\gamma}.$
+Together, the two bounds yield $\limsup=e^{\gamma}$. ∎
+Remark 7.5 (On effectivity).
+The asymptotic $\min_{p\leq x}c(p)\asymp 1/\!\log\log x$ can be made fully effective by combining explicit Mertens estimates (Rosser–Schoenfeld [29, Theorems 4 and 7], Dusart [9, Theorem 5.9]) with the Linnik-type bound $L\leq 5.18$ of Xylouris [37, Theorem 1.1], whose implied constant is effectively computable (though no numerical value has been published).
+Remark 7.6 (Convergence rate of the Euler product).
+Direct computation gives $P(5)\approx 0.7603$, $P(101)\approx 0.9900$, and $P(p)>1-10^{-9}$ for $p>2^{30}$; the correction factor $P(p)$ is thus negligible for all but the smallest primes.
+Define $B(\omega):=P_{3}\prod_{i=1}^{\omega}(1-1/p_{i})$; by Theorem 7.1, $c(p)\geq B(\omega(p-1))$ for every prime $p\geq 3$. Table 1 lists representative values of $B(K)$ for small values of $K$.
 $\omega(p-1)\leq K$ $B(K)$
@@
 2020 $\approx 0.0715828$
-Table 1: Representative values of the explicit lower bound $B(K)=P_{3}\prod_{i=1}^{K}(1-1/p_{i})$ appearing in Theorem 9. Thus every prime $p\geq 3$ with $\omega(p-1)\leq K$ satisfies $c(p)\geq B(K)$.
-Remark 6 (On the case $\omega(p-1)=1$).
-If $\omega(p-1)=1$ and $p$ is an odd prime, then p1p-1 is an even prime power, hence $p-1=2^{m}$ for some $m\geq 1$. Equivalently, $p$ is a Fermat prime. In this case
-$\frac{\varphi(p-1)}{p-1}=\frac{1}{2},\qquad c(p)=\frac{1}{2}\,P(p),\qquad P(p):=\prod_{j\geq 1}\Bigl(1-\frac{1}{p^{j}}\Bigr).$
-Hence $c(p)\to 1/2$ along any infinite subsequence of Fermat primes, so $B(1)$ is not tight for Fermat primes.
-Corollary 9 (Explicit lower bounds for large primes).
+Table 1.
+Representative values of the explicit lower bound $B(K)=P_{3}\prod_{i=1}^{K}(1-1/p_{i})$ appearing in Theorem 7.1.
+Thus every prime $p\geq 3$ with $\omega(p-1)\leq K$ satisfies $c(p)\geq B(K)$.
+Remark 7.7 (On the case $\omega(p-1)=1$).
+If $\omega(p-1)=1$ then $p$ is a Fermat prime, $\varphi(p-1)/(p-1)=1/2$, and $c(p)=P(p)/2$; since $P(p)>P(3)$ for every prime $p\geq 5$, each known Fermat prime $p\geq 5$ satisfies $c(p)>B(1)$, so the bound is not tight.
+Along any unbounded sequence of Fermat primes, should one exist, $P(p)\to 1$ and thus $c(p)\to 1/2$.
+Corollary 7.8 (Fully explicit coarse lower bound for large primes).
 Let $p$ be a prime with $p>2^{30}$. Then
@@
 Proof.
-The claim under $\omega(p-1)\leq 10$ follows from the bound $c(p)\geq B(\omega(p-1))$ (Theorem 9), since
+The claim under $\omega(p-1)\leq 10$ follows from the bound $c(p)\geq B(\omega(p-1))$ (Theorem 7.1), since
 $B(10)=P_{3}\prod_{i=1}^{10}\Bigl(1-\frac{1}{p_{i}}\Bigr)\approx 0.08847>0.088.$
@@
 $\frac{\varphi(p-1)}{p-1}=\prod_{r\mid(p-1)}\Bigl(1-\frac{1}{r}\Bigr)\;\geq\;\prod_{\begin{subarray}{c}r\leq p-1\\ r\text{ prime}\end{subarray}}\Bigl(1-\frac{1}{r}\Bigr).$
-By Dusart’s explicit product estimate [12, Theorem 5.9], which we use with threshold $x\geq 2\,278\,382$ (noting that $2\,278\,382<2^{22}\ll 2^{30}$, so the threshold is satisfied for all primes $p>2^{30}$),
-$\prod_{\begin{subarray}{c}r\leq x\\ r\text{ prime}\end{subarray}}\Bigl(1-\frac{1}{r}\Bigr)\;=\;\frac{e^{-\gamma}}{\log x}\,\Bigl(1+O^{*}\!\Bigl(\frac{0.2}{\log^{3}x}\Bigr)\Bigr).$
+By Dusart’s explicit product estimate [9, Theorem 5.9], which we use with threshold $x\geq 2\,278\,382$ (noting that $2\,278\,382<2^{22}\ll 2^{30}$, so the threshold is satisfied for all primes $p>2^{30}$),
+$\prod_{\begin{subarray}{c}r\leq x\\ r\text{ prime}\end{subarray}}\Bigl(1-\frac{1}{r}\Bigr)\;=\;\frac{e^{-\gamma}}{\log x}\,\Bigl(1+O^{*}\!\Bigl(\frac{0.2}{\log^{3}x}\Bigr)\Bigr),$
+where $O^{*}(\delta)$ denotes a quantity of absolute value at most $\delta$.
 Since $p>2^{30}$, we have p1>2 278 382p-1>2\,278\,382 and $\log(p-1)>20$, so the correction factor lies in $(1-2.5\times 10^{-5},\;1+2.5\times 10^{-5})$.
 Also,
-$\prod_{j\geq 1}\Bigl(1-\frac{1}{p^{j}}\Bigr)\geq 1-\sum_{j\geq 1}\frac{1}{p^{j}}=1-\frac{1}{p-1}>0.9999.$
+$\prod_{j\geq 1}\Bigl(1-\frac{1}{p^{j}}\Bigr)\geq 1-\sum_{j\geq 1}\frac{1}{p^{j}}=1-\frac{1}{p-1}>1-10^{-9}.$
 Therefore
-$c(p)=\frac{\varphi(p-1)}{p-1}\prod_{j\geq 1}\Bigl(1-\frac{1}{p^{j}}\Bigr)\;>\;\Bigl(1-\frac{0.2}{\log^{3}(p-1)}\Bigr)\cdot 0.9999\cdot\frac{e^{-\gamma}}{\log(p-1)}\;>\;\frac{0.56}{\log p},$
-using $e^{-\gamma}>0.5614$, $\log(p-1)<\log p$, and $0.2/\log^{3}(p-1)<2.5\times 10^{-5}$ for $p>2^{30}$. Equivalently, for $p>2^{30}$,
+$c(p)$ $=\frac{\varphi(p-1)}{p-1}\prod_{j\geq 1}\Bigl(1-\frac{1}{p^{j}}\Bigr)\;>\;\Bigl(1-\frac{0.2}{\log^{3}(p-1)}\Bigr)(1-10^{-9})\,\frac{e^{-\gamma}}{\log(p-1)}$
+$>\;\frac{0.56}{\log p},$
+using $e^{-\gamma}>0.5614$, $\log(p-1)<\log p$, $0.2/\log^{3}(p-1)<2.5\times 10^{-5}$, and $1/(p-1)<10^{-9}$ for $p>2^{30}$. Equivalently, for $p>2^{30}$,
 $\frac{1}{c(p)}\;<\;\frac{\log p}{0.56}\;<\;1.79\log p.$
-Remark 7.
-By the shifted-prime Erdős–Kac theorem (Barban [3]; see also Halberstam [20] and Elliott [13, Ch.
-12]), extending the classical result of Erdős and Kac [15], $(\omega(p-1)-\log\log x)/\sqrt{\log\log x}$ converges in distribution to the standard normal over primes $p\leq x$.
-For $p\approx 2^{128}$, $\log\log p\approx 4.5$, so $\omega(p-1)\leq 10$ holds for the vast majority of such primes.
-Standard $p$ generic $B(\omega(p-1))$ bound sharper factorization bound
-ML-KEM $3329$ $B(2)=P_{3}/3\approx 0.1867$ $\approx 0.4614$
-ML-DSA $8380417$ $B(4)=8P_{3}/35\approx 0.1280$ $\approx 0.2933$
-Table 2: The prime moduli appearing in FIPS 203 and FIPS 204 [31, 32].
-The third column gives the generic bound $B(\omega(p-1))$ from Theorem 9, depending only on the number of distinct prime factors of p1p-1, while the fourth gives the sharper lower bound obtained from the exact factorization of p1p-1 together with $P(p)\geq 1-\frac{1}{p-1}$.
-For the NIST-standardized moduli [31, 32], the exact factorizations $\varphi(3328)/3328=6/13$ and $\varphi(8380416)/8380416=100/341$ together with $P(p)\geq 1-1/(p-1)$ give $c(3329)\geq 0.4613$ and $c(8380417)\geq 0.2932$.
-The certified evaluation of Proposition 8 yields $c(3329)=0.461399\ldots$ and $c(8380417)=0.293255\ldots$, so the sharper factorization bounds agree with the exact values to $10$ and $21$ decimal places, respectively; the gap is $O(p^{-3})$, since replacing $P(p)$ by 11/(p1)1-1/(p-1) incurs an error of order $p^{-1}\cdot p^{-2}$ from the leading cross-term in the Euler product expansion.
-Context $p$ $\omega(p-1)$ $1/c(p)$
-ML-KEM (FIPS 203) $3329$ 22 $2.17$
-ML-DSA (FIPS 204) $8380417$ 44 $3.41$
-NewHope / FALCON [2, 19] $12289$ 22 $3.00$
-BFV/BGV FHE (small) [5, 18, 6] $786433$ 22 $3.00$
-BFV/BGV FHE (medium) [5, 18, 6] $2013265921$ 33 $3.75$
-NTTRU / BLISS [27, 11] $7681$ 33 $3.75$
-64-bit NTT prime [37] $2^{64}-2^{32}+1$ 66 $4.00$
-Fermat prime ($2^{16}+1$) $65537$ 11 $2.00$
-Table 3: Representative NTT-friendly primes from lattice-based cryptography and FHE implementations, together with $\omega(p-1)$ and the dimension-uniform expected rejection-sampling overhead $1/c(p)$.
-The displayed overheads are rounded to two decimal places; the exact values differ by less than $3\times 10^{-3}$.
-Remark 8 (Concrete example).
-For the 64-bit NTT prime $p=2^{64}-2^{32}+1$ (sometimes called the “Goldilocks prime”; not to be confused with the Ed448-Goldilocks prime [21]), used in lattice cryptography [37], one has $\omega(p-1)=6$, $\varphi(p-1)/(p-1)\approx 0.2500$, $P(p)>1-10^{-19}$, and $1/c(p)\approx 4$.
-Proposition 8 (Certified evaluation of $c(p)$).
-Given a prime $p\geq 3$ and the complete factorization $p-1=r_{1}^{e_{1}}\cdots r_{k}^{e_{k}}$, the following procedure computes $c(p)$ to $D$ correct decimal digits using $O(k+D\log 10/\!\log p)$ rational-arithmetic operations in the unit-cost rational-arithmetic model:
+The $1/c(p)<1.79\log p$ bound is explicit and valid uniformly for every prime $p>2^{30}$ without any factorisation data.
+The sharp asymptotic scale for the extremal order is the much slower $1/c(p)\leq(e^{\gamma}+o(1))\log\log p$ of Theorem 7.4.
+Remark 7.9 (Evaluation of $c(p)$).
+Given the factorisation of p1p-1, the value $c(p)$ can be computed to $D$ correct decimal places (absolute error < $10^{-D}$) using $O(k+D\log 10/\!\log p)$ rational-arithmetic operations, where $k=\omega(p-1)$ is the number of distinct prime factors of p1p-1, by computing the totient ratio $\varphi(p-1)/(p-1)$ exactly and truncating the Euler product $P(p)$ at depth $J=\lceil D\log 10/\!\log p\rceil$; writing $P_{J}(p):=\prod_{j=1}^{J}(1-p^{-j})$, the tail satisfies $0\leq P_{J}(p)-P(p)\leq p^{-J}/(p-1)$.
+8. Cryptographic Consequences
+The density $c(p)$ arises directly as a reduction constant in the PRIM-LWE variant of the learning with errors (LWE) problem [47], introduced by Sehrawat, Yeo, and Desmedt [31] to construct verifiable secret sharing schemes with “free” verification.
+The LWE-to-PRIM-LWE reduction proceeds by rejection-sampling: a uniformly random $n\times n$ matrix over $\mathbb{F}_{p}$ is accepted if and only if its determinant is a primitive root modulo $p$, and the acceptance probability in dimension $n$ is $c_{n}(p)$ (see Section 2).
+Since $c_{n}(p)$ is decreasing in $n$ with limit $c(p)$, the dimension-uniform expected rejection-sampling overhead is $1/c(p)$.
+This overhead represents a tightness loss in the security reduction: by the analysis of [31], the hardness guarantee for PRIM-LWE degrades by a factor of $c(p)$ relative to standard LWE, making tight bounds on $c(p)$ essential for concrete security estimates.
+Since $\inf_{p\textup{ prime}}c(p)=0$ (Theorem 3.3), the overhead is unbounded in general; the explicit bounds of Section 7 quantify it for any given prime.
+The results of the preceding sections translate immediately into cryptographic bounds. We collect the principal consequences.
+Corollary 8.1 (Compact overhead bound).
+Let $P_{3}$ be as in Theorem 7.1. If $p\geq 3$ is a prime modulus with $\omega(p-1)\leq K$, then the dimension-uniform expected overhead in the LWE-to-PRIM-LWE reduction satisfies
+$\frac{1}{c(p)}\;\leq\;\frac{1}{P_{3}}\prod_{i=1}^{K}\frac{p_{i}}{p_{i}-1}.$
+Proof.
+Immediate from Theorem 7.1 upon taking reciprocals. ∎
+For concrete deployment, it is useful to evaluate $c(p)$ exactly.
+The following proposition upgrades the evaluation procedure described in Remark 7.9 to a certified algorithm with explicit error control.
+Proposition 8.2 (Certified evaluation of $c(p)$).
+Given a prime $p\geq 3$ and the complete factorisation $p-1=r_{1}^{e_{1}}\cdots r_{k}^{e_{k}}$, the following procedure computes $c(p)$ to $D$ correct decimal digits using $O(k+D\log 10/\!\log p)$ rational-arithmetic operations in the unit-cost rational-arithmetic model:
   1. (i)
@@
 ∎
-8 Concluding Remarks
-This article studies the PRIM-LWE density constant $c(p)$ unconditionally from extremal, distributional, and computational viewpoints.
-We proved that $\inf_{p\text{ prime}}c(p)=0$ and that $\min_{p\leq x}c(p)\asymp 1/\log\log x$, so the worst-case rejection-sampling overhead grows only on a doubly logarithmic scale.
-We also showed that $c(p)$ admits a continuous, purely singular limiting distribution with support exactly $0,\tfrac{1}{2}]$ and Hausdorff dimension 0; in particular, the limiting measure is supported on a metrically negligible subset of $0,\tfrac{1}{2}]$ despite having full topological support.
-Finally, the explicit $\omega(p-1)$-based bounds and the certified evaluation algorithm of Proposition 8 make these asymptotic results effective for concrete cryptographic moduli.
-Several questions remain open. The squared modulus of the characteristic function admits an explicit Euler product (Proposition 5), but whether
+Proposition 8.2 and the exact factorisations of p1p-1 for the NIST-standardised moduli [45, 46] yield the bounds recorded in Table 2.
+The factorisations $\varphi(3328)/3328=6/13$ and $\varphi(8380416)/8380416=100/341$, together with $P(p)\geq 1-1/(p-1)$, give $c(3329)\geq 0.4613$ and $c(8380417)\geq 0.2932$.
+The certified evaluation yields $c(3329)=0.461399\ldots$ and $c(8380417)=0.293255\ldots$, so the sharper factorisation bounds agree with the exact values to $10$ and $21$ decimal places, respectively; the gap is $O(p^{-3})$, since replacing $P(p)$ by 11/(p1)1-1/(p-1) incurs an error of order $p^{-1}\cdot p^{-2}$ from the leading cross-term in the Euler product expansion.
+Standard $p$ generic $B(\omega(p-1))$ bound sharper factorisation bound
+ML-KEM (FIPS 203) $3329$ $B(2)=P_{3}/3\approx 0.1867$ $\approx 0.4614$
+ML-DSA (FIPS 204) $8380417$ $B(4)=8P_{3}/35\approx 0.1280$ $\approx 0.2933$
+Table 2.
+The prime moduli appearing in FIPS 203 and FIPS 204 [45, 46].
+The third column gives the generic bound $B(\omega(p-1))$ from Theorem 7.1, depending only on the number of distinct prime factors of p1p-1, while the fourth gives the sharper lower bound obtained from the exact factorisation of p1p-1 together with $P(p)\geq 1-1/(p-1)$.
+Table 3 collects the overhead $1/c(p)$ for representative NTT-friendly primes from lattice-based cryptography and fully homomorphic encryption implementations.
+Context $p$ $\omega(p-1)$ $1/c(p)$
+ML-KEM (FIPS 203) $3329$ 22 $2.17$
+ML-DSA (FIPS 204) $8380417$ 44 $3.41$
+NewHope / FALCON [38, 42] $12289$ 22 $3.00$
+BFV/BGV FHE (small) [40, 43, 39] $786433$ 22 $3.00$
+BFV/BGV FHE (medium) [40, 43, 39] $2013265921$ 33 $3.75$
+NTTRU / BLISS [44, 41] $7681$ 33 $3.75$
+64-bit NTT prime [48] $2^{64}-2^{32}+1$ 66 $4.00$
+Fermat prime ($2^{16}+1$) $65537$ 11 $2.00$
+Table 3.
+Representative NTT-friendly primes from lattice-based cryptography and FHE implementations, together with $\omega(p-1)$ and the dimension-uniform expected rejection-sampling overhead $1/c(p)$.
+The displayed overheads are rounded to two decimal places; the exact values differ by less than $3\times 10^{-3}$.
+Remark 8.3 (Typical overhead at cryptographic scales).
+By the shifted-prime Erdős–Kac theorem (Section 7), $\omega(p-1)$ is typically of order $\log\log p$.
+For $p\approx 2^{128}$, $\log\log p\approx 4.5$, so $\omega(p-1)\leq 10$ holds for the vast majority of such primes, and Table 1 gives $c(p)>0.088$.
+In practice, Table 3 shows $1/c(p)\leq 4$ for all surveyed moduli.
+By Corollary 7.8, the overhead satisfies $1/c(p)<1.79\log p$ uniformly for every prime $p>2^{30}$ without any factorisation data; the sharp asymptotic scale is $1/c(p)\leq(e^{\gamma}+o(1))\log\log p$ (Theorem 7.4), and the order $\min_{p\leq x}c(p)\asymp 1/\!\log\log x$ is tight (Corollary 7.2).
+Remark 8.4 (64-bit NTT prime).
+For the 64-bit NTT prime $p=2^{64}-2^{32}+1$ used in lattice cryptography [48], one has $\omega(p-1)=6$, $\varphi(p-1)/(p-1)\approx 0.2500$, $P(p)>1-10^{-19}$, and $1/c(p)\approx 4$.
+9. Concluding Remarks and Open Questions
+This article studies the primitive-root determinant density $c(p)$ from extremal, distributional, and metric viewpoints.
+We proved that $\inf_{p\text{ prime}}c(p)=0$ and that $\min_{p\leq x}c(p)\asymp 1/\log\log x$, so the extremal order of the reciprocal grows only on a doubly logarithmic scale, with sharp leading coefficient $e^{\gamma}$.
+We also showed that $c(p)$ admits a continuous, purely singular limiting distribution with support exactly $0,\tfrac{1}{2}]$ and Hausdorff dimension 0; in particular, the limiting measure is supported on a Lebesgue-null subset of $0,\tfrac{1}{2}]$ despite having full topological support.
+The explicit $\omega(p-1)$-based bounds make these asymptotic results effective.
+Section 8 translates the explicit bounds into concrete overhead estimates for the PRIM-LWE reduction in lattice-based cryptography.
+We conclude with two open questions.
+1. Fourier decay and the Rajchman property. The squared modulus of the characteristic function admits an explicit Euler product (Proposition 6.6), but whether
 $|\widehat{\mu_{f}}(\tau)|\to 0\qquad\text{or}\qquad\limsup_{|\tau|\to\infty}|\widehat{\mu_{f}}(\tau)|>0$
-remains open; by Proposition 6, the latter would give an independent Fourier-analytic proof of pure singularity.
-The numerical evidence of Remark 2 and Figure 3 suggests slow decay, but the oscillatory structure makes rigorous conclusions difficult.
-Determining a global modulus of continuity of $G$ is also open.
-Table 3 shows $1/c(p)\leq 4$ for all surveyed moduli; Corollary 8 gives $1/c(p)=O(\log\log p)$ pointwise.
-Closing the gap between the empirical bound $4$ and the asymptotic order $\log\log p$ remains an interesting open problem.
-Declaration of competing interests
-The author declares that there are no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
-Funding
-This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors.
-Research Data Statement
-No external datasets were used; all numerical values are reproducible from the explicit formulas and Monte Carlo procedures described in the paper.
-All empirical values of $c(p)$ in Figures 12 and Table 3 were computed by evaluating the totient ratio $\varphi(p-1)/(p-1)$ from the factorization of p1p-1, followed by the truncated Euler product $P_{J}(p)$ of Proposition 8 with $J=20$; the resulting error is below $10^{-9}$ for every prime $p\geq 3$.
-The Monte Carlo approximation of $G$ in Figure 1 samples the product distribution of Lemma 10 truncated at the 10,00010{,}000th odd prime $\ell_{N}$.
-Write $T_{N}:=\sum_{\ell>\ell_{N}}Z_{\ell}$ for the omitted tail and $S_{N}:=\sum_{3\leq\ell\leq\ell_{N}}Z_{\ell}$ for the retained partial sum, so that $X_{N}:=\tfrac{1}{2}e^{-S_{N}}$ approximates $X=\tfrac{1}{2}e^{-S}$ with $S=S_{N}+T_{N}$.
-The tail mean satisfies $\mathbb{E}[T_{N}]=\sum_{\ell>\ell_{N}}\frac{1}{\ell-1}\log\frac{\ell}{\ell-1}<10^{-5}$.
-Since $T_{N}\geq 0$ and $X_{N}\leq\tfrac{1}{2}$, the pointwise coupling bound $0\leq X_{N}-X=X_{N}(1-e^{-T_{N}})\leq T_{N}/2$ holds almost surely, giving $W_{1}(G_{N},G)\leq\tfrac{1}{2}\,\mathbb{E}[T_{N}]<5\times 10^{-6}$ for the Wasserstein-$1$ distance between the truncated and full distribution functions.
-Markov’s inequality further gives $\mathbb{P}(T_{N}>\varepsilon)<10^{-5}/\varepsilon$ for every $\varepsilon>0$.
-Figure 3 was produced by evaluating the closed-form Euler product of Proposition 5 over all primes $\ell\leq 10^{7}$.
-Tables 1 and 2 were computed from the explicit formula $B(K)=P_{3}\prod_{i=1}^{K}(1-1/p_{i})$ of Theorem 9 and the exact factorizations of p1p-1 for the listed moduli, respectively, using only rational arithmetic.
+remains open.
+If $\widehat{\mu_{f}}(\tau)\to 0$, the measure $\mu_{f}$ would be a Rajchman measure, placing the shifted-prime distribution of $\log(n/\varphi(n))$ in the same class as many classical singular measures arising from Bernoulli convolutions.
+A positive $\limsup$ would instead give an independent Fourier-analytic proof of pure singularity via the Riemann–Lebesgue lemma.
+Numerical evidence (Remark 6.8) suggests slow decay, but the oscillatory structure—driven by the mutual incommensurability of the values $\log(\ell/(\ell-1))$ across primes $\ell$—makes rigorous conclusions difficult.
+Standard techniques for Bernoulli convolutions (see, e.g., Peres, Schlag, and Solomyak [26]) do not apply directly, since the contraction ratios vary with $\ell$ and do not arise from a single algebraic parameter.
+2.
+Global modulus of continuity.
+Proposition 6.9 determines the behaviour of $G$ near its right endpoint: $1-G(1/2-\varepsilon)\sim\kappa/\!\log(1/\varepsilon)$, and Corollary 6.11 shows that $G$ is not Hölder of any positive order at $1/2$.
+Whether a logarithmic modulus of continuity holds globally—that is, whether there exists a constant $C>0$ such that $|G(\alpha)-G(\beta)|\leq C/\!\log(1/|\alpha-\beta|)$ for all $\alpha,\beta\in[0,\tfrac{1}{2}]$ with $|\alpha-\beta|$ small—remains open.
+For the analogous distribution of $\varphi(n)/n$ over the integers, the modulus of continuity is likewise unknown; see Schoenberg [30], Deshouillers–Hassani [7], and Toulmonde [35] for partial results in the integer case.
+Data availability statement
+All numerical computations are fully described by the explicit formulae in the paper and are reproducible from these descriptions alone—see Remark 5.15 for truncation and sampling parameters, and Remark 6.8 for the Fourier-decay evaluation.
+Python scripts reproducing Figures 13 and independently verifying Tables 13 are publicly available at https://github.com/Crypto-Vipin/primitive-root-ratio.
+No experimental data beyond the computations described in the text were used.